### v2

Abstract
We explore the relationship between two popular concepts in the allocation of divisible items: competitive equilibrium (CE) and allocations that maximize Nash welfare, i.e., allocations where the weighted geometric mean of the utilities is maximal. When agents have homogeneous concave utility functions, these two concepts coincide: the classical Eisenberg–Gale convex program that maximizes Nash welfare over feasible allocations yields a competitive equilibrium. However, these two concepts diverge for non–homogeneous utilities. From a computational perspective, maximizing Nash welfare amounts to solving a convex program for any concave utility functions, whereas computing CE becomes PPAD-hard already for separable piecewise linear concave (SPLC) utilities.
We introduce the concept of Gale-substitute utility functions, which is an analogue of the weak gross substitutes (WGS) property for the so-called Gale demand system. For Gale-substitutes utilities, we show that any allocation maximizing Nash welfare provides an approximate-CE with surprisingly strong guarantees, where every agent gets at least half the maximum utility they can get at any CE, and is approximately envy-free. Gale-substitutes include utility functions where computing CE is PPAD hard, such as all separable concave utilities and the previously studied non-separable class of Leontief-free utilities. We introduce a broad new class of utility functions called generalized network utilities based on the generalized flow model. This class includes SPLC and Leontief-free utilities, and we show that all such utilities are Gale-substitutes.
Conversely, although some agents may get much higher utility at a Nash welfare maximizing allocation than at a CE, we show a price of anarchy type result: for general concave utilities, every CE achieves at least $(1/\mathrm{e})^{1/\mathrm{e}}>0.69$ fraction of the maximum Nash welfare, and this factor is tight.
1 Introduction
Competitive (market) equilibrium is a central solution concept in economics to analyze and predict the outcomes of market interactions. The Fisher model is a fundamental market model where a set of $n$ agents ${\mathcal{A}}$ arrive at the market with a fixed supply of divisible goods ${\mathcal{M}}$. Each agent $i$ has a fixed budget $b_{i}$ and preferences over bundles of goods given by a utility function $u_{i}(.)$. At competitive equilibrium (CE), prices $\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ of goods and fractional allocation $\mathbf{x}\in\mathbb{R}_{\geq 0}^{{\mathcal{A}}\times{\mathcal{M}}}$ of goods to agents are such that each agent $i$ gets an optimal bundle of goods, i.e., $x_{i}$ maximizes $u_{i}(x_{i})$ subject to $\langle\mathbf{p},x_{i}\rangle\leq b_{i}$, and the market clears, i.e., the demand for each good meets its supply.
Equilibrium allocations satisfy many desirable properties, including Pareto optimality and the fairness guarantee of (weighted) envy-freeness. These properties make this an attractive concept also in other settings where no direct trade and money are involved. For example, competitive equilibrium with equal incomes (CEEI) is a popular mechanism in fair division, where a fictitious market is created by giving each agent the same budget Varian (1974). We note that CEEI is a special case of the Fisher model, where $b_{i}=1$ for each agent $i$.
It is standard to assume that the utility functions $u_{i}(.)$ are monotone non-decreasing, concave, and satisfy $u_{i}(\mathbf{0})=0$. Classical results in mathematical economics established the existence of an equilibrium in the Fisher model as well as more general exchange market models, by Arrow and Debreu (1954) and by McKenzie (1954). These results are based on fixed point theorems and thus do not lead to efficient computational methods. Finding an equilibrium by efficient computation has been an important topic in economics, optimization, and computer science over the past decades.
Eisenberg and Gale (1959) gave a convex programming formulation for CE with linear utility functions $u_{i}(x)=\sum_{j\in{\mathcal{M}}}v_{ij}x_{j}$. This program amounts to maximizing the weighted Nash welfare, i.e., the weighted geometric mean of agents’ utilities over feasible allocations. That is, any allocation maximizing $\prod_{i}u_{i}(x_{i})^{b_{i}}$ (equivalently, $\sum_{i}b_{i}\log{u_{i}(x_{i})}$) over feasible allocations gives a CE, where the dual solution yields the corresponding prices $\mathbf{p}$. Eisenberg (1961) showed that the same convex program yields competitive equilibria more generally for concave utility functions that are positively homogeneous of degree one, i.e., $u_{i}(\alpha x_{i})=\alpha u_{i}(x_{i})$ for any $\alpha>0$. This includes classical examples of constant elasticity of substitution (CES), Cobb–Douglas, and Leontief utility functions.
The Nash welfare takes its name due to its connection with the bargaining problem introduced by Nash (1950). Nash studied general cooperative bargaining problems where the set of possible outcomes is a convex set $K\subseteq\mathbb{R}^{n}$ with a disagreement point $d\in K$ and showed that maximizing the geometric mean of the excess utilities above the disagreement utility amounts to the only possible bargaining outcome that satisfies certain natural requirements. Taking $K$ as the set of feasible allocations and $d=\mathbf{0}$, the Nash bargaining solution is precisely the Nash welfare maximizing allocation in this setting Vazirani (2012).
Weak gross substitutes (WGS) utility functions are another broad class where a CE can be computed in polynomial time. Several fast algorithms have been designed for computing equilibria with WGS utilities Bei et al. (2019a), Codenotti et al. (2005b, c), Garg et al. (2023a). While the WGS class includes many homogeneous utilities, there are also non-homogeneous examples for which the Eisenberg–Gale solution may not correspond to a CE.
WGS and homogeneous utility functions cover most known cases where a CE can be computed in polynomial time. Already for the seemingly simple case of separable piecewise concave utilities, computing a CE is PPAD-hard Chen and Teng (2009). In fact, it is PPAD-hard even for bounded linear utilities, when there is an upper bound on the utility from each good, i.e., $u_{i}(x)=\sum_{j\in{\mathcal{M}}}\min\{v_{ij}x_{j},\ell_{ij}\}$ (Bei et al., 2019b). Example A.1 shows that the set of CE can be non-convex and disconnected already for bounded linear utilities.
1.1 Our contributions
In this paper, we study the relationship between two fundamental concepts in the allocation of divisible goods: competitive equilibrium and maximum Nash welfare allocations. We exhibit a close relationship between these solution concepts well beyond the homogeneous case. First, we show a ‘price of anarchy’ type result Dubey (1986), Koutsoupias and Papadimitriou (1999). One can interpret Nash welfare as a global measure of social utility. In contrast, a CE can be seen as an outcome reached by autonomous agents trying to improve their utilities individually. It turns out that, in the general setting of concave utilities, any CE recovers at least $(1/\mathrm{e})^{1/\mathrm{e}}>0.69$ fraction of the maximum Nash welfare, and this factor is tight (see Example A.4).
Theorem 1.1.
For Fisher markets under concave utility functions, the Nash welfare at any CE is at least $(1/\mathrm{e})^{1/\mathrm{e}}$ times the maximum Nash welfare.
We then study guarantees of the maximum Nash welfare solution in the context of CE. A major advantage of maximum Nash welfare is efficient computability: it can be arbitrarily well approximated for any concave utility functions. In contrast, computing CE becomes PPAD-hard already in very simple cases. In this context, our first simple result shows that the maximum Nash welfare yields an approximate CE for all concave utility functions. By an $\alpha$-demand-approximate CE, we mean allocations and prices such that every agent receives at least $1/\alpha$-fraction of the maximum possible utility (see Definition 2.8). As in the case of the Eisenberg–Gale program, one can interpret the optimal Lagrangian multipliers as prices. The result also implies approximate envy-freeness.
Theorem 1.2.
For Fisher markets under concave utility functions, any Nash welfare maximizing allocation $\mathbf{y}$, together with the dual prices $\mathbf{q}$, form a 2-demand-approximate CE. This also implies that $\mathbf{y}$ is half envy-free when $b_{i}$’s are identical, i.e., $u_{i}(y_{i})\geq\tfrac{1}{2}u_{i}(y_{k})$, for all $i,k\in{\mathcal{A}}$.
Theorem 1.2 already shows that maximum Nash welfare provides remarkable fairness and efficiency guarantees for all concave utility functions. However, there can be multiple equilibria, giving vastly different utilities to an agent. In Example A.5, we show that there could be a gap of $\Omega(n)$ in the utility of some agent between two different competitive equilibria. Hence, an agent could be much better off in some CE than in the maximum Nash welfare solution.
As the main contribution of the paper, we show that maximum Nash welfare provides surprisingly strong guarantees for a broad class of utility functions. For this, we explore the Gale demand correspondence Garg et al. (2023a), Nesterov and Shikhman (2018). For an agent $i\in{\mathcal{A}}$ with budget $b_{i}$ and utility function $u_{i}(.)$ at prices $\mathbf{q}$, the Gale demand correspondence comprises of the allocations $y_{i}$ that maximize $b_{i}\log u_{i}(y_{i})-\langle\mathbf{q},y_{i}\rangle$. It follows from Lagrangian duality that in a Nash welfare maximizing solution, every agent receives an allocation from their Gale demand correspondence.
The reason behind the equivalence of CE and maximum Nash welfare for homogeneous utilities is that, for this class, the standard demand correspondence and the Gale demand correspondence coincide. Nevertheless, these two demands can behave very differently for non-homogeneous utility functions, e.g., the Gale demand can be non-monotonic in the prices, whereas the standard demand is always monotonic (see Example A.3).
We identify a critical property of the Gale demand correspondence that we call Gale-substitutes property (Definition 2.14). This is analogous to the weak gross substitutes (WGS) property of the standard demand correspondence. In particular, it asserts that if some prices decrease then the demand for goods with unchanged prices may only decrease.
Competitive equilibrium becomes more subtle when the utility functions are satiable: there is an absolute maximum value of the utility that can be reached, regardless of the prices—bounded linear utilities mentioned above are such an example. To accommodate satiable utilities, we also introduce a stronger variant of the definition, called the $\Sigma$-Gale-substitutes property. We note that for non-satiable utilities, the Gale-substitutes and $\Sigma$-Gale-substitutes properties coincide. Figure 1 illustrates the containment relationships among various utility classes discussed in this paper, with those in which the Fisher market equilibrium can be efficiently computed highlighted in grey.
Our main result shows that for this class of utility functions, the utility of any agent at the Nash welfare maximizing allocation is at least half the maximum utility they may get at any CE.
Theorem 1.3.
Consider a Fisher market instance where all utility functions are $\Sigma$-Gale-substitutes. Then, for any competitive equilibrium $(\mathbf{x},\mathbf{p})$ and any Nash welfare maximizing allocation $\mathbf{y}$, $u_{i}(y_{i})\geq\frac{1}{2}u_{i}(x_{i})$ holds for every $i\in{\mathcal{A}}$.
As a counterpart to the ‘price of anarchy’ result in Theorem 1.1, one could call this a ‘price of authority’ result. The agents can either accept an allocation proposed by a central authority on the principle of maximizing Nash welfare or converge to a market solution (using some dynamics). By accepting the centrally proposed solution, each agent may be worse off by at most a factor 2 compared to the best possible equilibrium outcome. On the other hand, outcomes at different competitive equilibria can be very different, and the possible outcomes can be difficult to predict given the computational hardness of equilibria. Hence, for ($\Sigma$-)Gale-substitutes utility functions, the Nash welfare maximizing allocation provides a guaranteed good outcome to all agents simultaneously while also being efficiently computable. Thus, Theorem 1.3 establishes a strong relationship between CE and maximum Nash welfare beyond homogeneous functions and provides a natural alternative to the equilibrium selection problem Arrow and Hahn (1971) in this setting where multiple equilibria are possible.
We call a utility function $u(.)$ separable if $u(x)=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$. This class gives an important example of $\Sigma$-Gale-substitutes. Note that such utilities, in general, are neither homogeneous nor WGS. As mentioned above, competitive equilibrium computation is already PPAD-hard for the separable piecewise linear concave (SPLC) case when each $v_{j}$ is piecewise linear. This class has attracted significant attention on its own Anari et al. (2018), Chen et al. (2009, 2013), Garg et al. (2015), Vazirani and Yannakakis (2011).
Theorem 1.4.
Every separable utility function is $\Sigma$-Gale-substitutes.
We consider two further examples of piecewise linear Gale-substitutes utilities:
  • Leontief-free (LF) utilities Filos-Ratsikas et al. (2024), Garg et al. (2014), Goyal et al. (2023): Consider a set $K$ of segments, and let $u(x)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\max\sum_{k\in K}u_{k}$: $u_{k}=\sum_{j\in{\mathcal{M}}}a_{jk}z_{jk},\ \ u_{k}\leq\ell_{k};\ \forall k\in K$; $\sum_{k\in K}z_{jk}\leq x_{j}\ \forall j\in{\mathcal{M}}$; $\mathbf{z}\geq 0$. That is, an LF function is given by a set of linear functions $u_{k}$’s with a maximum utility limit $\ell_{k}$ on each $k\in K$. LF captures interesting non-separable PLC functions, and includes SPLC as a special case where each segment $k\in K$ has exactly one $j$ with $a_{jk}>0$.
  • Generalized network utilities: We introduce a new and broad class of utility functions defined by generalized flow instances. Here, the goods correspond to source nodes in a flow network with supply $x_{j}$, $j\in{\mathcal{M}}$. Each arc has a capacity and a gain factor, and the utility $u(x)$ is the maximum flow value that can reach a designated sink node $t$ from supplies $x$. The precise definitions are given in Section 2.1. LF can be formulated as a special case of a network with two layers, the middle layer between the sources and the sink corresponding to the segments $K$.
We show the following:
Theorem 1.5.
Generalized network utility functions are Gale-substitutes.
Generalized flows is a classical network flow model that can model commodity transportations with losses, as well as conversions between different types of entities such as in financial networks; see Chapter 15 in Ahuja et al. (1993) for a description and various applications.
Theorem 1.3 already shows that in Fisher markets with non-satiable generalized network utilities, Nash welfare maximizing allocations give a good approximation to utilities at competitive equilibrium. However, satiable generalized network utilities may not satisfy the stronger $\Sigma$-Gale-substitutes property required in Theorem 1.3. Example A.10 shows that the theorem does not hold if only assuming the Gale substitutes property. For satiable utility functions, it is natural to look at thrifty (strong) equilibria, which are guaranteed to exist Garg et al. (2022b). In such equilibria, each agent receives an optimal bundle at the cheapest possible price (see Definition 2.11). An appealing property of this concept is that thrifty equilibria are Pareto efficient, whereas this may not hold for (weak) competitive equilibrium. We also show that the statement of Theorem 1.3 remains true for generalized network utilities and for thrifty equilibria.
Theorem 1.6.
Consider a Fisher market instance where all utility functions are generalized network utilities. Then, for any thrifty competitive equilibrium $(\mathbf{x},\mathbf{p})$ and any Nash welfare maximizing allocation $\mathbf{y}$, $u_{i}(y_{i})\geq\frac{1}{2}u_{i}(x_{i})$ holds for every $i\in{\mathcal{A}}$.
Finally, we note that the statement of Theorem 1.3 does not hold for arbitrary concave utilities. In Example A.8, we show that without the Gale-substitutes property, agents may get much higher utilities in CE than in the maximum Nash welfare allocations.
Remark 1.7.
We note that no polynomial time approximation scheme (PTAS) can exist that, for some $\varepsilon>0$, provides any agent at least $(1-\varepsilon)$ times their maximum CE utility. This is due to the fact that, as shown in Example A.1, multiple thrifty equilibria may exist, each providing different utilities to the agents. The existence of an allocation that $\varepsilon$-approximates the best CE utility for each agent would, by compactness, imply the existence of an allocation where every agent achieves their best CE utility. However, this would contradict the fact that thrifty equilibria are Pareto efficient.
Figure 1: The graph illustrates the containment relationships among different utility classes, with the classes where the Fisher market equilibrium can be computed efficiently highlighted in grey.
1.2 Related work
This paper strengthens the relationship between the two fundamental concepts of competitive equilibrium and maximum Nash welfare. Both are intensely studied concepts with a variety of applications, so we briefly mention further related results. As mentioned earlier, computing equilibria in Fisher markets is polynomial time for homogeneous and WGS utility functions, and beyond that, the problem is essentially PPAD-hard. For SPLC utilities, achieving a $\frac{1}{11}$-approximate-CE is PPAD-hard Deligkas et al. (2024). We note that Chen and Teng (2009), Deligkas et al. (2024), Vazirani and Yannakakis (2011) use a different definition of approximate-CE, where every agent gets an optimal bundle, but the market clears approximately. Our approximate-CE definition also appears in several previous works, e.g., in Codenotti et al. (2005b, c), Codenotti and Varadarajan (2007).
A variant of the Gale substitutes property also appeared in Garg et al. (2022a) [^1] The technical report version of Garg et al. (2023a).. This was used to obtain an auction algorithm for the maximum Nash welfare problem with capped SPLC utilities, and the proof uses arguments similar to Section 7. Similar auction algorithms could be developed more generally for Gale substitutes functions.
Fisher markets are a special case of more general Arrow-Debreu (exchange) markets. Computing equilibria in exchange markets is polynomial time for WGS utility functions (see, e.g., Bei et al. (2019a), Codenotti et al. (2005c), Garg et al. (2023a)) and for CES utilities with parameter $-1\leq\rho<0$ Codenotti et al. (2005a), and beyond that the problem is essentially PPAD-hard Chen et al. (2009), even for homogeneous functions, such as Leontief and CES with parameter $\rho<-1$ Chen et al. (2013), Codenotti et al. (2006).
The maximum Nash welfare is a classic objective for allocating goods to agents. It has received significant attention in the literature on social choice and fair division; see, e.g., Moulin (2003). It provides remarkable fairness guarantees also for indivisible goods: under additive valuations, maximum Nash welfare allocations attain the “unreasonable” fairness guarantee of envy-freeness up to one good (EF1) Caragiannis et al. (2019). Furthermore, for more general subadditive valuations, the fairness guarantee of $\tfrac{1}{2}$-approximate envy-freeness up to any good (EFX) can be achieved at a $\tfrac{2}{3}$-approximate maximum Nash welfare allocation Feldman et al. (2024).
Another interesting line of research proposes Nash-bargaining-based solutions for allocation problems, offering efficient computation and other desirable structural properties; see, e.g., Vazirani (2012), Hosseini and Vazirani (2022), Garg et al. (2023b). Recently, Tröbst and Vazirani (2024) showed that computing a Pareto-efficient and envy-free allocation for matching utilities is PPAD-hard, and that the NSW solution gives a PO an $2$-envy-free allocation. This latter result can be viewed as a special case of our Theorem 1.2.
Overview
The remainder of the paper is organized as follows. Section 2 defines all models and definitions. Section 3 provides a brief overview of the techniques used. Section 4 proves Theorem 1.1, showing the tight bound of $(1/\mathrm{e})^{1/\mathrm{e}}$ on the efficiency of competitive equilibria. Section 5 presents several properties of the Gale demand correspondence and proves Theorem 1.2, showing that Nash welfare maximizing allocations yield $2$-demand approximate-CE. Section 6 proves Theorem 1.3, showing the stronger guarantees of the maximum Nash welfare allocations for Gale-substitutes utility functions. Section 7 proves Theorem 1.4, showing that separable utilities are $\Sigma$-Gale substitutes. Section 8 considers generalized network utility functions and proves Theorems 1.5 and 1.6. Several examples are deferred to the Appendix, and Appendix B presents an omitted proof from Section 8.
2 Preliminaries
Notation
For a function $f\,:\,X\to Y$, let $\operatorname{dom}(f)\subseteq X$ denote its domain. For $x\in\mathbb{R}^{n}$, let $\operatorname{supp}(x)=\{i\,:\,x_{i}\neq 0\}$ denote its support. For $x,y\in\mathbb{R}^{n}$, we use $\left\langle x,y\right\rangle=x^{\top}y$ for the standard scalar product.
Subgradients and supergradients
The functions we consider in this paper are not always differentiable. We will use sub- and supergradients, defined as follows.
Definition 2.1.
Given a function $f\,:\,\mathbb{R}^{n}\to\mathbb{R}$, a vector $\mathbf{g}\in\mathbb{R}^{n}$ is a supergradient of $f$ at the point $\mathbf{x}\in\operatorname{dom}(f)$ if for all $\mathbf{y}\in\operatorname{dom}(f)$, $f(\mathbf{x})+\left\langle\mathbf{g},\mathbf{y}-\mathbf{x}\right\rangle\geq f(\mathbf{y})$. Similarly, $\mathbf{g}\in\mathbb{R}^{n}$ is a subgradient of $f$ at $\mathbf{x}\in\operatorname{dom}(f)$ if $f(\mathbf{x})+\left\langle\mathbf{g},\mathbf{y}-\mathbf{x}\right\rangle\leq f(\mathbf{y})$ for all $\mathbf{y}\in\operatorname{dom}(f)$. We denote the set of all supergradients of $f$ at $\mathbf{x}$ as $\partial^{\wedge}{f}(\mathbf{x})$, set of all subgradients as $\partial^{\vee}{f}(\mathbf{x})$.
Thus, $\mathbf{g}$ if a supergradient of $f$ if and only if $-\mathbf{g}$ is a subgradient of $-f$. We recall that a function $f$ is convex if and only if $\partial^{\vee}{f}(\mathbf{x})\neq\emptyset$ for every $\mathbf{x}\in\operatorname{dom}(f)$ and concave if and only if $\partial^{\wedge}{f}(\mathbf{x})\neq\emptyset$ for every $\mathbf{x}\in\operatorname{dom}(f)$. For a differentiable convex (concave) function, the unique subgradient (supergradient) at every point is the gradient $\nabla f(\mathbf{x})$.
2.1 Utility functions
Given a set ${\mathcal{M}}$ of divisible goods, we will focus on utility functions with the following properties.
Definition 2.2.
A function $u(\cdot)\,,\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}$ is a utility function if $u(\mathbf{0})=0$, $u(\cdot)$ is concave, monotone non-decreasing, and there exists an $x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $u(x)>0$.
We say that a utility function is separable if $u(x)=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$. We will also distinguish between satiable and non-satiatiable utilities.
Definition 2.3.
A utility function $u(\cdot)$ is non-satiable if and only if, for any $x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, one can find $x^{\prime}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, such that $u(x^{\prime})>u(x)$. Otherwise, we say the utility function is satiable.
Demand correspondence and Gale demand correspondence
By a price vector, we mean a vector $\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$. Given prices $\mathbf{p}$ and a budget $b\in\mathbb{R}_{\geq 0}$, we define the demand correspondence as the set of maximum utility bundles achievable at $\mathbf{p}$ and $b$:
${\mathcal{D}}^{u}(\mathbf{p},b)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\arg\max_{x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\left\{u(x)\,:\,\left\langle\mathbf{p},x\right\rangle\leq b\right\}\,.$ (1)
Note that if $u$ is a non-satiable utility function, then $\left\langle\mathbf{p},x\right\rangle=b$ must hold for every $x\in{\mathcal{D}}^{u}(\mathbf{p},b)$.
${\mathcal{D}}^{u}(\mathbf{p},b)$ is always a non-empty compact set when $\mathbf{p}>0$. However, in cases when some prices may be 00, the supremum of utility over the budget constraint might be $\infty$, resulting in unbounded demand. This may happen even when the supremum value is finite, and causes technical difficulties. To circumvent these issues, we adopt the following assumption regarding utility functions $u(\cdot)$:
Assumption 1.
For any $\mathbf{p}\geq\mathbf{0}$ and $b>0$, if there is a sequence of allocations $\{x^{(k)}\}_{k=1}^{\infty}$ such that $\left\langle\mathbf{p},x^{(k)}\right\rangle$ converges to $b$ and $u(x^{(k)})$ converges to some finite value $M<\infty$, then there exists a finite allocation $x^{*}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $u(x^{*})=M$ and $\left\langle\mathbf{p},x\right\rangle\leq b$.
This assumption always holds if the set of possible allocations is restricted to a compact domain $X\subseteq\mathbb{R}_{\geq 0}^{\mathcal{M}}$; note Nesterov and Shikhman (2018) makes such a compactness assumption. For the unrestricted domain $\mathbb{R}_{\geq 0}^{\mathcal{M}}$, it can also be shown for arbitrary piecewise linear utility functions, in particular, for generalized network utilities defined below.
The Gale demand correspondence is defined as
$\mathcal{GD}^{u}(\mathbf{p},b)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\arg\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{p},y\right\rangle\,.$ (2)
We define the domain of $\mathcal{GD}^{u}$ with respect to prices as
$\operatorname{pdom}(u)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\left\{\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}\leavevmode\nobreak\ :\leavevmode\nobreak\ \exists b>0,\sup_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{p},y\right\rangle<+\infty\right\}\,.$ (3)
Under Assumption 1, $\mathbf{0}\in\operatorname{pdom}(u)$ if and only if $u(.)$ is satiable. Every strictly positive $\mathbf{p}>\mathbf{0}$ is in $\operatorname{pdom}(u)$, and if $u$ is satiable, then $\operatorname{pdom}(u)=\mathbb{R}_{\geq 0}^{\mathcal{M}}$ (see Lemma 5.2). In Section 5, we show the following.
Lemma 2.4.
Under Assumption 1, for any $\mathbf{p}\in\operatorname{pdom}(u)$ and any $b>0$, $\mathcal{GD}^{u}(\mathbf{p},b)\neq\emptyset$.
Gale demands may behave very differently from the standard demand correspondence. Example A.3 shows the counter-intuitive property increasing the prices $\mathbf{p}$ may actually increase the Gale utility value $\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{p},y\right\rangle$.
Recall the notions of bounded linear, separable piecewise linear (SPLC) and Leontief-free (LF) utilities from the Introduction. We now define generalized network utility functions.
Generalized network utilities
First, we introduce the generalized flow model. A generalized flow instance is defined as follows. Let $G=(V,E)$ be a directed graph with a sink node $t\in V$, a supply set $S\subseteq V\setminus\{t\}$, a capacity vector $h\in\mathbb{R}_{\geq 0}^{E}$, and positive gain factors $\gamma\in\mathbb{R}_{>0}^{E}$. We let $\partial^{\mathrm{in}}(S)$ and $\partial^{\mathrm{out}}(S)$ denote the set of arcs entering and leaving a subset $S\subseteq V$, respectively. For a node $v\in V$, we use $\partial^{\mathrm{in}}(v)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\partial^{\mathrm{in}}(\{v\})$ and $\partial^{\mathrm{out}}(v)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\partial^{\mathrm{out}}(\{v\})$. For a vector $f\in\mathbb{R}_{\geq 0}^{E}$, for every arc $e=(v,w)$, we interpret $f_{e}$ as the flow entering $e$ at node $v$. The flow gets multiplied by the gain factor $\gamma_{e}$ while traversing the arc, with $\gamma_{e}f_{e}$ amount reaching node $w$. Thus, the net flow at node $v\in V$ is defined as
$\operatorname{net}_{f}(v)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{e\in\partial^{\mathrm{in}}(v)}\gamma_{e}f_{e}-\sum_{e\in\partial^{\mathrm{out}}(v)}f_{e}\,.$
Given a supply vector $x\in\mathbb{R}_{\geq 0}^{S}$, we say that $f\in\mathbb{R}_{\geq 0}^{E}$ is a feasible generalized flow with supply $x$ and capacity $h$ if $0\leq f\leq h$, and $\operatorname{net}_{f}(v)\geq-x_{v}$ for all $v\in S$, and $\operatorname{net}_{f}(v)\geq 0$ for all $v\in V\setminus(S\cup\{t\})$. The maximum generalized flow value from supply $x$ is the maximum amount of $\operatorname{net}_{f}(t)$ among all feasible flows with supply $x$. It corresponds to the optimum value of the following linear program:
$\max\leavevmode\nobreak$ $\operatorname{net}_{f}(t)$ ($\operatorname{GenflowLP}(x)$)
$\operatorname{net}_{f}(v)$ $\geq-x_{v}\quad\forall v\in S$
$\operatorname{net}_{f}(v)$ $\geq 0\quad\forall v\in V\setminus(S\cup\{t\})$
00 0 $\leq f\leq h\,.$
Definition 2.5.
The function $u:\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ is a generalized network utility function if $u(x)$ is the maximum generalized flow value from supply $x$ in some generalized flow instance where the supply set is $S={\mathcal{M}}$, the maximum flow amount is finite for any supply vector, $u(\mathbf{0})=0$, and there exists $x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ with $u(x)>0$.
The following properties are immediate from the definition, and show that these are indeed utility functions in accordance with Definition 2.2.
Proposition 2.6.
If $u:\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ is a generalized network utility function, then $u(x)$ is monotone non-decreasing and concave, and satisfies Assumption 1.
SPLC and Leontief-free as special cases
An SPLC function can be represented as a generalized network utility function by setting $V={\mathcal{M}}\cup\{t\}$, and for each $j\in{\mathcal{M}}$, adding a capacitated arcs $e=(j,t)$ for each linear segment of $v_{j}$, where $\gamma_{e}$ is the marginal utility on this segment and $h_{e}$ is the length of the segment. Similary, we can represent a Leontief-free utility as a generalized network utility function by setting $V={\mathcal{M}}\cup K\cup\{t\}$; recall $K$ is the set of segments. For each $j\in{\mathcal{M}}$ and $k\in K$, we add an arc $e=(j,k)$ with gain factor $\gamma_{e}=a_{jk}$ and capacity $h_{e}=\infty$. For each segment $k\in K$, we add an arc $e=(k,t)$ with $\gamma_{e}=1$ and $h_{e}=\ell_{k}$.
Utilities from production
Assume that agents are competing for access to $m$ machines ${\mathcal{M}}$ that can produce $\ell$ different products. Machine $j$ can produce $a_{jk}$ units of product $k$ per time unit. The revenue of product $k$ is $r_{k}$. There are additionally limits $t_{jk}$ on how much of each product can be produced on each machine by the agent, as well a limit $d_{k}$ on the maximum possible amount producible from $k$. Given $x_{j}$ time on machine $j\in{\mathcal{M}}$, the utility of the agent is the maximum possible revenue subject to these constraints. This can be formulated using a two layer generalized flow network similarly to Leontief-free as above (albeit with additional capacity bounds $t_{jk}$.) The model allows for more complex production networks with possibly multiple layers corresponding to raw materials, intermediary and final products.
2.2 Competitive equilibrium
In a Fisher market instance, given is a set of agents ${\mathcal{A}}$ and a set of divisible goods ${\mathcal{M}}$, with one unit available from each good. Agents have utility functions $u_{i}\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ and budgets $b_{i}\in\mathbb{R}_{>0}$, $i\in{\mathcal{A}}$; we denote the budget vector as $\mathbf{b}=(b_{i})_{i\in{\mathcal{A}}}$. An allocation vector is $\mathbf{x}\in\mathbb{R}_{\geq 0}^{{\mathcal{A}}\times{\mathcal{M}}}$, where $x_{ij}$ denotes the amount of good $j$ allocated to agent $i$. A price vector is a vector $\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$. The demand correspondence of agent $i$ will be denoted as ${\mathcal{D}_{i}}(\mathbf{p},b_{i})$.
Definition 2.7 (Competitive equilibrium).
Given a Fisher market instance as above, the allocations and prices $(\mathbf{x},\mathbf{p})$ form a competitive (market) equilibrium if the following hold:
  1. (i)
    every agent gets an optimal utility at these prices: $x_{i}\in{\mathcal{D}_{i}}(\mathbf{p},b_{i})$ for every agent $i\in{\mathcal{A}}$
  2. (ii)
    no good is oversold: $\sum_{i}x_{ij}\leq 1$ for all $j\in{\mathcal{M}}$.
  3. (iii)
    every item with positive price is fully sold: $\sum_{i}x_{ij}=1$ if $p_{j}>0$.
In a demand-approximate competitive equilibrium, agents may have suboptimal bundles:
Definition 2.8 (Approximate competitive equilibrium).
Given a Fisher market instance and $\alpha\geq 1$, the allocations and prices $(\mathbf{x},\mathbf{p})$ form a $\alpha$-demand approximate competitive equilibrium if
  1. (i)
    $\left\langle\mathbf{p},x_{i}\right\rangle\leq b_{i}$ for each $i\in{\mathcal{A}}$, and $i$ gets at least $\alpha$-fraction of the optimal utility at these prices: $u_{i}(x_{i})\geq\frac{1}{\alpha}\max_{x_{i}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\{u_{i}(x_{i})\,:\,\left\langle\mathbf{p},x_{i}\right\rangle\leq b_{i}\}$.
  2. (ii)
    no good is oversold: $\sum_{i}x_{ij}\leq 1$ for all $j\in{\mathcal{M}}$.
  3. (iii)
    every item with positive price is fully sold: $\sum_{i}x_{ij}=1$ if $p_{j}>0$.
It is a well-established fact that when all $b_{i}$’s are identical, the competitive equilibrium attains exact envy-freeness. Here, we demonstrate that our definition of approximate competitive equilibrium also results in approximate envy-freeness.
Definition 2.9.
For a given allocation $\mathbf{x}\in\mathbb{R}_{\geq 0}^{{\mathcal{A}}\times{\mathcal{M}}}$, we say it is $\beta$-envy-free if, for any agents $i$ and $j$, the condition $u_{i}(x_{i})\geq\frac{1}{\beta}u_{i}(x_{j})$ holds.
Lemma 2.10.
When $b_{i}$’s are identical, the allocation $\mathbf{x}$ of an $\alpha$-approximate competitive equilibrium is also $\alpha$-envy-free.
Proof.
This follows directly from $u_{i}(x_{i})\geq\frac{1}{\alpha}\max_{x_{i}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\left\{u_{i}(x_{i})\,:\,\left\langle\mathbf{p},x_{i}\right\rangle\leq b_{i}\right\}\geq\frac{1}{\alpha}u_{i}(x_{j})$. ∎
Thrifty equilibrium
In case there are satiable utility functions, $\left\langle\mathbf{p},x_{i}\right\rangle<b_{i}$ may be possible in an (exact) competitive equilibrium. Hylland and Zeckhauser (1979) showed that competitive equilibrium may not be Pareto efficient due to satiable utilities. Mas-Colell (1992) strengthened the definition of a competitive equilibrium to require that each agent choose a least-cost optimal plan. This is also known as a strong equilibrium.
Definition 2.11 (Thrifty competitive equilibrium).
Given a Fisher market instance, the allocations and prices $(\mathbf{x},\mathbf{p})$ form a thrifty competitive equilibrium if $(\mathbf{x},\mathbf{p})$ is a competitive equilibrium, and moreover, for every $i\in{\mathcal{A}}$, the allocation $x_{i}$ minimizes $\left\langle\mathbf{p},x_{i}\right\rangle$ among all allocations ${\mathcal{D}_{i}}(\mathbf{p},b_{i})$.
Lemma 2.12 (Garg et al. (2022b), Gul et al. (2022)).
Thrifty competitive equilibria always exist and are Pareto efficient.
Note that in case all utilities are non-satiable, competitive equilibria and thrifty competitive equilibria coincide.
2.3 Nash welfare
Definition 2.13.
Given a Fisher market instance, the Nash welfare of an allocation $\mathbf{y}\in\mathbb{R}_{\geq 0}^{{\mathcal{A}}\times{\mathcal{M}}}$ is the weighted geometric mean of their utilities, namely,
$\left(\prod_{i\in{\mathcal{A}}}u_{i}(x_{i})^{b_{i}}\right)^{\frac{1}{\sum_{i\in{\mathcal{A}}}b_{i}}}\,.$
The allocation $\mathbf{y}\in\mathbb{R}_{\geq 0}^{{\mathcal{A}}\times{\mathcal{M}}}$ maximizes the (weighted) Nash welfare, if its Nash welfare is maximal along all allocations with $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1$ for all $j\in{\mathcal{M}}$, $\mathbf{y}\geq\mathbf{0}$.
Note that the Nash welfare maximizing allocations are precisely the optimal solutions to Eisenberg–Gale convex program Eisenberg and Gale (1959).
[…]
The Lagrangian dual of (EG) can be written as
$\varphi(\mathbf{q})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{i\in{\mathcal{A}}}\max_{y_{i}\in\mathbb{R}_{\geq 0}^{{\mathcal{M}}}}\left(b_{i}\log u_{i}(y_{i})-\left\langle\mathbf{q},y_{i}\right\rangle\right)+\sum_{j\in{\mathcal{M}}}q_{j}.$ (4)
We define the domain $\operatorname{dom}(\varphi)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\cap_{i\in{\mathcal{A}}}\operatorname{pdom}(u_{i})$ (see (3)). Thus, $\operatorname{dom}(\varphi)$ is the set of nonnegative prices $\mathbf{q}$ such that $\varphi(\mathbf{q})<\infty$. Strong duality holds and the dual optimum is taken according since Slater’s condition holds. Consequently, the optimal Nash welfare value equals the dual optimum
$\min_{\mathbf{q}\in\operatorname{dom}(\varphi)}\varphi(\mathbf{q})\,.$ (5)
This optimum is taken (see Lemma 6.2), and we refer to the optimal $\mathbf{q}\in\operatorname{dom}(\varphi)$ as Gale prices. We note that these may not be unique. To justify this term, let us recall the definition of the Gale demand correspondence (2). The term corresponding to agent $i$ in the definition of $\varphi(\mathbf{q})$ is the value of the assignments in the Gale demand correspondence of this agent at prices $\mathbf{q}$ and budget $b_{i}$. We use the notation $\mathcal{GD}_{i}(\mathbf{q},b_{i})$ for the Gale demand for the utility $u_{i}$, $i\in{\mathcal{A}}$.
2.4 Gale substitute utilities
We now define Gale substitute and $\Sigma$-Gale substitute utilities, the crucial classes where we show that the Nash welfare maximizing allocation recovers at least half of the utility of every agent in any possible competitive equilibrium (see Theorem 1.3).
Definition 2.14.
The utility function $u(\cdot)$ is Gale-substitutes if it satisfies the following properties:
  1. (i)
    for any $b>0$, any $\mathbf{q},\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$, and any $y\in\mathcal{GD}^{u}(\mathbf{q},b)$, if $\mathbf{q}\geq\mathbf{q}^{\prime}$, then there exists a $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ such that $y^{\prime}_{j}\leq y_{j}$ whenever $q_{j}^{\prime}=q_{j}$;
  2. (ii)
    for any $b\geq b^{\prime}>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$, and for any $y\in\mathcal{GD}^{u}(\mathbf{q},b)$, there exists a $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$ such that $y^{\prime}\leq y$.
We say that $u(\cdot)$ is $\Sigma$-Gale-substitutes, if it is Gale-substitutes and the following additional property holds:
$\mbox{if }\mathbf{0}\in\operatorname{pdom}(u)\,,\,b>0\mbox{ and }\mathbf{q}\in\mathbb{R}_{\geq 0}^{\mathcal{M}},\mbox{ then for any }y\in\mathcal{GD}(\mathbf{0},b),\exists y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b)\,:\,y^{\prime}\leq y\,.$ ($\Sigma$)
Remark 2.15.
As noted above, $\mathbf{0}\in\operatorname{pdom}(u)$ if and only if $u$ is satiable; thus, ($\Sigma$) is vacuously true for non-satiable utilities. In the satiable case, it asserts that if $y$ is a bundle of maximum possible utility, then at any prices there is an optimal bundle below $y$. Observe that this is immediate for separable utilities. On the other hand, ($\Sigma$) fails for the simple Leontief-free utility $u(x)=\max\{x_{1}+x_{2},1\}$: $y=(1,0)$ is an optimal bundle, whereas for $b=1$, $\mathbf{q}=(2,1)$, the unique solution in $\mathcal{GD}^{u}(\mathbf{q},b)$ is $y^{\prime}=(0,1)$.
Next, we demonstrate a utility function that exhibits Gale-substitutes but not substitutes properties.
Example 2.16.
Consider an agent with budget $b$ and a utility function over two goods given by $u(x)=\min\{x_{1},0.4\}+x_{2}$. This utility is separable and, as such, Gale-substitutes according to Lemma 7.1 shown later. However, we show that this is not weak gross substitutes (WGS): there exist prices $\mathbf{p}\leq\mathbf{p}^{\prime}$ and $x\in{\mathcal{D}}^{u}(\mathbf{p},b)$ such that for some good $j$ with $p_{j}=p^{\prime}_{j}$, every $x^{\prime}\in{\mathcal{D}}^{u}(\mathbf{p},b)$ has $x^{\prime}_{j}<x_{j}$. That is, let $\mathbf{p}=(0.8,1)$ and $\mathbf{p}^{\prime}=(0.9,1)$, and $b=1$. It is easy to see that $x=(0.4,0.68)$ is the unique optimal bundle at prices $\mathbf{p}$, and $x^{\prime}=(0.4,0.64)$ is the unique optimal bundle at prices $\mathbf{p}^{\prime}$ in contradiction to WGS property.
3 Overview of techniques
The ‘price of anarchy’ result Theorem 1.1 is proved using an elementary argument on concavity and the log-sum inequality. The proof is inspired by an argument in Theorem 1 in Brânzei et al. (2022). Example A.4 shows that the bound $(1/\mathrm{e})^{1/\mathrm{e}}$ is tight.
To prove Theorem 1.2 showing that the Nash welfare maximizing allocation is 2-demand-approximate CE, the key ingredient is Lemma 5.1, showing that the solutions in the Gale demand system are within budget, that is, for $y\in\mathcal{GD}^{u}(\mathbf{q},b)$ we have $\left\langle\mathbf{q},y\right\rangle\leq b$. Since at the Nash welfare maximizing solution, every agent gets an allocation from the Gale demand system at the dual prices $\mathbf{q}$, this lemma guarantees that the allocation is feasible. The 2-demand-approximate property then follows by a supergradient argument.
The proof of Theorem 1.3 relies on two key lemmas derived using Lagrangian duality and supergradient arguments. The first is a simple observation, showing that the Gale demand solution for lower prices yields at least half the utility of the utility maximizing bundle at the current prices.
Lemma 3.1.
Consider a utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$, prices $\mathbf{p},\mathbf{q}\in\operatorname{pdom}(u)$ such that $\mathbf{q}\leq\mathbf{p}$. Then, $u(y)\geq\frac{1}{2}u(x)$ for any $y\in\mathcal{GD}^{u}(\mathbf{q},b)$ and $x\in{\mathcal{D}}^{u}(\mathbf{p},b)$.
The second lemma shows that for every allocation of goods in the demand correspondence, there is a smaller or equal allocation of goods in the Gale demand correspondence. Although the first lemma was true for any utilities, this one critically relies on assuming the $\Sigma$-Gale-substitutes property.
Lemma 3.2.
Let $u(\cdot)$ be a $\Sigma$-Gale-substitutes utility and let $x\in{\mathcal{D}}^{u}(\mathbf{p},b)$ for $\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ and $b>0$. Then, $\mathbf{p}\in\operatorname{pdom}(u)$, and there exists $y\in\mathcal{GD}^{u}(\mathbf{p},b)$ such that $y\leq x$.
To derive Theorem 1.3, we start with a competitive equilibrium $(\mathbf{x},\mathbf{p})$, and show that the dual of the Eisenberg–Gale program (5), $\varphi(\mathbf{q})$, has an optimal solution $\mathbf{q}^{\star}$ such that $\mathbf{q}^{\star}\leq\mathbf{p}$. The statement then follows from Lemma 3.1.
An intuitive argument would be the following: we start from $\mathbf{q}=\mathbf{p}$ where the total Gale demand is less than the supply (by Lemma 3.2), and use a tâtonnement-like procedure using the Gale-substitutes property. As long as $\mathbf{q}$ is not a minimizer of $\varphi(.)$, there will be an underdemanded item $j$ with respect to Gale demands. Reducing $p_{j}$ for such a good $j$ may only decrease the Gale demand for all agents on all other goods. This iterative process converges to an equilibrium point and, since the prices are only decreasing, we can conclude that the optimal limit price vector satisfies $\mathbf{q}^{\star}\leq\mathbf{p}$.
[…]
In Section 7, we prove Theorem 1.4, i.e., every separable utility function is $\Sigma$-Gale-substitutes. We show the theorem for general, not necessarily differentiable utility functions using subgradient arguments. We next outline the argument under the simplying assumption that utilities are differentiable and strictly concave. Let $u(y)=\sum_{j\in{\mathcal{M}}}v_{j}(y_{j})$, and let $\nabla_{j}u(y)=\partial v_{j}(y_{j})/\partial y_{j}$ denote the $j$-th component of the gradient. Then, the KKT conditions yield $\nabla_{j}u(y)\leq q_{j}u(y)/b$, with equality whenever $y_{j}>0$ (see Lemma 7.2). Let $y\in\mathcal{GD}^{u}(\mathbf{q},b)$ and $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ for $\mathbf{q}^{\prime}\leq\mathbf{q}$. Using the optimality conditions, it is easy to derive that $u(y^{\prime})\geq u(y)$. Consequently, for every good $j$ with $q_{j}=q^{\prime}_{j}$ and $y_{j}>0$, $\nabla_{j}u(y^{\prime})\geq\nabla_{j}u(y)$ follows. By strict concavity, we must have $y^{\prime}_{j}\leq y_{j}$ as required. The proof without the differentiability and strict concavity assumptions poses additional difficulties: for example, we would not be able to conclude $y^{\prime}_{j}\leq y_{j}$ at this point, and might need a specific choice of $y^{\prime}_{j}$ from $\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$.
Theorem 1.5 shows that generalized network utilities are Gale substitutes. To verify the Gale substitutes property, we need to take two price vectors $\mathbf{q}\geq\mathbf{q}^{\prime}$ and $y\in\mathcal{GD}^{u}(\mathbf{q},b)$, and construct a $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q^{\prime}},b)$ such that $y^{\prime}_{j}\leq y_{j}$ whenever $q^{\prime}_{j}=q_{j}$; we have a second, analogous construction for budget changes. The proof is a finite algorithm that transforms the optimal generalized flow for demands $y$ to $y^{\prime}$ such that we throughout maintain primal-dual slackness condition corresponding to the Gale demand correspondence. The algorithm builds on the literature on maximum generalized flow algorithms. In particular, we use ideas from the classical highest gain augmenting path algorithm by Onaga Onaga (1966). While this does not give a polynomial-time algorithm, it was the starting point of further such algorithms, including Goldberg et al. (1991), Olver and Végh (2020), Végh (2014).
4 Efficiency of competitive equilibria
We now show the ‘price of anarchy’ result Theorem 1.1.
Theorem 4.1 (Restatement of Theorem 1.1).
Given a Fisher market instance, consider any competitive equilibrium $(\mathbf{x},\mathbf{p})$ and a Nash welfare maximizing allocation $\mathbf{y}$. Then,
$\left(\prod_{i\in{\mathcal{A}}}u_{i}(x_{i})^{b_{i}}\right)^{\frac{1}{\sum_{i\in{\mathcal{A}}}b_{i}}}\geq\left(\frac{1}{\mathrm{e}}\right)^{\frac{1}{\mathrm{e}}}\left(\prod_{i\in{\mathcal{A}}}u_{i}(y_{i})^{b_{i}}\right)^{\frac{1}{\sum_{i\in{\mathcal{A}}}b_{i}}}.$
Proof.
Without loss of generality, we can assume $\sum_{i\in{\mathcal{A}}}b_{i}=1$. Note that since $(\mathbf{x},\mathbf{p})$ is a competitive equilibrium, then $u_{i}(x_{i})\geq\min\left\{1,\frac{b_{i}}{\left\langle\mathbf{p},y_{i}\right\rangle}\right\}u_{i}(y_{i})$, as agent $i$ can use their budget, $b_{i}$, to buy at least $\frac{b_{i}}{\left\langle\mathbf{p},y_{i}\right\rangle}$ fraction of $y_{i}$. The lower bound holds because if $b_{i}<\left\langle\mathbf{p},y_{i}\right\rangle$, then $u_{i}\left(\frac{b_{i}}{\left\langle\mathbf{p},y_{i}\right\rangle}y_{i}\right)\geq\frac{b_{i}}{\left\langle\mathbf{p},y_{i}\right\rangle}u_{i}(y_{i})$ using concavity and $u_{i}(\mathbf{0})=0$. Thus,
$\sum_{i\in{\mathcal{A}}}{b_{i}}\log\frac{u_{i}(y_{i})}{u_{i}(x_{i})}\leq\sum_{i\in{\mathcal{A}}}{b_{i}}\log\max\left\{1,\frac{\left\langle\mathbf{p},y_{i}\right\rangle}{b_{i}}\right\}.$
Let $b^{\prime}_{i}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}{\left\langle\mathbf{p},y_{i}\right\rangle}$, and let ${\mathcal{A}}^{\prime}$ be the set of agents such that $b^{\prime}_{i}\geq b_{i}$, $B\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{i\in\mathcal{A}^{\prime}}b_{i}$, and $B^{\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{i\in\mathcal{A}^{\prime}}b^{\prime}_{i}$. Therefore, using the log sum inequality,
$\sum_{i\in{\mathcal{A}}}{b_{i}}\log\frac{u_{i}(y_{i})}{u_{i}(x_{i})}\leq\sum_{i\in{\mathcal{A}}^{\prime}}{b_{i}}\log\frac{b^{\prime}_{i}}{b_{i}}\leq B\log\frac{B^{\prime}}{B}\,.$
By definition, $B\leq\sum_{i\in{\mathcal{A}}}b_{i}=1$. We also note that $B^{\prime}\leq\sum_{i\in{\mathcal{A}}}b_{i}^{\prime}=\sum_{i\in{\mathcal{A}}}\left\langle\mathbf{p},y_{i}\right\rangle\leq\sum_{j\in{\mathcal{M}}}p_{j}\leq\sum_{i\in{\mathcal{A}}}b_{i}=1$.
$\sum_{i\in{\mathcal{A}}}{b_{i}}\log\frac{u_{i}(y_{i})}{u_{i}(x_{i})}\leq\max_{B\leq 1}B\log\frac{1}{B}\leq\frac{1}{\mathrm{e}}\,.$
The statement follows. ∎
5 Properties of the Gale demand correspondence
In this section, we show some properties of the Gale demand correspondence, and also derive Theorem 1.2.
Lemma 5.1.
Given any utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$, $b>0$ and $\mathbf{q}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, for any $y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $\left\langle\mathbf{q},y\right\rangle>b$, there exists $y^{\prime}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $\left\langle\mathbf{q},y^{\prime}\right\rangle\leq b$ and $b\log u(y^{\prime})-\left\langle\mathbf{q},y^{\prime}\right\rangle>b\log u(y)-\left\langle\mathbf{q},y\right\rangle$. In particular, if $y\in\mathcal{GD}^{u}(\mathbf{q},b)$, then, $\left\langle\mathbf{q},y\right\rangle\leq b$.
Proof.
Let $y^{\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}y\cdot\frac{b}{\left\langle\mathbf{q},y\right\rangle}$. By concavity and $u(\mathbf{0})=0$, we get $u(y^{\prime})\geq\frac{b}{\left\langle\mathbf{q},y\right\rangle}u(y)$. Thus,
$b\log u(y^{\prime})-\left\langle\mathbf{q},y^{\prime}\right\rangle$ $\geq b\log u(y)-b\log\frac{\left\langle\mathbf{q},y\right\rangle}{b}-b>b\log u(y)-\left\langle\mathbf{q},y\right\rangle\,,$
where the second inequality used $\log(\alpha)+1<\alpha$ for $\alpha>1$. ∎
Next, we derive some technical properties of the price domain $\operatorname{pdom}(u)$.
Lemma 5.2.
For any utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ and $b>0$, the following properties of the price domain $\operatorname{pdom}(u)$ hold:
  1. (i)
    for any $\mathbf{q}\in\operatorname{pdom}(u)$, $\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{q},y\right\rangle<+\infty$;
  2. (ii)
    $\{\mathbf{q}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{q}>0\}\subseteq\operatorname{pdom}(u)$;
  3. (iii)
    for any $\mathbf{q}\geq\mathbf{q}^{\prime}\geq 0$ such that $\{j:q^{\prime}_{j}=0\}=\{j:q_{j}=0\}$, $\mathbf{q}\in\operatorname{pdom}(u)$ implies $\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$;
  4. (iv)
    for any $\mathbf{p}\geq 0$, if $\max_{\left\langle\mathbf{p},y\right\rangle\leq b}\log u(y)<+\infty$, then $\mathbf{p}\in\operatorname{pdom}(u)$.
  5. (v)
    for any $M>0$, if there is a convergent sequence of prices $\mathbf{p}^{(k)}\in\operatorname{pdom}(u)$ with $\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{p},y\right\rangle<M$, $k=1,2,\ldots$ with a limit point $\mathbf{p}^{(k)}\to\bar{\mathbf{p}}$, then $\bar{\mathbf{p}}\in\operatorname{pdom}(u)$.
Proof.
We first prove property (i). According to the definition of $\operatorname{pdom}(u)$, there exists $b^{\prime}>0$ such that $\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b^{\prime}\log u(y)-\left\langle\mathbf{q},y^{\prime}\right\rangle<+\infty$. For any $b>0$ and for any $y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, if $\left\langle\mathbf{q},y\right\rangle\leq b$, then $b\log u(y)-\left\langle\mathbf{q},y\right\rangle<\frac{b}{b^{\prime}}\left(b^{\prime}\log u(y)-\left\langle\mathbf{q},y\right\rangle\right)+\frac{b-b^{\prime}}{b^{\prime}}\ \left\langle\mathbf{q},y\right\rangle$, which is bounded. Then, the result follows using Lemma 5.1.
Property (ii) holds as, by Lemma 5.1, $\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{q},y\right\rangle\leq\max_{\left\langle\mathbf{q},x\right\rangle\leq b}b\log u(x)-\left\langle\mathbf{q},x\right\rangle$, which is bounded as $x$ is bounded: $x_{j}\leq\frac{b}{q_{j}}<\infty$.
To prove property (iii), we let $\alpha=\min_{q_{j}>0}\{q^{\prime}_{j}/q_{j}\}$. Then, $\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}(\alpha b_{i})\log u_{i}(y_{i})-\left\langle\mathbf{q}^{\prime},y_{i}\right\rangle\leq\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}(\alpha b_{i})\log u_{i}(y_{i})-\alpha\left\langle\mathbf{q},y_{i}\right\rangle<+\infty$.
Property (iv) holds as, for any $y$ such that $\left\langle\mathbf{p},y\right\rangle\leq b$, $b\log u(y)-\left\langle\mathbf{q},y\right\rangle\leq b\log u(y)<+\infty$.
Finally, let us prove (v). For a contradiction, assume $\bar{\mathbf{p}}\notin\operatorname{pdom}(u)$, that is, $\sup_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\bar{\mathbf{p}},y\right\rangle=\infty$. Consider a vector $y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ with $b\log u(y)-\left\langle\bar{\mathbf{p}},y\right\rangle>2M$. Clearly, $y\neq 0$. Let us select $k$ such that $\|\mathbf{p}^{(k)}-\bar{\mathbf{p}}\|<M/\|y\|$. Then, $b\log u(y)-\left\langle\mathbf{p}^{(k)},y\right\rangle=b\log u(y)-\left\langle\bar{\mathbf{p}},y\right\rangle+\left\langle\bar{\mathbf{p}}-\mathbf{p}^{(k)},y\right\rangle>M$, in contradiction with the assumption on the sequence $\mathbf{p}^{(k)}$. ∎
We next prove Lemma 2.4. See 2.4
Proof.
Let $M\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sup_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{p},y\right\rangle$, which implies the existence of a sequence of allocation $\{y^{(k)}\}_{k=1}^{\infty}$ such that $b\log u(y^{(k)})-\left\langle\mathbf{p},y^{(k)}\right\rangle$ converges to $M$. We show the existence of an allocation $y^{*}$ satisfying $b\log u(y^{*})-\left\langle\mathbf{p},y^{*}\right\rangle=M$.
According to Lemma 5.1, there exists a sequence of $\bar{y}^{(k)}$ such that $b\log u(\bar{y}^{(k)})-\left\langle\mathbf{p},\bar{y}^{(k)}\right\rangle$ converges to $M$, with $0\leq\left\langle\mathbf{p},\bar{y}^{(k)}\right\rangle\leq b$ for all $k$. Thus, we can select a subsequence $\bar{y}^{(i_{k})}$ such that $b\log u(\bar{y}^{(i_{k})})-\left\langle\mathbf{p},\bar{y}^{(i_{k})}\right\rangle$ converges to $M$ and $\left\langle\mathbf{p},\bar{y}^{(i_{k})}\right\rangle$ converges to $b^{\prime}$ for some $b^{\prime}\in[0,b]$. Consequently, $b\log u(x^{(i_{k})})$ converges to $M+b^{\prime}$ as $\left\langle\mathbf{p},x^{(i_{k})}\right\rangle$ converges to $b^{\prime}$. By Assumption 1, there exists an allocation $y^{*}$ such that $b\log u(y^{*})=M+b^{\prime}$ and $\left\langle\mathbf{p},y^{*}\right\rangle\leq b^{\prime}$. This completes the proof. ∎
We can use the subgradient Karush–Kuhn–Tucker (KKT) conditions to characterize solutions in the Gale demand correspondence. We refer the reader to Lagrangian duality and KKT conditions of nondifferentiable functions to Chapter 3 in Ruszczyński (2011).
Lemma 5.3.
For a utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$, $b>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$, let $y\in\mathcal{GD}^{u}(\mathbf{q},b)$. Then, there exists a supergradient $\mathbf{g}\in\partial^{\wedge}u(y)$ such that
$g_{j}\leq{q_{j}}\cdot\frac{u(y)}{b}\,,\quad\forall j\in{\mathcal{M}}\,,\,\mbox{with equality if }y_{j}>0\,,$
Conversely, if such a $\mathbf{g}\in\partial^{\wedge}u(y)$ exists then $y\in\mathcal{GD}^{u}(\mathbf{q},b)$.
We are ready to prove Lemma 3.1
See 3.1
Proof.
By Lemma 5.3, there exists a supergradient $\mathbf{g}\in\partial^{\wedge}u(y)$ such that $g_{j}\leq q_{j}u(y)/b$ with equality whenever $y_{i}>0$. Therefore,
$u(x)$ $\leq u(y)+\left\langle\mathbf{g},x-y\right\rangle\leq u(y)+\frac{u(y)}{b}\left\langle\mathbf{q},x-y\right\rangle$
$\leq u(y)\left(1+\frac{\left\langle\mathbf{q},x\right\rangle}{b}\right)\leq u(y)\left(1+\frac{\left\langle\mathbf{p},x\right\rangle}{b}\right)\leq 2u(y)\,.$
5.1 Nash welfare as approximate competitive equilibrium
We now derive Theorem 1.2, showing that Nash welfare maximizing allocations can be seen as approximate competitive equilibria.
Theorem 5.4 (Restatement of Theorem 1.2).
Consider a Fisher market instance, and let $\mathbf{y}$ be an allocation maximizing Nash welfare along with Gale prices $\mathbf{q}$. Then, $(\mathbf{y},\mathbf{q})$ forms a $2$-demand-approximate competitive equilibrium. Additionally, when all $b_{i}$’s are identical, the allocation is $2$-envy-free.
Proof.
By Lemma 5.1, all agents satisfy the budget constraints. Additionally, the market clears by the KKT conditions for (EG). For the approximate competitive equilibrium, we only need to show $u_{i}(y_{i})\geq\frac{1}{2}\max_{\left\langle\mathbf{q},x_{i}\right\rangle\leq b_{i}}u_{i}(x_{i})$ for all agents.
Note that $y_{i}$ maximizes $b_{i}\log u_{i}(y_{i})-\left\langle\mathbf{q},y_{i}\right\rangle$. By Lemma 5.3, there is a supergradient $g_{i}$ such that $g_{ij}\leq q_{j}u_{i}(y_{i})/b_{i}$ with equality whenever $y_{ij}>0$. Then, for any $x_{i}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $\left\langle\mathbf{q},x_{i}\right\rangle\leq b_{i}$,
$u_{i}(x_{i})\leq u_{i}(y_{i})+\left\langle g_{i},x_{i}-y_{i}\right\rangle\leq u_{i}(y_{i})+\frac{u_{i}(y_{i})}{b_{i}}\left\langle\mathbf{q},x_{i}-y_{i}\right\rangle\leq\left(2-\frac{\left\langle\mathbf{q},y_{i}\right\rangle}{b_{i}}\right)u_{i}(y_{i})\leq 2u_{i}(y_{i})\,.$
This proves that $(\mathbf{y},\mathbf{q})$ forms a $2$-approximate competitive equilibrium. Additionally, allocation is $2$-envy-free follows directly from Lemma 2.10. ∎
6 Efficiency of Nash welfare for Gale-substitutes
In this section, we prove Theorem 1.3. We show that for $\Sigma$-Gale subsitutes utility functions, given a competitive equilibrium $(\mathbf{x},\mathbf{p})$, there are Gale prices $\mathbf{q}^{\star}$ such that $\mathbf{q}^{\star}\leq\mathbf{p}$. The statement then follows by Lemma 3.1. We start by re-stating and proving Lemma 3.2.
See 3.2
Proof.
According to the KKT conditions, there exists $\lambda\geq 0$ and supergradient $\mathbf{g}\in\partial^{\wedge}u(x)$ such that
$g_{j}\leq\lambda p_{j}\,,\quad\forall j\in{\mathcal{M}}\,,\,\mbox{with equality if }x_{j}>0\,,$ (6)
and $\left\langle\mathbf{p},x\right\rangle=b$ if $\lambda>0$.
Let us first consider the case $\lambda=0$. Then, $g_{j}\leq 0$ with equality if $x_{j}>0$, implying that $x=\arg\max_{x\in\mathbb{R}^{\mathcal{M}}}u(x)$. Consequently, $u(.)$ is satiable; therefore, $\mathbf{0}\in\operatorname{pdom}(u)$ and $x\in\mathcal{GD}^{u}(\mathbf{0},b)$. Now, ($\Sigma$) guarantees the existence of $y\in\mathcal{GD}^{u}(\mathbf{p},b)$, $y\leq x$. Clearly, $\operatorname{pdom}(u)=\mathbb{R}_{\geq 0}^{\mathcal{M}}$ in this case.
Next, assume $\lambda>0$. Let us define
$\mathbf{q}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\frac{\lambda b}{u(x)}\cdot\mathbf{p}\,.$
Note that $\mathbf{q}\leq\mathbf{p}$. This follows since $\lambda b=\lambda\left\langle\mathbf{p},x\right\rangle=\left\langle\mathbf{g},x\right\rangle\leq u(x)$, where the final inequality is by concavity and $u(\mathbf{0})=0$. Thus,
$g_{j}\leq\frac{u(x)}{b}q_{j}\,,\quad\forall j\in{\mathcal{M}}\,,\,\mbox{with equality if }x_{j}>0\,.$
According to Lemma 5.3, $x\in\mathcal{GD}^{u}(\mathbf{q},b)$.
Note that $\mathcal{GD}^{u}(\mathbf{q},b)=\mathcal{GD}^{u}(\mathbf{p},u(x)/\lambda)$. Utilizing Definition 2.14(ii) for $x\in\mathcal{GD}^{u}(\mathbf{p},u(x)/\lambda)$ and $b\leq u(x)/\lambda$, there exists $y\in\mathcal{GD}^{u}(\mathbf{p},b)$ such that $y\leq x$ as required. ∎
We next show that for any competitive equilibrium $(\mathbf{x},\mathbf{p})$, and any prices $\mathbf{q}\leq\mathbf{p}$, there exist an aggregate Gale demand such that $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1$ whenever $q_{j}=p_{j}$.
Lemma 6.1.
Consider a Fisher market instance such that $u_{i}(\cdot)$ is $\Sigma$-Gale-substitute utilities for all agents. For any competitive equilibrium $(\mathbf{x},\mathbf{p})$, for any $\mathbf{q}\in\cap_{i\in{\mathcal{A}}}\operatorname{pdom}(u_{i})$ such that $\mathbf{q}\leq\mathbf{p}$, there exists $y_{i}\in\mathcal{GD}_{i}(\mathbf{q},b_{i})$ for all $i\in{\mathcal{A}}$ and $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1$ for all $j\in{\mathcal{M}}$ such that $q_{j}=p_{j}$.
Proof.
By Lemma 3.2, there exists $\bar{y}_{i}\in\mathcal{GD}_{i}(\mathbf{p},b_{i})$ for every $i\in{\mathcal{A}}$ such that $\bar{y}_{i}\leq x_{i}$. Using Definition 2.14(i), there exists $y_{i}\in\mathcal{GD}_{i}(\mathbf{q},b_{i})$ and $y_{ij}\leq\bar{y}_{ij}$ if $q_{j}=p_{j}$. Thus, if $q_{j}=p_{j}$ then $\sum_{i\in{\mathcal{A}}}y_{ij}\leq\sum_{i\in{\mathcal{A}}}\bar{y}_{ij}\leq\sum_{i\in{\mathcal{A}}}x_{ij}\leq 1$. ∎
Recall the Lagrangian dual function $\varphi$ defined in (4) on the domain $\operatorname{dom}(\varphi)=\cap_{i\in{\mathcal{A}}}\operatorname{pdom}(u_{i})$.
$\varphi(\mathbf{q})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{i\in{\mathcal{A}}}\max_{y_{i}\in\mathbb{R}_{\geq 0}^{{\mathcal{M}}}}\left(b_{i}\log u_{i}(y_{i})-\left\langle\mathbf{q},y_{i}\right\rangle\right)+\sum_{j\in{\mathcal{M}}}q_{j}.$
The following properties are easy to verify.
Lemma 6.2.
The following hold for the function $\varphi(\cdot)$.
  1. (i)
    $\varphi(\mathbf{q})$ is convex.
  2. (ii)
    Given $\mathbf{q}\in\operatorname{dom}(\varphi)$, let $y_{i}\in\mathcal{GD}_{i}(\mathbf{q},b_{i})$ for all $i\in{\mathcal{A}}$. Then, $\mathbf{1}-\sum_{i\in{\mathcal{A}}}y_{i}\in\partial^{\vee}\varphi(\mathbf{q})$.
  3. (iii)
    The optimum value of (5), that is, the infimum of $\varphi(\mathbf{q})$ is taken in $\operatorname{dom}(\varphi)$.
Proof of Theorem 1.3.
Let $(\mathbf{x},\mathbf{p})$ be a competitive equilibrium solution. Consider the dual (5) of the Eisenberg-Gale program (EG), i.e.,
$\min_{\mathbf{q}\in\operatorname{dom}(\varphi)}\varphi(\mathbf{q})\,.$
We show that this has an optimal solution with $\mathbf{q}\leq\mathbf{p}$. The theorem then follows by Lemma 3.1. Let us take any optimal $q^{\star}\in\operatorname{dom}(\varphi)$; we are done if $\mathbf{q}^{\star}\leq\mathbf{p}$. Otherwise, let $J\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\{j\in{\mathcal{M}}\,:\,\mathbf{q}^{\star}_{j}>p_{j}\}$, and let us define $\bar{\mathbf{q}}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ as
$\bar{q}_{j}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\begin{cases}p_{j}\,,&\mbox{if }j\in J\,,\\ q^{\star}_{j}\,,&\mbox{if }j\notin J\,.\\ \end{cases}$
First, assuming $\bar{\mathbf{q}}\in\operatorname{dom}(\varphi)$, we show that $\bar{\mathbf{q}}$ is also optimal to (5) as claimed.
Let us pick an allocation $\mathbf{y}$ for the prices $\bar{\mathbf{q}}$ as guaranteed by Lemma 6.1: $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1$ for all $j\in{\mathcal{M}}$ with $\bar{q}_{j}=p_{j}$; in particular, for all $j\in J$. According to Lemma 6.2(ii), we get a subgradient $\bar{\mathbf{g}}\in\partial^{\vee}\varphi(\bar{\mathbf{q}})$ defined as $\bar{\mathbf{g}}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\mathbf{1}-\sum_{i\in{\mathcal{A}}}y_{i}$; this has the property $\bar{g}_{j}\geq 0$ for all $j\in J$. We then have
$\varphi(\mathbf{q}^{\star})\geq\varphi(\bar{\mathbf{q}})+\left\langle\bar{\mathbf{g}},\mathbf{q}^{\star}-\bar{\mathbf{q}}\right\rangle=\varphi(\bar{\mathbf{q}})+\sum_{j\in J}\bar{g}_{j}(q^{\star}_{j}-\bar{q}_{j})\geq\varphi(\bar{\mathbf{q}})\,.$
Here, the first inequality is by the definition of subgradients; the equality uses that $q^{\star}_{j}=\bar{q}_{j}$ if $j\notin J$, and the final inequality follows since $\bar{g}_{j}\geq 0$ and $q^{\star}_{j}>\bar{q}_{j}$ for $j\in J$. This shows that $\bar{\mathbf{q}}$ is also an optimal solution to (5), completing the proof assuming $\bar{\mathbf{q}}\in\operatorname{dom}(\varphi)$.
It is left to show that $\bar{\mathbf{q}}\in\operatorname{dom}(\varphi)$. Let $K\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\{j\in{\mathcal{M}}\,:\,\mathbf{q}^{\star}_{j}=0<p_{j}\}$. Clearly, $K\cap J=\emptyset$. If $K=\emptyset$, then Lemma 5.2(iii) already implies $\bar{\mathbf{q}}\in\operatorname{pdom}(u)$, since $\mathbf{p}\geq\bar{\mathbf{q}}$ and $\operatorname{supp}(\bar{\mathbf{q}})=\operatorname{supp}(\mathbf{p})$. Otherwise, for $k=1,2,\ldots$, let us define $\mathbf{q}^{(k)}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ as
$q_{j}^{(k)}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\begin{cases}p_{j}\,,&\mbox{if }j\in J\,,\\ \frac{p_{j}}{k}\,,&\mbox{if }j\in K\,,\\ q^{\star}_{j}\,,&\mbox{if }j\notin J\cup K\,.\\ \end{cases}$
We have $\mathbf{q}^{(k)}\in\operatorname{pdom}(u)$ again by Lemma 5.2(iii). As in the first part, we take a subgradient $\mathbf{g}^{(k)}$ as guaranteed by Lemma 6.2(ii), with the properties that $\bar{g}_{j}\geq 0$ for all $j\in J$ and also $\mathbf{g}^{(k)}\leq\mathbf{1}$. We can write
$\varphi(\mathbf{q}^{\star})\geq\varphi(\mathbf{q}^{(k)})+\left\langle\mathbf{g}^{(k)},\mathbf{q}^{\star}-\mathbf{q}^{(k)}\right\rangle=\varphi(\mathbf{q}^{(k)})+\sum_{j\in J}\bar{g}_{j}(q^{\star}_{j}-q^{(k)}_{j})-\sum_{j\in K}\bar{g}_{j}q^{(k)}_{j}\geq\varphi(\mathbf{q}^{(k)})-\frac{\|p\|_{\infty}}{k}\,.$
Here, we use in the first equality that $q^{\star}_{j}=q^{(k)}_{j}$ for $j\notin J\cup K$, and $q^{\star}_{j}=0$ for $j\in K$. The last inequality uses $\bar{g}_{j}\leq 1$ and the definition $q^{(k)}_{j}=p_{j}/(k+1)$. Thus, Lemma 5.2(v) is applicable with $M=\varphi(\mathbf{q}^{\star})+\|p\|_{\infty}$, using that for each $j\in{\mathcal{M}}$, $\max_{y_{i}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b_{i}\log u_{i}(y_{i})-\left\langle\mathbf{q}^{(k)},y_{i}\right\rangle\leq\varphi(\mathbf{q}^{(k)})$. This shows that $\bar{\mathbf{q}}\in\operatorname{pdom}(u_{i})$ for all $i\in{\mathcal{M}}$, and therefore $\bar{\mathbf{q}}\in\operatorname{dom}(\varphi)=\cap_{i\in{\mathcal{M}}}\operatorname{pdom}(u_{i})$. ∎
7 Separable utilities are $\Sigma$-Gale-substitutes
In this section, we prove Theorem 1.4. See 1.4 Recall the following simple monotonicity property of supergradients of univariate concave functions.
Lemma 7.1.
Let $v\,:\,\mathbb{R}\to\mathbb{R}$ be a univariate concave function, $\alpha,\alpha^{\prime}\in\operatorname{dom}(f)$, and $g\in\partial^{\wedge}v(\alpha)$, $g^{\prime}\in\partial^{\wedge}v(\alpha^{\prime})$. Then, $\alpha<\alpha^{\prime}$ implies $g\geq g^{\prime}$. Conversely, if $g>g^{\prime}$, then $\alpha\leq\alpha^{\prime}$.
We now adapt Lemma 5.3 to separable utilities.
Lemma 7.2.
Let $u(x)=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$ be a separable utility function. For any $b>0$, $\mathbf{q}\in\operatorname{pdom}(u)$, $y\in\mathcal{GD}^{u}(\mathbf{q},b)$, we have $u(y)>0$. Further, $y\in\mathcal{GD}^{u}(b,\mathbf{q})$ if and only if there exist supergradients $g_{j}\in\partial^{\wedge}v_{j}(y_{j})$, $j\in{\mathcal{M}}$ satisfying
  1. (i)
    $g_{j}\geq 0$; and
  2. (ii)
    $g_{j}\leq q_{j}u(y)/b$ with equality whenever $y_{j}>0$.
The proof of Theorem 1.4 follows by the next three lemmas, that prove the three properties in Definition 2.14 for separable utilities: property (i) in Lemma 7.3, property (ii) in Lemma 7.4, and property ($\Sigma$) in Lemma 7.5. A similar argument was used in Lemma 7.1 in Garg et al. (2022a) to show a variant of the Gale substitutes property for capped SPLC utilities.
Lemma 7.3.
Let $u(x)=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$ be a separable utility function, and let $b>0$, $\mathbf{q},\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$ and $y\in\mathcal{GD}^{u}(\mathbf{q},b)$. If $\mathbf{q}\geq\mathbf{q}^{\prime}$, then there exists a $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ such that $y^{\prime}_{j}\leq y_{j}$ whenever $q_{j}^{\prime}=q_{j}$.
Proof.
Let $J\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\{j\in{\mathcal{M}}\,:\,q^{\prime}_{j}=q_{j}\}$. Let us select $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ minimizing the potential $\Psi(y^{\prime})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{j\in J}(y^{\prime}_{j}-y_{j})^{+}$. Our goal is to show $\Psi(y^{\prime})=0$. Throughout the proof, let $g^{\prime}_{j}\in\partial^{\wedge}v_{j}(y^{\prime}_{j})$ and $g_{j}\in\partial^{\wedge}v_{j}(y_{j})$ as in Lemma 7.2.
First, we show $u(y^{\prime})\geq u(y)$. For a contradiction, assume $u(y^{\prime})<u(y)$. Then, according to Lemma 7.2, for every $j\in{\mathcal{M}}$ with $y_{j}>0$ and $q_{j}>0$ we have
$g^{\prime}_{j}\leq q^{\prime}_{j}u(y^{\prime})/b<q_{j}u(y)/b=g_{j}\,,$
while for $j\in{\mathcal{M}}$ with $q_{j}=0$ we have $g^{\prime}_{j}=g_{j}=0$. Consequently, for each $y_{j}>0$, if $q_{j}>0$ then Lemma 7.1 implies $y^{\prime}_{j}\geq y_{j}$, and if $q_{j}=0$ then $v_{j}(y_{j})=v_{j}(y_{j}^{\prime})$, both being equal to the maximum value of $v_{j}(.)$. These imply $u(y^{\prime})\geq u(y)$, a contradiction.
Using the above, for any $j\in J$ where $y^{\prime}_{j}>0$,
$g^{\prime}_{j}=\frac{q^{\prime}_{j}}{b}u(y^{\prime})\geq\frac{q_{j}}{b}u(y)\geq g_{j}\,.$ (7)
Let us first consider the case $u(y^{\prime})>u(y)$. For all $j\in J$ with $q^{\prime}_{j}=q_{j}>0$, we get $g^{\prime}_{j}>g_{j}$. By Lemma 7.1, $y^{\prime}_{j}\leq y_{j}$ follows for such goods $j$. On the other hand, if $q^{\prime}_{j}=q_{j}=0$, then this implies $g^{\prime}_{j}=g_{j}=0$. For such indices, if $y^{\prime}_{j}>y_{j}$, then replacing $y^{\prime}_{j}$ by $y_{j}$ results in another bundle $y^{\prime\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ with $\Psi(y^{\prime\prime})<\Psi(y^{\prime})$, a contradiction.
For the rest of the proof, let us assume that $u(y^{\prime})=u(y)$. We next show $y^{\prime}_{j}\geq y_{j}$ for all $j\in{\mathcal{M}}\setminus J$, that is, for all goods with $q^{\prime}_{j}<q_{j}$. This is clearly true if $y_{j}=0$. Whenever $y_{j}>0$ for such an index, we have
$g^{\prime}_{j}\leq\frac{q^{\prime}_{j}}{b}u(y^{\prime})<\frac{q_{j}}{b}u(y)=g_{j}\,,$
and therefore $y^{\prime}_{j}\geq y_{j}$ by Lemma 7.1. By monotonicity, $v_{j}(y^{\prime}_{j})\geq v_{j}(y_{j})$ for all $j\in{\mathcal{M}}\setminus J$. Since $u(y^{\prime})=u(y)$, we have $\sum_{j\in J}v_{j}(y^{\prime}_{j})\leq\sum_{j\in J}v_{j}(y_{j})$.
From (7), it follows that $g^{\prime}_{j}=g_{j}$ for all $j\in J$ with $y^{\prime}_{j}>0$. This is because the first inequality is now equality, and the second inequality may only be strict if $y_{j}=0<y^{\prime}_{j}$. By Lemma 7.1, $g_{j}\geq g^{\prime}_{j}$ in this case, leading to a contradiction.
Assume for a contradiction $\Psi(y^{\prime})>0$, and let us select an index $k\in J$ with $y^{\prime}_{k}>y_{k}$. If $v_{k}(y^{\prime}_{k})=v_{k}(y_{k})$, then we get a contradiction, because changing $y^{\prime}_{k}$ to $y_{k}$ would preserve the the optimality conditions in Lemma 7.2, while decreasing the potential $\Psi(y^{\prime})$. Thus, $v_{k}(y^{\prime}_{k})>v_{k}(y_{k})$. By $\sum_{j\in J}v_{j}(y^{\prime}_{j})\leq\sum_{j\in J}v_{j}(y_{j})$, there must be an index $j\in J$ with $v_{j}(y^{\prime}_{j})<v_{j}(y_{j})$. We can now find suitable $\varepsilon_{k},\varepsilon_{j}>0$ such that $v_{k}(y^{\prime}_{k}-\varepsilon_{k})+v_{j}(y^{\prime}_{j}+\varepsilon_{j})=v_{k}(y^{\prime}_{k})+v_{j}(y^{\prime}_{j})$, and $y^{\prime}_{j}+\varepsilon_{j}\leq y_{j}$. This modified solution $y^{\prime\prime}$ satisfies the optimality conditions but $\Psi(y^{\prime\prime})<\Psi(y^{\prime})$, again leading to a contradiction. ∎
Lemma 7.4.
Let $u(x)=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$ be a separable utility function, $b>b^{\prime}>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$. Then, for any $y\in\mathcal{GD}^{u}(\mathbf{q},b)$, there exists a $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$ such that $y^{\prime}\leq y$.
Proof.
We argue similarly as in the previous proof: we pick $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$ which minimizes $\Psi(y^{\prime})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{j\in J}(y^{\prime}_{j}-y_{j})^{+}$, and show that $\Psi(y^{\prime})=0$.
First, we show that ${u(y)}/{b}\leq{u(y^{\prime})}/{b^{\prime}}$. For a contradiction, suppose ${u(y)}/{b}>{u(y^{\prime})}/{b^{\prime}}$. If $y_{j}>0$, then, by Lemma 7.2
$g_{j}=\frac{q_{j}u(y)}{b}>\frac{q_{j}u(y^{\prime})}{b^{\prime}}\geq g^{\prime}_{j},$
which implies $v_{j}(y_{j})\leq v_{j}(y^{\prime}_{j})$ whenever $y_{j}>0$ (Lemma 7.1), and therefore, $u(y)\leq u(y^{\prime})$. Together with $b>b^{\prime}$, this implies ${u(y)}/{b}\leq{u(y^{\prime})}/{b^{\prime}}$, a contradiction.
Whenever $y^{\prime}_{j}>0$, we now have
$g^{\prime}_{j}=\frac{q_{j}u(y^{\prime})}{b^{\prime}}\geq\frac{q_{j}u(y)}{b}\geq g_{j}\,.$ (8)
Assume first $u(y^{\prime})/b^{\prime}>u(y)/b$. If $q_{j}>0$, then this inequality is strict, and therefore $y^{\prime}_{j}\leq y_{j}$ for all $j\in{\mathcal{M}}$ follows by Lemma 7.1. If $q_{j}=0$, then $g^{\prime}_{j}=g_{j}=0$. If $y^{\prime}_{j}>y_{j}$ for such a good $j$, we can decrease $y^{\prime}_{j}$ to $y_{j}$ and thereby get an optimal allocation with strictly smaller $\Psi(y^{\prime})$ value, leading to a contradiction. This completes the proof for $u(y^{\prime})/b^{\prime}>u(y)/b$.
For the rest, we may assume $u(y^{\prime})/b^{\prime}=u(y)/b$. Now, $g^{\prime}_{j}=g_{j}$ must hold whenever $y^{\prime}_{j}>0$. This is because the first inequality in (8) is equality, and if the second is strict, then $y_{j}=0<y^{\prime}_{j}$ would give a contradiction to $g^{\prime}_{j}>g_{j}$.
Assume for a contradiction $\Psi(y^{\prime})>0$, and let us pick $k\in{\mathcal{M}}$ with $y^{\prime}_{k}>y_{k}$. The equality $u(y^{\prime})/b^{\prime}=u(y)/b$ and $b^{\prime}<b$ gives $u(y^{\prime})<u(y)$. We can derive a contradiction by decreasing $y^{\prime}_{k}$ and potentially increasing an $y^{\prime}_{j}$ similarly as in the proof of Lemma 7.3. ∎
Lemma 7.5.
Let $u(x)=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$ be a separable utility function such that $\mathbf{0}\in\operatorname{pdom}(u)$. Then, for any $\mathbf{q}\geq\mathbf{0}$, $b>0$ and any $y\in\mathcal{GD}^{u}(\mathbf{0},b)$, there exists a $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b)$ such that $y^{\prime}\leq y$.
Proof.
$\mathbf{0}\in\operatorname{pdom}(u)$ implies $\sup_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}u(y)<\infty$. Therefore, $\operatorname{pdom}(u)=\mathbb{R}_{\geq 0}^{\mathcal{M}}$; in particular, $\mathbf{q}\in\operatorname{pdom}(u)$. Let us pick $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b)$ such that $\sum_{j\in{\mathcal{M}}}y^{\prime}_{j}$ is minimal.
Note that $u(y)=\sup_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}u(y)$. In particular, for any $\alpha\geq y_{j}$, $v_{j}(\alpha)=v_{j}(y_{j})$. For a contradiction, assume there exists a $k\in{\mathcal{M}}$ with $y_{k}<y^{\prime}_{k}$. Then, reducing $y^{\prime}_{k}$ to $y_{k}$ may only increase $b\log u(y^{\prime})-\left\langle\mathbf{q},y^{\prime}\right\rangle$, while decreasing $\sum_{j\in{\mathcal{M}}}y^{\prime}_{j}$. This contradicts the choice of $y^{\prime}$. ∎
8 Generalized network utilities
In this section, we prove Theorem 1.5. We will first need a characterization of subgradients by linear programming duality. We state this in a more general form. The proof based on LP duality is deferred to Appendix B.
Lemma 8.1.
Let $\mathbf{A}\in\mathbb{R}^{k\times\ell}$, $\mathbf{B}\in\mathbb{R}^{k\times m}$, $c\in\mathbb{R}^{\ell}$, $b\in\mathbb{R}^{k}$. Let $\mathcal{D}\subseteq\mathbb{R}^{m}$ be the set of vectors $x\in\mathbb{R}^{m}$ such that the following parametric linear program is feasible and bounded, and let us define $\varphi:\mathcal{D}\to\mathbb{R}$ such that $\varphi(x)$ is the optimum value of this program:
$\max\leavevmode\nobreak$ $c^{\top}z$ ($\operatorname{Primal}(x)$)
$\mathbf{A}z$ $\leq\mathbf{B}x+b$
$z$ $\geq 0$
Then, for every $x\in\mathcal{D}$, $g\in\partial^{\wedge}\varphi(x)$ if and only if $g=\mathbf{B}^{\top}\pi$ for an optimal solution $\pi\in\mathbb{R}_{\geq 0}^{k}$ to the dual linear program
$\min\leavevmode\nobreak$ $x^{\top}\mathbf{B}^{\top}\pi+b^{\top}\pi$ ($\operatorname{Dual}(x)$)
$\mathbf{A}^{\top}\pi$ $\geq c$
$\pi$ $\geq 0\,.$
Consider a generalized flow instance defined in a graph $G=(V,E)$, sink $t\in V$, capacities $h\in\mathbb{R}_{\geq 0}^{E}$, gain factors $\gamma\in\mathbb{R}_{>0}^{E}$ and supply set $S={\mathcal{M}}$. For the rest of this section, let $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ be a generalized network utility function defined in the form $\operatorname{GenflowLP}(x)$ in Section 2.
Using Lemma 8.1 and complementary slackness, we get the following characterization of subgradients of generalized network utility functions. Note that for a generalized network utility function, the domain is $\mathcal{D}=\mathbb{R}_{\geq 0}^{\mathcal{M}}$, because $f=0$ is always a feasible solution to $\operatorname{GenflowLP}(x)$, and the definition assumes that the objective is always bounded.
Lemma 8.2.
[…]
$\gamma_{e}\pi_{w}-\pi_{v}$ $=0\,,\quad\forall e=(v,w)\in E:h_{e}>f_{e}>0\,,$ (9)
$\gamma_{e}\pi_{w}-\pi_{v}$ $\leq 0\,,\quad\forall e=(v,w)\in E:f_{e}=0\,,$
$\gamma_{e}\pi_{w}-\pi_{v}$ $\geq 0\,,\quad\forall e=(v,w)\in E:f_{e}=h_{e}\,,$
$\pi_{v}$ $=0\,,\quad\forall v\in{\mathcal{M}}:\operatorname{net}_{f}(v)>-x_{v}\mbox{ or }v\in V\setminus({\mathcal{M}}\cup t):\operatorname{net}_{f}(v)>0\,,$
$\pi_{t}$ $=1\,,$
$\pi$ $\geq 0\,.$
In accordance with this lemma, we say that $f\in\mathbb{R}^{E}$ and $\pi\in\mathbb{R}^{V}$ form a fitting pair with respect to $x$, if $f$ is a feasible generalized flow with supply $x$, and $\pi$ satisfies (9). Note that this implies that $f$ and $\pi$ are primal and dual optimal solutions to $\operatorname{GenflowLP}(x)$. Further, $g\in\partial^{\wedge}u(x)$ if and only if $g$ is the restriction of $\pi$ to ${\mathcal{M}}$ for a fitting pair $(f,\pi)$ with respect to $x$.
For every arc $e=(v,w)\in E$, let us define the reverse arc $\overleftarrow{e}=(v,w)$ with gain factor $\gamma_{\overleftarrow{e}}=1/\gamma_{e}$. By sending $\alpha$ units of flow on $\overleftarrow{e}$ we mean decreasing the flow value on $e$ by $\gamma_{e}\alpha$. Given a fitting pair $(f,\pi)$ with respect to $x$, we define the auxiliary network $G^{\textrm{aux}}_{f,\pi}=(V,E^{\textrm{aux}}_{f,\pi})$ by
$E^{\textrm{aux}}_{f,\pi}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\left\{e=(v,w)\in E\,:\,\gamma_{e}\pi_{w}-\pi_{v}=0\mbox{ and }f_{e}<h_{e}\right\}\cup\left\{\overleftarrow{e}\,:\,e\in E\,:\,\gamma_{e}\pi_{w}-\pi_{v}=0\mbox{ and }0<f_{e}\right\}\,.$
In the proof of Theorem 1.5, we use two flow subroutines in the generalized flow network. The first one starts with a fitting pair $(f,\pi)$ with respect to a demand vector $x$, and outputs a flow $f^{\prime}$ for a decreased demand vector $x^{\prime}$ such that the flow value $\operatorname{net}_{f^{\prime}}(t)$ is decreased to the smallest possible amount while maintaining $(f^{\prime},\pi)$ as a fitting pair.
Input: $x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, and a fitting pair $(f,\pi)$ with respect to $x$
Output: A supply vector $0\leq x^{\prime}\leq x$ and a feasible generalized flow $f^{\prime}$ such that $(f^{\prime},\pi)$ is a fitting pair with respect to $x^{\prime}$, and subject to this, $\operatorname{net}_{f^{\prime}}(t)$ is minimal.
Algorithm 1 Pullback-Flow
Lemma 8.3.
Pullback-Flow can be formulated as a linear program. For the output flow $f^{\prime}$, the auxiliary graph $E^{\textrm{aux}}_{f^{\prime},\pi}$ does not contain any directed paths from the sink $t$ to $\operatorname{supp}(x^{\prime})$.
Proof.
When fixing $\pi$, the requirement that $(f^{\prime},\pi)$ be a fitting pair can be formulated by linear constraints on $f^{\prime}$. For example, for every arc $e=(v,w)$, if $\gamma_{e}\pi_{w}-\pi_{v}<0$ then we require $f^{\prime}_{e}=0$, and if $\gamma_{e}\pi_{w}-\pi_{v}>0$ then $f^{\prime}_{e}=h_{e}$; if $\gamma_{e}\pi_{w}-\pi_{v}=0$ then any value $0\leq f^{\prime}_{e}\leq h_{e}$ is allowed. For the second part, assume for a contradiction that $E^{\textrm{aux}}_{f^{\prime},\pi}$ contains a directed path $P$ from $t$ to a node $j\in{\mathcal{M}}$ with $x^{\prime}_{j}>0$. Then, we can send a small amount of flow along this path by maintaining the flow balance on every intermediate node, and decreasing $\operatorname{net}_{f^{\prime}}(t)$ and $x^{\prime}_{j}$ by positive amounts. This contradicts the minimality of $\operatorname{net}_{f^{\prime}}(t)$. ∎
Remark 8.4.
Pullback-Flow can be in fact formulated as a (classical) maximum flow problem. To see this, one can use the standard generalized flow relabelling technique, first introduced by Glover Glover and Klingman (1973), see also e.g., Goldberg et al. (1991), Olver and Végh (2020)). That is, we multiply the flow at node $v$ by $\pi_{v}$ and replacing $\gamma_{e}$ by $\gamma_{e}\pi_{w}/\pi_{v}$ for every $e=(v,w)\in E$. On every arc, we fix $f^{\prime}_{e}$ to the upper or lower bound unless $\gamma_{e}\pi_{w}-\pi_{v}=0$, in which case the relabelled gain factor is 1.
Our second subroutine aims to increase the supply at a designated node $k\in{\mathcal{M}}$ while the supply at all other goods is non-increasing, and the flow value $\operatorname{net}_{f}(t)$ is non-decreasing. The goal is again to maintain a fitting pair $(f^{\prime},\pi)$ for a demand vector $x^{\prime}$.
Input: $x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, and a fitting pair $(f,\pi)$ with respect to $x$, and a good $k\in{\mathcal{M}}$ such that $\pi_{k}>0$.
Output: A supply vector $x^{\prime}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $x^{\prime}_{k}\geq x_{k}$ and $x^{\prime}_{j}\leq x_{j}$ if $j\in{\mathcal{M}}\setminus\{k\}$, and a feasible generalized flow $f^{\prime}$ such that $(f^{\prime},\pi)$ is a fitting pair with respect to $x^{\prime}$, $\operatorname{net}_{f}(t^{\prime})\geq\operatorname{net}_{f}(t)$, and subject to this, $x^{\prime}_{k}$ is maximal.
Algorithm 2 Reroute-Flow
The following lemma can be derived similarly to Lemma 8.3.
Lemma 8.5.
Reroute-Flow can be formulated as a linear program. For the output flow $f^{\prime}$, the auxiliary graph $E^{\textrm{aux}}_{f^{\prime},\pi}$ does not contain any directed path from $k$ to the set $(\operatorname{supp}(x^{\prime})\setminus\{k\})\cup\{t\}$.
From Lemma 5.3, we get the following.
Lemma 8.6.
Consider a generalized network utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$, $b>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$, and $x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$. Then, $x\in\mathcal{GD}^{u}(b,\mathbf{q})$ if and only if there exists a fitting pair $(f,\pi)$ with respect to $x$ such that $u(x)=\operatorname{net}_{f}(t)$ and
$\pi_{j}\leq\frac{q_{j}}{b}\cdot u(x)\,\quad\forall j\in{\mathcal{M}}\,,\,\text{ with equality for }j\in\operatorname{supp}(x)\,.$ (10)
Theorem 1.5 follows by showing the two properties in Definition 2.14 in Lemmas 8.7 and Lemma 8.8 below. We start by property (ii).
Lemma 8.7.
Let $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ be a generalized network utility function, $b>b^{\prime}>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$, and let $x\in\mathcal{GD}^{u}(b,\mathbf{q})$. Then, there exists $x^{\prime}\in\mathcal{GD}^{u}(b^{\prime},\mathbf{q})$ such that $x^{\prime}\leq x$.
Proof.
The proof is given by a finite algorithm, shown as Algorithm 3. We start from fitting pair $(f,\pi)$ with respect to $x$ as in Lemma 8.6. Thus, the inequality (10) holds for $\pi$, $x$, $\mathbf{q}$ and $b$. If $u(x)=0$, then we must have $\pi_{j}=0$ for all $j\in{\mathcal{M}}$ and thus (10) holds also when replacing $b$ by $b^{\prime}$, and therefore we can return $x^{\prime}=x$. For the rest of the proof, let us assume $u(x)>0$. Similarly, we can assume that $\mathbf{q}\neq\mathbf{0}$. Note that $\pi_{j}>0$ for every $j\in\operatorname{supp}(x)\cap\operatorname{supp}(\mathbf{q})$. When replacing $b$ by $b^{\prime}$, all inequalities with in (10) with $q_{j}>0$ become strict; the ratio of the right and left hand sides will be $\alpha\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}b/b^{\prime}>1$ for all $j\in\operatorname{supp}(x)\cap\operatorname{supp}(\mathbf{q})$. Also, recall $u(x)=\operatorname{net}_{f}(t)$.
Algorithm 3 alternates between two operations: changing $x$ and $f$ while keeping $\pi$ unchanged; and changing the dual $\pi$ while keeping $x$ and $f$ unchanged. Throughout, it maintains a fitting pair with respect to a decreasing vector $x^{\prime}$, such that the required equalities in (10) are gradually restored.
[…]
Case I:
If $\alpha^{\prime}\leq 1$, then we choose $\lambda\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}(\alpha-1)/(\alpha-\alpha^{\prime})\in[0,1]$ such that $\lambda\alpha+(1-\lambda)\alpha^{\prime}=1$. Then, $x^{\prime\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\lambda x^{\prime}+(1-\lambda)x$ and the flow $f^{\prime\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\lambda f^{\prime}+(1-\lambda)f$ satisfy that $(f^{\prime\prime},\pi)$ is a fitting pair with respect to $x^{\prime\prime}$, and (10) holds for $\pi$, $x^{\prime\prime}$ and $b^{\prime}$. Thus, we can terminate by outputting $x^{\prime\prime}\in\mathcal{GD}^{u}(b^{\prime},\mathbf{q})$.
Case II:
If $\alpha^{\prime}>1$, then we update the dual $\pi$ to $\pi^{\prime}$ as follows. Let $T\supseteq\operatorname{supp}(x^{\prime})$ be the set of nodes that can reach $\operatorname{supp}(x^{\prime})$ in the auxiliary graph $(G,E^{\textrm{aux}}_{f^{\prime},\pi})$. By Lemma 8.3, $t\notin T$. Let us define $\pi^{\prime}$ such that $\pi^{\prime}_{v}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\pi_{v}$ for $v\notin T$ and $\pi^{\prime}_{v}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\delta\pi_{v}$ for $v\in T$ for the largest value $1\leq\delta\leq\alpha^{\prime}$ such that $(f^{\prime},\pi^{\prime})$ is a fitting pair.
We claim that $\delta^{\prime}>1$. Recall that fitting pairs are required to satisfy (9). The sign of $\gamma_{e}\pi^{\prime}_{w}-\pi^{\prime}_{v}$ is the same as the sign of $\gamma_{e}\pi_{w}-\pi_{v}$ for all arcs $e=(v,w)$ with both endpoints inside $T$ or outside $T$. There is no arc $e=(v,w)$ entering or leaving $T$ with $0<f^{\prime}_{e}<u_{e}$, since $e,\overleftarrow{e}\in E^{\textrm{aux}}_{f^{\prime},\pi}$ for all such arcs. For every arc $e=(v,w)$ entering $T$ with $f^{\prime}_{e}=0$, we must have $\gamma_{e}\pi_{w}-\pi_{v}<0$ so that $e\notin E^{\textrm{aux}}_{f^{\prime},\pi}$. Similarly, for every arc $e=(v,w)$ leaving $T$ with $f^{\prime}_{e}=u_{e}$, we must have $\gamma_{e}\pi_{w}-\pi_{v}>0$. Hence, we can always choose $\delta^{\prime}>1$. If $\delta^{\prime}=\alpha^{\prime}$, then $\pi^{\prime}$, $x^{\prime}$, $\mathbf{q}$ and $b^{\prime}$ satisfy (10), since the right hand side for all $j\in\operatorname{supp}(x^{\prime})$ has increased by a factor $\alpha^{\prime}$, and therefore equality holds for all these indices. We can therefore conclude $x^{\prime}\in\mathcal{GD}^{u}(b,\mathbf{q})$ and terminate by outputting $x^{\prime}$.
In case $\delta^{\prime}<\alpha^{\prime}$, then (10) holds with strict inequality for all $j$ for $\pi^{\prime}$, $x^{\prime}$, $\mathbf{q}$, and $b^{\prime}$. Moreover, the ratio of the right hand side and the left hand side is the same for all $j\in\operatorname{supp}(x^{\prime})\cap\operatorname{supp}(\mathbf{q})$, namely, $\alpha^{\prime}/\delta^{\prime}<\alpha$. We now restart the above procedure with $(f^{\prime},\pi^{\prime})$ and $x^{\prime}$ in place of $(f,\pi)$ and $x$.
We claim that the algorithm terminates in a finite number of iterations. This holds because the linear program Pullback-Flow only depends on the sign pattern of $\gamma_{e}\pi_{w}-\pi_{v}$ for the arcs $e=(w,v)$ and on $\operatorname{supp}(x)$. There are $3^{|E|}$ possible sign patterns; the sest $\operatorname{supp}(x)$ is monotone decreasing and therefore there are at most $|{\mathcal{M}}|$ different possibilities. The same configuration of arc sign patterns and $\operatorname{supp}(x)$ cannot occur twice, because $u(x^{\prime})=\operatorname{net}_{f^{\prime}}(t)$ is strictly decreasing in every iteration of the algorithm. Hence, after a finite number of iterations we arrive at $\pi^{\prime}$ and $x^{\prime}$ satisfying (10) for $b^{\prime}$. Since $x^{\prime}$ is decreasing throughout, we have $x^{\prime}\leq x$ for the input supply vector $x$, as required. ∎
Input: $\mathbf{q}\in\operatorname{pdom}(u)$, $b,b^{\prime}>0$, $x\in\mathcal{GD}^{u}(b,\mathbf{q})$.
Output: $0\leq x^{\prime}\leq x$ such that $x^{\prime}\in\mathcal{GD}^{u}(b^{\prime},\mathbf{q})$.
1 Find a fitting pair $(f,\pi)$ with respect to $x$ such that $u(x)=\operatorname{net}_{t}(x)$ and (10) holds;
2 if $u(x)=0$ or $\mathbf{q}=\mathbf{0}$ then return $x$;
3 $\bar{x}\leftarrow x$ ; $\alpha\leftarrow b/b^{\prime}$ ;
4 while $\alpha>1$ do
5       $(x^{\prime},f^{\prime})\leftarrow\textnormal{{Pullback-Flow}}(\bar{x},f,\pi)$ ;
6       $\alpha^{\prime}\leftarrow\alpha\frac{u(x^{\prime})}{u(x)}$ ;
7       if $\alpha^{\prime}\leq 1$ then
8             $\lambda\leftarrow\frac{\alpha-1}{\alpha-\alpha^{\prime}}$ ;
9             $x^{\prime\prime}\leftarrow\frac{\alpha-1}{\alpha-\alpha^{\prime}}x^{\prime}+\frac{1-\alpha^{\prime}}{\alpha-\alpha^{\prime}}\bar{x}$ ;
10             return $x^{\prime\prime}$
11      else
12             $T\leftarrow\left\{v\in V\,:\,v\text{ can reach }\operatorname{supp}(x^{\prime})\text{ in }(G,E^{\textrm{aux}}_{f^{\prime},\pi})\right\}$  ;
13             $\delta\leftarrow\min\left\{\alpha,\min\left\{\frac{\gamma_{e}\pi_{w}}{\pi_{v}}\,:\,e=(v,w)\in\partial^{\mathrm{in}}(T)\right\},\min\left\{\frac{\pi_{v}}{\gamma_{e}\pi_{w}}\,:\,e=(v,w)\in\partial^{\mathrm{out}}(T)\right\}\right\}$;
14             if $\delta^{\prime}=\alpha^{\prime}$ then return $x^{\prime}$;
15             for $v\in T$ do $\pi_{v}\leftarrow\delta^{\prime}\pi_{v}$;
16             $\alpha\leftarrow\alpha^{\prime}/\delta^{\prime}$ ; $\bar{x}\leftarrow x^{\prime}$ ;
17
18
Algorithm 3 Budget-Decrease
We next prove property (i) in Definition 2.14.
Lemma 8.8.
Let $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ be a generalized network utility function and $\mathbf{q},\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$ such that $0\leq\mathbf{q}^{\prime}\leq\mathbf{q}$, and let $x\in\mathcal{GD}^{u}(b,\mathbf{q})$. Then, there exists $x^{\prime}\in\mathcal{GD}^{u}(b,\mathbf{q}^{\prime})$ such that $x^{\prime}_{j}\leq x_{j}$ whenever $q^{\prime}_{j}=q_{j}$.
Proof.
Clearly, it suffices to show that the statement holds when $q^{\prime}_{k}<q_{k}$ for a single good $k\in{\mathcal{M}}$. Again, we start from fitting pair $(f,\pi)$ with respect to $x$ as in Lemma 8.6. Similarly to Lemma 8.7, we may assume that $u(x)>0$ and $\mathbf{q}\neq\mathbf{0}$. Let us now replace $\mathbf{q}$ by $\mathbf{q}^{\prime}$. The condition (10) can only be violated for $j=k$. If it is not violated, then we may return $x^{\prime}=x$; this is always the case when $\pi_{k}=0$. For the rest of the proof, assume that the (10) is violated for $k$ with the cost $q^{\prime}_{k}$. Note that $\pi_{j}=\beta q^{\prime}_{j}$ for every $j\neq k$, where $\beta\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}u(x)/b$; we now have $\pi_{k}>\beta q^{\prime}_{k}\geq 0$. This value of $\beta$ will be fixed throughout the argument.
[…]
We can again argue that this algorithm terminates in a finite number of iterations. Note that $\operatorname{supp}(x^{\prime})$ is monotone decreasing throughout, and $\pi^{\prime}_{j}=\beta q^{\prime}_{j}$ is maintained for all $j\in\operatorname{supp}(x^{\prime})\setminus\{k\}$. For goods with $x^{\prime}_{j}=0$, $\pi^{\prime}_{j}$ may decrease, and therefore $\pi^{\prime}_{j}\leq\beta q_{j}$ for these goods. At termination, we have $\pi^{\prime}_{j}=\beta q^{\prime}_{j}$ for all $j\in\operatorname{supp}(x^{\prime})$ — including $j=k$ — and $\pi^{\prime}_{j}\leq\beta q^{\prime}_{j}$ whenever $x^{\prime}_{j}=0$. Moreover, note that $x^{\prime}_{j}\leq x_{j}$ for all $j\neq k$.
Throughout the algorithm, $u(x^{\prime})$ is non-decreasing. Hence, $u(x^{\prime})/b\geq\beta$ holds. By setting $\hat{b}>b$ such that $u(x^{\prime})/\hat{b}=\beta$, (10) holds for $x^{\prime}$, $\pi^{\prime}$, $q^{\prime}$ and $\hat{b}$. Hence, $x^{\prime}\in\mathcal{GD}_{u}(\hat{b},\mathbf{q}^{\prime})$. Applying now Lemma 8.7 for $x^{\prime}$, the prices $\mathbf{q}^{\prime}$ and budgets $\hat{b}\geq b$, we can find $x^{\prime\prime}\in\mathcal{GD}_{u}(b,\mathbf{q}^{\prime})$ such that $x^{\prime\prime}\leq x^{\prime}$. Note that $x^{\prime\prime}_{j}\leq x^{\prime}_{j}\leq x_{j}$ for all $j\neq k$, and therefore $x^{\prime\prime}$ satisfies the requirements of the Lemma.
[…]
This concludes the proof of Theorem 1.5, namely, that all generalized network utilities are Gale-substitutes. Nevertheless, they are not $\Sigma$-Gale-substitutes, and so Theorem 1.3 is only applicable in the case when the generalized network utilities are non-satiable. Here, we also show Theorem 1.6, namely, that the desired utility bound holds for all generalized network utilities if we consider thrifty competitive equilibria.
See 1.6
Proof sketch.
We only need to slightly modify the argument in Section 6. Property ($\Sigma$) is only used in the proof of Lemma 3.2. We show that for generalized network utilities, if $x\in{\mathcal{D}}^{u}(\mathbf{p},b)$ for $\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ and further $x$ minimizes $\left\langle\mathbf{p},x\right\rangle$ in ${\mathcal{D}}^{u}(\mathbf{p},b)$, then there exists $y\in\mathcal{GD}^{u}(\mathbf{p},b)$ such that $y\leq x$. From this property, we can derive Lemma 6.1 for thrifty competitive equilibria; the rest of the proof remains identical to the proof of Theorem 1.3.
Consider now (6) in the proof of Lemma 3.2. If $\lambda>0$, the same argument is applicable. Consider now the case $\lambda=0$. This means that the constraint $\left\langle\mathbf{p},x\right\rangle\leq b$ is not active in the program defining ${\mathcal{D}}^{u}(\mathbf{p},b)$; that is, $\max_{z\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\{u(z)\,:\,\left\langle\mathbf{p},z\right\rangle\leq b\}=\max_{z\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}u(z)=v^{\star}$. In particular, $u(x)=v^{\star}$. The budget minimization property means that $x$ is an optimal solution to the program
$\min\leavevmode\nobreak$ $\mathbf{p}^{\top}x$ (11)
$\operatorname{net}_{f}(t)$ $\geq v^{\star}$
$\operatorname{net}_{f}(v)$ $\geq-x_{v}\quad\forall v\in S$
$\operatorname{net}_{f}(v)$ $\geq 0\quad\forall v\in V\setminus(S\cup\{t\})$
00 0 $\leq f\leq h\,.$
From here, using LP duality and Lemma 8.1, we can see that there exists a supergradient $\mathbf{g}\in\partial^{\wedge}u(x)$ and $\mu\in\mathbb{R}_{\geq 0}$ such that
$\mu g_{j}\leq p_{j}\,,\quad\forall j\in{\mathcal{M}}\,,\,\mbox{with equality if }x_{j}>0\,.$
If $\mu>0$, then we can use the same argument for $\lambda=1/\mu$. If $\mu=0$, given that $\mu$ acts as the dual variable of $\operatorname{net}_{f}(t)\geq v^{\star}$, which is no longer active, the optimal value of (11) becomes 00 since $\mathbf{0}$ is a feasible allocation of $u(\cdot)$ according to our definition. This implies the existence of a zero-cost optimal allocation $x$: $u(x)=v^{*}$ and $\mathbf{p}^{\top}x=0$, which indicates that $x$ belongs to both ${\mathcal{D}}^{u}(\mathbf{p},b)$ and $\mathcal{GD}^{u}(\mathbf{p},b)$. Thus, the result follows. ∎
9 Discussion
In this paper, we compared competitive equilibrium and Nash welfare maximizing solutions in the Fisher market setting. We identified Gale-substitutes as a broad family of utility functions where the Nash welfare uniformly guarantees every agent at least half of their maximum utility at any competitive equilibrium. We also introduced generalized network utilities. This is a very general class with a rich combinatorial structure and may be a relevant example in various market and fair division settings.
Gale-substitutes may be a class of concave functions worth investigating on its own. In the context of indivisible goods, the analogous class of gross substitute utilities, introduced by Keslo and Crawford (1982), turned out to be the same as the important class of $M^{\natural}$-concave utilities in discrete convex analysis Fujishige and Yang (2003). In particular, one may wonder whether the Gale-substitutes property implies submodularity, that is, $u(x)+u(y)\geq u(x\vee y)+u(x\wedge y)$ for any $x,y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, where $x\vee y$ and $x\wedge y$ denote the pointwise maximum and minimum of the two vectors, respectively. This property trivially holds for separable utilities and can also be shown for generalized network utilities.
We note that one can construct an example of a concave quadratic function that is submodular but not Gale-substitutes. On the other hand, it is unclear whether Theorem 1.6 would remain true for submodular utilities instead of Gale-substitutes. One would also be interested in whether Theorem 1.6 holds for matching utilities, which, like submodular utilities, are not necessarily Gale-substitutesl(see Example A.11). Another natural question related to the well-studied notion of weak gross substitutes (WGS) utilities: Gale-substitutes includes utility functions that are not WGS, such as concave separable; but it is not clear whether all WGS utilities are also Gale-substitutes.
Gale-substitutes utilities may also have potential algorithmic applications. The WGS property leads to natural tâtonnement and auction mechanisms for finding competitive equilibria, see, e.g., Arrow et al. (1959), Arrow and Hurwicz (1960), Bei et al. (2019a), Codenotti et al. (2005b, c), Garg et al. (2023a). The Gale-substitutes property may be useful to design simple such mechanisms and combinatorial algorithms for solving the convex program (EG). For the linear case, Devanur et al. (2008) gave a primal-dual algorithm that was followed by a series of works on combinatorial algorithms for various market settings; see, e.g., Chaudhury et al. (2022), Cole and Gkatzelis (2015), Duan and Mehlhorn (2015), Garg and Végh (2019).
Appendix A Examples
In this section, we present multiple examples mentioned in the paper.
A.1 Connectivity of equilibria with bounded linear utilities
We first investigate the (dis)connectivity of competitive equilibria with bounded linear utilities. We begin by demonstrating that the thrifty competitive equilibria with bounded linear utilities might not be connected.
Example A.1 (Disconnectivity of thrifty competitive equilibrium).
There are two agents and two goods. Each agent has a unit budget. The first agent’s utility is $u_{1}(x_{1})=1.3\min\{x_{11},0.8\}+0.45x_{12}$ and the second agent’s utility is $u_{2}(x_{2})=0.01\min\{x_{21},0.3\}+2\min\{x_{22},0.8\}$.
First, we show there are at least two equilibria. The first equilibrium is with $\mathbf{p}=(1,1)$. Under this price, $x_{11}=0.8$, $x_{12}=0.2$, $x_{21}=0.2$, $x_{22}=0.8$. Another competitive equilibrium is with $\mathbf{p}=(1.3,0.45)$. Under this price, $x_{11}=0.7$, $x_{12}=0.2$, $x_{21}=0.3$, $x_{22}=0.8$.
Now, we show the second competitive equilibrium is isolated under the thrifty assumption.
First, we show the price is isolated. Suppose there is a nearby price $\mathbf{p}^{\prime}=(1.3+\epsilon_{1},0.45+\epsilon_{2})$. For this price, the thrifty demand of the second agent remains the same: $x_{21}=0.3$, $x_{22}=0.8$. To maintain equilibrium, the first agent should receive the same allocation. As $1.3\times 0.7+0.45\times 0.2=1$, $\epsilon_{1}$ and $\epsilon_{2}$ can not be both positive or negative. However, if one $\epsilon$ is negative, the first agent will prefer this item more at $\mathbf{p}^{\prime}$. This implies the allocation of the other item of the first agent is 00. This leads to a contradiction.
Second, we show the allocation is isolated. Suppose there is a nearby equilibrium allocation of the second agent $x^{\prime}_{2}=(x_{21}-\epsilon_{1},x_{22}-\epsilon_{2})$ with at least one of $\epsilon_{1}$ and $\epsilon_{2}$ is strictly positive. In this case, both agents should spend all their budgets as they are non-satiable. Therefore, the equilibrium prices satisfy $p^{\prime}_{1}+p^{\prime}_{2}=2$. For the first agent, the allocation is $x^{\prime}_{11}=0.7+\epsilon_{1}$, $x^{\prime}_{12}=0.2+\epsilon_{2}$. Note that both $\epsilon_{1}$ and $\epsilon_{2}$ are very tiny. This implies $p^{\prime}_{1}:p^{\prime}_{2}=1.3:0.45$. Therefore, $p^{\prime}_{1}=\frac{2}{1.3+0.45}\times 1.3$ and $p^{\prime}_{2}=\frac{2}{1.3+0.45}\times 0.45$. However, the first agent cannot afford these allocations under these prices.
The next example shows that competitive equilibria are not connected without the thrifty assumption.
Example A.2 (Disconnectivity of competitive equilibria).
There are three agents and three goods. Each agent has a unit budget. The first agent’s utility is $u_{1}(x_{1})=1.3\min\{x_{11},0.8\}+0.45x_{12}$, the second agent’s utility is $u_{2}(x_{2})=0.01\min\{x_{21},0.3\}+2\min\{x_{22},0.8\}+10^{-10}x_{23}$, and the third agent’s utility is $u_{3}(x_{3})=x_{33}$.
First, we show there are at least two equilibrium. The first equilibrium is with $\mathbf{p}=(1,1,1)$, and the allocation is $x_{1}=(0.8,0.2,0)$, $x_{2}=(0.2,0.8,0)$, and $x_{3}=(0,0,1)$. The second equilibrium is with $\mathbf{p}=(1.3,0.45,1.25)$ and the allocation is $x_{1}=(0.7,0.2,0)$, $x_{2}=(0.3,0.8,0.2)$, and $x_{3}=(0,0,0.8)$.
In order to show our results, we first state the following properties: for any equilibrium,
  1. 1.
    $p_{3}>0$ as the third agent is only interested in the third good;
  2. 2.
    $x_{13}=0$ as the first agent shows no interest in the third good and its price is strictly positive;
  3. 3.
    $p_{1}>0$ and $p_{2}>0$ as, otherwise, these goods will be over-demanded;
  4. 4.
    $x_{31}=x_{32}=0$ as the third agent shows no interest in the first two goods and their prices are strictly positive;
  5. 5.
    $p_{1}+p_{2}+p_{3}=3$ as all agents are non-satiated.
Now, we show the second equilibrium is isolated.
First, we demonstrate that the price is isolated. Suppose there is a nearby equilibrium price $\mathbf{p}^{\prime}=(1.3+\epsilon_{1},0.45+\epsilon_{2},1.25+\epsilon_{3})$. For this price, the allocation of the second agent on the first two items is the same. Therefore, the first agent should receive the same allocation. As $1.3\times 0.7+0.45\times 0.2=1$, $\epsilon_{1}$ and $\epsilon_{2}$ can not be both positive or negative. Similar to the previous example, if one of $\epsilon_{1}$ and $\epsilon_{2}$ is negative, then the first agent will prefer this item more at this nearby price. This implies the allocation of the other item of the first agent is 00. This leads to the conclusion $\epsilon_{1}=\epsilon_{2}=\epsilon_{3}=0$.
Second, we show the allocation is isolated. Suppose there is a nearby equilibrium allocation $\mathbf{x}^{\prime}$ with the equilibrium $\mathbf{p}^{\prime}$. As $x_{11}<0.8$, $p^{\prime}_{1}:p^{\prime}_{2}=1.3:0.45$.  [^2] Otherwise, the allocation of the first agent will be changed significantly. Additionally, $p^{\prime}_{1}\times x^{\prime}_{11}+p^{\prime}_{2}\times x^{\prime}_{12}=1$. This implies $p_{1}^{\prime}$ is very close to $p_{1}$ and $p_{2}^{\prime}$ is very close to $p_{2}$: if $\|x^{\prime}-x\|_{\infty}\leq\epsilon$, then $|p_{1}^{\prime}-p_{1}|\leq O(\epsilon)$ and $|p_{2}^{\prime}-p_{2}|\leq O(\epsilon)$.  [^3] Note that $p^{\prime}_{1}:p^{\prime}_{2}=1.3:0.45=p_{1}:p_{2}$. By budget constraint, $\frac{1.3}{0.45}p^{\prime}_{2}x^{\prime}_{11}+p^{\prime}_{2}x^{\prime}_{12}=\frac{1.3}{0.45}p_{2}x_{11}+p_{2}x_{12}$. Therefore, $p_{2}^{\prime}-p_{2}=\frac{p^{\prime}_{2}(\frac{1.3}{0.45}(x_{11}-x^{\prime}_{11})+(x_{12}-x^{\prime}_{12}))}{\frac{1.3}{0.45}x_{11}+x_{12}}$. We have a similar result for $p_{1}^{\prime}-p_{1}$. This also implies $|p_{3}^{\prime}-p_{3}|\leq O(\epsilon)$ as $p^{\prime}_{1}+p^{\prime}_{2}+p^{\prime}_{3}=3$. Since price is isolated, $p^{\prime}_{1}=p_{1}$, $p^{\prime}_{2}=p_{2}$, and $p^{\prime}_{3}=p_{3}$. Under this price, $\mathbf{x}$ is the only allocation.
A.2 Non-monotonicity of Gale Demands
The next example shows the somewhat counterintuitive property that increasing the prices may lead to an increase in the Gale utility.
Example A.3 (Non-monotonicity of Gale demands).
There are two goods and a single agent $i$ with $b_{i}=1$. The agent’s utility function is $u_{i}(x_{i})=\min\{x_{i1}+x_{i2},2+0.1\cdot x_{i1}+0.2\cdot x_{i2}\}$. When $\mathbf{q}=\left(\frac{0.1}{2/0.9+0.2\cdot 0.01},\frac{0.2}{2/0.9+0.2\cdot 0.01}\right)$, the associated Gale demand is $\mathcal{GD}_{i}(b_{i},\mathbf{q})=\{(2/0.9,0.01)\}$ and the utility is $u_{i}(\mathcal{GD}_{i}(b_{i},\mathbf{q}))=2/0.9+0.2\cdot 0.01$. When $\mathbf{q}^{\prime}=\left(\frac{1}{0.01+2/0.85},\frac{0.2}{0.01+2/0.85}\right)$, the Gale demand becomes $\mathcal{GD}_{i}(b_{i},\mathbf{q}^{\prime})=\{(0.01,2/0.85)\}$ and $u_{i}(\mathcal{GD}_{i}(b_{i},\mathbf{q}^{\prime}))=2/0.85+0.01$. Clearly, $\mathbf{q}<\mathbf{q}^{\prime}$, yet $u_{i}(\mathcal{GD}_{i}(b_{i},\mathbf{q}))<u_{i}(\mathcal{GD}_{i}(b_{i},\mathbf{q}^{\prime}))$.
A.3 Tight Example on Price of Anarchy
The next example shows that the bound in Theorem 1.1 is tight.
Example A.4 (Tight Example on Price of Anarchy).
We consider the case with two agents and two goods. The first agent has a budget of $1$, and desires the first good: $u_{1}(x_{1})=x_{11}$. The second agent has a budget of $\mathrm{e}-1$ and has a utility function $u_{2}(x_{2})=\varepsilon x_{21}+\min\{x_{22},1\}$, where $\varepsilon>0$ is sufficiently small. There is a competitive equilibrium with prices $p_{1}=\mathrm{e}$ and $p_{2}=0$, and allocations $x_{1}=(\frac{1}{\mathrm{e}},0)$ and $x_{2}=(\frac{\mathrm{e}-1}{\mathrm{e}},1)$. The Nash welfare objective for this allocation is $\left(\frac{1}{\mathrm{e}}(1+\varepsilon\cdot\frac{\mathrm{e}-1}{\mathrm{e}})^{\mathrm{e}-1}\right)^{\frac{1}{\mathrm{e}}}$. In the allocation maximizing Nash welfare, we have $y_{1}=(1,0)$ and $y_{2}=(0,1)$, where the Nash welfare objective is $1$. When $\varepsilon\rightarrow 0$, the difference of these two scenarios converges to $\left(\frac{1}{\mathrm{e}}\right)^{\frac{1}{\mathrm{e}}}$.
A.4 Utility Difference in Different Equilibria
The following example shows agents might receive significantly different utilities in different equilibria.
Example A.5 (Utility Difference in Different Equilibria).
Consider a scenario involving $n$ agents and $n$ goods, where each agent has a unit budget. For $11\leq i\leq n-$ 1, agent $i$’s utility is defined as $u_{i}(x_{i})=\min\{x_{ii},1\}+\epsilon x_{in}$, while agent $n$’s utility is given by $u_{i}(x_{i})=x_{nn}$. There exist at least two competitive equilibria. In the first competitive equilibrium, all goods are priced at $1$, and each good $i$ is allocated to agent $i$ for $1\leq i\leq n$, resulting in a utility of $1$ for agent $n$. However, in the second competitive equilibrium, prices are set to 00 for the first $n-1$ goods and the price of the $n$-th good is $n$. Consequently, goods $1$ to $n-1$ are allocated to agents $1$ to $n-1$ respectively, while the $n$-th good is uniformly allocated to all agents. In this equilibrium, agent $n$ can only attain a utility of $\approx\frac{1}{n}$.
A.5 Tight Example on Demand-Approximate CE
The following example shows Theorem 1.2 is tight.
Example A.6 (Tight Example on Demand-Approximate CE).
Consider the case with two agents and one good. Both agents have unit budgets. The first agent’s utility on this good is linear: $u_{1}(x_{1})=x_{1}$. The second agent’s utility is defined as follows: $u_{2}(x_{2})=\alpha x_{2}$ for $x_{2}\leq\frac{1}{\alpha}$ and $u_{2}(x_{2})=1+\beta(x_{2}-\frac{1}{\alpha})$ for $x_{2}>\frac{1}{\alpha}$, where $\alpha$ is a large number and $\beta\leq\alpha$. When $\beta=\frac{1-\frac{1}{\alpha}}{1-\frac{2}{\alpha}}$, the Nash welfare maximizing allocation is $x_{1}=1-\frac{1}{\alpha}$ and $x_{2}=\frac{1}{\alpha}$, with the dual variable $q$ equal to $\frac{1}{x_{1}}=\frac{1}{1-\frac{1}{\alpha}}$. Under this price, agent $2$ can obtain $x^{\prime}_{2}=1-\frac{1}{\alpha}$ with his/her budget, resulting in a utility of $1+\beta(x^{\prime}_{2}-\frac{1}{\alpha})=2-\frac{1}{\alpha}$. In contrast, in the Nash welfare maximizing allocation, agent $2$ only receives a utility of $1$.
A.6 (In)-approximate Nash Welfare utility by Equilibrium utility
In this paper, we show for $\Sigma$-Gale-substitues utilities, the utility of any agent at the Nash welfare maximizing allocation is at least half the maximum utility they may get at any CE. It’s natural to ask, inversely, if the utility agent can attain at any CE can approximate the utility they receive at Nash welfare maximizing allocation. We show a constant approximation is not achievable by providing an example in which the CE is unique and there exists an agent whose utility at this CE is much worse than his/her utility at Nash welfare maximizing allocation.
Example A.7.
Consider a scenario with $n+2$ agents and $n+1$ goods. Set $\epsilon<1$. For $1\leq i\leq n$, let $u_{i}(x_{i})=\min\{x_{ii},0.5\}+\epsilon x_{i(n+1)}$. Additionally, set $u_{n+1}(x_{n+1})=x_{(n+1)(n+1)}$ and $u_{n+2}(x_{n+2})=\sum_{j=1}^{n}x_{ij}$. We first demonstrate the unique competitive equilibrium is achieved when $p_{j}=\frac{2}{n}$ for $1\leq j\leq n$ and $p_{n+1}=n$. In this equilibrium, $x_{ii}=0.5$ and $x_{i(n+1)}=\frac{n-1}{n^{2}}$ for $1\leq i\leq n$; $x_{(n+1)(n+1)}=\frac{1}{n}$; and $x_{(n+2)j}=0.5$ for $1\leq j\leq n$.
We now show this is the only competitive equilibrium. First, for all $j:1\leq j\leq n$, $p_{j}>0$. Otherwise, agent $n+2$ would receive infinite goods. Given $p_{j}>0$ for any $1\leq j\leq n$, $x_{(n+2)j}\geq 0.5$. Consequently, the first $n$ goods share the same prices and $10.5p_{j}n\leq\sum_{j}x_{(n+2)j}p_{j}=$ 1 which implies $p_{j}\leq\frac{2}{n}$. If $p_{j}<\frac{2}{n}$ for $1\leq j\leq n$, then $p_{n+1}=\sum_{i}b_{i}-\sum_{j}p_{j}>n$. In this case, $x_{ii}=0.5$ for $1\leq i\leq n$ and $10.5p_{j}n=$ 1 for $1\leq j\leq n$, which implies the result.
Therefore, agent $n+1$ receives utility $\frac{1}{n}$ in the unique competitive equilibrium. However, for a small enough $\epsilon$, this agent receives $1$ in the maximal Nash welfare solution.
A.7 (In)-approximation for general utilities
Our next example demonstrates if the utility function is not Gale substitutes, then agents’ competitive equilibrium utilities can be much higher than the Nash welfare maximizing solution.
Example A.8.
There are three types of agents (A, B, C): $1$ type-A agent; $n^{2}$ type-B agents; $n$ type-C agents; and four types of goods (O, P, R, S): $1$ type-O good; $n$ type-P goods; $n^{2}$ type-R goods; $n$ type-S goods. In total, there are $N=n^{2}+n+1$ agents and $M=n^{2}+2n+1$ goods.
The type-A agent is interested in the type-O good and type-P goods. Type-A agent has a bounded linear utility with equal preferences over these type-O and type-P goods: $v_{ij}=1$ for $i=A$ and $j\in O\cup P$. This agent has an upper bound of 1 on the utility of each good: $u_{A}(x_{A})=\sum_{j\in O\cup P}\min\{x_{Aj},1\}$.
For type-B agents, let’s label each of these agents by $(s,t)$ for $s\in[1,n]$ and $t\in[1,n]$ and also label each of type-R goods by $(s,t)$ for $s\in[1,n]$ and $t\in[1,n]$. Agent $(s,t)$ is interested in the corresponding $(s,t)$ good in R and the $s$-th good in P. The utility function of agent $i=(s,t)$ is $\min\{x_{iP_{s}}+\frac{n}{n+1},x_{iR_{s,t}}\}$, where $x_{iP_{s}}$ is agent $i$’s allocation of $P_{s}$ and $x_{iR_{s,t}}$ is agent $i$’s allocation of $R_{s,t}$.
For type-C agents, the $s$-th type-C agent is only interested in the $s$-th type-S good and all type-R good with label $(s,\cdot)$. This agent also has a bounded linear utility with equal preference ($v_{ij}=1$) over these goods and an upper bound, $l_{ij}=1$, on the utility of each good.
We first consider the maximal Nash welfare solution. Here, $k$ is a parameter, which will be specified later.
The type-A agent receives type-O good and $k+\frac{1}{n+1}$ portion of type-P goods, his/her utility will be $1+nk+\frac{n}{n+1}$.
Each type-B agent obtains $1-\frac{k}{n}$ portion of corresponding type-R good and $\frac{1}{n+1}-\frac{k}{n}$ portion of corresponding type-P good, yielding a utility of $1-\frac{k}{n}$.
The type-C agents receive the remaining goods, resulting in a utility of $1+k$.
The goal is to maximize $(1+nk+\frac{n}{n+1})(1-\frac{k}{n})^{n^{2}}(1+k)^{n}$. By calculation, the optimal $k$ is found to be no larger than $n^{-\frac{1}{2}}$.  [^4] First, we take the log and calculate the gradient, which is $\frac{n}{1+nk+\frac{n}{n+1}}+\frac{n}{1+k}-\frac{n}{1-\frac{k}{n}}$. Note that this is a decreasing function with $k$. So, in order to show the optimal $k$ is no larger than $n^{-\frac{1}{2}}$, we only need to show when $k=n^{-\frac{1}{2}}$, $\frac{n}{1+nk+\frac{n}{n+1}}+\frac{n}{1+k}\leq\frac{n}{1-\frac{k}{n}}$. This is true as $\frac{n}{1-\frac{k}{n}}-\frac{n}{1+k}=\frac{(n+1)k}{1+k-\frac{k}{n}-\frac{k^{2}}{n}}\geq nk=\frac{1}{k}\geq\frac{n}{1+nk+\frac{n}{n+1}}$. Therefore, the type-A agent receives at most $1+\sqrt{n}+\frac{n}{n+1}$ utility in the maximal Nash welfare solution.
On the other hand, let’s consider the (thrifty) competitive equilibrium where all agents have unit budgets.
The prices of type-O, P, and S goods are 00. and prices of type-R goods are $\frac{n+1}{n}$. In this case, the type-A agent will get the type-O good and type-P goods, resulting in a utility of $n+1$. Each type-B agent receives $\frac{n}{n+1}$ portion of the corresponding type-R good, leading to a utility of $\frac{n}{n+1}$. The type-C agent acquires $1$ portion of the corresponding type-S good and the rest of type-R goods, resulting in a utility of $1+\frac{n}{n+1}$.
Comparing the type-A agent’s utility in the two scenarios, the type-A agent achieves at most $1+\sqrt{n}+\frac{n}{n+1}$ utility in the maximal Nash welfare solution, whereas the agent attains $n+1$ utility in the (thrifty) competitive equilibrium.
A.8 Tight Example of Theorem 1.3
The following example shows the bound in Theorem 1.3 is tight.
Example A.9.
Consider the case that there are $n$ agents and $n$ goods. All agents have bounded linear utilities. Additionally, all agents have the same preferences ($v_{ij}=1$) on all goods except for the first good. For the first good, only the first agent likes the first good ($v_{11}=0$) and all other agents don’t like this good $v_{i1}=0$ for $i\neq 1$. Additionally, we assume the upper bound on the utility of each good of each agent is $1$, $l_{ij}=1$ for all $i$ and $j$.
In this case, for the maximal Nash welfare solution, the utilities of all agents equal $1$. However, there exists a competitive equilibrium, at which the price of the first good equals 00 and the prices for all other goods equal $\frac{n}{n-1}$. Then, the first agent can get the first good and $\frac{n-1}{n}$ fraction of the second good, and achieve a utility of $1+\frac{n-1}{n}$ which tends to $2$ in the case $n\rightarrow\infty$.
A.9 Example with Leontief-free utility functions
Not all Leontief-free utility functions satisfy property ($\Sigma$). The subsequent example illustrates a scenario where an agent can attain significantly more utility in competitive equilibrium compared to the Nash welfare maximizing allocation when agents have Leontief-free utilities.
Example A.10.
Consider the scenario with $n$ goods and $n+1$ agents. Agent $i=1,2,\cdots,n$ only likes good $i$, has budget $b_{i}=1$ and a linear utility function $u_{i}(x_{i})=x_{ii}$. Agent $n+1$ has a substantially larger budget $b_{k+1}=B$, and a utility function $\min\{\sum_{i=1}^{k}x_{(n+1)i},(n-1)(1-\epsilon)\}$ with a small $\epsilon>0$.
There is a competitive equilibrium where the prices are $p_{1}=1$ and $p_{2}=p_{3}=\cdot=p_{n}=\frac{1}{\epsilon}$. Agent $1$ obtains $x_{11}=1$, the other agents get $x_{ii}=\epsilon$ for $i=2,\cdots,n$, and agent $n+1$ gets $x_{(n+1)1}=0$ and $x_{(n+1)j}=1-\epsilon$ for $j=2,\cdots,n$. The competitive equilibrium holds if $B>\frac{(n-1)(1-\epsilon)}{\epsilon}$.
On the other hand, in the Nash welfare maximizing allocation, when $B$ is big enough, agent $n+1$ receives (approximately) $\frac{(1-\epsilon)(n-1)}{n}$ of each good, with the remaining goods shared among the remaining agents. Agent $1$ receives a utility of $\epsilon+\frac{1-\epsilon}{n}$, which is significantly less than the utility received in the competitive equilibrium.
In this case, agent $n+1$’s utility doesn’t satisfy property ($\Sigma$). When $\mathbf{p}=\mathbf{0}$, $x_{n+1}=(0,1-\epsilon,1-\epsilon,\cdots,1-\epsilon)\in\mathcal{GD}_{n+1}(\mathbf{p},B)$. However, when $\mathbf{p}=(1,\frac{1}{\epsilon},\frac{1}{\epsilon},\cdots,\frac{1}{\epsilon})$, for any $x_{n+1}\in\mathcal{GD}_{n+1}(\mathbf{p},B)$, $x_{(n+1)1}>0$, which violates property ($\Sigma$).
A.10 Example with matching utility functions
The following example illustrates that matching utilities are not Gale-substitutes in general.
Example A.11.
Consider the scenario with $3$ goods and an agent with unit budget. The agent’s utility is a matching utilities, with preferences for the goods being $1$, $2.5$, and $0.7$ respectively. The initial prices for the goods are $0.3$, $1.2$, and $0.1$. In this case, the Gale demand is
${\arg\max}_{x=(x_{1},x_{2},x_{3}):x_{1}+x_{2}+x_{3}\leq 1}\{\log(x_{1}+2.5x_{2}+0.7x_{3})-(0.3x_{1}+1.2x_{2}+0.1x_{3})\},$
which gives the solution $(0,103/198,95/198)$. Now we keep the prices of the first two goods unchanged and increase the price of the third good from $0.1$ to $0.6$. By calculation, under the new prices, the new Gale demand is $(5/9,4/9,0)$. The Gale demand of good $2$ decreases, which implies that it is not substitutes.
Appendix B Proof of Lemma 8.1
See 8.1
Proof.
First, assume that $\pi\in\mathbb{R}_{\geq 0}^{\ell}$ is an optimal solution to $\operatorname{Dual}(x)$ for a given $x\in\mathbb{R}^{m}$. We need to show that $g=B^{\top}\pi$ is a supergradient at $x$, that is, $\varphi(t)\leq\varphi(x)+\left\langle g,t-x\right\rangle$ for any $t\in\mathcal{D}$.
The programs $\operatorname{Dual}(x)$ and $\operatorname{Dual}(t)$ have the same feasible region, with linear objective functions $\mathbf{B}x+b$ and $\mathbf{B}t+b$, respectively. By weak duality, the cost of every feasible dual solution is at least the primal optimum value, thus, $\varphi(t)\leq t^{\top}\mathbf{B}\pi+b^{\top}\pi=g^{\top}t+b^{\top}\pi$. The statement follows as
$\varphi(t)\leq g^{\top}t+b^{\top}\pi=g^{\top}(t-x)+x^{\top}\mathbf{B}\pi+b^{\top}\pi=\varphi(x)+g^{\top}(t-x)\,,$
noting also that $\varphi(x)=x^{\top}\mathbf{B}\pi+b^{\top}\pi$ since $\pi$ is an optimal solution to $\operatorname{Dual}(x)$, and $\varphi(x)$ is the optimum value.
Conversely, we need to show that if $g\in\partial^{\wedge}\varphi(x)$, then there exists a dual optimal solution $\pi$ such that $g=\mathbf{B}^{\top}\pi$. Let $z$ be a primal optimal solution to $\operatorname{Primal}(x)$. Using complementary slackness, we can turn $\operatorname{Dual}(x)$ into a feasibility system. Namely, $g=\mathbf{B}^{\top}\pi$ for a dual optimal solution if and only if
$\mathbf{B}^{\top}\pi$ $=g$ (12)
$\mathbf{A}_{i}^{\top}\pi$ $=c_{i}\,,\quad\forall i:z_{i}>0$
$\mathbf{A}_{i}^{\top}\pi$ $\geq c_{i}\,,\quad\forall i:z_{i}=0$
$\pi_{j}$ $=0\,,\quad\forall j:(\mathbf{A}z)_{j}<(\mathbf{B}x+b)_{j}$
$\pi_{j}$ $\geq 0,\quad\forall j:(\mathbf{A}z)_{j}=(\mathbf{B}x+b)_{j}\,.$
Here, $\mathbf{A}_{i}$ denotes the $i$-th column of $\mathbf{A}$. By Farkas’s lemma, if (12) is not feasible, then the following system has a feasible solution:
$(\mathbf{A}w)_{j}+(\mathbf{B}r)_{j}$ $\leq 0\,,\quad\forall j:(\mathbf{A}z)_{j}=(\mathbf{B}x+b)_{j}$
$w_{i}$ $\geq 0\,,\quad\forall i:z_{i}=0$
$c^{\top}w+g^{\top}r$ $>0\,.$
For a small enough $\varepsilon>0$, $z+\varepsilon w$ is feasible to the system $\operatorname{Primal}(x-\varepsilon r)$. Hence, for $t=x-\varepsilon r$, $\varphi(t)\geq c^{\top}(z+\varepsilon w)=c^{\top}z+\varepsilon c^{\top}w$. Using $c^{\top}z=\varphi(x)$ by the choice of $z$ as an optimal solution to $\operatorname{Primal}(x)$, and the bound $c^{\top}w>-g^{\top}r$, we get $\varphi(t)>\varphi(x)-\varepsilon g^{\top}r=\varphi(x)+\left\langle g,t-x\right\rangle$, a contradiction to $g\in\partial^{\wedge}\varphi(x)$. ∎

### v3

@@
 Abstract
-We explore the relationship between two popular concepts in the allocation of divisible items: competitive equilibrium (CE) and allocations that maximize Nash welfare, i.e., allocations where the weighted geometric mean of the utilities is maximal.
-When agents have homogeneous concave utility functions, these two concepts coincide: the classical Eisenberg–Gale convex program that maximizes Nash welfare over feasible allocations yields a competitive equilibrium.
-However, these two concepts diverge for non–homogeneous utilities.
+We study the relationship between two central concepts in the allocation of divisible goods: competitive equilibrium (CE) and allocations that maximize Nash welfare, i.e., allocations where the weighted geometric mean of the utilities is maximal.
+When agents have homogeneous concave utility functions, these concepts coincide: the classical Eisenberg–Gale convex program that maximizes Nash welfare over feasible allocations yields a competitive equilibrium.
+However, they diverge for non–homogeneous utilities.
 From a computational perspective, maximizing Nash welfare amounts to solving a convex program for any concave utility functions, whereas computing CE becomes PPAD-hard already for separable piecewise linear concave (SPLC) utilities.
-We introduce the concept of Gale-substitute utility functions, which is an analogue of the weak gross substitutes (WGS) property for the so-called Gale demand system.
+We introduce the concept of Gale-substitute utility functions, an analogue of the weak gross substitutes (WGS) property for the so-called Gale demand system.
 For Gale-substitutes utilities, we show that any allocation maximizing Nash welfare provides an approximate-CE with surprisingly strong guarantees, where every agent gets at least half the maximum utility they can get at any CE, and is approximately envy-free.
@@
 This class includes SPLC and Leontief-free utilities, and we show that all such utilities are Gale-substitutes.
-Conversely, although some agents may get much higher utility at a Nash welfare maximizing allocation than at a CE, we show a price of anarchy type result: for general concave utilities, every CE achieves at least $(1/\mathrm{e})^{1/\mathrm{e}}>0.69$ fraction of the maximum Nash welfare, and this factor is tight.
+Conversely, although some agents may get much higher utility at a Nash welfare maximizing allocation than at a CE, we show a ‘price of anarchy’ type result: for general concave utilities, every CE achieves at least $(1/\mathrm{e})^{1/\mathrm{e}}>0.69$ fraction of the maximum Nash welfare, and this factor is tight.
 1 Introduction
-Competitive (market) equilibrium is a central solution concept in economics to analyze and predict the outcomes of market interactions.
-The Fisher model is a fundamental market model where a set of $n$ agents ${\mathcal{A}}$ arrive at the market with a fixed supply of divisible goods ${\mathcal{M}}$.
-Each agent $i$ has a fixed budget $b_{i}$ and preferences over bundles of goods given by a utility function $u_{i}(.)$.
-At competitive equilibrium (CE), prices $\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ of goods and fractional allocation $\mathbf{x}\in\mathbb{R}_{\geq 0}^{{\mathcal{A}}\times{\mathcal{M}}}$ of goods to agents are such that each agent $i$ gets an optimal bundle of goods, i.e., $x_{i}$ maximizes $u_{i}(x_{i})$ subject to $\langle\mathbf{p},x_{i}\rangle\leq b_{i}$, and the market clears, i.e., the demand for each good meets its supply.
-Equilibrium allocations satisfy many desirable properties, including Pareto optimality and the fairness guarantee of (weighted) envy-freeness.
-These properties make this an attractive concept also in other settings where no direct trade and money are involved.
-For example, competitive equilibrium with equal incomes (CEEI) is a popular mechanism in fair division, where a fictitious market is created by giving each agent the same budget Varian (1974).
-We note that CEEI is a special case of the Fisher model, where $b_{i}=1$ for each agent $i$.
-It is standard to assume that the utility functions $u_{i}(.)$ are monotone non-decreasing, concave, and satisfy $u_{i}(\mathbf{0})=0$.
-Classical results in mathematical economics established the existence of an equilibrium in the Fisher model as well as more general exchange market models, by Arrow and Debreu (1954) and by McKenzie (1954).
-These results are based on fixed point theorems and thus do not lead to efficient computational methods.
-Finding an equilibrium by efficient computation has been an important topic in economics, optimization, and computer science over the past decades.
-Eisenberg and Gale (1959) gave a convex programming formulation for CE with linear utility functions $u_{i}(x)=\sum_{j\in{\mathcal{M}}}v_{ij}x_{j}$.
-This program amounts to maximizing the weighted Nash welfare, i.e., the weighted geometric mean of agents’ utilities over feasible allocations.
-That is, any allocation maximizing $\prod_{i}u_{i}(x_{i})^{b_{i}}$ (equivalently, $\sum_{i}b_{i}\log{u_{i}(x_{i})}$) over feasible allocations gives a CE, where the dual solution yields the corresponding prices $\mathbf{p}$.
-Eisenberg (1961) showed that the same convex program yields competitive equilibria more generally for concave utility functions that are positively homogeneous of degree one, i.e., $u_{i}(\alpha x_{i})=\alpha u_{i}(x_{i})$ for any $\alpha>0$.
-This includes classical examples of constant elasticity of substitution (CES), Cobb–Douglas, and Leontief utility functions.
-The Nash welfare takes its name due to its connection with the bargaining problem introduced by Nash (1950).
-Nash studied general cooperative bargaining problems where the set of possible outcomes is a convex set $K\subseteq\mathbb{R}^{n}$ with a disagreement point $d\in K$ and showed that maximizing the geometric mean of the excess utilities above the disagreement utility amounts to the only possible bargaining outcome that satisfies certain natural requirements.
-Taking $K$ as the set of feasible allocations and $d=\mathbf{0}$, the Nash bargaining solution is precisely the Nash welfare maximizing allocation in this setting Vazirani (2012).
-Weak gross substitutes (WGS) utility functions are another broad class where a CE can be computed in polynomial time.
+We investigate a setting where a set $\mathcal{M}$} of divisible goods needs to be fairly allocated among a set $\mathcal{A}$} of $n$ agents in a cardinal utility framework.
+Each agent $i\in{\mathcal{A}}$ has preferences given by a utility function $u_{i}(.)\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$.
+We make the standard assumptions that $u_{i}(.)$’s are monotone non-decreasing, concave, and satisfy $u_{i}(\mathbf{0})=0$.
+[^1] All concepts discussed in the Introduction will be formally defined in Section 2.
+Two classical allocation concepts are Nash welfare and competitive equilibrium from equal incomes (CEEI).
+In Nash welfare, we select an allocation $(\mathbf{x}_{i})_{i\in{\mathcal{A}}}$ that maximizes $\prod_{i\in{\mathcal{A}}}u_{i}(\mathbf{x}_{i})$.
+In CEEI, each agent is initially given an equal share of all goods and allowed to trade with others; the resulting allocation $(\mathbf{x}_{i})_{i\in{\mathcal{A}}}$ corresponds to a competitive equilibrium (CE) outcome.
+Both these concepts have desirable properties that have made them widely popular in fair division research.
+Nash welfare derives its name from its roots in cooperative bargaining, introduced by Nash (1950).
+In this model, a convex set $K\subseteq\mathbb{R}^{n}$ of feasible outcomes is given, along with a disagreement point $\mathbf{d}\in K$ and concave utility functions $u_{i}\,:\,K\to\mathbb{R}_{\geq 0}$.
+Nash showed that the unique bargaining solution satisfying natural axioms—Pareto optimality, invariance under affine transformations, symmetry, and independence of irrelevant alternatives—is the one that maximizes the geometric mean of the agents’ gains over the disagreement point.
+Taking $K$ as the set of feasible allocations and $\mathbf{d}=\mathbf{0}$, the Nash bargaining solution exactly corresponds to the Nash welfare maximizing allocation in this setting (see also Vazirani (2012)).
+CEEI, introduced by Varian (1974), also satisfies desirable properties such as Pareto optimality, scale invariance, and symmetry.
+In addition, it guarantees envy-freeness, which may not be the case for Nash welfare.
+For further discussion and references on these concepts, see Moulin (2003, Chapters 3 and 7).
+Nash welfare maximization and CEEI remarkably coincide for degree one positively homogeneous utilities, i.e., if $u(\alpha\mathbf{x})=\alpha u(\mathbf{x})$ for any $\alpha>0$.
+This was first shown by Eisenberg and Gale (1959) for linear utilities (i.e., $u(\mathbf{x})=\sum_{j\in{\mathcal{M}}}v_{j}x_{j}$), and later extended by Eisenberg (1961) to the homogeneous case that includes classical examples of constant elasticity of substitution (CES), Cobb–Douglas, and Leontief utility functions.
+In general, Nash welfare maximization and CEEI can lead to different outcomes.
+Moreover, while the Nash welfare solution yields a unique utility profile for the agents, CEEI may admit multiple equilibria with significantly different utilities across agents.
+This raises a nontrivial equilibrium selection problem Arrow and Hahn (1971), Harsanyi and Selten (1998).
+The problem of efficiently computing equilibria has been an important topic in economics, optimization, and computer science over the past decades.
+The existence of CEEI follows from classical results in general equilibrium theory by Arrow and Debreu (1954) and by McKenzie (1954), based on fixed point theorems.
+While these results apply to more general exchange market models, they do not yield efficient computational methods.
+In fact, the computational complexity of finding equilibria is generally negative: even for the seemingly simple case of separable piecewise concave (SPLC) utilities, computing a CE falls into the class of PPAD-hard Chen and Teng (2009) problems, suggesting that an efficient algorithm is unlikely to exist.
+The hardness persists even under bounded linear utilities, where the utility from each good is capped, i.e., $u(\mathbf{x})=\sum_{j\in{\mathcal{M}}}\min\{v_{j}x_{j},\ell_{j}\}$ (Bei et al., 2019b).
+Examples E.1 and E.2 show that the set of CE can be non-convex and disconnected even under such preferences.
+In contrast, Nash welfare maximization corresponds to solving a convex program and can therefore be computed efficiently for any concave valuations.
+Besides homogeneous utilities, another important class where CEEI can be computed in polynomial time is weak gross substitutes (WGS) utility functions.
+[^2] A utility function satisfies the weak gross substitutes property when raising the price of one good does not decrease the demand for any other good.
 Several fast algorithms have been designed for computing equilibria with WGS utilities Bei et al.
@@
 (2023a).
-While the WGS class includes many homogeneous utilities, there are also non-homogeneous examples for which the Eisenberg–Gale solution may not correspond to a CE.
-WGS and homogeneous utility functions cover most known cases where a CE can be computed in polynomial time.
-Already for the seemingly simple case of separable piecewise concave utilities, computing a CE is PPAD-hard Chen and Teng (2009).
-In fact, it is PPAD-hard even for bounded linear utilities, when there is an upper bound on the utility from each good, i.e., $u_{i}(x)=\sum_{j\in{\mathcal{M}}}\min\{v_{ij}x_{j},\ell_{ij}\}$ (Bei et al., 2019b).
-Example A.1 shows that the set of CE can be non-convex and disconnected already for bounded linear utilities.
+While the WGS class includes many homogeneous utilities, there are also non-homogeneous examples where the Eisenberg–Gale solution may not correspond to a CE.
+Together, WGS and homogeneous utility functions cover the majority of known settings in which a CE can be computed efficiently.
+In this paper, we study the relationship between two fundamental concepts in the allocation of divisible goods: competitive equilibrium and maximum Nash welfare allocations.
+We reveal a strong connection between these solution concepts that extend well beyond the classical homogeneous case.
+While the classical fair division concepts are CEEI and Nash welfare maximizing allocations, our results apply more broadly to the Fisher market model and weighted Nash welfare maximization.
+In the Fisher model, a fixed supply of goods $\mathcal{M}$} is to be allocated among agents $\mathcal{A}$}, where each agent $i\in{\mathcal{A}}$ has cardinal preferences given by a utility function $u_{i}(.)$.
+Each agent also has a fixed budget $b_{i}>0$.
+At competitive equilibrium (CE), prices $\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ of goods and fractional allocation $(\mathbf{x}_{i})_{i\in{\mathcal{A}}}$ satisfy two conditions: (i) each agent $i$ gets an optimal bundle $\mathbf{x}_{i}$ that maximizes $u_{i}(\mathbf{x}_{i})$ subject to the budget constraint $\langle\mathbf{p},\mathbf{x}_{i}\rangle\leq b_{i}$, and (ii) the market clears, meaning demand meets supply for every good.
+The CEEI setting corresponds to the special case where all budgets are equal, i.e., $b_{i}=1$ for all $i\in{\mathcal{A}}$.
+Similarly, for given weights $b_{i}>0$, the weighted Nash welfare maximizing allocation is a feasible allocation of goods that maximizes $\prod_{i}u_{i}(\mathbf{x}_{i})^{b_{i}}$, or equivalently, $\sum_{i}b_{i}\log{u_{i}(\mathbf{x}_{i})}$.
+Classical results in Eisenberg and Gale (1959), Eisenberg (1961) establish that these allocations coincide with CE outcomes for degree one positively homogeneous utilities and arbitrary budgets $b_{i}>0$.
+Henceforth, by Nash welfare maximizing allocation we refer to this weighted concept.
 1.1 Our contributions
-In this paper, we study the relationship between two fundamental concepts in the allocation of divisible goods: competitive equilibrium and maximum Nash welfare allocations.
-We exhibit a close relationship between these solution concepts well beyond the homogeneous case.
-First, we show a ‘price of anarchy’ type result Dubey (1986), Koutsoupias and Papadimitriou (1999).
+Our first result is a ‘price of anarchy’ type bound Dubey (1986), Koutsoupias and Papadimitriou (1999).
 One can interpret Nash welfare as a global measure of social utility.
 In contrast, a CE can be seen as an outcome reached by autonomous agents trying to improve their utilities individually.
-It turns out that, in the general setting of concave utilities, any CE recovers at least $(1/\mathrm{e})^{1/\mathrm{e}}>0.69$ fraction of the maximum Nash welfare, and this factor is tight (see Example A.4).
+It turns out that, in the general setting of concave utilities, any CE recovers at least $(1/\mathrm{e})^{1/\mathrm{e}}>0.69$ fraction of the maximum Nash welfare, and this factor is tight (see Example 3.2).
 Theorem 1.1.
@@
 We then study guarantees of the maximum Nash welfare solution in the context of CE.
-A major advantage of maximum Nash welfare is efficient computability: it can be arbitrarily well approximated for any concave utility functions.
-In contrast, computing CE becomes PPAD-hard already in very simple cases.
+As noted above, a major advantage of maximum Nash welfare is efficient computability: it can be arbitrarily well approximated for any concave utility functions, whereas computing CE becomes PPAD-hard already in very simple cases.
 In this context, our first simple result shows that the maximum Nash welfare yields an approximate CE for all concave utility functions.
-By an $\alpha$-demand-approximate CE, we mean allocations and prices such that every agent receives at least $1/\alpha$-fraction of the maximum possible utility (see Definition 2.8).
 As in the case of the Eisenberg–Gale program, one can interpret the optimal Lagrangian multipliers as prices.
+By an $\alpha$-demand-approximate CE, we mean allocations and prices such that every agent receives at least $1/\alpha$-fraction of the maximum possible utility at those prices (see Definition 2.8).
 The result also implies approximate envy-freeness.
 Theorem 1.2.
-For Fisher markets under concave utility functions, any Nash welfare maximizing allocation $\mathbf{y}$, together with the dual prices $\mathbf{q}$, form a 2-demand-approximate CE.
-This also implies that $\mathbf{y}$ is half envy-free when $b_{i}$’s are identical, i.e., $u_{i}(y_{i})\geq\tfrac{1}{2}u_{i}(y_{k})$, for all $i,k\in{\mathcal{A}}$.
+For Fisher markets under concave utility functions, any Nash welfare maximizing allocation $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$, together with the dual prices $\mathbf{q}$, form a 2-demand-approximate CE.
+This also implies that $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$ is half envy-free when $b_{i}$’s are identical, i.e., $u_{i}(\mathbf{y}_{i})\geq\tfrac{1}{2}u_{i}(\mathbf{y}_{k})$, for all $i,k\in{\mathcal{A}}$.
 Theorem 1.2 already shows that maximum Nash welfare provides remarkable fairness and efficiency guarantees for all concave utility functions.
 However, there can be multiple equilibria, giving vastly different utilities to an agent.
-In Example A.5, we show that there could be a gap of $\Omega(n)$ in the utility of some agent between two different competitive equilibria.
-Hence, an agent could be much better off in some CE than in the maximum Nash welfare solution.
+In Example 4.4, we show that there could be a gap of $\Omega(n)$ in the utility of some agent between two different competitive equilibria.
+Hence, some agent could be much better off in some CE than in the maximum Nash welfare solution.
 As the main contribution of the paper, we show that maximum Nash welfare provides surprisingly strong guarantees for a broad class of utility functions.
@@
 (2023a), Nesterov and Shikhman (2018).
-For an agent $i\in{\mathcal{A}}$ with budget $b_{i}$ and utility function $u_{i}(.)$ at prices $\mathbf{q}$, the Gale demand correspondence comprises of the allocations $y_{i}$ that maximize $b_{i}\log u_{i}(y_{i})-\langle\mathbf{q},y_{i}\rangle$.
+For an agent $i\in{\mathcal{A}}$ with budget $b_{i}$ and utility function $u_{i}(.)$ at prices $\mathbf{q}$, the Gale demand correspondence [^3] In contrast, the standard demand correspondence comprises of the allocations $\mathbf{x}_{i}$ that maximize $\{u_{i}(\mathbf{x}_{i})\ |\ \langle\mathbf{q},\mathbf{x}_{i}\rangle\leq b_{i}\,\}.$ comprises of the allocations $\mathbf{y}_{i}$ that maximize
+$b_{i}\log u_{i}(\mathbf{y}_{i})-\langle\mathbf{q},\mathbf{y}_{i}\rangle\,.$
 It follows from Lagrangian duality that in a Nash welfare maximizing solution, every agent receives an allocation from their Gale demand correspondence.
 The reason behind the equivalence of CE and maximum Nash welfare for homogeneous utilities is that, for this class, the standard demand correspondence and the Gale demand correspondence coincide.
-Nevertheless, these two demands can behave very differently for non-homogeneous utility functions, e.g., the Gale demand can be non-monotonic in the prices, whereas the standard demand is always monotonic (see Example A.3).
+Nevertheless, these two demands can behave very differently for non-homogeneous utility functions, e.g., the utility value from the standard demand is always monotone non-increasing, whereas the utility under Gale demand can be non-monotonic (see Example E.3).
 We identify a critical property of the Gale demand correspondence that we call Gale-substitutes property (Definition 2.14).
@@
 We note that for non-satiable utilities, the Gale-substitutes and $\Sigma$-Gale-substitutes properties coincide.
-Figure 1 illustrates the containment relationships among various utility classes discussed in this paper, with those in which the Fisher market equilibrium can be efficiently computed highlighted in grey.
+Figure 1 illustrates the containment relationships among various utility classes discussed in this paper.
+Figure 1: The graph illustrates the containment relationships among different utility classes, with the classes where the Fisher market equilibrium can be computed efficiently highlighted in grey.
 Our main result shows that for this class of utility functions, the utility of any agent at the Nash welfare maximizing allocation is at least half the maximum utility they may get at any CE.
@@
 Consider a Fisher market instance where all utility functions are $\Sigma$-Gale-substitutes.
-Then, for any competitive equilibrium $(\mathbf{x},\mathbf{p})$ and any Nash welfare maximizing allocation $\mathbf{y}$, $u_{i}(y_{i})\geq\frac{1}{2}u_{i}(x_{i})$ holds for every $i\in{\mathcal{A}}$.
+Then, for any competitive equilibrium $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$ and any Nash welfare maximizing allocation $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$, $u_{i}(\mathbf{y}_{i})\geq\frac{1}{2}u_{i}(\mathbf{x}_{i})$ holds for every $i\in{\mathcal{A}}$.
 As a counterpart to the ‘price of anarchy’ result in Theorem 1.1, one could call this a ‘price of authority’ result.
@@
 Hence, for ($\Sigma$-)Gale-substitutes utility functions, the Nash welfare maximizing allocation provides a guaranteed good outcome to all agents simultaneously while also being efficiently computable.
-Thus, Theorem 1.3 establishes a strong relationship between CE and maximum Nash welfare beyond homogeneous functions and provides a natural alternative to the equilibrium selection problem Arrow and Hahn (1971) in this setting where multiple equilibria are possible.
-We call a utility function $u(.)$ separable if $u(x)=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$.
+Thus, Theorem 1.3 establishes a strong relationship between CE and maximum Nash welfare beyond homogeneous functions and provides a natural alternative to the equilibrium selection problem Arrow and Hahn (1971), Harsanyi and Selten (1998) in this setting where multiple equilibria are possible.
+We note that the statement of Theorem 1.3 does not hold for arbitrary concave utilities.
+In Example E.5, we show that without the Gale-substitutes property, agents may get much higher utilities in CE than in the maximum Nash welfare allocations.
+Examples of Gale-substitute utilities
+It turns out that ($\Sigma$-)Gale-substitutes capture a wide range of natural utility functions.
+We call a utility function $u(.)$ separable if $u(\mathbf{x})=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$.
 This class gives an important example of $\Sigma$-Gale-substitutes.
@@
 Every separable utility function is $\Sigma$-Gale-substitutes.
-We consider two further examples of piecewise linear Gale-substitutes utilities:
+We consider two further examples of non-separable, piecewise linear Gale-substitutes utilities:
   • Leontief-free (LF) utilities Filos-Ratsikas et al.
@@
 (2014), Goyal et al.
-(2023): Consider a set $K$ of segments, and let $u(x)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\max\sum_{k\in K}u_{k}$: $u_{k}=\sum_{j\in{\mathcal{M}}}a_{jk}z_{jk},\ \ u_{k}\leq\ell_{k};\ \forall k\in K$; $\sum_{k\in K}z_{jk}\leq x_{j}\ \forall j\in{\mathcal{M}}$; $\mathbf{z}\geq 0$.
-That is, an LF function is given by a set of linear functions $u_{k}$’s with a maximum utility limit $\ell_{k}$ on each $k\in K$.
+(2023): Consider a set $K$ of segments, and let $u(\mathbf{x})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\max\sum_{k\in K}u_{k}$: $u_{k}=\sum_{j\in{\mathcal{M}}}a_{jk}z_{jk},\ \ u_{k}\leq\ell_{k};\ \forall k\in K$; $\sum_{k\in K}z_{jk}\leq x_{j}\ \forall j\in{\mathcal{M}}$; $\mathbf{z}\geq 0$.
+That is, an LF function is given by a set of linear functions $u_{k}$, with a maximum utility limit $\ell_{k}$ on each $k\in K$.
 LF captures interesting non-separable PLC functions, and includes SPLC as a special case where each segment $k\in K$ has exactly one $j$ with $a_{jk}>0$.
@@
 Here, the goods correspond to source nodes in a flow network with supply $x_{j}$, $j\in{\mathcal{M}}$.
-Each arc has a capacity and a gain factor, and the utility $u(x)$ is the maximum flow value that can reach a designated sink node $t$ from supplies $x$.
+Each arc has a capacity and a gain factor, and the utility $u(\mathbf{x})$ is the maximum flow value that can reach a designated sink node $t$ from supplies $\mathbf{x}$.
+This value can be computed in polynomial time.
 The precise definitions are given in Section 2.1.
-LF can be formulated as a special case of a network with two layers, the middle layer between the sources and the sink corresponding to the segments $K$.
+Additionally, LF can be formulated as a special case of a network with two layers, the middle layer between the sources and the sink corresponding to the segments $K$.
+We note that generalized flows is a classical network flow model that can model commodity transportation with losses, as well as conversions between different types of entities such as in financial networks; see Ahuja et al.
+(1993, Chapter 15) for a description and various applications.
 We show the following:
@@
 Generalized network utility functions are Gale-substitutes.
-Generalized flows is a classical network flow model that can model commodity transportations with losses, as well as conversions between different types of entities such as in financial networks; see Chapter 15 in Ahuja et al.
-(1993) for a description and various applications.
 Theorem 1.3 already shows that in Fisher markets with non-satiable generalized network utilities, Nash welfare maximizing allocations give a good approximation to utilities at competitive equilibrium.
 However, satiable generalized network utilities may not satisfy the stronger $\Sigma$-Gale-substitutes property required in Theorem 1.3.
-Example A.10 shows that the theorem does not hold if only assuming the Gale substitutes property.
-For satiable utility functions, it is natural to look at thrifty (strong) equilibria, which are guaranteed to exist Garg et al.
+Example E.6 shows that the theorem does not hold if only assuming the Gale substitutes property.
+For satiable utility functions, it is natural to look at thrifty (strong) equilibria, a strengthening of CE, which are also guaranteed to exist, see, e.g., Garg et al.
 (2022b).
@@
 Consider a Fisher market instance where all utility functions are generalized network utilities.
-Then, for any thrifty competitive equilibrium $(\mathbf{x},\mathbf{p})$ and any Nash welfare maximizing allocation $\mathbf{y}$, $u_{i}(y_{i})\geq\frac{1}{2}u_{i}(x_{i})$ holds for every $i\in{\mathcal{A}}$.
-Finally, we note that the statement of Theorem 1.3 does not hold for arbitrary concave utilities.
-In Example A.8, we show that without the Gale-substitutes property, agents may get much higher utilities in CE than in the maximum Nash welfare allocations.
+Then, for any thrifty competitive equilibrium $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$ and any Nash welfare maximizing allocation $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$, $u_{i}(\mathbf{y}_{i})\geq\frac{1}{2}u_{i}(\mathbf{x}_{i})$ holds for every $i\in{\mathcal{A}}$.
 Remark 1.7.
-We note that no polynomial time approximation scheme (PTAS) can exist that, for some $\varepsilon>0$, provides any agent at least $(1-\varepsilon)$ times their maximum CE utility.
-This is due to the fact that, as shown in Example A.1, multiple thrifty equilibria may exist, each providing different utilities to the agents.
+We note that no algorithm can guarantee agents at least $(1-\varepsilon)$ times their maximum CE utility for any $\varepsilon>0$.
+This is due to the fact that, as shown in Example E.1, multiple thrifty equilibria may exist, each providing different utilities to the agents.
 The existence of an allocation that $\varepsilon$-approximates the best CE utility for each agent would, by compactness, imply the existence of an allocation where every agent achieves their best CE utility.
 However, this would contradict the fact that thrifty equilibria are Pareto efficient.
-Figure 1: The graph illustrates the containment relationships among different utility classes, with the classes where the Fisher market equilibrium can be computed efficiently highlighted in grey.
 1.2 Related work
@@
 (2023a)..
-This was used to obtain an auction algorithm for the maximum Nash welfare problem with capped SPLC utilities, and the proof uses arguments similar to Section 7.
+This was used to obtain an auction algorithm for the maximum Nash welfare problem with capped SPLC utilities, and the proof uses arguments similar to those in Appendix C.
 Similar auction algorithms could be developed more generally for Gale substitutes functions.
@@
 (2005c), Garg et al.
-(2023a)) and for CES utilities with parameter $-1\leq\rho<0$ Codenotti et al.
+(2023a)) and for CES utilities with parameter $1\leq\rho<0$ Codenotti et al.
 (2005a), and beyond that the problem is essentially PPAD-hard Chen et al.
@@
 (2006).
-The maximum Nash welfare is a classic objective for allocating goods to agents.
+The maximum Nash welfare is a classical objective for allocating goods to agents.
 It has received significant attention in the literature on social choice and fair division; see, e.g., Moulin (2003).
@@
 Section 2 defines all models and definitions.
-Section 3 provides a brief overview of the techniques used.
-Section 4 proves Theorem 1.1, showing the tight bound of $(1/\mathrm{e})^{1/\mathrm{e}}$ on the efficiency of competitive equilibria.
-Section 5 presents several properties of the Gale demand correspondence and proves Theorem 1.2, showing that Nash welfare maximizing allocations yield $2$-demand approximate-CE.
+Section 3 proves Theorem 1.1, establishing the tight bound of $(1/\mathrm{e})^{1/\mathrm{e}}$ on the efficiency of competitive equilibria.
+In Section 4, we show that Nash welfare maximizing allocations yield $2$-demand approximate-CE.
 Section 5 proves Theorem 1.3, showing the stronger guarantees of the maximum Nash welfare allocations for Gale-substitutes utility functions.
-Section 7 proves Theorem 1.4, showing that separable utilities are $\Sigma$-Gale substitutes.
-Section 8 considers generalized network utility functions and proves Theorems 1.5 and 1.6.
-Several examples are deferred to the Appendix, and Appendix B presents an omitted proof from Section 8.
-2 Preliminaries
-Notation
+Section 6 proves Theorem 1.4, showing that separable utilities are $\Sigma$-Gale substitutes, and considers generalized network utility functions.
+Appendix A presents several properties of the Gale demand correspondence.
+Appendix B provides the complete proof of Theorem 1.3.
+Appendix C proves Theorem 1.4 for the general case, and Appendix D contains the proofs of Theorems 1.5 and 1.6.
+Additional examples are presented in Appendix E.
+2 Models and definitions
 For a function $f\,:\,X\to Y$, let $\operatorname{dom}(f)\subseteq X$ denote its domain.
-For $x\in\mathbb{R}^{n}$, let $\operatorname{supp}(x)=\{i\,:\,x_{i}\neq 0\}$ denote its support.
-For $x,y\in\mathbb{R}^{n}$, we use $\left\langle x,y\right\rangle=x^{\top}y$ for the standard scalar product.
+For $\mathbf{x}\in\mathbb{R}^{n}$, let $\operatorname{supp}(\mathbf{x})=\{i\,:\,x_{i}\neq 0\}$ denote its support.
+For $\mathbf{x},\mathbf{y}\in\mathbb{R}^{n}$, we use $\left\langle\mathbf{x},\mathbf{y}\right\rangle=\mathbf{x}^{\top}\mathbf{y}$ for the standard scalar product.
 Subgradients and supergradients
@@
 We denote the set of all supergradients of $f$ at $\mathbf{x}$ as $\partial^{\wedge}{f}(\mathbf{x})$, set of all subgradients as $\partial^{\vee}{f}(\mathbf{x})$.
-Thus, $\mathbf{g}$ if a supergradient of $f$ if and only if $-\mathbf{g}$ is a subgradient of $-f$.
+Thus, $\mathbf{g}$ if a supergradient of $f$ if and only if $\mathbf{g}$ is a subgradient of f-f.
 We recall that a function $f$ is convex if and only if $\partial^{\vee}{f}(\mathbf{x})\neq\emptyset$ for every $\mathbf{x}\in\operatorname{dom}(f)$ and concave if and only if $\partial^{\wedge}{f}(\mathbf{x})\neq\emptyset$ for every $\mathbf{x}\in\operatorname{dom}(f)$.
-For a differentiable convex (concave) function, the unique subgradient (supergradient) at every point is the gradient $\nabla f(\mathbf{x})$.
+For a differentiable convex (concave) function $f$, the unique subgradient (supergradient) at every point is the gradient $\nabla f(\mathbf{x})$.
 2.1 Utility functions
@@
 Definition 2.2.
-A function $u(\cdot)\,,\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}$ is a utility function if $u(\mathbf{0})=0$, $u(\cdot)$ is concave, monotone non-decreasing, and there exists an $x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $u(x)>0$.
-We say that a utility function is separable if $u(x)=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$. We will also distinguish between satiable and non-satiatiable utilities.
+A function $u(\cdot)\,,\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}$ is a utility function if $u(\mathbf{0})=0$, $u(\cdot)$ is concave, continuous, monotone non-decreasing, and there exists an $\mathbf{x}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $u(\mathbf{x})>0$.
+We say that a utility function is separable if $u(\mathbf{x})=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$. We will also distinguish between satiable and non-satiable utilities.
 Definition 2.3.
-A utility function $u(\cdot)$ is non-satiable if and only if, for any $x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, one can find $x^{\prime}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, such that $u(x^{\prime})>u(x)$.
+A utility function $u(\cdot)$ is non-satiable if and only if, for any $\mathbf{x}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, there exists $\mathbf{x}^{\prime}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, such that $u(\mathbf{x}^{\prime})>u(\mathbf{x})$.
 Otherwise, we say the utility function is satiable.
@@
 Given prices $\mathbf{p}$ and a budget $b\in\mathbb{R}_{\geq 0}$, we define the demand correspondence as the set of maximum utility bundles achievable at $\mathbf{p}$ and $b$:
-${\mathcal{D}}^{u}(\mathbf{p},b)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\arg\max_{x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\left\{u(x)\,:\,\left\langle\mathbf{p},x\right\rangle\leq b\right\}\,.$ (1)
-Note that if $u$ is a non-satiable utility function, then $\left\langle\mathbf{p},x\right\rangle=b$ must hold for every $x\in{\mathcal{D}}^{u}(\mathbf{p},b)$.
+${\mathcal{D}}^{u}(\mathbf{p},b)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\arg\max_{\mathbf{x}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\left\{u(\mathbf{x})\,:\,\left\langle\mathbf{p},\mathbf{x}\right\rangle\leq b\right\}\,.$ (1)
+Note that if $u$ is a non-satiable utility function, then $\left\langle\mathbf{p},\mathbf{x}\right\rangle=b$ must hold for every $\mathbf{x}\in{\mathcal{D}}^{u}(\mathbf{p},b)$.
 ${\mathcal{D}}^{u}(\mathbf{p},b)$ is always a non-empty compact set when $\mathbf{p}>0$.
-However, in cases when some prices may be 00, the supremum of utility over the budget constraint might be $\infty$, resulting in unbounded demand.
-This may happen even when the supremum value is finite, and causes technical difficulties.
+However, in cases when some prices may be 0, the supremum of utility over the budget constraint might be $\infty$, resulting in unbounded demand.
+This may occur even when the supremum is finite, causing technical difficulties.
+[^5] For example, given $u(x_{1},x_{2})=1-\frac{1}{x_{1}+x_{2}+1}$, $\mathbf{p}=(1,0)$, and $b=1$, $\sup_{\mathbf{x}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\left\{u(\mathbf{x})\,:\,\left\langle\mathbf{p},\mathbf{x}\right\rangle\leq b\right\}=1$.
+This value is approached as $x_{1}=1$ and $x_{2}\rightarrow\infty$, but it is never achieved by any finite bundle $\mathbf{x}$.
 To circumvent these issues, we adopt the following assumption regarding utility functions $u(\cdot)$:
 Assumption 1.
-For any $\mathbf{p}\geq\mathbf{0}$ and $b>0$, if there is a sequence of allocations $\{x^{(k)}\}_{k=1}^{\infty}$ such that $\left\langle\mathbf{p},x^{(k)}\right\rangle$ converges to $b$ and $u(x^{(k)})$ converges to some finite value $M<\infty$, then there exists a finite allocation $x^{*}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $u(x^{*})=M$ and $\left\langle\mathbf{p},x\right\rangle\leq b$.
-This assumption always holds if the set of possible allocations is restricted to a compact domain $X\subseteq\mathbb{R}_{\geq 0}^{\mathcal{M}}$; note Nesterov and Shikhman (2018) makes such a compactness assumption.
-For the unrestricted domain $\mathbb{R}_{\geq 0}^{\mathcal{M}}$, it can also be shown for arbitrary piecewise linear utility functions, in particular, for generalized network utilities defined below.
+For any $\mathbf{p}\geq\mathbf{0}$ and $b>0$, if there is a sequence of allocations $\{\mathbf{x}^{(k)}\}_{k=1}^{\infty}$ such that $\left\langle\mathbf{p},\mathbf{x}^{(k)}\right\rangle$ converges to $b$ and $u(\mathbf{x}^{(k)})$ converges to some finite value $M<\infty$, then there exists a finite allocation $\mathbf{x}^{*}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $u(\mathbf{x}^{*})=M$ and $\left\langle\mathbf{p},\mathbf{x}^{*}\right\rangle\leq b$.
+This assumption always holds if the set of possible allocations is restricted to a compact domain $Z\subseteq\mathbb{R}_{\geq 0}^{\mathcal{M}}$; note that Nesterov and Shikhman (2018) make such a compactness assumption.
+For the unrestricted domain $\mathbb{R}_{\geq 0}^{\mathcal{M}}$, Assumption 1 can be shown to hold for arbitrary piecewise linear utility functions, in particular, for generalized network utilities defined below.
 The Gale demand correspondence is defined as
-$\mathcal{GD}^{u}(\mathbf{p},b)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\arg\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{p},y\right\rangle\,.$ (2)
-We define the domain of $\mathcal{GD}^{u}$ with respect to prices as
-$\operatorname{pdom}(u)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\left\{\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}\leavevmode\nobreak\ :\leavevmode\nobreak\ \exists b>0,\sup_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{p},y\right\rangle<+\infty\right\}\,.$ (3)
+$\mathcal{GD}^{u}(\mathbf{p},b)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\arg\max_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(\mathbf{y})-\left\langle\mathbf{p},\mathbf{y}\right\rangle\,.$ (2)
+We define the domain of $\mathcal{GD}^{u}$ with respect to prices as  [^6] An equivalent formulation of the definition is given by: $\operatorname{pdom}(u)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\left\{\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}\penalty 10000\ :\penalty 10000\ \forall b>0,\sup_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}} b\log u(\mathbf{y})-\left\langle\mathbf{p},\mathbf{y}\right\rangle<+\infty\right\}.$ This definition is obtained by substituting the universal quantifier ($\forall$) for the existential quantifier ($\exists$).
+Lemma A.1 (i) provides further details on this equivalence.
+$\operatorname{pdom}(u)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\left\{\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}\penalty 10000\ :\penalty 10000\ \exists b>0,\sup_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(\mathbf{y})-\left\langle\mathbf{p},\mathbf{y}\right\rangle<+\infty\right\}\,.$ (3)
 Under Assumption 1, $\mathbf{0}\in\operatorname{pdom}(u)$ if and only if $u(.)$ is satiable.
-Every strictly positive $\mathbf{p}>\mathbf{0}$ is in $\operatorname{pdom}(u)$, and if $u$ is satiable, then $\operatorname{pdom}(u)=\mathbb{R}_{\geq 0}^{\mathcal{M}}$ (see Lemma 5.2).
-In Section 5, we show the following.
+Every strictly positive $\mathbf{p}>\mathbf{0}$ is in $\operatorname{pdom}(u)$, and if $u$ is satiable, then $\operatorname{pdom}(u)=\mathbb{R}_{\geq 0}^{\mathcal{M}}$ (see Lemma A.1).
+In Appendix A, we show the following.
 Lemma 2.4.
-Under Assumption 1, for any $\mathbf{p}\in\operatorname{pdom}(u)$ and any $b>0$, $\mathcal{GD}^{u}(\mathbf{p},b)\neq\emptyset$.
+Under Assumption 1, for any $\mathbf{p}\in\operatorname{pdom}(u)$ and for any $b>0$, it holds that $\mathcal{GD}^{u}(\mathbf{p},b)\neq\emptyset$.
 Gale demands may behave very differently from the standard demand correspondence.
-Example A.3 shows the counter-intuitive property increasing the prices $\mathbf{p}$ may actually increase the Gale utility value $\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{p},y\right\rangle$.
-Recall the notions of bounded linear, separable piecewise linear (SPLC) and Leontief-free (LF) utilities from the Introduction. We now define generalized network utility functions.
+Example E.3 shows the counter-intuitive property increasing the prices $\mathbf{p}$ may actually increase the Gale utility value $\max_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(\mathbf{y})-\left\langle\mathbf{p},\mathbf{y}\right\rangle$.
+Recall the notions of bounded linear, separable piecewise linear (SPLC) and Leontief-free (LF) utilities from the Introduction. We now define the class of generalized network utility functions.
 Generalized network utilities
-First, we introduce the generalized flow model.
-A generalized flow instance is defined as follows.
-Let $G=(V,E)$ be a directed graph with a sink node $t\in V$, a supply set $S\subseteq V\setminus\{t\}$, a capacity vector $h\in\mathbb{R}_{\geq 0}^{E}$, and positive gain factors $\gamma\in\mathbb{R}_{>0}^{E}$.
+We now formally introduce the generalized flow model.
+A generalized flow instance is given by a directed graph $G=(V,E)$ with a sink node $t\in V$, a supply set $S\subseteq V\setminus\{t\}$, a capacity vector $\mathbf{h}\in\mathbb{R}_{\geq 0}^{E}$, and positive gain factors $\bm{\gamma}\in\mathbb{R}_{>0}^{E}$.
 We let $\partial^{\mathrm{in}}(S)$ and $\partial^{\mathrm{out}}(S)$ denote the set of arcs entering and leaving a subset $S\subseteq V$, respectively.
 For a node $v\in V$, we use $\partial^{\mathrm{in}}(v)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\partial^{\mathrm{in}}(\{v\})$ and $\partial^{\mathrm{out}}(v)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\partial^{\mathrm{out}}(\{v\})$.
-For a vector $f\in\mathbb{R}_{\geq 0}^{E}$, for every arc $e=(v,w)$, we interpret $f_{e}$ as the flow entering $e$ at node $v$.
+For a vector $\mathbf{f}\in\mathbb{R}_{\geq 0}^{E}$, for every arc $e=(v,w)$, we interpret $f_{e}$ as the flow entering $e$ at node $v$.
 The flow gets multiplied by the gain factor $\gamma_{e}$ while traversing the arc, with $\gamma_{e}f_{e}$ amount reaching node $w$.
 Thus, the net flow at node $v\in V$ is defined as
-$\operatorname{net}_{f}(v)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{e\in\partial^{\mathrm{in}}(v)}\gamma_{e}f_{e}-\sum_{e\in\partial^{\mathrm{out}}(v)}f_{e}\,.$
-Given a supply vector $x\in\mathbb{R}_{\geq 0}^{S}$, we say that $f\in\mathbb{R}_{\geq 0}^{E}$ is a feasible generalized flow with supply $x$ and capacity $h$ if $0\leq f\leq h$, and $\operatorname{net}_{f}(v)\geq-x_{v}$ for all $v\in S$, and $\operatorname{net}_{f}(v)\geq 0$ for all $v\in V\setminus(S\cup\{t\})$.
-The maximum generalized flow value from supply $x$ is the maximum amount of $\operatorname{net}_{f}(t)$ among all feasible flows with supply $x$.
+$\operatorname{net}_{\mathbf{f}}(v)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{e\in\partial^{\mathrm{in}}(v)}\gamma_{e}f_{e}-\sum_{e\in\partial^{\mathrm{out}}(v)}f_{e}\,.$
+Given a supply vector $\mathbf{x}\in\mathbb{R}_{\geq 0}^{S}$, we say that $\mathbf{f}\in\mathbb{R}_{\geq 0}^{E}$ is a feasible generalized flow with supply $\mathbf{x}$ and capacity $\mathbf{h}$ if $\mathbf{0}\leq\mathbf{f}\leq\mathbf{h}$, and $\operatorname{net}_{\mathbf{f}}(v)\geq-x_{v}$ for all $v\in S$, and $\operatorname{net}_{\mathbf{f}}(v)\geq 0$ for all $v\in V\setminus(S\cup\{t\})$.
+The maximum generalized flow value from supply $\mathbf{x}$ is the maximum amount of $\operatorname{net}_{\mathbf{f}}(t)$ among all feasible flows with supply $\mathbf{x}$.
 It corresponds to the optimum value of the following linear program:
-$\max\leavevmode\nobreak$ $\operatorname{net}_{f}(t)$ ($\operatorname{GenflowLP}(x)$)
-$\operatorname{net}_{f}(v)$ $\geq-x_{v}\quad\forall v\in S$
-$\operatorname{net}_{f}(v)$ $\geq 0\quad\forall v\in V\setminus(S\cup\{t\})$
-00 0 $\leq f\leq h\,.$
+[…]
+$\operatorname{net}_{\mathbf{f}}(v)$ $\geq-x_{v}\quad\forall v\in S$
+$\operatorname{net}_{\mathbf{f}}(v)$ $\geq 0\quad\forall v\in V\setminus(S\cup\{t\})$
+$\mathbf{0}$ $\leq\mathbf{f}\leq\mathbf{h}\,.$
 Definition 2.5.
-The function $u:\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ is a generalized network utility function if $u(x)$ is the maximum generalized flow value from supply $x$ in some generalized flow instance where the supply set is $S={\mathcal{M}}$, the maximum flow amount is finite for any supply vector, $u(\mathbf{0})=0$, and there exists $x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ with $u(x)>0$.
-The following properties are immediate from the definition, and show that these are indeed utility functions in accordance with Definition 2.2.
+The function $u:\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ is a generalized network utility function if $u(\mathbf{x})$ is the maximum generalized flow value from supply $\mathbf{x}$ in some generalized flow instance where the supply set is $S={\mathcal{M}}$, the maximum flow amount is finite for any supply vector, $u(\mathbf{0})=0$, and there exists $\mathbf{x}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ with $u(\mathbf{x})>0$.
+The following properties are immediate from the definition and show that these are indeed utility functions in accordance with Definition 2.2.
 Proposition 2.6.
-If $u:\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ is a generalized network utility function, then $u(x)$ is monotone non-decreasing and concave, and satisfies Assumption 1.
+If $u:\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ is a generalized network utility function, then $u(\mathbf{x})$ is monotone non-decreasing and concave, and satisfies Assumption 1.
 SPLC and Leontief-free as special cases
-An SPLC function can be represented as a generalized network utility function by setting $V={\mathcal{M}}\cup\{t\}$, and for each $j\in{\mathcal{M}}$, adding a capacitated arcs $e=(j,t)$ for each linear segment of $v_{j}$, where $\gamma_{e}$ is the marginal utility on this segment and $h_{e}$ is the length of the segment.
-Similary, we can represent a Leontief-free utility as a generalized network utility function by setting $V={\mathcal{M}}\cup K\cup\{t\}$; recall $K$ is the set of segments.
+An SPLC function can be represented as a generalized network utility function by setting $V={\mathcal{M}}\cup\{t\}$, and for each $j\in{\mathcal{M}}$, adding a capacitated arc $e=(j,t)$ for each linear segment of $v_{j}$, where $\gamma_{e}$ is the marginal utility on this segment and $h_{e}$ is the length of the segment.
+Similarly, we can represent a Leontief-free utility as a generalized network utility function by setting $V={\mathcal{M}}\cup K\cup\{t\}$; recall $K$ is the set of segments.
 For each $j\in{\mathcal{M}}$ and $k\in K$, we add an arc $e=(j,k)$ with gain factor $\gamma_{e}=a_{jk}$ and capacity $h_{e}=\infty$.
@@
 Assume that agents are competing for access to $m$ machines $\mathcal{M}$} that can produce $\ell$ different products.
-Machine $j$ can produce $a_{jk}$ units of product $k$ per time unit.
-The revenue of product $k$ is $r_{k}$.
-There are additionally limits $t_{jk}$ on how much of each product can be produced on each machine by the agent, as well a limit $d_{k}$ on the maximum possible amount producible from $k$.
+The machine $j$ can produce $a_{jk}$ units of product $k$ per time unit.
+The unit revenue of product $k$ is $r_{k}$.
+There are also limits $t_{jk}$ on the quantity of each product $k$ that the agent can produce on machine $j$, as well a limit $d_{k}$ on the maximum possible amount producible from $k$.
 Given $x_{j}$ time on machine $j\in{\mathcal{M}}$, the utility of the agent is the maximum possible revenue subject to these constraints.
-This can be formulated using a two layer generalized flow network similarly to Leontief-free as above (albeit with additional capacity bounds $t_{jk}$.) The model allows for more complex production networks with possibly multiple layers corresponding to raw materials, intermediary and final products.
+This can be formulated using a two-layer generalized flow network, similar to Leontief-free case described above, but with additional capacity bounds $t_{jk}$.
+See Figure 2 for an illustration.
+The model allows for more complex production networks with possibly multiple layers corresponding to raw materials, intermediary and final products.
+$x_{1}\bullet$$1\bullet$$x_{2}\bullet$$\vdots$$\vdots$$k\bullet$$\bullet$$x_{i}\bullet$$\vdots$$\vdots$$\ell\bullet$$x_{m}\bullet$$\scriptstyle{(r_{k},d_{k})}$$\scriptstyle{\quad\quad\quad(a_{jk},t_{jk})}\scriptstyle{\vdots}$
+Figure 2: Utilities from production.
 2.2 Competitive equilibrium
-In a Fisher market instance, given is a set of agents ${\mathcal{A}}$ and a set of divisible goods ${\mathcal{M}}$, with one unit available from each good.
+In a Fisher market instance, we are given a set of agents $\mathcal{A}$} and a set of divisible goods $\mathcal{M}$}, with one unit available of each good.
 Agents have utility functions $u_{i}\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ and budgets $b_{i}\in\mathbb{R}_{>0}$, $i\in{\mathcal{A}}$; we denote the budget vector as $\mathbf{b}=(b_{i})_{i\in{\mathcal{A}}}$.
-An allocation vector is $\mathbf{x}\in\mathbb{R}_{\geq 0}^{{\mathcal{A}}\times{\mathcal{M}}}$, where $x_{ij}$ denotes the amount of good $j$ allocated to agent $i$.
+An allocation is represented by $(\mathbf{x}_{i})_{i\in{\mathcal{A}}}$, where $x_{ij}$ denotes the amount of good $j$ allocated to agent $i$.
 A price vector is a vector $\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$.
@@
 Definition 2.7 (Competitive equilibrium).
-Given a Fisher market instance as above, the allocations and prices $(\mathbf{x},\mathbf{p})$ form a competitive (market) equilibrium if the following hold:
+Given a Fisher market instance as above, the allocations and prices $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$ form a competitive (market) equilibrium if the following hold:
   1. (i)
-    every agent gets an optimal utility at these prices: $x_{i}\in{\mathcal{D}_{i}}(\mathbf{p},b_{i})$ for every agent $i\in{\mathcal{A}}$
+    every agent gets an optimal utility at these prices: $\mathbf{x}_{i}\in{\mathcal{D}_{i}}(\mathbf{p},b_{i})$ for every agent $i\in{\mathcal{A}}$.
   2. (ii)
@@
   3. (iii)
-    every item with positive price is fully sold: $\sum_{i}x_{ij}=1$ if $p_{j}>0$.
+    every good with positive price is fully sold: $\sum_{i}x_{ij}=1$ if $p_{j}>0$.
 In a demand-approximate competitive equilibrium, agents may have suboptimal bundles:
 Definition 2.8 (Approximate competitive equilibrium).
-Given a Fisher market instance and $\alpha\geq 1$, the allocations and prices $(\mathbf{x},\mathbf{p})$ form a $\alpha$-demand approximate competitive equilibrium if
+Given a Fisher market instance and $\alpha\geq 1$, the allocations and prices $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$ form an $\alpha$-demand approximate competitive equilibrium if
   1. (i)
-    $\left\langle\mathbf{p},x_{i}\right\rangle\leq b_{i}$ for each $i\in{\mathcal{A}}$, and $i$ gets at least $\alpha$-fraction of the optimal utility at these prices: $u_{i}(x_{i})\geq\frac{1}{\alpha}\max_{x_{i}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\{u_{i}(x_{i})\,:\,\left\langle\mathbf{p},x_{i}\right\rangle\leq b_{i}\}$.
+    $\left\langle\mathbf{p},\mathbf{x}_{i}\right\rangle\leq b_{i}$ for each $i\in{\mathcal{A}}$, and $i$ gets at least $\alpha$-fraction of the optimal utility at these prices: $u_{i}(\mathbf{x}_{i})\geq\frac{1}{\alpha}\max_{\mathbf{x}_{i}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\{u_{i}(\mathbf{x}_{i})\,:\,\left\langle\mathbf{p},\mathbf{x}_{i}\right\rangle\leq b_{i}\}$.
   2. (ii)
@@
   3. (iii)
-    every item with positive price is fully sold: $\sum_{i}x_{ij}=1$ if $p_{j}>0$.
-It is a well-established fact that when all $b_{i}$’s are identical, the competitive equilibrium attains exact envy-freeness.
+    every good with positive price is fully sold: $\sum_{i}x_{ij}=1$ if $p_{j}>0$.
+It is well known that when all $b_{i}$’s are identical, competitive equilibria yield an allocation that is exactly envy-free.
 Here, we demonstrate that our definition of approximate competitive equilibrium also results in approximate envy-freeness.
 Definition 2.9.
-For a given allocation $\mathbf{x}\in\mathbb{R}_{\geq 0}^{{\mathcal{A}}\times{\mathcal{M}}}$, we say it is $\beta$-envy-free if, for any agents $i$ and $j$, the condition $u_{i}(x_{i})\geq\frac{1}{\beta}u_{i}(x_{j})$ holds.
+For a given allocation $(\mathbf{x}_{i})_{i\in{\mathcal{A}}}$, we say it is $\beta$-envy-free if, for any agents $i$ and $j$, the condition $u_{i}(\mathbf{x}_{i})\geq\frac{1}{\beta}u_{i}(\mathbf{x}_{j})$ holds.
 Lemma 2.10.
-When $b_{i}$’s are identical, the allocation $\mathbf{x}$ of an $\alpha$-approximate competitive equilibrium is also $\alpha$-envy-free.
+When $b_{i}$’s are identical, the allocation $(\mathbf{x}_{i})_{i\in{\mathcal{A}}}$ of an $\alpha$-approximate competitive equilibrium is also $\alpha$-envy-free.
 Proof.
-This follows directly from $u_{i}(x_{i})\geq\frac{1}{\alpha}\max_{x_{i}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\left\{u_{i}(x_{i})\,:\,\left\langle\mathbf{p},x_{i}\right\rangle\leq b_{i}\right\}\geq\frac{1}{\alpha}u_{i}(x_{j})$.
+This follows directly from $u_{i}(\mathbf{x}_{i})\geq\frac{1}{\alpha}\max_{\mathbf{x}_{i}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\left\{u_{i}(\mathbf{x}_{i})\,:\,\left\langle\mathbf{p},\mathbf{x}_{i}\right\rangle\leq b_{i}\right\}\geq\frac{1}{\alpha}u_{i}(\mathbf{x}_{j})$.
 ∎
 Thrifty equilibrium
-In case there are satiable utility functions, $\left\langle\mathbf{p},x_{i}\right\rangle<b_{i}$ may be possible in an (exact) competitive equilibrium.
-Hylland and Zeckhauser (1979) showed that competitive equilibrium may not be Pareto efficient due to satiable utilities.
-Mas-Colell (1992) strengthened the definition of a competitive equilibrium to require that each agent choose a least-cost optimal plan.
+In the case of satiable utility functions, it is possible for an (exact) equilibrium to satisfy $\left\langle\mathbf{p},\mathbf{x}_{i}\right\rangle<b_{i}$.
+Hylland and Zeckhauser (1979) showed that competitive equilibria may not be Pareto efficient in the presence of satiable utilities.
+Mas-Colell (1992) strengthened the definition of competitive equilibrium by requiring that each agent select a least-cost optimal bundle.
 This is also known as a strong equilibrium.
 Definition 2.11 (Thrifty competitive equilibrium).
-Given a Fisher market instance, the allocations and prices $(\mathbf{x},\mathbf{p})$ form a thrifty competitive equilibrium if $(\mathbf{x},\mathbf{p})$ is a competitive equilibrium, and moreover, for every $i\in{\mathcal{A}}$, the allocation $x_{i}$ minimizes $\left\langle\mathbf{p},x_{i}\right\rangle$ among all allocations ${\mathcal{D}_{i}}(\mathbf{p},b_{i})$.
+Given a Fisher market instance, the allocations and prices $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$ form a thrifty competitive equilibrium if $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$ is a competitive equilibrium, and moreover, for every $i\in{\mathcal{A}}$, the allocation $\mathbf{x}_{i}$ minimizes $\left\langle\mathbf{p},\mathbf{x}_{i}\right\rangle$ among all allocations ${\mathcal{D}_{i}}(\mathbf{p},b_{i})$.
 Lemma 2.12 (Garg et al. (2022b), Gul et al. (2024)).
 Thrifty competitive equilibria always exist and are Pareto efficient.
-Note that in case all utilities are non-satiable, competitive equilibria and thrifty competitive equilibria coincide.
+Note that when all utility functions are non-satiable, every competitive equilibrium is thrifty.
 2.3 Nash welfare
 Definition 2.13.
-Given a Fisher market instance, the Nash welfare of an allocation $\mathbf{y}\in\mathbb{R}_{\geq 0}^{{\mathcal{A}}\times{\mathcal{M}}}$ is the weighted geometric mean of their utilities, namely,
-$\left(\prod_{i\in{\mathcal{A}}}u_{i}(x_{i})^{b_{i}}\right)^{\frac{1}{\sum_{i\in{\mathcal{A}}}b_{i}}}\,.$
-The allocation $\mathbf{y}\in\mathbb{R}_{\geq 0}^{{\mathcal{A}}\times{\mathcal{M}}}$ maximizes the (weighted) Nash welfare, if its Nash welfare is maximal along all allocations with $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1$ for all $j\in{\mathcal{M}}$, $\mathbf{y}\geq\mathbf{0}$.
-Note that the Nash welfare maximizing allocations are precisely the optimal solutions to Eisenberg–Gale convex program Eisenberg and Gale (1959).
-[…]
+Given a Fisher market instance, the Nash welfare of an allocation $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$ is the weighted geometric mean of their utilities, namely,
+$\left(\prod_{i\in{\mathcal{A}}}u_{i}(\mathbf{y}_{i})^{b_{i}}\right)^{\frac{1}{\sum_{i\in{\mathcal{A}}}b_{i}}}\,.$
+The allocation $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$ maximizes the (weighted) Nash welfare, if it achieves the highest Nash welfare among all allocations satisfying $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1$ for all $j\in{\mathcal{M}}$.
+It is well known that the Nash welfare maximizing allocations correspond exactly to the optimal solutions of the convex program introduced by Eisenberg and Gale (1959).
+$\max$ $\sum_{i\in{\mathcal{A}}}b_{i}\log u_{i}(\mathbf{y}_{i})\ \ \text{s.t.}$ $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1,\ \forall j\in{\mathcal{M}}\ \ \text{and}$ $\penalty 10000\ \penalty 10000\ \penalty 10000\ y_{ij}\geq 0,\ \forall i\in{\mathcal{A}},j\in{\mathcal{M}}\,.$ (EG)
 The Lagrangian dual of (EG) can be written as
-$\varphi(\mathbf{q})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{i\in{\mathcal{A}}}\max_{y_{i}\in\mathbb{R}_{\geq 0}^{{\mathcal{M}}}}\left(b_{i}\log u_{i}(y_{i})-\left\langle\mathbf{q},y_{i}\right\rangle\right)+\sum_{j\in{\mathcal{M}}}q_{j}.$ (4)
+$\varphi(\mathbf{q})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{i\in{\mathcal{A}}}\max_{\mathbf{y}_{i}\in\mathbb{R}_{\geq 0}^{{\mathcal{M}}}}\left(b_{i}\log u_{i}(\mathbf{y}_{i})-\left\langle\mathbf{q},\mathbf{y}_{i}\right\rangle\right)+\sum_{j\in{\mathcal{M}}}q_{j}.$ (4)
 We define the domain $\operatorname{dom}(\varphi)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\cap_{i\in{\mathcal{A}}}\operatorname{pdom}(u_{i})$ (see (3)).
 Thus, $\operatorname{dom}(\varphi)$ is the set of nonnegative prices $\mathbf{q}$ such that $\varphi(\mathbf{q})<\infty$.
-Strong duality holds and the dual optimum is taken according since Slater’s condition holds.
-Consequently, the optimal Nash welfare value equals the dual optimum
-$\min_{\mathbf{q}\in\operatorname{dom}(\varphi)}\varphi(\mathbf{q})\,.$ (5)
-This optimum is taken (see Lemma 6.2), and we refer to the optimal $\mathbf{q}\in\operatorname{dom}(\varphi)$ as Gale prices.
+Slater’s condition is satisfied; therefore strong duality holds, and the dual optimum is attained.
+Consequently, the optimal Nash welfare value is equal to
+$\min_{\mathbf{q}\in\operatorname{dom}(\varphi)}\varphi(\mathbf{q})\,$, (5)
+and we refer to the optimal $\mathbf{q}\in\operatorname{dom}(\varphi)$ as Gale prices.
 We note that these may not be unique.
@@
 The term corresponding to agent $i$ in the definition of $\varphi(\mathbf{q})$ is the value of the assignments in the Gale demand correspondence of this agent at prices $\mathbf{q}$ and budget $b_{i}$.
-We use the notation $\mathcal{GD}_{i}(\mathbf{q},b_{i})$ for the Gale demand for the utility $u_{i}$, $i\in{\mathcal{A}}$.
+We denote by $\mathcal{GD}_{i}(\mathbf{q},b_{i})$ the Gale demand corresponding to the utility function $u_{i}$, $i\in{\mathcal{A}}$.
 2.4 Gale substitute utilities
@@
   1. (i)
-    for any $b>0$, any $\mathbf{q},\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$, and any $y\in\mathcal{GD}^{u}(\mathbf{q},b)$, if $\mathbf{q}\geq\mathbf{q}^{\prime}$, then there exists a $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ such that $y^{\prime}_{j}\leq y_{j}$ whenever $q_{j}^{\prime}=q_{j}$;
+    for any $b>0$, any $\mathbf{q},\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$, and any $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$, if $\mathbf{q}\geq\mathbf{q}^{\prime}$, then there exists a $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ such that $y^{\prime}_{j}\leq y_{j}$ whenever $q_{j}^{\prime}=q_{j}$;
   2. (ii)
-    for any $b\geq b^{\prime}>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$, and for any $y\in\mathcal{GD}^{u}(\mathbf{q},b)$, there exists a $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$ such that $y^{\prime}\leq y$.
-We say that $u(\cdot)$ is $\Sigma$-Gale-substitutes, if it is Gale-substitutes and the following additional property holds:
-$\mbox{if }\mathbf{0}\in\operatorname{pdom}(u)\,,\,b>0\mbox{ and }\mathbf{q}\in\mathbb{R}_{\geq 0}^{\mathcal{M}},\mbox{ then for any }y\in\mathcal{GD}(\mathbf{0},b),\exists y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b)\,:\,y^{\prime}\leq y\,.$ ($\Sigma$)
+    for any $b\geq b^{\prime}>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$, and for any $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$, there exists a $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$ such that $\mathbf{y}^{\prime}\leq\mathbf{y}$.
+We say that $u(\cdot)$ is $\Sigma$-Gale-substitutes, if it satisfies the Gale-substitutes condition along with the following additional property:
+$\mbox{if }\mathbf{0}\in\operatorname{pdom}(u)\,,\,b>0\mbox{ and }\mathbf{q}\in\mathbb{R}_{\geq 0}^{\mathcal{M}},\mbox{ then for any }\mathbf{y}\in\mathcal{GD}(\mathbf{0},b),\exists\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b)\,:\,\mathbf{y}^{\prime}\leq\mathbf{y}\,.$ ($\Sigma$)
 Remark 2.15.
 As noted above, $\mathbf{0}\in\operatorname{pdom}(u)$ if and only if $u$ is satiable; thus, ($\Sigma$) is vacuously true for non-satiable utilities.
-In the satiable case, it asserts that if $y$ is a bundle of maximum possible utility, then at any prices there is an optimal bundle below $y$.
+In the satiable case, it asserts that if $\mathbf{y}$ is a bundle of maximum possible utility, then at any prices there is an optimal bundle below $\mathbf{y}$.
 Observe that this is immediate for separable utilities.
-On the other hand, ($\Sigma$) fails for the simple Leontief-free utility $u(x)=\max\{x_{1}+x_{2},1\}$: $y=(1,0)$ is an optimal bundle, whereas for $b=1$, $\mathbf{q}=(2,1)$, the unique solution in $\mathcal{GD}^{u}(\mathbf{q},b)$ is $y^{\prime}=(0,1)$.
-Next, we demonstrate a utility function that exhibits Gale-substitutes but not substitutes properties.
+On the other hand, ($\Sigma$) fails for the simple Leontief-free utility $u(\mathbf{x})=\max\{x_{1}+x_{2},1\}$: $\mathbf{y}=(1,0)$ is an optimal bundle at $\mathbf{q}=\mathbf{0}$, whereas for $b=1$, $\mathbf{q}=(2,1)$, the unique solution in $\mathcal{GD}^{u}(\mathbf{q},b)$ is $\mathbf{y}^{\prime}=(0,1)$.
+Next, we present a utility function that satisfies the Gale-substitutes property but does not exhibit the weak gross substitutes (WGS) property.
 Example 2.16.
-Consider an agent with budget $b$ and a utility function over two goods given by $u(x)=\min\{x_{1},0.4\}+x_{2}$.
-This utility is separable and, as such, Gale-substitutes according to Lemma 7.1 shown later.
-However, we show that this is not weak gross substitutes (WGS): there exist prices $\mathbf{p}\leq\mathbf{p}^{\prime}$ and $x\in{\mathcal{D}}^{u}(\mathbf{p},b)$ such that for some good $j$ with $p_{j}=p^{\prime}_{j}$, every $x^{\prime}\in{\mathcal{D}}^{u}(\mathbf{p},b)$ has $x^{\prime}_{j}<x_{j}$.
+Consider an agent with budget $b$ and a utility function over two goods given by $u(\mathbf{x})=\min\{x_{1},0.4\}+x_{2}$.
+This utility is separable and, as such, Gale-substitutes (see Theorem 1.4).
+However, we show that this is not WGS: there exist prices $\mathbf{p}\leq\mathbf{p}^{\prime}$ and $\mathbf{x}\in{\mathcal{D}}^{u}(\mathbf{p},b)$ such that for some good $j$ with $p_{j}=p^{\prime}_{j}$, every $\mathbf{x}^{\prime}\in{\mathcal{D}}^{u}(\mathbf{p},b)$ has $x^{\prime}_{j}<x_{j}$.
 That is, let $\mathbf{p}=(0.8,1)$ and $\mathbf{p}^{\prime}=(0.9,1)$, and $b=1$.
-It is easy to see that $x=(0.4,0.68)$ is the unique optimal bundle at prices $\mathbf{p}$, and $x^{\prime}=(0.4,0.64)$ is the unique optimal bundle at prices $\mathbf{p}^{\prime}$ in contradiction to WGS property.
-3 Overview of techniques
-The ‘price of anarchy’ result Theorem 1.1 is proved using an elementary argument on concavity and the log-sum inequality.
-The proof is inspired by an argument in Theorem 1 in Brânzei et al.
+It is easy to see that $\mathbf{x}=(0.4,0.68)$ is the unique optimal bundle at prices $\mathbf{p}$, and $\mathbf{x}^{\prime}=(0.4,0.64)$ is the unique optimal bundle at prices $\mathbf{p}^{\prime}$ in contradiction to the WGS property.
+We note that the gross substitutes property is one of the key conditions that allows for the efficient computation of market equilibria.
+However, SPLC utilities do not satisfy the gross substitutes property, as illustrated in the above example.
+In a later section, we show that SPLC preferences are Gale-substitutes, which forms the foundation of our main results.
+3 The price of anarchy bound
+We now show the ‘price of anarchy’ result stated in Theorem 1.1.
+The proof employs an argument based on concavity and the log-sum inequality, drawing inspiration from Theorem 1 in Brânzei et al.
 (2022).
-Example A.4 shows that the bound $(1/\mathrm{e})^{1/\mathrm{e}}$ is tight.
-To prove Theorem 1.2 showing that the Nash welfare maximizing allocation is 2-demand-approximate CE, the key ingredient is Lemma 5.1, showing that the solutions in the Gale demand system are within budget, that is, for $y\in\mathcal{GD}^{u}(\mathbf{q},b)$ we have $\left\langle\mathbf{q},y\right\rangle\leq b$.
-Since at the Nash welfare maximizing solution, every agent gets an allocation from the Gale demand system at the dual prices $\mathbf{q}$, this lemma guarantees that the allocation is feasible.
-The 2-demand-approximate property then follows by a supergradient argument.
-The proof of Theorem 1.3 relies on two key lemmas derived using Lagrangian duality and supergradient arguments.
+Theorem 3.1 (Restatement of Theorem 1.1).
+Given a Fisher market instance with concave utility functions, consider any competitive equilibrium $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$ and a Nash welfare maximizing allocation $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$.
+Then,
+$\left(\prod_{i\in{\mathcal{A}}}u_{i}(\mathbf{x}_{i})^{b_{i}}\right)^{\frac{1}{\sum_{i\in{\mathcal{A}}}b_{i}}}\geq\left(\frac{1}{\mathrm{e}}\right)^{\frac{1}{\mathrm{e}}}\left(\prod_{i\in{\mathcal{A}}}u_{i}(\mathbf{y}_{i})^{b_{i}}\right)^{\frac{1}{\sum_{i\in{\mathcal{A}}}b_{i}}}.$
+Proof.
+Without loss of generality, we can assume $\sum_{i\in{\mathcal{A}}}b_{i}=1$.
+Note that since $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$ is a competitive equilibrium, then $u_{i}(\mathbf{x}_{i})\geq\min\left\{1,\frac{b_{i}}{\left\langle\mathbf{p},\mathbf{y}_{i}\right\rangle}\right\}u_{i}(\mathbf{y}_{i})$, as agent $i$ can use their budget, $b_{i}$, to buy at least $\frac{b_{i}}{\left\langle\mathbf{p},\mathbf{y}_{i}\right\rangle}$ fraction of $\mathbf{y}_{i}$.
+The lower bound holds because if $b_{i}<\left\langle\mathbf{p},\mathbf{y}_{i}\right\rangle$, then $u_{i}\left(\frac{b_{i}}{\left\langle\mathbf{p},\mathbf{y}_{i}\right\rangle}\mathbf{y}_{i}\right)\geq\frac{b_{i}}{\left\langle\mathbf{p},\mathbf{y}_{i}\right\rangle}u_{i}(\mathbf{y}_{i})$ using concavity and $u_{i}(\mathbf{0})=0$.
+Thus,
+$\sum_{i\in{\mathcal{A}}}{b_{i}}\log\frac{u_{i}(\mathbf{y}_{i})}{u_{i}(\mathbf{x}_{i})}\leq\sum_{i\in{\mathcal{A}}}{b_{i}}\log\max\left\{1,\frac{\left\langle\mathbf{p},\mathbf{y}_{i}\right\rangle}{b_{i}}\right\}.$
+Let $b^{\prime}_{i}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}{\left\langle\mathbf{p},\mathbf{y}_{i}\right\rangle}$, and let $\mathcal{A}$$^{\prime}$ be the set of agents such that $b^{\prime}_{i}\geq b_{i}$, $B\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{i\in\mathcal{A}^{\prime}}b_{i}$, and $B^{\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{i\in\mathcal{A}^{\prime}}b^{\prime}_{i}$.
+Therefore, using the log sum inequality,
+$\sum_{i\in{\mathcal{A}}}{b_{i}}\log\frac{u_{i}(\mathbf{y}_{i})}{u_{i}(\mathbf{x}_{i})}\leq\sum_{i\in{\mathcal{A}}^{\prime}}{b_{i}}\log\frac{b^{\prime}_{i}}{b_{i}}\leq B\log\frac{B^{\prime}}{B}\,.$
+By definition, $B\leq\sum_{i\in{\mathcal{A}}}b_{i}=1$.
+We also note that $B^{\prime}\leq\sum_{i\in{\mathcal{A}}}b_{i}^{\prime}=\sum_{i\in{\mathcal{A}}}\left\langle\mathbf{p},\mathbf{y}_{i}\right\rangle\leq\sum_{j\in{\mathcal{M}}}p_{j}\leq\sum_{i\in{\mathcal{A}}}b_{i}\leq 1$.
+$\sum_{i\in{\mathcal{A}}}{b_{i}}\log\frac{u_{i}(\mathbf{y}_{i})}{u_{i}(\mathbf{x}_{i})}\leq\max_{B\leq 1}B\log\frac{1}{B}\leq\frac{1}{\mathrm{e}}\,.$
+The statement follows. ∎
+The following example demonstrates that the bound $(1/\mathrm{e})^{1/\mathrm{e}}$ is tight.
+Example 3.2 (Tight example on price of anarchy).
+Consider an instance with two agents and two goods. The first agent has a budget of $1$ and has a utility function:
+$u_{1}(\mathbf{x}_{1})=x_{11}+\varepsilon x_{12},$
+and the second agent has a budget of $\mathrm{e}-1$ and has a utility function
+$u_{2}(\mathbf{x}_{2})=\varepsilon x_{21}+\min\{x_{22},1-\varepsilon\},$
+where $\varepsilon>0$ is sufficiently small.
+The maximum Nash welfare is at least $(1-\varepsilon)^{(\mathrm{e}-1)/\mathrm{e}}$, taken at $\mathbf{y}_{1}=(1,0)$ and $\mathbf{y}_{2}=(0,1)$.
+Conversely, the unique competitive equilibrium is characterized by prices $p_{1}=\mathrm{e}(\frac{1}{1+\varepsilon})$ and $p_{2}=\mathrm{e}(\frac{\varepsilon}{1+\varepsilon})$, with corresponding allocations given by:
+$\mathbf{x}_{1}=\left(\frac{1+\varepsilon-\mathrm{e}\varepsilon^{2}}{\mathrm{e}},\varepsilon\right),\quad\mathbf{x}_{2}=\left(1-\frac{1+\varepsilon-\mathrm{e}\varepsilon^{2}}{\mathrm{e}},1-\varepsilon\right).$
+The Nash welfare objective for this allocation is $\left(\frac{1+\varepsilon}{\mathrm{e}}\cdot\left[1-\varepsilon+\varepsilon\cdot\left(1-\frac{1+\varepsilon-\mathrm{e}\varepsilon^{2}}{\mathrm{e}}\right)\right]^{\mathrm{e}-1}\right)^{\frac{1}{\mathrm{e}}}$.
+When $\varepsilon\to 0$, the Nash welfare ratio between these two allocations converges to $(1/\mathrm{e})^{1/\mathrm{e}}$.
+The above example also shows that the price of stability is the same as the price of anarchy, as there is a unique CE.
+4 Nash welfare approximates competitive equilibrium
+We now derive Theorem 1.2, showing that Nash welfare maximizing allocations can be seen as approximate competitive equilibria.
+Our first simple lemma shows that the Gale demand solutions are feasible with respect to the budget:
+Lemma 4.1.
+Given any utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$, $b>0$ and $\mathbf{q}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, for any $\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $\left\langle\mathbf{q},\mathbf{y}\right\rangle>b$, there exists $\mathbf{y}^{\prime}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $\left\langle\mathbf{q},\mathbf{y}^{\prime}\right\rangle\leq b$ and $b\log u(\mathbf{y}^{\prime})-\left\langle\mathbf{q},\mathbf{y}^{\prime}\right\rangle>b\log u(\mathbf{y})-\left\langle\mathbf{q},\mathbf{y}\right\rangle$.
+In particular, if $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$, then, $\left\langle\mathbf{q},\mathbf{y}\right\rangle\leq b$.
+Proof.
+Let $\mathbf{y}^{\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\mathbf{y}\cdot\frac{b}{\left\langle\mathbf{q},\mathbf{y}\right\rangle}$.
+By concavity and $u(\mathbf{0})=0$, we get $u(\mathbf{y}^{\prime})\geq\frac{b}{\left\langle\mathbf{q},\mathbf{y}\right\rangle}u(\mathbf{y})$.
+Thus,
+$b\log u(\mathbf{y}^{\prime})-\left\langle\mathbf{q},\mathbf{y}^{\prime}\right\rangle$ $\geq b\log u(\mathbf{y})-b\log\frac{\left\langle\mathbf{q},\mathbf{y}\right\rangle}{b}-b>b\log u(\mathbf{y})-\left\langle\mathbf{q},\mathbf{y}\right\rangle\,$,
+where the second inequality used $\log(\alpha)+1<\alpha$ for $\alpha>1$. ∎
+The next lemma follows directly from the subgradient Karush–Kuhn–Tucker (KKT) conditions used to characterize solutions in the Gale demand correspondence.
+For a detailed treatment of Lagrangian duality and KKT conditions for nondifferentiable functions, we refer the reader to Chapter 3 of Ruszczyński (2011).
+Lemma 4.2.
+For a utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$, $b>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$, let $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$.
+Then, there exists a supergradient $\mathbf{g}\in\partial^{\wedge}u(\mathbf{y})$ such that
+$g_{j}\leq{q_{j}}\cdot\frac{u(\mathbf{y})}{b}\,,\quad\forall j\in{\mathcal{M}}\,,\,\mbox{with equality if }y_{j}>0\,$,
+Conversely, if such a $\mathbf{g}\in\partial^{\wedge}u(\mathbf{y})$ exists then $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$.
+See 1.2
+Proof.
+By Lemma 4.1, all agents satisfy the budget constraints, $\left\langle\mathbf{q},\mathbf{y}_{i}\right\rangle\leq b_{i}$ for all $i$.
+Additionally, the market clears by the KKT conditions for (EG).
+For the approximate competitive equilibrium, we only need to show $u_{i}(\mathbf{y}_{i})\geq\frac{1}{2}\max_{\left\langle\mathbf{q},\mathbf{x}_{i}\right\rangle\leq b_{i}}u_{i}(\mathbf{x}_{i})$ for all agents.
+Note that $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$ is the allocation that maximizes Nash welfare, where each $\mathbf{y}_{i}$ maximizes $b_{i}\log u_{i}(\mathbf{y}_{i})-\left\langle\mathbf{q},\mathbf{y}_{i}\right\rangle$.
+By Lemma 4.2, there is a supergradient $\mathbf{g}_{i}$ such that $g_{ij}\leq q_{j}u_{i}(\mathbf{y}_{i})/b_{i}$ with equality whenever $y_{ij}>0$.
+Then, for any $\mathbf{x}_{i}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $\left\langle\mathbf{q},\mathbf{x}_{i}\right\rangle\leq b_{i}$,
+$u_{i}(\mathbf{x}_{i})$ $\leq u_{i}(\mathbf{y}_{i})+\left\langle\mathbf{g}_{i},\mathbf{x}_{i}-\mathbf{y}_{i}\right\rangle$ (convexity of $u_{i}$)
+$\leq u_{i}(\mathbf{y}_{i})+\frac{u_{i}(\mathbf{y}_{i})}{b_{i}}\left\langle\mathbf{q},\mathbf{x}_{i}-\mathbf{y}_{i}\right\rangle$ ($g_{ij}\leq q_{j}u_{i}(\mathbf{y}_{i})/b_{i}$ with equality whenever $y_{ij}>0$)
+$\leq\left(2-\frac{\left\langle\mathbf{q},\mathbf{y}_{i}\right\rangle}{b_{i}}\right)u_{i}(\mathbf{y}_{i})$ ($\left\langle\mathbf{q},\mathbf{x}_{i}\right\rangle\leq b_{i}$)
+$\leq 2u_{i}(\mathbf{y}_{i})\,.$
+This proves that $((\mathbf{y}_{i})_{i\in{\mathcal{A}}},\mathbf{q})$ forms a $2$-approximate competitive equilibrium.
+Additionally, the $2$-envy-freeness of the allocation follows directly from Lemma 2.10.
+∎
+One may wonder if the $2$-demand-approximate CE can be improved. The following example demonstrates that this bound of $2$ is tight.
+Example 4.3 (Tight example on demand-approximate CE).
+Consider the case with two agents and a single good. Both agents have unit budgets. The first agent’s utility for this good is linear:
+$u_{1}(x_{1})=x_{1}.$
+The second agent’s utility is defined as:
+$u_{2}(x_{2})=\begin{cases}\alpha x_{2}&\text{for }x_{2}\leq\frac{1}{\alpha},\\ 1+\beta\left(x_{2}-\frac{1}{\alpha}\right)&\text{for }x_{2}>\frac{1}{\alpha},\end{cases}$
+where $\alpha$ is a large number and $\beta\leq\alpha$.
+When $\beta=\frac{1-\frac{1}{\alpha}}{1-\frac{2}{\alpha}}$, with the optimal dual variable (price) $q$ equal to $q=\frac{1}{x_{1}}=\frac{1}{1-\frac{1}{\alpha}}$, the Nash welfare maximizing allocation is
+$x_{1}=1-\frac{1}{\alpha},\quad x_{2}=\frac{1}{\alpha}.$
+Under this price, agent 2 can afford $x_{2}^{\prime}=1-\frac{1}{\alpha}$ with the budget, resulting in a utility of
+$u_{2}(x_{2}^{\prime})=1+\beta\left(x_{2}^{\prime}-\frac{1}{\alpha}\right)=2-\frac{1}{\alpha}.$
+In contrast, in the Nash welfare maximizing allocation, agent 2 only receives a utility of $1$.
+Although Theorem 1.1 guarantees that the Nash welfare of any competitive equilibrium is at least a constant fraction of the optimum, the following example illustrates that agents can receive significantly different utilities across different equilibria.
+Example 4.4 (Utility discrepancy across competitive equilibria).
+Consider a market with $n$ agents and $n$ goods, where each agent has a unit budget. For $1\leq i\leq n-1$, agent $i$’s utility is given by
+$u_{i}(\mathbf{x}_{i})=\min\{x_{ii},1\}+\varepsilon x_{in},$
+for a small $\varepsilon>0$, and agent $n$’s utility is
+$u_{n}(\mathbf{x}_{n})=x_{nn}.$
+There exist at least two distinct competitive equilibria:
+  • In the first equilibrium, all goods are priced at $1$, and each good $i$ is allocated entirely to agent $i$ for $1\leq i\leq n$.
+Agent $n$ thus attains a utility of $1$.
+This equilibrium is also a Nash welfare maximizing allocation.
+  • In the second equilibrium, the first n1n-1 goods are priced at 0, while the $n$-th good is priced at $n$.
+Goods $1$ to n1n-1 are allocated to agents $1$ through n1n-1 respectively, and the $n$-th good is shared equally among all $n$ agents.
+As a result, agent $n$ obtains utility equal to $\frac{1}{n}$.
+5 Agent-wise utility guarantee
+We now turn to the proof of the main result of this paper, Theorem 1.3, asserting that for Gale-substitutes utility functions, each agent’s utility in the Nash welfare maximizing allocation is at least half of the maximum utility they could achieve in any competitive equilibrium.
+The proof relies on two key lemmas derived using Lagrangian duality and supergradient arguments.
 The first is a simple observation, showing that the Gale demand solution for lower prices yields at least half the utility of the utility maximizing bundle at the current prices.
@@
 Consider a utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$, prices $\mathbf{p},\mathbf{q}\in\operatorname{pdom}(u)$ such that $\mathbf{q}\leq\mathbf{p}$.
-Then, $u(y)\geq\frac{1}{2}u(x)$ for any $y\in\mathcal{GD}^{u}(\mathbf{q},b)$ and $x\in{\mathcal{D}}^{u}(\mathbf{p},b)$.
-The second lemma shows that for every allocation of goods in the demand correspondence, there is a smaller or equal allocation of goods in the Gale demand correspondence.
-Although the first lemma was true for any utilities, this one critically relies on assuming the $\Sigma$-Gale-substitutes property.
+Then, for any $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$ and $\mathbf{x}\in{\mathcal{D}}^{u}(\mathbf{p},b)$, it holds that $u(\mathbf{y})\geq\frac{1}{2}u(\mathbf{x})$.
+Proof.
+By Lemma 4.2, there exists a supergradient $\mathbf{g}\in\partial^{\wedge}u(\mathbf{y})$ such that $g_{j}\leq q_{j}u(\mathbf{y})/b$ with equality whenever $y_{j}>0$. Therefore,
+$u(\mathbf{x})$ $\leq u(\mathbf{y})+\left\langle\mathbf{g},\mathbf{x}-\mathbf{y}\right\rangle\leq u(\mathbf{y})+\frac{u(\mathbf{y})}{b}\left\langle\mathbf{q},\mathbf{x}-\mathbf{y}\right\rangle$
+$\leq u(\mathbf{y})\left(1+\frac{\left\langle\mathbf{q},\mathbf{x}\right\rangle}{b}\right)\leq u(\mathbf{y})\left(1+\frac{\left\langle\mathbf{p},\mathbf{x}\right\rangle}{b}\right)\leq 2u(\mathbf{y})\,.$
+The second lemma establishes that for every allocation of goods in the standard demand correspondence, there exists a smaller or equal allocation of goods in the Gale demand correspondence.
+While Lemma 5.1 holds for any utility functions, this result critically depends on the assumption that the utility function satisfies the $\Sigma$-Gale-substitutes property.
+The proof is given in Appendix B, transforming the KKT-conditions of the demand correspondence ${\mathcal{D}}^{u}(\mathbf{p},b)$ to the KKT-conditions of $\mathcal{GD}^{u}(\mathbf{p},b)$ (Lemma 4.2), by making use of property ($\Sigma$).
 Lemma 5.2.
-Let $u(\cdot)$ be a $\Sigma$-Gale-substitutes utility and let $x\in{\mathcal{D}}^{u}(\mathbf{p},b)$ for $\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ and $b>0$.
-Then, $\mathbf{p}\in\operatorname{pdom}(u)$, and there exists $y\in\mathcal{GD}^{u}(\mathbf{p},b)$ such that $y\leq x$.
-To derive Theorem 1.3, we start with a competitive equilibrium $(\mathbf{x},\mathbf{p})$, and show that the dual of the Eisenberg–Gale program (5), $\varphi(\mathbf{q})$, has an optimal solution $\mathbf{q}^{\star}$ such that $\mathbf{q}^{\star}\leq\mathbf{p}$.
+Let $u(\cdot)$ be a $\Sigma$-Gale-substitutes utility and let $\mathbf{x}\in{\mathcal{D}}^{u}(\mathbf{p},b)$ for $\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ and $b>0$.
+Then, $\mathbf{p}\in\operatorname{pdom}(u)$, and there exists $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{p},b)$ such that $\mathbf{y}\leq\mathbf{x}$.
+We now restate Theorem 1.3, and sketch the proof. The full proof is given in Appendix B. See 1.3
+Proof sketch.
+We start with a competitive equilibrium $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$, and show that the dual of the Eisenberg–Gale program (5), $\varphi(\mathbf{q})$, has an optimal solution $\mathbf{q}^{\star}$ such that $\mathbf{q}^{\star}\leq\mathbf{p}$.
 The statement then follows from Lemma 5.1.
-An intuitive argument would be the following: we start from $\mathbf{q}=\mathbf{p}$ where the total Gale demand is less than the supply (by Lemma 3.2), and use a tâtonnement-like procedure using the Gale-substitutes property.
-As long as $\mathbf{q}$ is not a minimizer of $\varphi(.)$, there will be an underdemanded item $j$ with respect to Gale demands.
-Reducing $p_{j}$ for such a good $j$ may only decrease the Gale demand for all agents on all other goods.
+An intuitive argument is as follows: starting from $\mathbf{q}=\mathbf{p}$, where the total Gale demand is less than the supply (by Lemma 5.2), we decrease the prices in a tâtonnement-like procedure Cheung et al.
+(2020) using the Gale-substitutes property.
+As long as $\mathbf{q}$ is not a minimizer of $\varphi(.)$, there will be an underdemanded good $j$ with respect to Gale demands.
+Reducing $q_{j}$ for such a good $j$ may only decrease the Gale demand for all agents on all other goods.
 This iterative process converges to an equilibrium point and, since the prices are only decreasing, we can conclude that the optimal limit price vector satisfies $\mathbf{q}^{\star}\leq\mathbf{p}$.
-[…]
-In Section 7, we prove Theorem 1.4, i.e., every separable utility function is $\Sigma$-Gale-substitutes.
-We show the theorem for general, not necessarily differentiable utility functions using subgradient arguments.
-We next outline the argument under the simplying assumption that utilities are differentiable and strictly concave.
-Let $u(y)=\sum_{j\in{\mathcal{M}}}v_{j}(y_{j})$, and let $\nabla_{j}u(y)=\partial v_{j}(y_{j})/\partial y_{j}$ denote the $j$-th component of the gradient.
-Then, the KKT conditions yield $\nabla_{j}u(y)\leq q_{j}u(y)/b$, with equality whenever $y_{j}>0$ (see Lemma 7.2).
-Let $y\in\mathcal{GD}^{u}(\mathbf{q},b)$ and $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ for $\mathbf{q}^{\prime}\leq\mathbf{q}$.
-Using the optimality conditions, it is easy to derive that $u(y^{\prime})\geq u(y)$.
-Consequently, for every good $j$ with $q_{j}=q^{\prime}_{j}$ and $y_{j}>0$, $\nabla_{j}u(y^{\prime})\geq\nabla_{j}u(y)$ follows.
+Rather than analysing this continuous procedure, the proof in Appendix B establishes $\mathbf{q}^{\star}\leq\mathbf{p}$ directly, using Lagrangian duality.
+For any price vector $\mathbf{q}$, we get a subgradient in $\partial^{\vee}\varphi(\mathbf{q})$ from the difference between the supply and the total Gale demand.
+Further, for $\mathbf{q}\leq\mathbf{p}$ we get that if $q_{j}=p_{j}$ then the $j$-th component of this subgradient at $\mathbf{q}$ is non-negative.
+This is shown in Lemma B.1, by making use of an allocation $\mathbf{y}$ derived using Lemma 5.2 for the prices $\mathbf{q}$.
+Thus, in the region $\{\mathbf{q}\penalty 10000\ |\penalty 10000\ \mathbf{0}\leq\mathbf{q}\leq\mathbf{p}\}$, the subgradients on the boundary all point outward.
+Consequently, if $\mathbf{q}^{\star}\not\leq\mathbf{p}$, then we can take the projection $\mathbf{q}$ of $\mathbf{q}^{\star}$ to the boundary of the region, where this subgradient shows that $\mathbf{q}$ is also optimal.
+∎
+The following example shows the bound in Theorem 1.3 is tight.
+Example 5.3.
+Consider a market with $n$ agents and $n$ goods, where all agents have bounded linear utilities. Agent 1’s utility is
+$u_{1}(\mathbf{x}_{1})=\sum_{j}\min\{1,x_{1j}\},$
+while, for all other agents ($2\leq i\leq n$), they all have identical preferences on all goods except for the first one
+$u_{i}(\mathbf{x}_{i})=\sum_{j>1}\min\{1,x_{ij}\}.$
+In the Nash welfare maximizing allocation, each agent receives a bundle yielding utility exactly equal to $1$.
+However, consider a competitive equilibrium in which the price of the first good is 0, and the prices of all other goods are set to $\frac{n}{n-1}$.
+In this equilibrium, agent $1$ obtains the entire first good and a $\frac{n-1}{n}$ fraction of the other goods, leading to a total utility of
+$u_{1}=1+\frac{n-1}{n},$
+which approaches $2$ as $n\to\infty$.
+Remark 5.4.
+The assumption that the utilities are $\Sigma$-Gale-substitutes is important in Theorem 1.3.
+Example E.5 demonstrates that for non-Gale substitute utilities, some agents may achieve significantly higher utility in a competitive equilibrium than in the Nash welfare maximizing allocation.
+Remark 5.5.
+Theorem 1.3 naturally raises a question about the converse: for any agent $i$, does there always exist a CE $(\mathbf{x}_{i})_{i\in{\mathcal{A}}}$ such that $u_{i}(\mathbf{x}_{i})\geq C\cdot u_{i}(\mathbf{y}_{i})$ for some constant $C>0$, where $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$ is a Nash welfare maximizing allocation?
+We show that such a constant-factor guarantee is not always possible by presenting an example (see Example E.4) in which the CE is unique, yet there exists an agent whose utility under this equilibrium is significantly lower than in the Nash welfare maximizing allocation.
+6 Gale substitute utilities
+First, we show that separable utilities are $\Sigma$-Gale-substitutes. See 1.4 For this, we first adapt Lemma 4.2 to the setting of separable utilities.
+Lemma 6.1.
+Let $u(\mathbf{x})=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$ be a separable utility function.
+For any $b>0$, $\mathbf{q}\in\operatorname{pdom}(u)$, $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$, we have $u(\mathbf{y})>0$.
+Further, $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$ if and only if there exist supergradients $g_{j}\in\partial^{\wedge}v_{j}(y_{j})$, $j\in{\mathcal{M}}$ satisfying
+  1. (i)
+    $g_{j}\geq 0$; and
+  2. (ii)
+    $g_{j}\leq q_{j}u(\mathbf{y})/b$ with equality whenever $y_{j}>0$.
+We now show that strictly concave and differentiable utility functions satisfy Definition 2.14 (i) : for any $b>0$, any $\mathbf{q},\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$, and any $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$, if $\mathbf{q}\geq\mathbf{q}^{\prime}$, then there exists a $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ such that $y^{\prime}_{j}\leq y_{j}$ whenever $q_{j}^{\prime}=q_{j}$.
+Proof of Theorem 1.4 for strictly concave and differentiable utilities.
+Let $u(\mathbf{y})=\sum_{j\in{\mathcal{M}}}v_{j}(y_{j})$ be a separable and differentiable concave utility function, and let $\nabla_{\!j}u(\mathbf{y})=\partial v_{j}(y_{j})/\partial y_{j}$ denote the $j$-th component of the gradient.
+Recall that for differentiable utilities, the only supergradient of $v_{j}(y_{j})$ is $\nabla_{\!j}u(\mathbf{y})$.
+By Lemma 6.1, $\nabla_{\!j}u(\mathbf{y})\leq q_{j}u(\mathbf{y})/b$, with equality whenever $y_{j}>0$.
+Let $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$ and $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ for $\mathbf{q}^{\prime}\leq\mathbf{q}$.
+We first claim that $u(\mathbf{y}^{\prime})\geq u(\mathbf{y})$. For a contradiction, assume $u(\mathbf{y}^{\prime})<u(\mathbf{y})$. From the above, whenever $y_{j}>0$, we have
+$\nabla_{\!j}u(\mathbf{y})=q_{j}u(\mathbf{y})/b\geq q^{\prime}_{j}u(\mathbf{y}^{\prime})/b\geq\nabla_{\!j}u(\mathbf{y}^{\prime})\,.$
+By strict concavity, this means $y^{\prime}_{j}\geq y_{j}$.
+Thus, we see that $\mathbf{y}^{\prime}\geq\mathbf{y}$, a contradiction to the monotonicity of $u(.)$ and the assumption $u(\mathbf{y}^{\prime})<u(\mathbf{y})$.
+From $u(\mathbf{y}^{\prime})\geq u(\mathbf{y})$, for every good $j$ with $q_{j}=q^{\prime}_{j}$ and $y_{j}>0$, $\nabla_{\!j}u(\mathbf{y}^{\prime})\geq\nabla_{\!j}u(\mathbf{y})$ follows.
 By strict concavity, we must have $y^{\prime}_{j}\leq y_{j}$ as required.
-The proof without the differentiability and strict concavity assumptions poses additional difficulties: for example, we would not be able to conclude $y^{\prime}_{j}\leq y_{j}$ at this point, and might need a specific choice of $y^{\prime}_{j}$ from $\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$.
-Theorem 1.5 shows that generalized network utilities are Gale substitutes.
-To verify the Gale substitutes property, we need to take two price vectors $\mathbf{q}\geq\mathbf{q}^{\prime}$ and $y\in\mathcal{GD}^{u}(\mathbf{q},b)$, and construct a $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q^{\prime}},b)$ such that $y^{\prime}_{j}\leq y_{j}$ whenever $q^{\prime}_{j}=q_{j}$; we have a second, analogous construction for budget changes.
-The proof is a finite algorithm that transforms the optimal generalized flow for demands $y$ to $y^{\prime}$ such that we throughout maintain primal-dual slackness condition corresponding to the Gale demand correspondence.
+∎
+The proof without the differentiability and strict concavity assumptions poses additional difficulties: for example, we would not be able to conclude $y^{\prime}_{j}\leq y_{j}$ in the last step, and might need a specific choice of $y^{\prime}_{j}$ from $\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$.
+The proof of the general case is given in Appendix C.
+Our final theorem is the Gale substitutability of generalized network utility functions. This is deferred to Appendix D, and we only present a high level summary here. See 1.5
+To verify the Gale substitutes property, we need to take two price vectors $\mathbf{q}\geq\mathbf{q}^{\prime}$ and $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$, and construct a $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q^{\prime}},b)$ such that $y^{\prime}_{j}\leq y_{j}$ whenever $q^{\prime}_{j}=q_{j}$; we have a second, analogous construction for budget changes.
+The proof is a finite algorithm that transforms the optimal generalized flow for demands $\mathbf{y}$ to $\mathbf{y}^{\prime}$ such that we throughout maintain the primal-dual slackness condition corresponding to the Gale demand correspondence.
 The algorithm builds on the literature on maximum generalized flow algorithms.
@@
 (1991), Olver and Végh (2020), Végh (2014).
-4 Efficiency of competitive equilibria
-We now show the ‘price of anarchy’ result Theorem 1.1.
-Theorem 4.1 (Restatement of Theorem 1.1).
-Given a Fisher market instance, consider any competitive equilibrium $(\mathbf{x},\mathbf{p})$ and a Nash welfare maximizing allocation $\mathbf{y}$. Then,
-$\left(\prod_{i\in{\mathcal{A}}}u_{i}(x_{i})^{b_{i}}\right)^{\frac{1}{\sum_{i\in{\mathcal{A}}}b_{i}}}\geq\left(\frac{1}{\mathrm{e}}\right)^{\frac{1}{\mathrm{e}}}\left(\prod_{i\in{\mathcal{A}}}u_{i}(y_{i})^{b_{i}}\right)^{\frac{1}{\sum_{i\in{\mathcal{A}}}b_{i}}}.$
-Proof.
-Without loss of generality, we can assume $\sum_{i\in{\mathcal{A}}}b_{i}=1$.
-Note that since $(\mathbf{x},\mathbf{p})$ is a competitive equilibrium, then $u_{i}(x_{i})\geq\min\left\{1,\frac{b_{i}}{\left\langle\mathbf{p},y_{i}\right\rangle}\right\}u_{i}(y_{i})$, as agent $i$ can use their budget, $b_{i}$, to buy at least $\frac{b_{i}}{\left\langle\mathbf{p},y_{i}\right\rangle}$ fraction of $y_{i}$.
-The lower bound holds because if $b_{i}<\left\langle\mathbf{p},y_{i}\right\rangle$, then $u_{i}\left(\frac{b_{i}}{\left\langle\mathbf{p},y_{i}\right\rangle}y_{i}\right)\geq\frac{b_{i}}{\left\langle\mathbf{p},y_{i}\right\rangle}u_{i}(y_{i})$ using concavity and $u_{i}(\mathbf{0})=0$.
-Thus,
-$\sum_{i\in{\mathcal{A}}}{b_{i}}\log\frac{u_{i}(y_{i})}{u_{i}(x_{i})}\leq\sum_{i\in{\mathcal{A}}}{b_{i}}\log\max\left\{1,\frac{\left\langle\mathbf{p},y_{i}\right\rangle}{b_{i}}\right\}.$
-Let $b^{\prime}_{i}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}{\left\langle\mathbf{p},y_{i}\right\rangle}$, and let ${\mathcal{A}}^{\prime}$ be the set of agents such that $b^{\prime}_{i}\geq b_{i}$, $B\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{i\in\mathcal{A}^{\prime}}b_{i}$, and $B^{\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{i\in\mathcal{A}^{\prime}}b^{\prime}_{i}$.
-Therefore, using the log sum inequality,
-$\sum_{i\in{\mathcal{A}}}{b_{i}}\log\frac{u_{i}(y_{i})}{u_{i}(x_{i})}\leq\sum_{i\in{\mathcal{A}}^{\prime}}{b_{i}}\log\frac{b^{\prime}_{i}}{b_{i}}\leq B\log\frac{B^{\prime}}{B}\,.$
-By definition, $B\leq\sum_{i\in{\mathcal{A}}}b_{i}=1$.
-We also note that $B^{\prime}\leq\sum_{i\in{\mathcal{A}}}b_{i}^{\prime}=\sum_{i\in{\mathcal{A}}}\left\langle\mathbf{p},y_{i}\right\rangle\leq\sum_{j\in{\mathcal{M}}}p_{j}\leq\sum_{i\in{\mathcal{A}}}b_{i}=1$.
-$\sum_{i\in{\mathcal{A}}}{b_{i}}\log\frac{u_{i}(y_{i})}{u_{i}(x_{i})}\leq\max_{B\leq 1}B\log\frac{1}{B}\leq\frac{1}{\mathrm{e}}\,.$
-The statement follows. ∎
-5 Properties of the Gale demand correspondence
-In this section, we show some properties of the Gale demand correspondence, and also derive Theorem 1.2.
-Lemma 5.1.
-Given any utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$, $b>0$ and $\mathbf{q}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, for any $y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $\left\langle\mathbf{q},y\right\rangle>b$, there exists $y^{\prime}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $\left\langle\mathbf{q},y^{\prime}\right\rangle\leq b$ and $b\log u(y^{\prime})-\left\langle\mathbf{q},y^{\prime}\right\rangle>b\log u(y)-\left\langle\mathbf{q},y\right\rangle$.
-In particular, if $y\in\mathcal{GD}^{u}(\mathbf{q},b)$, then, $\left\langle\mathbf{q},y\right\rangle\leq b$.
-Proof.
-Let $y^{\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}y\cdot\frac{b}{\left\langle\mathbf{q},y\right\rangle}$.
-By concavity and $u(\mathbf{0})=0$, we get $u(y^{\prime})\geq\frac{b}{\left\langle\mathbf{q},y\right\rangle}u(y)$.
-Thus,
-$b\log u(y^{\prime})-\left\langle\mathbf{q},y^{\prime}\right\rangle$ $\geq b\log u(y)-b\log\frac{\left\langle\mathbf{q},y\right\rangle}{b}-b>b\log u(y)-\left\langle\mathbf{q},y\right\rangle\,,$
-where the second inequality used $\log(\alpha)+1<\alpha$ for $\alpha>1$. ∎
-Next, we derive some technical properties of the price domain $\operatorname{pdom}(u)$.
-Lemma 5.2.
-For any utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ and $b>0$, the following properties of the price domain $\operatorname{pdom}(u)$ hold:
-  1. (i)
-    for any $\mathbf{q}\in\operatorname{pdom}(u)$, $\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{q},y\right\rangle<+\infty$;
-  2. (ii)
-    $\{\mathbf{q}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{q}>0\}\subseteq\operatorname{pdom}(u)$;
-  3. (iii)
-    for any $\mathbf{q}\geq\mathbf{q}^{\prime}\geq 0$ such that $\{j:q^{\prime}_{j}=0\}=\{j:q_{j}=0\}$, $\mathbf{q}\in\operatorname{pdom}(u)$ implies $\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$;
-  4. (iv)
-    for any $\mathbf{p}\geq 0$, if $\max_{\left\langle\mathbf{p},y\right\rangle\leq b}\log u(y)<+\infty$, then $\mathbf{p}\in\operatorname{pdom}(u)$.
-  5. (v)
-    for any $M>0$, if there is a convergent sequence of prices $\mathbf{p}^{(k)}\in\operatorname{pdom}(u)$ with $\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{p},y\right\rangle<M$, $k=1,2,\ldots$ with a limit point $\mathbf{p}^{(k)}\to\bar{\mathbf{p}}$, then $\bar{\mathbf{p}}\in\operatorname{pdom}(u)$.
-Proof.
-We first prove property (i).
-According to the definition of $\operatorname{pdom}(u)$, there exists $b^{\prime}>0$ such that $\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b^{\prime}\log u(y)-\left\langle\mathbf{q},y^{\prime}\right\rangle<+\infty$.
-For any $b>0$ and for any $y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, if $\left\langle\mathbf{q},y\right\rangle\leq b$, then $b\log u(y)-\left\langle\mathbf{q},y\right\rangle<\frac{b}{b^{\prime}}\left(b^{\prime}\log u(y)-\left\langle\mathbf{q},y\right\rangle\right)+\frac{b-b^{\prime}}{b^{\prime}}\ \left\langle\mathbf{q},y\right\rangle$, which is bounded.
-Then, the result follows using Lemma 5.1.
-Property (ii) holds as, by Lemma 5.1, $\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{q},y\right\rangle\leq\max_{\left\langle\mathbf{q},x\right\rangle\leq b}b\log u(x)-\left\langle\mathbf{q},x\right\rangle$, which is bounded as $x$ is bounded: $x_{j}\leq\frac{b}{q_{j}}<\infty$.
-To prove property (iii), we let $\alpha=\min_{q_{j}>0}\{q^{\prime}_{j}/q_{j}\}$.
-Then, $\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}(\alpha b_{i})\log u_{i}(y_{i})-\left\langle\mathbf{q}^{\prime},y_{i}\right\rangle\leq\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}(\alpha b_{i})\log u_{i}(y_{i})-\alpha\left\langle\mathbf{q},y_{i}\right\rangle<+\infty$.
-Property (iv) holds as, for any $y$ such that $\left\langle\mathbf{p},y\right\rangle\leq b$, $b\log u(y)-\left\langle\mathbf{q},y\right\rangle\leq b\log u(y)<+\infty$.
-Finally, let us prove (v).
-For a contradiction, assume $\bar{\mathbf{p}}\notin\operatorname{pdom}(u)$, that is, $\sup_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\bar{\mathbf{p}},y\right\rangle=\infty$.
-Consider a vector $y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ with $b\log u(y)-\left\langle\bar{\mathbf{p}},y\right\rangle>2M$.
-Clearly, $y\neq 0$.
-Let us select $k$ such that $\|\mathbf{p}^{(k)}-\bar{\mathbf{p}}\|<M/\|y\|$.
-Then, $b\log u(y)-\left\langle\mathbf{p}^{(k)},y\right\rangle=b\log u(y)-\left\langle\bar{\mathbf{p}},y\right\rangle+\left\langle\bar{\mathbf{p}}-\mathbf{p}^{(k)},y\right\rangle>M$, in contradiction with the assumption on the sequence $\mathbf{p}^{(k)}$.
-∎
-We next prove Lemma 2.4. See 2.4
-Proof.
-Let $M\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sup_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{p},y\right\rangle$, which implies the existence of a sequence of allocation $\{y^{(k)}\}_{k=1}^{\infty}$ such that $b\log u(y^{(k)})-\left\langle\mathbf{p},y^{(k)}\right\rangle$ converges to $M$.
-We show the existence of an allocation $y^{*}$ satisfying $b\log u(y^{*})-\left\langle\mathbf{p},y^{*}\right\rangle=M$.
-According to Lemma 5.1, there exists a sequence of $\bar{y}^{(k)}$ such that $b\log u(\bar{y}^{(k)})-\left\langle\mathbf{p},\bar{y}^{(k)}\right\rangle$ converges to $M$, with $0\leq\left\langle\mathbf{p},\bar{y}^{(k)}\right\rangle\leq b$ for all $k$.
-Thus, we can select a subsequence $\bar{y}^{(i_{k})}$ such that $b\log u(\bar{y}^{(i_{k})})-\left\langle\mathbf{p},\bar{y}^{(i_{k})}\right\rangle$ converges to $M$ and $\left\langle\mathbf{p},\bar{y}^{(i_{k})}\right\rangle$ converges to $b^{\prime}$ for some $b^{\prime}\in[0,b]$.
-Consequently, $b\log u(x^{(i_{k})})$ converges to $M+b^{\prime}$ as $\left\langle\mathbf{p},x^{(i_{k})}\right\rangle$ converges to $b^{\prime}$.
-By Assumption 1, there exists an allocation $y^{*}$ such that $b\log u(y^{*})=M+b^{\prime}$ and $\left\langle\mathbf{p},y^{*}\right\rangle\leq b^{\prime}$.
-This completes the proof.
-∎
-We can use the subgradient Karush–Kuhn–Tucker (KKT) conditions to characterize solutions in the Gale demand correspondence.
-We refer the reader to Lagrangian duality and KKT conditions of nondifferentiable functions to Chapter 3 in Ruszczyński (2011).
-Lemma 5.3.
-For a utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$, $b>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$, let $y\in\mathcal{GD}^{u}(\mathbf{q},b)$.
-Then, there exists a supergradient $\mathbf{g}\in\partial^{\wedge}u(y)$ such that
-$g_{j}\leq{q_{j}}\cdot\frac{u(y)}{b}\,,\quad\forall j\in{\mathcal{M}}\,,\,\mbox{with equality if }y_{j}>0\,,$
-Conversely, if such a $\mathbf{g}\in\partial^{\wedge}u(y)$ exists then $y\in\mathcal{GD}^{u}(\mathbf{q},b)$.
-We are ready to prove Lemma 3.1
-See 3.1
-Proof.
-By Lemma 5.3, there exists a supergradient $\mathbf{g}\in\partial^{\wedge}u(y)$ such that $g_{j}\leq q_{j}u(y)/b$ with equality whenever $y_{i}>0$. Therefore,
-$u(x)$ $\leq u(y)+\left\langle\mathbf{g},x-y\right\rangle\leq u(y)+\frac{u(y)}{b}\left\langle\mathbf{q},x-y\right\rangle$
-$\leq u(y)\left(1+\frac{\left\langle\mathbf{q},x\right\rangle}{b}\right)\leq u(y)\left(1+\frac{\left\langle\mathbf{p},x\right\rangle}{b}\right)\leq 2u(y)\,.$
-5.1 Nash welfare as approximate competitive equilibrium
-We now derive Theorem 1.2, showing that Nash welfare maximizing allocations can be seen as approximate competitive equilibria.
-Theorem 5.4 (Restatement of Theorem 1.2).
-Consider a Fisher market instance, and let $\mathbf{y}$ be an allocation maximizing Nash welfare along with Gale prices $\mathbf{q}$.
-Then, $(\mathbf{y},\mathbf{q})$ forms a $2$-demand-approximate competitive equilibrium.
-Additionally, when all $b_{i}$’s are identical, the allocation is $2$-envy-free.
-Proof.
-By Lemma 5.1, all agents satisfy the budget constraints.
-Additionally, the market clears by the KKT conditions for (EG).
-For the approximate competitive equilibrium, we only need to show $u_{i}(y_{i})\geq\frac{1}{2}\max_{\left\langle\mathbf{q},x_{i}\right\rangle\leq b_{i}}u_{i}(x_{i})$ for all agents.
-Note that $y_{i}$ maximizes $b_{i}\log u_{i}(y_{i})-\left\langle\mathbf{q},y_{i}\right\rangle$.
-By Lemma 5.3, there is a supergradient $g_{i}$ such that $g_{ij}\leq q_{j}u_{i}(y_{i})/b_{i}$ with equality whenever $y_{ij}>0$.
-Then, for any $x_{i}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $\left\langle\mathbf{q},x_{i}\right\rangle\leq b_{i}$,
-$u_{i}(x_{i})\leq u_{i}(y_{i})+\left\langle g_{i},x_{i}-y_{i}\right\rangle\leq u_{i}(y_{i})+\frac{u_{i}(y_{i})}{b_{i}}\left\langle\mathbf{q},x_{i}-y_{i}\right\rangle\leq\left(2-\frac{\left\langle\mathbf{q},y_{i}\right\rangle}{b_{i}}\right)u_{i}(y_{i})\leq 2u_{i}(y_{i})\,.$
-This proves that $(\mathbf{y},\mathbf{q})$ forms a $2$-approximate competitive equilibrium. Additionally, allocation is $2$-envy-free follows directly from Lemma 2.10. ∎
-6 Efficiency of Nash welfare for Gale-substitutes
-In this section, we prove Theorem 1.3.
-We show that for $\Sigma$-Gale subsitutes utility functions, given a competitive equilibrium $(\mathbf{x},\mathbf{p})$, there are Gale prices $\mathbf{q}^{\star}$ such that $\mathbf{q}^{\star}\leq\mathbf{p}$.
-The statement then follows by Lemma 3.1.
-We start by re-stating and proving Lemma 3.2.
-See 3.2
-Proof.
-According to the KKT conditions, there exists $\lambda\geq 0$ and supergradient $\mathbf{g}\in\partial^{\wedge}u(x)$ such that
-$g_{j}\leq\lambda p_{j}\,,\quad\forall j\in{\mathcal{M}}\,,\,\mbox{with equality if }x_{j}>0\,,$ (6)
-and $\left\langle\mathbf{p},x\right\rangle=b$ if $\lambda>0$.
-Let us first consider the case $\lambda=0$.
-Then, $g_{j}\leq 0$ with equality if $x_{j}>0$, implying that $x=\arg\max_{x\in\mathbb{R}^{\mathcal{M}}}u(x)$.
-Consequently, $u(.)$ is satiable; therefore, $\mathbf{0}\in\operatorname{pdom}(u)$ and $x\in\mathcal{GD}^{u}(\mathbf{0},b)$.
-Now, ($\Sigma$) guarantees the existence of $y\in\mathcal{GD}^{u}(\mathbf{p},b)$, $y\leq x$.
-Clearly, $\operatorname{pdom}(u)=\mathbb{R}_{\geq 0}^{\mathcal{M}}$ in this case.
-Next, assume $\lambda>0$. Let us define
-$\mathbf{q}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\frac{\lambda b}{u(x)}\cdot\mathbf{p}\,.$
-Note that $\mathbf{q}\leq\mathbf{p}$.
-This follows since $\lambda b=\lambda\left\langle\mathbf{p},x\right\rangle=\left\langle\mathbf{g},x\right\rangle\leq u(x)$, where the final inequality is by concavity and $u(\mathbf{0})=0$.
-Thus,
-$g_{j}\leq\frac{u(x)}{b}q_{j}\,,\quad\forall j\in{\mathcal{M}}\,,\,\mbox{with equality if }x_{j}>0\,.$
-According to Lemma 5.3, $x\in\mathcal{GD}^{u}(\mathbf{q},b)$.
-Note that $\mathcal{GD}^{u}(\mathbf{q},b)=\mathcal{GD}^{u}(\mathbf{p},u(x)/\lambda)$.
-Utilizing Definition 2.14(ii) for $x\in\mathcal{GD}^{u}(\mathbf{p},u(x)/\lambda)$ and $b\leq u(x)/\lambda$, there exists $y\in\mathcal{GD}^{u}(\mathbf{p},b)$ such that $y\leq x$ as required.
-∎
-We next show that for any competitive equilibrium $(\mathbf{x},\mathbf{p})$, and any prices $\mathbf{q}\leq\mathbf{p}$, there exist an aggregate Gale demand such that $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1$ whenever $q_{j}=p_{j}$.
-Lemma 6.1.
-Consider a Fisher market instance such that $u_{i}(\cdot)$ is $\Sigma$-Gale-substitute utilities for all agents.
-For any competitive equilibrium $(\mathbf{x},\mathbf{p})$, for any $\mathbf{q}\in\cap_{i\in{\mathcal{A}}}\operatorname{pdom}(u_{i})$ such that $\mathbf{q}\leq\mathbf{p}$, there exists $y_{i}\in\mathcal{GD}_{i}(\mathbf{q},b_{i})$ for all $i\in{\mathcal{A}}$ and $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1$ for all $j\in{\mathcal{M}}$ such that $q_{j}=p_{j}$.
-Proof.
-By Lemma 3.2, there exists $\bar{y}_{i}\in\mathcal{GD}_{i}(\mathbf{p},b_{i})$ for every $i\in{\mathcal{A}}$ such that $\bar{y}_{i}\leq x_{i}$.
-Using Definition 2.14(i), there exists $y_{i}\in\mathcal{GD}_{i}(\mathbf{q},b_{i})$ and $y_{ij}\leq\bar{y}_{ij}$ if $q_{j}=p_{j}$.
-Thus, if $q_{j}=p_{j}$ then $\sum_{i\in{\mathcal{A}}}y_{ij}\leq\sum_{i\in{\mathcal{A}}}\bar{y}_{ij}\leq\sum_{i\in{\mathcal{A}}}x_{ij}\leq 1$.
-∎
-Recall the Lagrangian dual function $\varphi$ defined in (4) on the domain $\operatorname{dom}(\varphi)=\cap_{i\in{\mathcal{A}}}\operatorname{pdom}(u_{i})$.
-$\varphi(\mathbf{q})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{i\in{\mathcal{A}}}\max_{y_{i}\in\mathbb{R}_{\geq 0}^{{\mathcal{M}}}}\left(b_{i}\log u_{i}(y_{i})-\left\langle\mathbf{q},y_{i}\right\rangle\right)+\sum_{j\in{\mathcal{M}}}q_{j}.$
-The following properties are easy to verify.
-Lemma 6.2.
-The following hold for the function $\varphi(\cdot)$.
-  1. (i)
-    $\varphi(\mathbf{q})$ is convex.
-  2. (ii)
-    Given $\mathbf{q}\in\operatorname{dom}(\varphi)$, let $y_{i}\in\mathcal{GD}_{i}(\mathbf{q},b_{i})$ for all $i\in{\mathcal{A}}$.
-Then, $\mathbf{1}-\sum_{i\in{\mathcal{A}}}y_{i}\in\partial^{\vee}\varphi(\mathbf{q})$.
-  3. (iii)
-    The optimum value of (5), that is, the infimum of $\varphi(\mathbf{q})$ is taken in $\operatorname{dom}(\varphi)$.
-Proof of Theorem 1.3.
-Let $(\mathbf{x},\mathbf{p})$ be a competitive equilibrium solution. Consider the dual (5) of the Eisenberg-Gale program (EG), i.e.,
-$\min_{\mathbf{q}\in\operatorname{dom}(\varphi)}\varphi(\mathbf{q})\,.$
-We show that this has an optimal solution with $\mathbf{q}\leq\mathbf{p}$.
-The theorem then follows by Lemma 3.1.
-Let us take any optimal $q^{\star}\in\operatorname{dom}(\varphi)$; we are done if $\mathbf{q}^{\star}\leq\mathbf{p}$.
-Otherwise, let $J\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\{j\in{\mathcal{M}}\,:\,\mathbf{q}^{\star}_{j}>p_{j}\}$, and let us define $\bar{\mathbf{q}}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ as
-$\bar{q}_{j}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\begin{cases}p_{j}\,,&\mbox{if }j\in J\,,\\ q^{\star}_{j}\,,&\mbox{if }j\notin J\,.\\ \end{cases}$
-First, assuming $\bar{\mathbf{q}}\in\operatorname{dom}(\varphi)$, we show that $\bar{\mathbf{q}}$ is also optimal to (5) as claimed.
-Let us pick an allocation $\mathbf{y}$ for the prices $\bar{\mathbf{q}}$ as guaranteed by Lemma 6.1: $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1$ for all $j\in{\mathcal{M}}$ with $\bar{q}_{j}=p_{j}$; in particular, for all $j\in J$.
-According to Lemma 6.2(ii), we get a subgradient $\bar{\mathbf{g}}\in\partial^{\vee}\varphi(\bar{\mathbf{q}})$ defined as $\bar{\mathbf{g}}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\mathbf{1}-\sum_{i\in{\mathcal{A}}}y_{i}$; this has the property $\bar{g}_{j}\geq 0$ for all $j\in J$.
-We then have
-$\varphi(\mathbf{q}^{\star})\geq\varphi(\bar{\mathbf{q}})+\left\langle\bar{\mathbf{g}},\mathbf{q}^{\star}-\bar{\mathbf{q}}\right\rangle=\varphi(\bar{\mathbf{q}})+\sum_{j\in J}\bar{g}_{j}(q^{\star}_{j}-\bar{q}_{j})\geq\varphi(\bar{\mathbf{q}})\,.$
-Here, the first inequality is by the definition of subgradients; the equality uses that $q^{\star}_{j}=\bar{q}_{j}$ if $j\notin J$, and the final inequality follows since $\bar{g}_{j}\geq 0$ and $q^{\star}_{j}>\bar{q}_{j}$ for $j\in J$.
-This shows that $\bar{\mathbf{q}}$ is also an optimal solution to (5), completing the proof assuming $\bar{\mathbf{q}}\in\operatorname{dom}(\varphi)$.
-It is left to show that $\bar{\mathbf{q}}\in\operatorname{dom}(\varphi)$.
-Let $K\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\{j\in{\mathcal{M}}\,:\,\mathbf{q}^{\star}_{j}=0<p_{j}\}$.
-Clearly, $K\cap J=\emptyset$.
-If $K=\emptyset$, then Lemma 5.2(iii) already implies $\bar{\mathbf{q}}\in\operatorname{pdom}(u)$, since $\mathbf{p}\geq\bar{\mathbf{q}}$ and $\operatorname{supp}(\bar{\mathbf{q}})=\operatorname{supp}(\mathbf{p})$.
-Otherwise, for $k=1,2,\ldots$, let us define $\mathbf{q}^{(k)}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ as
-$q_{j}^{(k)}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\begin{cases}p_{j}\,,&\mbox{if }j\in J\,,\\ \frac{p_{j}}{k}\,,&\mbox{if }j\in K\,,\\ q^{\star}_{j}\,,&\mbox{if }j\notin J\cup K\,.\\ \end{cases}$
-We have $\mathbf{q}^{(k)}\in\operatorname{pdom}(u)$ again by Lemma 5.2(iii).
-As in the first part, we take a subgradient $\mathbf{g}^{(k)}$ as guaranteed by Lemma 6.2(ii), with the properties that $\bar{g}_{j}\geq 0$ for all $j\in J$ and also $\mathbf{g}^{(k)}\leq\mathbf{1}$.
-We can write
-$\varphi(\mathbf{q}^{\star})\geq\varphi(\mathbf{q}^{(k)})+\left\langle\mathbf{g}^{(k)},\mathbf{q}^{\star}-\mathbf{q}^{(k)}\right\rangle=\varphi(\mathbf{q}^{(k)})+\sum_{j\in J}\bar{g}_{j}(q^{\star}_{j}-q^{(k)}_{j})-\sum_{j\in K}\bar{g}_{j}q^{(k)}_{j}\geq\varphi(\mathbf{q}^{(k)})-\frac{\|p\|_{\infty}}{k}\,.$
-Here, we use in the first equality that $q^{\star}_{j}=q^{(k)}_{j}$ for $j\notin J\cup K$, and $q^{\star}_{j}=0$ for $j\in K$.
-The last inequality uses $\bar{g}_{j}\leq 1$ and the definition $q^{(k)}_{j}=p_{j}/(k+1)$.
-Thus, Lemma 5.2(v) is applicable with $M=\varphi(\mathbf{q}^{\star})+\|p\|_{\infty}$, using that for each $j\in{\mathcal{M}}$, $\max_{y_{i}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b_{i}\log u_{i}(y_{i})-\left\langle\mathbf{q}^{(k)},y_{i}\right\rangle\leq\varphi(\mathbf{q}^{(k)})$.
-This shows that $\bar{\mathbf{q}}\in\operatorname{pdom}(u_{i})$ for all $i\in{\mathcal{M}}$, and therefore $\bar{\mathbf{q}}\in\operatorname{dom}(\varphi)=\cap_{i\in{\mathcal{M}}}\operatorname{pdom}(u_{i})$.
-∎
-7 Separable utilities are $\Sigma$-Gale-substitutes
-In this section, we prove Theorem 1.4. See 1.4 Recall the following simple monotonicity property of supergradients of univariate concave functions.
-Lemma 7.1.
-Let $v\,:\,\mathbb{R}\to\mathbb{R}$ be a univariate concave function, $\alpha,\alpha^{\prime}\in\operatorname{dom}(f)$, and $g\in\partial^{\wedge}v(\alpha)$, $g^{\prime}\in\partial^{\wedge}v(\alpha^{\prime})$.
-Then, $\alpha<\alpha^{\prime}$ implies $g\geq g^{\prime}$.
-Conversely, if $g>g^{\prime}$, then $\alpha\leq\alpha^{\prime}$.
-We now adapt Lemma 5.3 to separable utilities.
-Lemma 7.2.
-Let $u(x)=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$ be a separable utility function.
-For any $b>0$, $\mathbf{q}\in\operatorname{pdom}(u)$, $y\in\mathcal{GD}^{u}(\mathbf{q},b)$, we have $u(y)>0$.
-Further, $y\in\mathcal{GD}^{u}(b,\mathbf{q})$ if and only if there exist supergradients $g_{j}\in\partial^{\wedge}v_{j}(y_{j})$, $j\in{\mathcal{M}}$ satisfying
-  1. (i)
-    $g_{j}\geq 0$; and
-  2. (ii)
-    $g_{j}\leq q_{j}u(y)/b$ with equality whenever $y_{j}>0$.
-The proof of Theorem 1.4 follows by the next three lemmas, that prove the three properties in Definition 2.14 for separable utilities: property (i) in Lemma 7.3, property (ii) in Lemma 7.4, and property ($\Sigma$) in Lemma 7.5.
-A similar argument was used in Lemma 7.1 in Garg et al.
-(2022a) to show a variant of the Gale substitutes property for capped SPLC utilities.
-Lemma 7.3.
-Let $u(x)=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$ be a separable utility function, and let $b>0$, $\mathbf{q},\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$ and $y\in\mathcal{GD}^{u}(\mathbf{q},b)$.
-If $\mathbf{q}\geq\mathbf{q}^{\prime}$, then there exists a $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ such that $y^{\prime}_{j}\leq y_{j}$ whenever $q_{j}^{\prime}=q_{j}$.
-Proof.
-Let $J\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\{j\in{\mathcal{M}}\,:\,q^{\prime}_{j}=q_{j}\}$.
-Let us select $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ minimizing the potential $\Psi(y^{\prime})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{j\in J}(y^{\prime}_{j}-y_{j})^{+}$.
-Our goal is to show $\Psi(y^{\prime})=0$.
-Throughout the proof, let $g^{\prime}_{j}\in\partial^{\wedge}v_{j}(y^{\prime}_{j})$ and $g_{j}\in\partial^{\wedge}v_{j}(y_{j})$ as in Lemma 7.2.
-First, we show $u(y^{\prime})\geq u(y)$. For a contradiction, assume $u(y^{\prime})<u(y)$. Then, according to Lemma 7.2, for every $j\in{\mathcal{M}}$ with $y_{j}>0$ and $q_{j}>0$ we have
-$g^{\prime}_{j}\leq q^{\prime}_{j}u(y^{\prime})/b<q_{j}u(y)/b=g_{j}\,,$
-while for $j\in{\mathcal{M}}$ with $q_{j}=0$ we have $g^{\prime}_{j}=g_{j}=0$.
-Consequently, for each $y_{j}>0$, if $q_{j}>0$ then Lemma 7.1 implies $y^{\prime}_{j}\geq y_{j}$, and if $q_{j}=0$ then $v_{j}(y_{j})=v_{j}(y_{j}^{\prime})$, both being equal to the maximum value of $v_{j}(.)$.
-These imply $u(y^{\prime})\geq u(y)$, a contradiction.
-Using the above, for any $j\in J$ where $y^{\prime}_{j}>0$,
-$g^{\prime}_{j}=\frac{q^{\prime}_{j}}{b}u(y^{\prime})\geq\frac{q_{j}}{b}u(y)\geq g_{j}\,.$ (7)
-Let us first consider the case $u(y^{\prime})>u(y)$.
-For all $j\in J$ with $q^{\prime}_{j}=q_{j}>0$, we get $g^{\prime}_{j}>g_{j}$.
-By Lemma 7.1, $y^{\prime}_{j}\leq y_{j}$ follows for such goods $j$.
-On the other hand, if $q^{\prime}_{j}=q_{j}=0$, then this implies $g^{\prime}_{j}=g_{j}=0$.
-For such indices, if $y^{\prime}_{j}>y_{j}$, then replacing $y^{\prime}_{j}$ by $y_{j}$ results in another bundle $y^{\prime\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ with $\Psi(y^{\prime\prime})<\Psi(y^{\prime})$, a contradiction.
-For the rest of the proof, let us assume that $u(y^{\prime})=u(y)$.
-We next show $y^{\prime}_{j}\geq y_{j}$ for all $j\in{\mathcal{M}}\setminus J$, that is, for all goods with $q^{\prime}_{j}<q_{j}$.
-This is clearly true if $y_{j}=0$.
-Whenever $y_{j}>0$ for such an index, we have
-$g^{\prime}_{j}\leq\frac{q^{\prime}_{j}}{b}u(y^{\prime})<\frac{q_{j}}{b}u(y)=g_{j}\,,$
-and therefore $y^{\prime}_{j}\geq y_{j}$ by Lemma 7.1.
-By monotonicity, $v_{j}(y^{\prime}_{j})\geq v_{j}(y_{j})$ for all $j\in{\mathcal{M}}\setminus J$.
-Since $u(y^{\prime})=u(y)$, we have $\sum_{j\in J}v_{j}(y^{\prime}_{j})\leq\sum_{j\in J}v_{j}(y_{j})$.
-From (7), it follows that $g^{\prime}_{j}=g_{j}$ for all $j\in J$ with $y^{\prime}_{j}>0$.
-This is because the first inequality is now equality, and the second inequality may only be strict if $y_{j}=0<y^{\prime}_{j}$.
-By Lemma 7.1, $g_{j}\geq g^{\prime}_{j}$ in this case, leading to a contradiction.
-Assume for a contradiction $\Psi(y^{\prime})>0$, and let us select an index $k\in J$ with $y^{\prime}_{k}>y_{k}$.
-If $v_{k}(y^{\prime}_{k})=v_{k}(y_{k})$, then we get a contradiction, because changing $y^{\prime}_{k}$ to $y_{k}$ would preserve the the optimality conditions in Lemma 7.2, while decreasing the potential $\Psi(y^{\prime})$.
-Thus, $v_{k}(y^{\prime}_{k})>v_{k}(y_{k})$.
-By $\sum_{j\in J}v_{j}(y^{\prime}_{j})\leq\sum_{j\in J}v_{j}(y_{j})$, there must be an index $j\in J$ with $v_{j}(y^{\prime}_{j})<v_{j}(y_{j})$.
-We can now find suitable $\varepsilon_{k},\varepsilon_{j}>0$ such that $v_{k}(y^{\prime}_{k}-\varepsilon_{k})+v_{j}(y^{\prime}_{j}+\varepsilon_{j})=v_{k}(y^{\prime}_{k})+v_{j}(y^{\prime}_{j})$, and $y^{\prime}_{j}+\varepsilon_{j}\leq y_{j}$.
-This modified solution $y^{\prime\prime}$ satisfies the optimality conditions but $\Psi(y^{\prime\prime})<\Psi(y^{\prime})$, again leading to a contradiction.
-∎
-Lemma 7.4.
-Let $u(x)=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$ be a separable utility function, $b>b^{\prime}>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$.
-Then, for any $y\in\mathcal{GD}^{u}(\mathbf{q},b)$, there exists a $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$ such that $y^{\prime}\leq y$.
-Proof.
-We argue similarly as in the previous proof: we pick $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$ which minimizes $\Psi(y^{\prime})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{j\in J}(y^{\prime}_{j}-y_{j})^{+}$, and show that $\Psi(y^{\prime})=0$.
-First, we show that ${u(y)}/{b}\leq{u(y^{\prime})}/{b^{\prime}}$. For a contradiction, suppose ${u(y)}/{b}>{u(y^{\prime})}/{b^{\prime}}$. If $y_{j}>0$, then, by Lemma 7.2
-$g_{j}=\frac{q_{j}u(y)}{b}>\frac{q_{j}u(y^{\prime})}{b^{\prime}}\geq g^{\prime}_{j},$
-which implies $v_{j}(y_{j})\leq v_{j}(y^{\prime}_{j})$ whenever $y_{j}>0$ (Lemma 7.1), and therefore, $u(y)\leq u(y^{\prime})$.
-Together with $b>b^{\prime}$, this implies ${u(y)}/{b}\leq{u(y^{\prime})}/{b^{\prime}}$, a contradiction.
-Whenever $y^{\prime}_{j}>0$, we now have
-$g^{\prime}_{j}=\frac{q_{j}u(y^{\prime})}{b^{\prime}}\geq\frac{q_{j}u(y)}{b}\geq g_{j}\,.$ (8)
-Assume first $u(y^{\prime})/b^{\prime}>u(y)/b$.
-If $q_{j}>0$, then this inequality is strict, and therefore $y^{\prime}_{j}\leq y_{j}$ for all $j\in{\mathcal{M}}$ follows by Lemma 7.1.
-If $q_{j}=0$, then $g^{\prime}_{j}=g_{j}=0$.
-If $y^{\prime}_{j}>y_{j}$ for such a good $j$, we can decrease $y^{\prime}_{j}$ to $y_{j}$ and thereby get an optimal allocation with strictly smaller $\Psi(y^{\prime})$ value, leading to a contradiction.
-This completes the proof for $u(y^{\prime})/b^{\prime}>u(y)/b$.
-For the rest, we may assume $u(y^{\prime})/b^{\prime}=u(y)/b$.
-Now, $g^{\prime}_{j}=g_{j}$ must hold whenever $y^{\prime}_{j}>0$.
-This is because the first inequality in (8) is equality, and if the second is strict, then $y_{j}=0<y^{\prime}_{j}$ would give a contradiction to $g^{\prime}_{j}>g_{j}$.
-Assume for a contradiction $\Psi(y^{\prime})>0$, and let us pick $k\in{\mathcal{M}}$ with $y^{\prime}_{k}>y_{k}$.
-The equality $u(y^{\prime})/b^{\prime}=u(y)/b$ and $b^{\prime}<b$ gives $u(y^{\prime})<u(y)$.
-We can derive a contradiction by decreasing $y^{\prime}_{k}$ and potentially increasing an $y^{\prime}_{j}$ similarly as in the proof of Lemma 7.3.
-∎
-Lemma 7.5.
-Let $u(x)=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$ be a separable utility function such that $\mathbf{0}\in\operatorname{pdom}(u)$.
-Then, for any $\mathbf{q}\geq\mathbf{0}$, $b>0$ and any $y\in\mathcal{GD}^{u}(\mathbf{0},b)$, there exists a $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b)$ such that $y^{\prime}\leq y$.
-Proof.
-$\mathbf{0}\in\operatorname{pdom}(u)$ implies $\sup_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}u(y)<\infty$.
-Therefore, $\operatorname{pdom}(u)=\mathbb{R}_{\geq 0}^{\mathcal{M}}$; in particular, $\mathbf{q}\in\operatorname{pdom}(u)$.
-Let us pick $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b)$ such that $\sum_{j\in{\mathcal{M}}}y^{\prime}_{j}$ is minimal.
-Note that $u(y)=\sup_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}u(y)$.
-In particular, for any $\alpha\geq y_{j}$, $v_{j}(\alpha)=v_{j}(y_{j})$.
-For a contradiction, assume there exists a $k\in{\mathcal{M}}$ with $y_{k}<y^{\prime}_{k}$.
-Then, reducing $y^{\prime}_{k}$ to $y_{k}$ may only increase $b\log u(y^{\prime})-\left\langle\mathbf{q},y^{\prime}\right\rangle$, while decreasing $\sum_{j\in{\mathcal{M}}}y^{\prime}_{j}$.
-This contradicts the choice of $y^{\prime}$.
-∎
-8 Generalized network utilities
-In this section, we prove Theorem 1.5.
-We will first need a characterization of subgradients by linear programming duality.
-We state this in a more general form.
-The proof based on LP duality is deferred to Appendix B.
-Lemma 8.1.
-Let $\mathbf{A}\in\mathbb{R}^{k\times\ell}$, $\mathbf{B}\in\mathbb{R}^{k\times m}$, $c\in\mathbb{R}^{\ell}$, $b\in\mathbb{R}^{k}$.
-Let $\mathcal{D}\subseteq\mathbb{R}^{m}$ be the set of vectors $x\in\mathbb{R}^{m}$ such that the following parametric linear program is feasible and bounded, and let us define $\varphi:\mathcal{D}\to\mathbb{R}$ such that $\varphi(x)$ is the optimum value of this program:
-$\max\leavevmode\nobreak$ $c^{\top}z$ ($\operatorname{Primal}(x)$)
-$\mathbf{A}z$ $\leq\mathbf{B}x+b$
-$z$ $\geq 0$
-Then, for every $x\in\mathcal{D}$, $g\in\partial^{\wedge}\varphi(x)$ if and only if $g=\mathbf{B}^{\top}\pi$ for an optimal solution $\pi\in\mathbb{R}_{\geq 0}^{k}$ to the dual linear program
-$\min\leavevmode\nobreak$ $x^{\top}\mathbf{B}^{\top}\pi+b^{\top}\pi$ ($\operatorname{Dual}(x)$)
-$\mathbf{A}^{\top}\pi$ $\geq c$
-$\pi$ $\geq 0\,.$
-Consider a generalized flow instance defined in a graph $G=(V,E)$, sink $t\in V$, capacities $h\in\mathbb{R}_{\geq 0}^{E}$, gain factors $\gamma\in\mathbb{R}_{>0}^{E}$ and supply set $S={\mathcal{M}}$.
-For the rest of this section, let $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ be a generalized network utility function defined in the form $\operatorname{GenflowLP}(x)$ in Section 2.
-Using Lemma 8.1 and complementary slackness, we get the following characterization of subgradients of generalized network utility functions.
-Note that for a generalized network utility function, the domain is $\mathcal{D}=\mathbb{R}_{\geq 0}^{\mathcal{M}}$, because $f=0$ is always a feasible solution to $\operatorname{GenflowLP}(x)$, and the definition assumes that the objective is always bounded.
-Lemma 8.2.
-[…]
-$\gamma_{e}\pi_{w}-\pi_{v}$ $=0\,,\quad\forall e=(v,w)\in E:h_{e}>f_{e}>0\,,$ (9)
-$\gamma_{e}\pi_{w}-\pi_{v}$ $\leq 0\,,\quad\forall e=(v,w)\in E:f_{e}=0\,,$
-$\gamma_{e}\pi_{w}-\pi_{v}$ $\geq 0\,,\quad\forall e=(v,w)\in E:f_{e}=h_{e}\,,$
-$\pi_{v}$ $=0\,,\quad\forall v\in{\mathcal{M}}:\operatorname{net}_{f}(v)>-x_{v}\mbox{ or }v\in V\setminus({\mathcal{M}}\cup t):\operatorname{net}_{f}(v)>0\,,$
-$\pi_{t}$ $=1\,,$
-$\pi$ $\geq 0\,.$
-In accordance with this lemma, we say that $f\in\mathbb{R}^{E}$ and $\pi\in\mathbb{R}^{V}$ form a fitting pair with respect to $x$, if $f$ is a feasible generalized flow with supply $x$, and $\pi$ satisfies (9).
-Note that this implies that $f$ and $\pi$ are primal and dual optimal solutions to $\operatorname{GenflowLP}(x)$.
-Further, $g\in\partial^{\wedge}u(x)$ if and only if $g$ is the restriction of $\pi$ to ${\mathcal{M}}$ for a fitting pair $(f,\pi)$ with respect to $x$.
-For every arc $e=(v,w)\in E$, let us define the reverse arc $\overleftarrow{e}=(v,w)$ with gain factor $\gamma_{\overleftarrow{e}}=1/\gamma_{e}$.
-By sending $\alpha$ units of flow on $\overleftarrow{e}$ we mean decreasing the flow value on $e$ by $\gamma_{e}\alpha$.
-Given a fitting pair $(f,\pi)$ with respect to $x$, we define the auxiliary network $G^{\textrm{aux}}_{f,\pi}=(V,E^{\textrm{aux}}_{f,\pi})$ by
-$E^{\textrm{aux}}_{f,\pi}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\left\{e=(v,w)\in E\,:\,\gamma_{e}\pi_{w}-\pi_{v}=0\mbox{ and }f_{e}<h_{e}\right\}\cup\left\{\overleftarrow{e}\,:\,e\in E\,:\,\gamma_{e}\pi_{w}-\pi_{v}=0\mbox{ and }0<f_{e}\right\}\,.$
-In the proof of Theorem 1.5, we use two flow subroutines in the generalized flow network.
-The first one starts with a fitting pair $(f,\pi)$ with respect to a demand vector $x$, and outputs a flow $f^{\prime}$ for a decreased demand vector $x^{\prime}$ such that the flow value $\operatorname{net}_{f^{\prime}}(t)$ is decreased to the smallest possible amount while maintaining $(f^{\prime},\pi)$ as a fitting pair.
-Input: $x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, and a fitting pair $(f,\pi)$ with respect to $x$
-Output: A supply vector $0\leq x^{\prime}\leq x$ and a feasible generalized flow $f^{\prime}$ such that $(f^{\prime},\pi)$ is a fitting pair with respect to $x^{\prime}$, and subject to this, $\operatorname{net}_{f^{\prime}}(t)$ is minimal.
-Algorithm 1 Pullback-Flow
-Lemma 8.3.
-Pullback-Flow can be formulated as a linear program.
-For the output flow $f^{\prime}$, the auxiliary graph $E^{\textrm{aux}}_{f^{\prime},\pi}$ does not contain any directed paths from the sink $t$ to $\operatorname{supp}(x^{\prime})$.
-Proof.
-When fixing $\pi$, the requirement that $(f^{\prime},\pi)$ be a fitting pair can be formulated by linear constraints on $f^{\prime}$.
-For example, for every arc $e=(v,w)$, if $\gamma_{e}\pi_{w}-\pi_{v}<0$ then we require $f^{\prime}_{e}=0$, and if $\gamma_{e}\pi_{w}-\pi_{v}>0$ then $f^{\prime}_{e}=h_{e}$; if $\gamma_{e}\pi_{w}-\pi_{v}=0$ then any value $0\leq f^{\prime}_{e}\leq h_{e}$ is allowed.
-For the second part, assume for a contradiction that $E^{\textrm{aux}}_{f^{\prime},\pi}$ contains a directed path $P$ from $t$ to a node $j\in{\mathcal{M}}$ with $x^{\prime}_{j}>0$.
-Then, we can send a small amount of flow along this path by maintaining the flow balance on every intermediate node, and decreasing $\operatorname{net}_{f^{\prime}}(t)$ and $x^{\prime}_{j}$ by positive amounts.
-This contradicts the minimality of $\operatorname{net}_{f^{\prime}}(t)$.
-∎
-Remark 8.4.
-Pullback-Flow can be in fact formulated as a (classical) maximum flow problem.
-To see this, one can use the standard generalized flow relabelling technique, first introduced by Glover Glover and Klingman (1973), see also e.g., Goldberg et al.
-(1991), Olver and Végh (2020)).
-That is, we multiply the flow at node $v$ by $\pi_{v}$ and replacing $\gamma_{e}$ by $\gamma_{e}\pi_{w}/\pi_{v}$ for every $e=(v,w)\in E$.
-On every arc, we fix $f^{\prime}_{e}$ to the upper or lower bound unless $\gamma_{e}\pi_{w}-\pi_{v}=0$, in which case the relabelled gain factor is 1.
-Our second subroutine aims to increase the supply at a designated node $k\in{\mathcal{M}}$ while the supply at all other goods is non-increasing, and the flow value $\operatorname{net}_{f}(t)$ is non-decreasing.
-The goal is again to maintain a fitting pair $(f^{\prime},\pi)$ for a demand vector $x^{\prime}$.
-Input: $x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, and a fitting pair $(f,\pi)$ with respect to $x$, and a good $k\in{\mathcal{M}}$ such that $\pi_{k}>0$.
-Output: A supply vector $x^{\prime}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $x^{\prime}_{k}\geq x_{k}$ and $x^{\prime}_{j}\leq x_{j}$ if $j\in{\mathcal{M}}\setminus\{k\}$, and a feasible generalized flow $f^{\prime}$ such that $(f^{\prime},\pi)$ is a fitting pair with respect to $x^{\prime}$, $\operatorname{net}_{f}(t^{\prime})\geq\operatorname{net}_{f}(t)$, and subject to this, $x^{\prime}_{k}$ is maximal.
-Algorithm 2 Reroute-Flow
-The following lemma can be derived similarly to Lemma 8.3.
-Lemma 8.5.
-Reroute-Flow can be formulated as a linear program.
-For the output flow $f^{\prime}$, the auxiliary graph $E^{\textrm{aux}}_{f^{\prime},\pi}$ does not contain any directed path from $k$ to the set $(\operatorname{supp}(x^{\prime})\setminus\{k\})\cup\{t\}$.
-From Lemma 5.3, we get the following.
-Lemma 8.6.
-Consider a generalized network utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$, $b>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$, and $x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$.
-Then, $x\in\mathcal{GD}^{u}(b,\mathbf{q})$ if and only if there exists a fitting pair $(f,\pi)$ with respect to $x$ such that $u(x)=\operatorname{net}_{f}(t)$ and
-$\pi_{j}\leq\frac{q_{j}}{b}\cdot u(x)\,\quad\forall j\in{\mathcal{M}}\,,\,\text{ with equality for }j\in\operatorname{supp}(x)\,.$ (10)
-Theorem 1.5 follows by showing the two properties in Definition 2.14 in Lemmas 8.7 and Lemma 8.8 below. We start by property (ii).
-Lemma 8.7.
-Let $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ be a generalized network utility function, $b>b^{\prime}>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$, and let $x\in\mathcal{GD}^{u}(b,\mathbf{q})$.
-Then, there exists $x^{\prime}\in\mathcal{GD}^{u}(b^{\prime},\mathbf{q})$ such that $x^{\prime}\leq x$.
-Proof.
-The proof is given by a finite algorithm, shown as Algorithm 3.
-We start from fitting pair $(f,\pi)$ with respect to $x$ as in Lemma 8.6.
-Thus, the inequality (10) holds for $\pi$, $x$, $\mathbf{q}$ and $b$.
-If $u(x)=0$, then we must have $\pi_{j}=0$ for all $j\in{\mathcal{M}}$ and thus (10) holds also when replacing $b$ by $b^{\prime}$, and therefore we can return $x^{\prime}=x$.
-For the rest of the proof, let us assume $u(x)>0$.
-Similarly, we can assume that $\mathbf{q}\neq\mathbf{0}$.
-Note that $\pi_{j}>0$ for every $j\in\operatorname{supp}(x)\cap\operatorname{supp}(\mathbf{q})$.
-When replacing $b$ by $b^{\prime}$, all inequalities with in (10) with $q_{j}>0$ become strict; the ratio of the right and left hand sides will be $\alpha\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}b/b^{\prime}>1$ for all $j\in\operatorname{supp}(x)\cap\operatorname{supp}(\mathbf{q})$.
-Also, recall $u(x)=\operatorname{net}_{f}(t)$.
-Algorithm 3 alternates between two operations: changing $x$ and $f$ while keeping $\pi$ unchanged; and changing the dual $\pi$ while keeping $x$ and $f$ unchanged.
-Throughout, it maintains a fitting pair with respect to a decreasing vector $x^{\prime}$, such that the required equalities in (10) are gradually restored.
-[…]
-Case I:
-If $\alpha^{\prime}\leq 1$, then we choose $\lambda\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}(\alpha-1)/(\alpha-\alpha^{\prime})\in[0,1]$ such that $\lambda\alpha+(1-\lambda)\alpha^{\prime}=1$.
-Then, $x^{\prime\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\lambda x^{\prime}+(1-\lambda)x$ and the flow $f^{\prime\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\lambda f^{\prime}+(1-\lambda)f$ satisfy that $(f^{\prime\prime},\pi)$ is a fitting pair with respect to $x^{\prime\prime}$, and (10) holds for $\pi$, $x^{\prime\prime}$ and $b^{\prime}$.
-Thus, we can terminate by outputting $x^{\prime\prime}\in\mathcal{GD}^{u}(b^{\prime},\mathbf{q})$.
-Case II:
-If $\alpha^{\prime}>1$, then we update the dual $\pi$ to $\pi^{\prime}$ as follows.
-Let $T\supseteq\operatorname{supp}(x^{\prime})$ be the set of nodes that can reach $\operatorname{supp}(x^{\prime})$ in the auxiliary graph $(G,E^{\textrm{aux}}_{f^{\prime},\pi})$.
-By Lemma 8.3, $t\notin T$.
-Let us define $\pi^{\prime}$ such that $\pi^{\prime}_{v}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\pi_{v}$ for $v\notin T$ and $\pi^{\prime}_{v}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\delta\pi_{v}$ for $v\in T$ for the largest value $1\leq\delta\leq\alpha^{\prime}$ such that $(f^{\prime},\pi^{\prime})$ is a fitting pair.
-We claim that $\delta^{\prime}>1$.
-Recall that fitting pairs are required to satisfy (9).
-The sign of $\gamma_{e}\pi^{\prime}_{w}-\pi^{\prime}_{v}$ is the same as the sign of $\gamma_{e}\pi_{w}-\pi_{v}$ for all arcs $e=(v,w)$ with both endpoints inside $T$ or outside $T$.
-There is no arc $e=(v,w)$ entering or leaving $T$ with $0<f^{\prime}_{e}<u_{e}$, since $e,\overleftarrow{e}\in E^{\textrm{aux}}_{f^{\prime},\pi}$ for all such arcs.
-For every arc $e=(v,w)$ entering $T$ with $f^{\prime}_{e}=0$, we must have $\gamma_{e}\pi_{w}-\pi_{v}<0$ so that $e\notin E^{\textrm{aux}}_{f^{\prime},\pi}$.
-Similarly, for every arc $e=(v,w)$ leaving $T$ with $f^{\prime}_{e}=u_{e}$, we must have $\gamma_{e}\pi_{w}-\pi_{v}>0$.
-Hence, we can always choose $\delta^{\prime}>1$.
-If $\delta^{\prime}=\alpha^{\prime}$, then $\pi^{\prime}$, $x^{\prime}$, $\mathbf{q}$ and $b^{\prime}$ satisfy (10), since the right hand side for all $j\in\operatorname{supp}(x^{\prime})$ has increased by a factor $\alpha^{\prime}$, and therefore equality holds for all these indices.
-We can therefore conclude $x^{\prime}\in\mathcal{GD}^{u}(b,\mathbf{q})$ and terminate by outputting $x^{\prime}$.
-In case $\delta^{\prime}<\alpha^{\prime}$, then (10) holds with strict inequality for all $j$ for $\pi^{\prime}$, $x^{\prime}$, $\mathbf{q}$, and $b^{\prime}$.
-Moreover, the ratio of the right hand side and the left hand side is the same for all $j\in\operatorname{supp}(x^{\prime})\cap\operatorname{supp}(\mathbf{q})$, namely, $\alpha^{\prime}/\delta^{\prime}<\alpha$.
-We now restart the above procedure with $(f^{\prime},\pi^{\prime})$ and $x^{\prime}$ in place of $(f,\pi)$ and $x$.
-We claim that the algorithm terminates in a finite number of iterations.
-This holds because the linear program Pullback-Flow only depends on the sign pattern of $\gamma_{e}\pi_{w}-\pi_{v}$ for the arcs $e=(w,v)$ and on $\operatorname{supp}(x)$.
-There are $3^{|E|}$ possible sign patterns; the sest $\operatorname{supp}(x)$ is monotone decreasing and therefore there are at most $|{\mathcal{M}}|$ different possibilities.
-The same configuration of arc sign patterns and $\operatorname{supp}(x)$ cannot occur twice, because $u(x^{\prime})=\operatorname{net}_{f^{\prime}}(t)$ is strictly decreasing in every iteration of the algorithm.
-Hence, after a finite number of iterations we arrive at $\pi^{\prime}$ and $x^{\prime}$ satisfying (10) for $b^{\prime}$.
-Since $x^{\prime}$ is decreasing throughout, we have $x^{\prime}\leq x$ for the input supply vector $x$, as required.
-∎
-Input: $\mathbf{q}\in\operatorname{pdom}(u)$, $b,b^{\prime}>0$, $x\in\mathcal{GD}^{u}(b,\mathbf{q})$.
-Output: $0\leq x^{\prime}\leq x$ such that $x^{\prime}\in\mathcal{GD}^{u}(b^{\prime},\mathbf{q})$.
-1 Find a fitting pair $(f,\pi)$ with respect to $x$ such that $u(x)=\operatorname{net}_{t}(x)$ and (10) holds;
-2 if $u(x)=0$ or $\mathbf{q}=\mathbf{0}$ then return $x$;
-3 $\bar{x}\leftarrow x$ ; $\alpha\leftarrow b/b^{\prime}$ ;
-4 while $\alpha>1$ do
-5       $(x^{\prime},f^{\prime})\leftarrow\textnormal{{Pullback-Flow}}(\bar{x},f,\pi)$ ;
-6       $\alpha^{\prime}\leftarrow\alpha\frac{u(x^{\prime})}{u(x)}$ ;
-7       if $\alpha^{\prime}\leq 1$ then
-8             $\lambda\leftarrow\frac{\alpha-1}{\alpha-\alpha^{\prime}}$ ;
-9             $x^{\prime\prime}\leftarrow\frac{\alpha-1}{\alpha-\alpha^{\prime}}x^{\prime}+\frac{1-\alpha^{\prime}}{\alpha-\alpha^{\prime}}\bar{x}$ ;
-10             return $x^{\prime\prime}$
-11      else
-12             $T\leftarrow\left\{v\in V\,:\,v\text{ can reach }\operatorname{supp}(x^{\prime})\text{ in }(G,E^{\textrm{aux}}_{f^{\prime},\pi})\right\}$  ;
-13             $\delta\leftarrow\min\left\{\alpha,\min\left\{\frac{\gamma_{e}\pi_{w}}{\pi_{v}}\,:\,e=(v,w)\in\partial^{\mathrm{in}}(T)\right\},\min\left\{\frac{\pi_{v}}{\gamma_{e}\pi_{w}}\,:\,e=(v,w)\in\partial^{\mathrm{out}}(T)\right\}\right\}$;
-14             if $\delta^{\prime}=\alpha^{\prime}$ then return $x^{\prime}$;
-15             for $v\in T$ do $\pi_{v}\leftarrow\delta^{\prime}\pi_{v}$;
-16             $\alpha\leftarrow\alpha^{\prime}/\delta^{\prime}$ ; $\bar{x}\leftarrow x^{\prime}$ ;
-17
-18
-Algorithm 3 Budget-Decrease
-We next prove property (i) in Definition 2.14.
-Lemma 8.8.
-Let $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ be a generalized network utility function and $\mathbf{q},\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$ such that $0\leq\mathbf{q}^{\prime}\leq\mathbf{q}$, and let $x\in\mathcal{GD}^{u}(b,\mathbf{q})$.
-Then, there exists $x^{\prime}\in\mathcal{GD}^{u}(b,\mathbf{q}^{\prime})$ such that $x^{\prime}_{j}\leq x_{j}$ whenever $q^{\prime}_{j}=q_{j}$.
-Proof.
-Clearly, it suffices to show that the statement holds when $q^{\prime}_{k}<q_{k}$ for a single good $k\in{\mathcal{M}}$.
-Again, we start from fitting pair $(f,\pi)$ with respect to $x$ as in Lemma 8.6.
-Similarly to Lemma 8.7, we may assume that $u(x)>0$ and $\mathbf{q}\neq\mathbf{0}$.
-Let us now replace $\mathbf{q}$ by $\mathbf{q}^{\prime}$.
-The condition (10) can only be violated for $j=k$.
-If it is not violated, then we may return $x^{\prime}=x$; this is always the case when $\pi_{k}=0$.
-For the rest of the proof, assume that the (10) is violated for $k$ with the cost $q^{\prime}_{k}$.
-Note that $\pi_{j}=\beta q^{\prime}_{j}$ for every $j\neq k$, where $\beta\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}u(x)/b$; we now have $\pi_{k}>\beta q^{\prime}_{k}\geq 0$.
-This value of $\beta$ will be fixed throughout the argument.
-[…]
-We can again argue that this algorithm terminates in a finite number of iterations.
-Note that $\operatorname{supp}(x^{\prime})$ is monotone decreasing throughout, and $\pi^{\prime}_{j}=\beta q^{\prime}_{j}$ is maintained for all $j\in\operatorname{supp}(x^{\prime})\setminus\{k\}$.
-For goods with $x^{\prime}_{j}=0$, $\pi^{\prime}_{j}$ may decrease, and therefore $\pi^{\prime}_{j}\leq\beta q_{j}$ for these goods.
-At termination, we have $\pi^{\prime}_{j}=\beta q^{\prime}_{j}$ for all $j\in\operatorname{supp}(x^{\prime})$ — including $j=k$ — and $\pi^{\prime}_{j}\leq\beta q^{\prime}_{j}$ whenever $x^{\prime}_{j}=0$.
-Moreover, note that $x^{\prime}_{j}\leq x_{j}$ for all $j\neq k$.
-Throughout the algorithm, $u(x^{\prime})$ is non-decreasing.
-Hence, $u(x^{\prime})/b\geq\beta$ holds.
-By setting $\hat{b}>b$ such that $u(x^{\prime})/\hat{b}=\beta$, (10) holds for $x^{\prime}$, $\pi^{\prime}$, $q^{\prime}$ and $\hat{b}$.
-Hence, $x^{\prime}\in\mathcal{GD}_{u}(\hat{b},\mathbf{q}^{\prime})$.
-Applying now Lemma 8.7 for $x^{\prime}$, the prices $\mathbf{q}^{\prime}$ and budgets $\hat{b}\geq b$, we can find $x^{\prime\prime}\in\mathcal{GD}_{u}(b,\mathbf{q}^{\prime})$ such that $x^{\prime\prime}\leq x^{\prime}$.
-Note that $x^{\prime\prime}_{j}\leq x^{\prime}_{j}\leq x_{j}$ for all $j\neq k$, and therefore $x^{\prime\prime}$ satisfies the requirements of the Lemma.
-[…]
-This concludes the proof of Theorem 1.5, namely, that all generalized network utilities are Gale-substitutes.
-Nevertheless, they are not $\Sigma$-Gale-substitutes, and so Theorem 1.3 is only applicable in the case when the generalized network utilities are non-satiable.
-Here, we also show Theorem 1.6, namely, that the desired utility bound holds for all generalized network utilities if we consider thrifty competitive equilibria.
+Although generalized network utilities are not $\Sigma$-Gale-substitutes in general (see Example E.6), we can still establish the following theorem with a slight modification to the proof of Theorem 1.3; the proof is also given in Appendix D.
 See 1.6
-Proof sketch.
-We only need to slightly modify the argument in Section 6.
-Property ($\Sigma$) is only used in the proof of Lemma 3.2.
-We show that for generalized network utilities, if $x\in{\mathcal{D}}^{u}(\mathbf{p},b)$ for $\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ and further $x$ minimizes $\left\langle\mathbf{p},x\right\rangle$ in ${\mathcal{D}}^{u}(\mathbf{p},b)$, then there exists $y\in\mathcal{GD}^{u}(\mathbf{p},b)$ such that $y\leq x$.
-From this property, we can derive Lemma 6.1 for thrifty competitive equilibria; the rest of the proof remains identical to the proof of Theorem 1.3.
-Consider now (6) in the proof of Lemma 3.2.
-If $\lambda>0$, the same argument is applicable.
-Consider now the case $\lambda=0$.
-This means that the constraint $\left\langle\mathbf{p},x\right\rangle\leq b$ is not active in the program defining ${\mathcal{D}}^{u}(\mathbf{p},b)$; that is, $\max_{z\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\{u(z)\,:\,\left\langle\mathbf{p},z\right\rangle\leq b\}=\max_{z\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}u(z)=v^{\star}$.
-In particular, $u(x)=v^{\star}$.
-The budget minimization property means that $x$ is an optimal solution to the program
-$\min\leavevmode\nobreak$ $\mathbf{p}^{\top}x$ (11)
-$\operatorname{net}_{f}(t)$ $\geq v^{\star}$
-$\operatorname{net}_{f}(v)$ $\geq-x_{v}\quad\forall v\in S$
-$\operatorname{net}_{f}(v)$ $\geq 0\quad\forall v\in V\setminus(S\cup\{t\})$
-00 0 $\leq f\leq h\,.$
-From here, using LP duality and Lemma 8.1, we can see that there exists a supergradient $\mathbf{g}\in\partial^{\wedge}u(x)$ and $\mu\in\mathbb{R}_{\geq 0}$ such that
-$\mu g_{j}\leq p_{j}\,,\quad\forall j\in{\mathcal{M}}\,,\,\mbox{with equality if }x_{j}>0\,.$
-If $\mu>0$, then we can use the same argument for $\lambda=1/\mu$.
-If $\mu=0$, given that $\mu$ acts as the dual variable of $\operatorname{net}_{f}(t)\geq v^{\star}$, which is no longer active, the optimal value of (11) becomes 00 since $\mathbf{0}$ is a feasible allocation of $u(\cdot)$ according to our definition.
-This implies the existence of a zero-cost optimal allocation $x$: $u(x)=v^{*}$ and $\mathbf{p}^{\top}x=0$, which indicates that $x$ belongs to both ${\mathcal{D}}^{u}(\mathbf{p},b)$ and $\mathcal{GD}^{u}(\mathbf{p},b)$.
-Thus, the result follows.
-∎
 7 Discussion
@@
 We also introduced generalized network utilities.
-This is a very general class with a rich combinatorial structure and may be a relevant example in various market and fair division settings.
-Gale-substitutes may be a class of concave functions worth investigating on its own.
-In the context of indivisible goods, the analogous class of gross substitute utilities, introduced by Keslo and Crawford (1982), turned out to be the same as the important class of $M^{\natural}$-concave utilities in discrete convex analysis Fujishige and Yang (2003).
-In particular, one may wonder whether the Gale-substitutes property implies submodularity, that is, $u(x)+u(y)\geq u(x\vee y)+u(x\wedge y)$ for any $x,y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, where $x\vee y$ and $x\wedge y$ denote the pointwise maximum and minimum of the two vectors, respectively.
+This is a very general construction with rich combinatorial structure and may be a relevant example in various market and fair division settings.
+Gale-substitutes may form a compelling subclass of concave functions worth investigating on its own.
+In the context of indivisible goods, the analogous class of gross substitute utilities, introduced by Kelso and Crawford (1982), turned out to be the same as the important class of $M^{\natural}$-concave utilities in discrete convex analysis Fujishige and Yang (2003).
+In particular, one may wonder whether the Gale-substitutes property implies submodularity, that is, $u(\mathbf{x})+u(\mathbf{y})\geq u(\mathbf{x}\vee\mathbf{y})+u(\mathbf{x}\wedge\mathbf{y})$ for any $\mathbf{x},\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, where $\mathbf{x}\vee\mathbf{y}$ and $\mathbf{x}\wedge\mathbf{y}$ denote the pointwise maximum and minimum of the two vectors, respectively.
 This property trivially holds for separable utilities and can also be shown for generalized network utilities.
+Another interesting question is whether computing competitive equilibrium in generalized network utilities lies in PPAD.
 We note that one can construct an example of a concave quadratic function that is submodular but not Gale-substitutes.
 On the other hand, it is unclear whether Theorem 1.6 would remain true for submodular utilities instead of Gale-substitutes.
-One would also be interested in whether Theorem 1.6 holds for matching utilities, which, like submodular utilities, are not necessarily Gale-substitutesl(see Example A.11).
+One would also be interested in whether Theorem 1.6 holds for matching utilities, which, like submodular utilities, are not necessarily Gale-substitutes (see Example E.7).
 Another natural question related to the well-studied notion of weak gross substitutes (WGS) utilities: Gale-substitutes includes utility functions that are not WGS, such as concave separable; but it is not clear whether all WGS utilities are also Gale-substitutes.
@@
 (2023a).
+In particular, Garg et al.
+(2023a) gives an auction algorithm for maximizing Nash welfare under capped SPLC utilities.
+Similar methods could be extended to the broader class of Gale substitutes utilities.
 The Gale-substitutes property may be useful to design simple such mechanisms and combinatorial algorithms for solving the convex program (EG).
@@
 (2022), Cole and Gkatzelis (2015), Duan and Mehlhorn (2015), Garg and Végh (2023).
-Appendix A Examples
+Appendix A Properties of the Gale demand correspondence
+In this Appendix, we prove Lemma 2.4. We also derive some technical properties of the price domain $\operatorname{pdom}(u)$ that will be used in the subsequent Appendices.
+See 2.4
+Proof.
+Let $M\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sup_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(\mathbf{y})-\left\langle\mathbf{p},\mathbf{y}\right\rangle$, which implies the existence of a sequence of allocations $\{\mathbf{y}^{(k)}\}_{k=1}^{\infty}$ such that $b\log u(\mathbf{y}^{(k)})-\left\langle\mathbf{p},\mathbf{y}^{(k)}\right\rangle$ converges to $M$.
+We note that the finiteness of $M$ follows from Lemma 4.1, which is explicitly established in Lemma A.1(i).
+We show the existence of an allocation $\mathbf{y}^{*}$ satisfying $b\log u(\mathbf{y}^{*})-\left\langle\mathbf{p},\mathbf{y}^{*}\right\rangle=M$.
+According to Lemma 4.1, there exists a sequence of $\bar{\mathbf{y}}^{(k)}$ such that $b\log u(\bar{\mathbf{y}}^{(k)})-\left\langle\mathbf{p},\bar{\mathbf{y}}^{(k)}\right\rangle$ converges to $M$, with $0\leq\left\langle\mathbf{p},\bar{\mathbf{y}}^{(k)}\right\rangle\leq b$ for all $k$.
+Thus, we can select a subsequence $\bar{\mathbf{y}}^{(i_{k})}$ such that $b\log u(\bar{\mathbf{y}}^{(i_{k})})-\left\langle\mathbf{p},\bar{\mathbf{y}}^{(i_{k})}\right\rangle$ converges to $M$ and $\left\langle\mathbf{p},\bar{\mathbf{y}}^{(i_{k})}\right\rangle$ converges to $b^{\prime}$ for some $b^{\prime}\in[0,b]$.
+Consequently, $b\log u(\bar{\mathbf{y}}^{(i_{k})})$ converges to $M+b^{\prime}$ as $\left\langle\mathbf{p},\bar{\mathbf{y}}^{(i_{k})}\right\rangle$ converges to $b^{\prime}$.
+By Assumption 1, there exists an allocation $\mathbf{y}^{*}$ such that $b\log u(\mathbf{y}^{*})=M+b^{\prime}$ and $\left\langle\mathbf{p},\mathbf{y}^{*}\right\rangle\leq b^{\prime}$.
+This completes the proof.
+∎
+Lemma A.1.
+For any utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ and $b>0$, the following properties of the price domain $\operatorname{pdom}(u)$ hold:
+  1. (i)
+    for any $\mathbf{q}\in\operatorname{pdom}(u)$, $\max_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(\mathbf{y})-\left\langle\mathbf{q},\mathbf{y}\right\rangle<+\infty$;
+  2. (ii)
+    $\{\mathbf{q}\penalty 10000\ |\penalty 10000\ \mathbf{q}>0\}\subseteq\operatorname{pdom}(u)$;
+  3. (iii)
+    for any $\mathbf{q}\geq\mathbf{q}^{\prime}\geq 0$ such that $\{j:q^{\prime}_{j}=0\}=\{j:q_{j}=0\}$, $\mathbf{q}\in\operatorname{pdom}(u)$ implies $\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$;
+  4. (iv)
+    for any $\mathbf{p}\geq 0$, if $\max_{\left\langle\mathbf{p},\mathbf{y}\right\rangle\leq b}\log u(\mathbf{y})<+\infty$, then $\mathbf{p}\in\operatorname{pdom}(u)$.
+  5. (v)
+    for any $M>0$, if there is a convergent sequence of prices $\mathbf{p}^{(k)}\in\operatorname{pdom}(u)$ with $M>\max_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(\mathbf{y})-\left\langle\mathbf{p}^{(k)},\mathbf{y}\right\rangle$, $k=1,2,\ldots$ with a limit point $\mathbf{p}^{(k)}\to\bar{\mathbf{p}}$, then $\bar{\mathbf{p}}\in\operatorname{pdom}(u)$.
+Proof.
+We first prove property (i).
+According to the definition of $\operatorname{pdom}(u)$, there exists $b^{\prime}>0$ such that $\max_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b^{\prime}\log u(\mathbf{y})-\left\langle\mathbf{q},\mathbf{y}\right\rangle<+\infty$.
+For any $b>0$ and for any $\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, if $\left\langle\mathbf{q},\mathbf{y}\right\rangle\leq b$, then $b\log u(\mathbf{y})-\left\langle\mathbf{q},\mathbf{y}\right\rangle<\frac{b}{b^{\prime}}\left(b^{\prime}\log u(\mathbf{y})-\left\langle\mathbf{q},\mathbf{y}\right\rangle\right)+\frac{b-b^{\prime}}{b^{\prime}}\ \left\langle\mathbf{q},\mathbf{y}\right\rangle$, which is bounded.
+Then, the result follows using Lemma 4.1.
+Property (ii) holds as, by Lemma 4.1, $\max_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(\mathbf{y})-\left\langle\mathbf{q},\mathbf{y}\right\rangle\leq\max_{\left\langle\mathbf{q},\mathbf{x}\right\rangle\leq b}b\log u(\mathbf{x})-\left\langle\mathbf{q},\mathbf{x}\right\rangle$, which is bounded as $\mathbf{x}$ is bounded: $x_{j}\leq\frac{b}{q_{j}}<\infty$.
+To prove property (iii), we let $\alpha=\min_{q_{j}>0}\{q^{\prime}_{j}/q_{j}\}$.
+Then, $\max_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}(\alpha b)\log u(\mathbf{y})-\left\langle\mathbf{q}^{\prime},\mathbf{y}\right\rangle\leq\max_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}(\alpha b)\log u(\mathbf{y})-\alpha\left\langle\mathbf{q},\mathbf{y}\right\rangle<+\infty$.
+Property (iv) holds as, for any $\mathbf{y}$ such that $\left\langle\mathbf{p},\mathbf{y}\right\rangle\leq b$, $b\log u(\mathbf{y})-\left\langle\mathbf{p},\mathbf{y}\right\rangle\leq b\log u(\mathbf{y})<+\infty$.
+Finally, let us prove (v).
+For a contradiction, assume $\bar{\mathbf{p}}\notin\operatorname{pdom}(u)$, that is, $\sup_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(\mathbf{y})-\left\langle\bar{\mathbf{p}},\mathbf{y}\right\rangle=\infty$.
+Consider a vector $\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ with $b\log u(\mathbf{y})-\left\langle\bar{\mathbf{p}},\mathbf{y}\right\rangle>2M$.
+Clearly, $\mathbf{y}\neq 0$.
+Let us select $k$ such that $\|\mathbf{p}^{(k)}-\bar{\mathbf{p}}\|<M/\|\mathbf{y}\|$.
+Then, $b\log u(\mathbf{y})-\left\langle\mathbf{p}^{(k)},\mathbf{y}\right\rangle=b\log u(\mathbf{y})-\left\langle\bar{\mathbf{p}},\mathbf{y}\right\rangle+\left\langle\bar{\mathbf{p}}-\mathbf{p}^{(k)},\mathbf{y}\right\rangle>M$, in contradiction with the assumption on the sequence $\mathbf{p}^{(k)}$.
+∎
+Appendix B Efficiency of Nash welfare for Gale-substitutes
+In this Appendix, we give the full proof of Theorem 1.3.
+We show that for $\Sigma$-Gale substitutes utility functions, any competitive equilibrium $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$ admits Gale prices $\mathbf{q}^{\star}$ such that $\mathbf{q}^{\star}\leq\mathbf{p}$.
+The theorem statement then follows from Lemma 5.1.
+We start by re-stating and proving Lemma 5.2.
+See 5.2
+Proof.
+According to the KKT conditions, there exists $\lambda\geq 0$ and supergradient $\mathbf{g}\in\partial^{\wedge}u(\mathbf{x})$ such that
+$g_{j}\leq\lambda p_{j}\,,\quad\forall j\in{\mathcal{M}}\,,\,\mbox{with equality if }x_{j}>0\,$, (6)
+and $\left\langle\mathbf{p},\mathbf{x}\right\rangle=b$ if $\lambda>0$.
+Let us first consider the case $\lambda=0$.
+Then, $g_{j}\leq 0$ with equality if $x_{j}>0$, implying that $\mathbf{x}=\arg\max_{\mathbf{x}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}u(\mathbf{x})$.
+Consequently, $u(.)$ is satiable; therefore, $\mathbf{0}\in\operatorname{pdom}(u)$ and $\mathbf{x}\in\mathcal{GD}^{u}(\mathbf{0},b)$.
+Now, ($\Sigma$) guarantees the existence of $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{p},b)$, $\mathbf{y}\leq\mathbf{x}$.
+Clearly, $\operatorname{pdom}(u)=\mathbb{R}_{\geq 0}^{\mathcal{M}}$ in this case.
+Next, assume $\lambda>0$. Let us define
+$\mathbf{q}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\frac{\lambda b}{u(\mathbf{x})}\cdot\mathbf{p}\,.$
+Note that $\mathbf{q}\leq\mathbf{p}$.
+This follows since $\lambda b=\lambda\left\langle\mathbf{p},\mathbf{x}\right\rangle=\left\langle\mathbf{g},\mathbf{x}\right\rangle\leq u(\mathbf{x})$, where the final inequality is by concavity and $u(\mathbf{0})=0$.
+Thus,
+$g_{j}\leq\frac{u(\mathbf{x})}{b}q_{j}\,,\quad\forall j\in{\mathcal{M}}\,,\,\mbox{with equality if }x_{j}>0\,.$
+According to Lemma 4.2, $\mathbf{x}\in\mathcal{GD}^{u}(\mathbf{q},b)$.
+Note that $\mathcal{GD}^{u}(\mathbf{q},b)=\mathcal{GD}^{u}(\mathbf{p},u(\mathbf{x})/\lambda)$.
+Utilizing Definition 2.14(ii) for $\mathbf{x}\in\mathcal{GD}^{u}(\mathbf{p},u(\mathbf{x})/\lambda)$ and $b\leq u(\mathbf{x})/\lambda$, there exists $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{p},b)$ such that $\mathbf{y}\leq\mathbf{x}$ as required.
+∎
+We next show that for any competitive equilibrium $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$, and any prices $\mathbf{q}\leq\mathbf{p}$, there exist an aggregate Gale demand such that $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1$ whenever $q_{j}=p_{j}$.
+Lemma B.1.
+Consider a Fisher market instance such that $u_{i}(\cdot)$ is $\Sigma$-Gale-substitute utilities for all agents.
+For any competitive equilibrium $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$, for any $\mathbf{q}\in\cap_{i\in{\mathcal{A}}}\operatorname{pdom}(u_{i})$ such that $\mathbf{q}\leq\mathbf{p}$, there exists $\mathbf{y}_{i}\in\mathcal{GD}_{i}(\mathbf{q},b_{i})$ for all $i\in{\mathcal{A}}$ and $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1$ for all $j\in{\mathcal{M}}$ such that $q_{j}=p_{j}$.
+Proof.
+By Lemma 5.2, there exists $\bar{\mathbf{y}}_{i}\in\mathcal{GD}_{i}(\mathbf{p},b_{i})$ for every $i\in{\mathcal{A}}$ such that $\bar{\mathbf{y}}_{i}\leq\mathbf{x}_{i}$.
+Using Definition 2.14(i), there exists $\mathbf{y}_{i}\in\mathcal{GD}_{i}(\mathbf{q},b_{i})$ and $y_{ij}\leq\bar{y}_{ij}$ if $q_{j}=p_{j}$.
+Thus, if $q_{j}=p_{j}$ then $\sum_{i\in{\mathcal{A}}}y_{ij}\leq\sum_{i\in{\mathcal{A}}}\bar{y}_{ij}\leq\sum_{i\in{\mathcal{A}}}x_{ij}\leq 1$.
+∎
+Recall the Lagrangian dual function $\varphi$ defined in (4) on the domain $\operatorname{dom}(\varphi)=\cap_{i\in{\mathcal{A}}}\operatorname{pdom}(u_{i})$.
+$\varphi(\mathbf{q})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{i\in{\mathcal{A}}}\max_{\mathbf{y}_{i}\in\mathbb{R}_{\geq 0}^{{\mathcal{M}}}}\left(b_{i}\log u_{i}(\mathbf{y}_{i})-\left\langle\mathbf{q},\mathbf{y}_{i}\right\rangle\right)+\sum_{j\in{\mathcal{M}}}q_{j}.$
+The properties stated in the next lemma are easy to verify.
+Lemma B.2.
+The following hold for the function $\varphi(\cdot)$.
+  1. (i)
+    $\varphi(\mathbf{q})$ is convex.
+  2. (ii)
+    Given $\mathbf{q}\in\operatorname{dom}(\varphi)$, let $\mathbf{y}_{i}\in\mathcal{GD}_{i}(\mathbf{q},b_{i})$ for all $i\in{\mathcal{A}}$.
+Then, $\mathbf{1}-\sum_{i\in{\mathcal{A}}}\mathbf{y}_{i}\in\partial^{\vee}\varphi(\mathbf{q})$.
+  3. (iii)
+    The optimum value of (5), that is, the infimum of $\varphi(\mathbf{q})$ is taken in $\operatorname{dom}(\varphi)$.
+Proof of Theorem 1.3.
+Let $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$ be a competitive equilibrium solution. Consider the dual (5) of the Eisenberg-Gale program (EG), i.e.,
+$\min_{\mathbf{q}\in\operatorname{dom}(\varphi)}\varphi(\mathbf{q})\,.$
+We show that this has an optimal solution with $\mathbf{q}\leq\mathbf{p}$.
+The theorem then follows by Lemma 5.1.
+Let us take any optimal $\mathbf{q}^{\star}\in\operatorname{dom}(\varphi)$; we are done if $\mathbf{q}^{\star}\leq\mathbf{p}$.
+Otherwise, let $J\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\{j\in{\mathcal{M}}\,:\,q^{\star}_{j}>p_{j}\}$, and let us define $\bar{\mathbf{q}}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ as
+[…]
+First, assuming $\bar{\mathbf{q}}\in\operatorname{dom}(\varphi)$, we show that $\bar{\mathbf{q}}$ is also optimal to (5) as claimed.
+Let us pick an allocation $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$ for the prices $\bar{\mathbf{q}}$ as guaranteed by Lemma B.1: $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1$ for all $j\in{\mathcal{M}}$ with $\bar{q}_{j}=p_{j}$; in particular, for all $j\in J$.
+According to Lemma B.2(ii), we get a subgradient $\bar{\mathbf{g}}\in\partial^{\vee}\varphi(\bar{\mathbf{q}})$ defined as $\bar{\mathbf{g}}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\mathbf{1}-\sum_{i\in{\mathcal{A}}}\mathbf{y}_{i}$; this has the property $\bar{g}_{j}\geq 0$ for all $j\in J$.
+We then have
+$\varphi(\mathbf{q}^{\star})\geq\varphi(\bar{\mathbf{q}})+\left\langle\bar{\mathbf{g}},\mathbf{q}^{\star}-\bar{\mathbf{q}}\right\rangle=\varphi(\bar{\mathbf{q}})+\sum_{j\in J}\bar{g}_{j}(q^{\star}_{j}-\bar{q}_{j})\geq\varphi(\bar{\mathbf{q}})\,.$
+Here, the first inequality is by the definition of subgradients; the equality uses that $q^{\star}_{j}=\bar{q}_{j}$ if $j\notin J$, and the final inequality follows since $\bar{g}_{j}\geq 0$ and $q^{\star}_{j}>\bar{q}_{j}$ for $j\in J$.
+This shows that $\bar{\mathbf{q}}$ is also an optimal solution to (5), completing the proof assuming $\bar{\mathbf{q}}\in\operatorname{dom}(\varphi)$.
+It is left to show that $\bar{\mathbf{q}}\in\operatorname{dom}(\varphi)$.
+Let $K\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\{j\in{\mathcal{M}}\,:\,q^{\star}_{j}=0<p_{j}\}$.
+Clearly, $K\cap J=\emptyset$.
+If $K=\emptyset$, then Lemma A.1(iii) already implies $\bar{\mathbf{q}}\in\operatorname{pdom}(u)$, since $\mathbf{p}\geq\bar{\mathbf{q}}$ and $\operatorname{supp}(\bar{\mathbf{q}})=\operatorname{supp}(\mathbf{p})$.
+Otherwise, for $k=1,2,\ldots$, let us define $\mathbf{q}^{(k)}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ as
+[…]
+We have $\mathbf{q}^{(k)}\in\operatorname{pdom}(u)$ again by Lemma A.1(iii).
+As in the first part, we take a subgradient $\mathbf{g}^{(k)}$ as guaranteed by Lemma B.2(ii), with the properties that $g^{(k)}_{j}\geq 0$ for all $j\in J$ and also $\mathbf{g}^{(k)}\leq\mathbf{1}$.
+We can write
+$\varphi(\mathbf{q}^{\star})\geq\varphi(\mathbf{q}^{(k)})+\left\langle\mathbf{g}^{(k)},\mathbf{q}^{\star}-\mathbf{q}^{(k)}\right\rangle=\varphi(\mathbf{q}^{(k)})+\sum_{j\in J}g^{(k)}_{j}(q^{\star}_{j}-q^{(k)}_{j})-\sum_{j\in K}g^{(k)}_{j}q^{(k)}_{j}\geq\varphi(\mathbf{q}^{(k)})-\frac{\|\mathbf{p}\|_{\infty}}{k}|K|\,.$
+Here, we use in the first equality that $q^{\star}_{j}=q^{(k)}_{j}$ for $j\notin J\cup K$, and $q^{\star}_{j}=0$ for $j\in K$.
+The last inequality uses $g^{(k)}_{j}\leq 1$ and the definition $q^{(k)}_{j}=p_{j}/k$.
+Thus, Lemma A.1(v) is applicable with $M=\varphi(\mathbf{q}^{\star})+|K|\cdot\|\mathbf{p}\|_{\infty}$, using that for each $j\in{\mathcal{M}}$, $\max_{\mathbf{y}_{i}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b_{i}\log u_{i}(\mathbf{y}_{i})-\left\langle\mathbf{q}^{(k)},\mathbf{y}_{i}\right\rangle\leq\varphi(\mathbf{q}^{(k)})$.
+This shows that $\bar{\mathbf{q}}\in\operatorname{pdom}(u_{i})$ for all $i\in{\mathcal{M}}$, and therefore $\bar{\mathbf{q}}\in\operatorname{dom}(\varphi)=\cap_{i\in{\mathcal{M}}}\operatorname{pdom}(u_{i})$.
+∎
+Appendix C Separable utilities are $\Sigma$-Gale-substitutes
+In this Appendix, we prove Theorem 1.4 for the general case. Recall the following simple monotonicity property of supergradients of univariate concave functions.
+Lemma C.1.
+Let $v\,:\,\mathbb{R}\to\mathbb{R}$ be a univariate concave function, $\alpha,\alpha^{\prime}\in\operatorname{dom}(v)$, and $g\in\partial^{\wedge}v(\alpha)$, $g^{\prime}\in\partial^{\wedge}v(\alpha^{\prime})$.
+Then, $\alpha<\alpha^{\prime}$ implies $g\geq g^{\prime}$.
+Conversely, if $g>g^{\prime}$, then $\alpha\leq\alpha^{\prime}$.
+The proof of Theorem 1.4 follows by the next three lemmas, that prove the three properties in Definition 2.14 for separable utilities: property (i) in Lemma C.2, property (ii) in Lemma C.3, and property ($\Sigma$) in Lemma C.4.
+A similar argument was used in Lemma 7.1 in Garg et al.
+(2022a) to show a variant of the Gale substitutes property for capped SPLC utilities.
+Lemma C.2.
+Let $u(\mathbf{x})=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$ be a separable utility function, and let $b>0$, $\mathbf{q},\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$ and $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$.
+If $\mathbf{q}\geq\mathbf{q}^{\prime}$, then there exists a $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ such that $y^{\prime}_{j}\leq y_{j}$ whenever $q_{j}^{\prime}=q_{j}$.
+Proof.
+Let $J\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\{j\in{\mathcal{M}}\,:\,q^{\prime}_{j}=q_{j}\}$.
+Let us select $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ minimizing the potential $\Psi(\mathbf{y}^{\prime})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{j\in J}(y^{\prime}_{j}-y_{j})^{+}$.
+Our goal is to show $\Psi(\mathbf{y}^{\prime})=0$.
+Throughout the proof, let $g^{\prime}_{j}\in\partial^{\wedge}v_{j}(y^{\prime}_{j})$ and $g_{j}\in\partial^{\wedge}v_{j}(y_{j})$ as in Lemma 6.1.
+First, we show $u(\mathbf{y}^{\prime})\geq u(\mathbf{y})$.
+For a contradiction, assume $u(\mathbf{y}^{\prime})<u(\mathbf{y})$.
+Then, according to Lemma 6.1, for every $j\in{\mathcal{M}}$ with $y_{j}>0$ and $q_{j}>0$ we have
+$g^{\prime}_{j}\leq q^{\prime}_{j}u(\mathbf{y}^{\prime})/b<q_{j}u(\mathbf{y})/b=g_{j}\,$,
+while for $j\in{\mathcal{M}}$ with $q_{j}=0$ we have $g^{\prime}_{j}=g_{j}=0$.
+Consequently, for each $y_{j}>0$, if $q_{j}>0$ then Lemma C.1 implies $y^{\prime}_{j}\geq y_{j}$, and if $q_{j}=0$ then $v_{j}(y_{j})=v_{j}(y_{j}^{\prime})$, both being equal to the maximum value of $v_{j}(.)$.
+These imply $u(\mathbf{y}^{\prime})\geq u(\mathbf{y})$, a contradiction.
+Using the above, for any $j\in J$ where $y^{\prime}_{j}>0$,
+$g^{\prime}_{j}=\frac{q^{\prime}_{j}}{b}u(\mathbf{y}^{\prime})\geq\frac{q_{j}}{b}u(\mathbf{y})\geq g_{j}\,.$ (7)
+Let us first consider the case $u(\mathbf{y}^{\prime})>u(\mathbf{y})$.
+For all $j\in J$ with $q^{\prime}_{j}=q_{j}>0$, we get $g^{\prime}_{j}>g_{j}$.
+By Lemma C.1, $y^{\prime}_{j}\leq y_{j}$ follows for such goods $j$.
+On the other hand, if $q^{\prime}_{j}=q_{j}=0$, then this implies $g^{\prime}_{j}=g_{j}=0$.
+For such indices, if $y^{\prime}_{j}>y_{j}$, then replacing $y^{\prime}_{j}$ by $y_{j}$ results in another bundle $\mathbf{y}^{\prime\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ with $\Psi(\mathbf{y}^{\prime\prime})<\Psi(\mathbf{y}^{\prime})$, a contradiction.
+For the rest of the proof, let us assume that $u(\mathbf{y}^{\prime})=u(\mathbf{y})$.
+We next show $y^{\prime}_{j}\geq y_{j}$ for all $j\in{\mathcal{M}}\setminus J$, that is, for all goods with $q^{\prime}_{j}<q_{j}$.
+This is clearly true if $y_{j}=0$.
+Whenever $y_{j}>0$ for such an index, we have
+$g^{\prime}_{j}\leq\frac{q^{\prime}_{j}}{b}u(\mathbf{y}^{\prime})<\frac{q_{j}}{b}u(\mathbf{y})=g_{j}\,$,
+and therefore $y^{\prime}_{j}\geq y_{j}$ by Lemma C.1.
+By monotonicity, $v_{j}(y^{\prime}_{j})\geq v_{j}(y_{j})$ for all $j\in{\mathcal{M}}\setminus J$.
+Since $u(\mathbf{y}^{\prime})=u(\mathbf{y})$, we have $\sum_{j\in J}v_{j}(y^{\prime}_{j})\leq\sum_{j\in J}v_{j}(y_{j})$.
+From (7), it follows that $g^{\prime}_{j}=g_{j}$ for all $j\in J$ with $y^{\prime}_{j}>0$.
+This is because the first inequality is now equality, and the second inequality may only be strict if $y_{j}=0<y^{\prime}_{j}$.
+By Lemma C.1, $g_{j}\geq g^{\prime}_{j}$ in this case, leading to a contradiction.
+It also follows that $g^{\prime}_{k}=g_{k}$ for all $k\in J$ with $y_{k}>0$ from the same argument.
+Assume for a contradiction $\Psi(\mathbf{y}^{\prime})>0$, and let us select an index $j\in J$ with $y^{\prime}_{j}>y_{j}$.
+If $v_{j}(y^{\prime}_{j})=v_{j}(y_{j})$, then we get a contradiction, because changing $y^{\prime}_{j}$ to $y_{j}$ would preserve the optimality conditions in Lemma 6.1, while decreasing the potential $\Psi(\mathbf{y}^{\prime})$.
+Thus, $v_{j}(y^{\prime}_{j})>v_{j}(y_{j})$.
+By $\sum_{j^{\prime}\in J}v_{j^{\prime}}(y^{\prime}_{j^{\prime}})\leq\sum_{j^{\prime}\in J}v_{j^{\prime}}(y_{j^{\prime}})$, there must be an index $k\in J$ with $v_{k}(y^{\prime}_{k})<v_{k}(y_{k})$.
+We note as $u(\mathbf{y})=u(\mathbf{y}^{\prime})$ and $j\in J$ and $k\in J$, this implies $g^{\prime}_{k}=g_{k}=\frac{q^{\prime}_{k}}{b}u(\mathbf{y}^{\prime})$ and $g^{\prime}_{j}=g_{j}=\frac{q^{\prime}_{j}}{b}u(\mathbf{y}^{\prime})$.
+We can now find small $\varepsilon_{k},\varepsilon_{j}>0$ such that $\varepsilon_{k}\cdot q^{\prime}_{k}=\varepsilon_{j}\cdot q^{\prime}_{j}$; $y^{\prime}_{j}-\varepsilon_{j}\geq y_{j}$ and $y^{\prime}_{k}+\varepsilon_{k}\leq y_{k}$.
+We note after making the update $y^{\prime\prime}_{j}=y^{\prime}_{j}-\varepsilon_{j}\geq y_{j}$ and $y^{\prime\prime}_{k}=y^{\prime}_{k}+\varepsilon_{k}\leq y_{k}$.
+Additionally, $v_{k}(y^{\prime}_{k}+\varepsilon_{k})+v_{j}(y^{\prime}_{j}-\varepsilon_{j})=v_{k}(y^{\prime}_{k})+v_{j}(y^{\prime}_{j})$ and $y^{\prime\prime}_{j}q^{\prime}_{j}+y^{\prime\prime}_{k}q^{\prime}_{k}=y^{\prime}_{j}q^{\prime}_{j}+y^{\prime}_{k}q^{\prime}_{k}$.
+Therefore, this modified solution $\mathbf{y}^{\prime\prime}$ satisfies the optimality conditions but $\Psi(\mathbf{y}^{\prime\prime})<\Psi(\mathbf{y}^{\prime})$, again leading to a contradiction.
+∎
+Lemma C.3.
+Let $u(\mathbf{x})=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$ be a separable utility function, $b>b^{\prime}>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$.
+Then, for any $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$, there exists a $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$ such that $\mathbf{y}^{\prime}\leq\mathbf{y}$.
+Proof.
+We argue similarly as in the previous proof: we pick $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$ which minimizes $\Psi(\mathbf{y}^{\prime})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{j\in{\mathcal{M}}}(y^{\prime}_{j}-y_{j})^{+}$, and show that $\Psi(\mathbf{y}^{\prime})=0$.
+First, we show that ${u(\mathbf{y})}/{b}\leq{u(\mathbf{y}^{\prime})}/{b^{\prime}}$.
+For a contradiction, suppose ${u(\mathbf{y})}/{b}>{u(\mathbf{y}^{\prime})}/{b^{\prime}}$.
+If $y_{j}>0$, then, by Lemma 6.1
+$g_{j}=\frac{q_{j}u(\mathbf{y})}{b}>\frac{q_{j}u(\mathbf{y}^{\prime})}{b^{\prime}}\geq g^{\prime}_{j},$
+which implies $v_{j}(y_{j})\leq v_{j}(y^{\prime}_{j})$ whenever $y_{j}>0$ (Lemma C.1), and therefore, $u(\mathbf{y})\leq u(\mathbf{y}^{\prime})$.
+Together with $b>b^{\prime}$, this implies ${u(\mathbf{y})}/{b}\leq{u(\mathbf{y}^{\prime})}/{b^{\prime}}$, a contradiction.
+Whenever $y^{\prime}_{j}>0$, we now have
+$g^{\prime}_{j}=\frac{q_{j}u(\mathbf{y}^{\prime})}{b^{\prime}}\geq\frac{q_{j}u(\mathbf{y})}{b}\geq g_{j}\,.$ (8)
+Assume first $u(\mathbf{y}^{\prime})/b^{\prime}>u(\mathbf{y})/b$.
+If $q_{j}>0$, then inequality (8) is strict, and therefore $y^{\prime}_{j}\leq y_{j}$ for all $j\in{\mathcal{M}}$ follows by Lemma C.1.
+If $q_{j}=0$, then $g^{\prime}_{j}=g_{j}=0$.
+If $y^{\prime}_{j}>y_{j}$ for such a good $j$, we can decrease $y^{\prime}_{j}$ to $y_{j}$ and thereby get an optimal allocation with strictly smaller $\Psi(\mathbf{y}^{\prime})$ value, leading to a contradiction.
+This completes the proof for $u(\mathbf{y}^{\prime})/b^{\prime}>u(\mathbf{y})/b$.
+For the rest, we may assume $u(\mathbf{y}^{\prime})/b^{\prime}=u(\mathbf{y})/b$.
+Now, $g^{\prime}_{j}=g_{j}$ must hold whenever $y^{\prime}_{j}>0$.
+This is because the first inequality in (8) is equality, and if the second is strict, then $y_{j}=0<y^{\prime}_{j}$ would give a contradiction to $g^{\prime}_{j}>g_{j}$.
+It also follows from the same argument that $g^{\prime}_{k}=g_{k}$ for all $k\in{\mathcal{M}}$ with $y_{k}>0$.
+Assume for a contradiction that $\Psi(\mathbf{y}^{\prime})>0$, and let us select an index $j\in{\mathcal{M}}$ such that $y^{\prime}_{j}>y_{j}$.
+If $v_{j}(y^{\prime}_{j})=v_{j}(y_{j})$, we obtain a contradiction because changing $y^{\prime}_{j}$ to $y_{j}$ would preserve the optimality conditions in Lemma 6.1 while decreasing the potential $\Psi(\mathbf{y}^{\prime})$.
+Hence, we must have $v_{j}(y^{\prime}_{j})>v_{j}(y_{j})$.
+Since $\sum_{j^{\prime}\in{\mathcal{M}}}v_{j^{\prime}}(y^{\prime}_{j^{\prime}})\leq\sum_{j^{\prime}\in{\mathcal{M}}}v_{j^{\prime}}(y_{j^{\prime}})$, there must exist an index $k\in{\mathcal{M}}$ with $v_{k}(y^{\prime}_{k})<v_{k}(y_{k})$.
+Noting that $u(\mathbf{y})=u(\mathbf{y}^{\prime})$ and $j,k\in{\mathcal{M}}$, this implies $g^{\prime}_{k}=g_{k}=\frac{q^{\prime}_{k}}{b}u(\mathbf{y}^{\prime})$ and $g^{\prime}_{j}=g_{j}=\frac{q^{\prime}_{j}}{b}u(\mathbf{y}^{\prime})$.
+We can now choose small $\varepsilon_{k},\varepsilon_{j}>0$ such that $\varepsilon_{k}\cdot q^{\prime}_{k}=\varepsilon_{j}\cdot q^{\prime}_{j}$; $y^{\prime}_{j}-\varepsilon_{j}\geq y_{j}$ and $y^{\prime}_{k}+\varepsilon_{k}\leq y_{k}$.
+After making this update, define $y^{\prime\prime}_{j}=y^{\prime}_{j}-\varepsilon_{j}\geq y_{j}$ and $y^{\prime\prime}_{k}=y^{\prime}_{k}+\varepsilon_{k}\leq y_{k}$.
+Moreover, we have $v_{k}(y^{\prime}_{k}+\varepsilon_{k})+v_{j}(y^{\prime}_{j}-\varepsilon_{j})=v_{k}(y^{\prime}_{k})+v_{j}(y^{\prime}_{j})$ and $y^{\prime\prime}_{j}q^{\prime}_{j}+y^{\prime\prime}_{k}q^{\prime}_{k}=y^{\prime}_{j}q^{\prime}_{j}+y^{\prime}_{k}q^{\prime}_{k}$.
+Thus, the modified solution $\mathbf{y}^{\prime\prime}$ satisfies the optimality conditions, but $\Psi(\mathbf{y}^{\prime\prime})<\Psi(\mathbf{y}^{\prime})$, again leading to a contradiction.
+∎
+Lemma C.4.
+Let $u(\mathbf{x})=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$ be a separable utility function such that $\mathbf{0}\in\operatorname{pdom}(u)$.
+Then, for any $\mathbf{q}\geq\mathbf{0}$, $b>0$ and any $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{0},b)$, there exists a $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b)$ such that $\mathbf{y}^{\prime}\leq\mathbf{y}$.
+Proof.
+$\mathbf{0}\in\operatorname{pdom}(u)$ implies $\sup_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}u(\mathbf{y})<\infty$.
+Therefore, $\operatorname{pdom}(u)=\mathbb{R}_{\geq 0}^{\mathcal{M}}$; in particular, $\mathbf{q}\in\operatorname{pdom}(u)$.
+Let us pick $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b)$ such that $\sum_{j\in{\mathcal{M}}}y^{\prime}_{j}$ is minimal.
+Note that $u(\mathbf{y})=\sup_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}u(\mathbf{y})$.
+In particular, for any $\alpha\geq y_{j}$, $v_{j}(\alpha)=v_{j}(y_{j})$.
+For a contradiction, assume there exists a $k\in{\mathcal{M}}$ with $y_{k}<y^{\prime}_{k}$.
+Then, reducing $y^{\prime}_{k}$ to $y_{k}$ may only increase $b\log u(\mathbf{y}^{\prime})-\left\langle\mathbf{q},\mathbf{y}^{\prime}\right\rangle$, while decreasing $\sum_{j\in{\mathcal{M}}}y^{\prime}_{j}$.
+This contradicts the choice of $\mathbf{y}^{\prime}$.
+∎
+Appendix D Generalized network utilities
+In this Appendix, we prove Theorem 1.5 and Theorem 1.6. We will first need a characterization of subgradients by linear programming duality. We state this in a more general form.
+Lemma D.1.
+Let $A\in\mathbb{R}^{k\times\ell}$, $B\in\mathbb{R}^{k\times m}$, $\mathbf{c}\in\mathbb{R}^{\ell}$, $\mathbf{b}\in\mathbb{R}^{k}$.
+Let $\mathcal{D}\subseteq\mathbb{R}^{m}$ be the set of vectors $\mathbf{x}\in\mathbb{R}^{m}$ such that the following parametric linear program is feasible and bounded, and let us define $\varphi:\mathcal{D}\to\mathbb{R}$ such that $\varphi(\mathbf{x})$ is the optimum value of this program:
+[…]
+$A\mathbf{z}$ $\leq B\mathbf{x}+\mathbf{b}$
+$\mathbf{z}$ $\geq\mathbf{0}$
+Then, for every $\mathbf{x}\in\mathcal{D}$, $\mathbf{g}\in\partial^{\wedge}\varphi(\mathbf{x})$ if and only if $\mathbf{g}=B^{\top}\bm{\pi}$ for an optimal solution $\bm{\pi}\in\mathbb{R}_{\geq 0}^{k}$ to the dual linear program
+[…]
+$A^{\top}\bm{\pi}$ $\geq\mathbf{c}$
+$\bm{\pi}$ $\geq\mathbf{0}\,.$
+Proof.
+First, assume that $\bm{\pi}\in\mathbb{R}_{\geq 0}^{\ell}$ is an optimal solution to $\operatorname{Dual}(\mathbf{x})$ for a given $\mathbf{x}\in\mathbb{R}^{m}$.
+We need to show that $\mathbf{g}=B^{\top}\bm{\pi}$ is a supergradient at $\mathbf{x}$, that is, $\varphi(\mathbf{t})\leq\varphi(\mathbf{x})+\left\langle\mathbf{g},\mathbf{t}-\mathbf{x}\right\rangle$ for any $\mathbf{t}\in\mathcal{D}$.
+The programs $\operatorname{Dual}(\mathbf{x})$ and $\operatorname{Dual}(\mathbf{t})$ have the same feasible region, with linear objective functions $B\mathbf{x}+\mathbf{b}$ and $B\mathbf{t}+\mathbf{b}$, respectively.
+By weak duality, the cost of every feasible dual solution is at least the primal optimum value, thus, $\varphi(\mathbf{t})\leq\mathbf{t}^{\top}B\bm{\pi}+\mathbf{b}^{\top}\bm{\pi}=\mathbf{g}^{\top}\mathbf{t}+\mathbf{b}^{\top}\bm{\pi}$.
+The statement follows as
+$\varphi(\mathbf{t})\leq\mathbf{g}^{\top}\mathbf{t}+\mathbf{b}^{\top}\bm{\pi}=\mathbf{g}^{\top}(\mathbf{t}-\mathbf{x})+\mathbf{x}^{\top}B\bm{\pi}+\mathbf{b}^{\top}\bm{\pi}=\varphi(\mathbf{x})+\left\langle\mathbf{g},\mathbf{t}-\mathbf{x}\right\rangle\,$,
+noting also that $\varphi(\mathbf{x})=\mathbf{x}^{\top}B\bm{\pi}+\mathbf{b}^{\top}\bm{\pi}$ since $\bm{\pi}$ is an optimal solution to $\operatorname{Dual}(\mathbf{x})$, and $\varphi(\mathbf{x})$ is the optimum value.
+Conversely, we need to show that if $\mathbf{g}\in\partial^{\wedge}\varphi(\mathbf{x})$, then there exists a dual optimal solution $\bm{\pi}$ such that $\mathbf{g}=B^{\top}\bm{\pi}$.
+Let $\mathbf{z}$ be a primal optimal solution to $\operatorname{Primal}(\mathbf{x})$.
+Using complementary slackness, we can turn $\operatorname{Dual}(\mathbf{x})$ into a feasibility system.
+Namely, $\mathbf{g}=B^{\top}\bm{\pi}$ for a dual optimal solution if and only if
+$B^{\top}\bm{\pi}$ $=\mathbf{g}$ (9)
+$A_{i}^{\top}\bm{\pi}$ $=c_{i}\,,\quad\forall i:z_{i}>0$
+$A_{i}^{\top}\bm{\pi}$ $\geq c_{i}\,,\quad\forall i:z_{i}=0$
+$\pi_{j}$ $=0\,,\quad\forall j:(A\mathbf{z})_{j}<(B\mathbf{x}+\mathbf{b})_{j}$
+$\pi_{j}$ $\geq 0,\quad\forall j:(A\mathbf{z})_{j}=(B\mathbf{x}+\mathbf{b})_{j}\,.$
+Here, $A_{i}$ denotes the $i$-th column of $A$. By Farkas’s lemma, if (9) is not feasible, then the following system has a feasible solution:
+$(A\mathbf{w})_{j}+(B\mathbf{r})_{j}$ $\leq 0\,,\quad\forall j:(A\mathbf{z})_{j}=(B\mathbf{x}+\mathbf{b})_{j}$
+$w_{i}$ $\geq 0\,,\quad\forall i:z_{i}=0$
+$\mathbf{c}^{\top}\mathbf{w}+\mathbf{g}^{\top}\mathbf{r}$ $>0\,.$
+For a small enough $\varepsilon>0$, $\mathbf{z}+\varepsilon\mathbf{w}$ is feasible to the system $\operatorname{Primal}(\mathbf{x}-\varepsilon\mathbf{r})$.
+Hence, for $\mathbf{t}=\mathbf{x}-\varepsilon\mathbf{r}$, $\varphi(\mathbf{t})\geq\mathbf{c}^{\top}(\mathbf{z}+\varepsilon\mathbf{w})=\mathbf{c}^{\top}\mathbf{z}+\varepsilon\mathbf{c}^{\top}\mathbf{w}$.
+Using $\mathbf{c}^{\top}\mathbf{z}=\varphi(\mathbf{x})$ by the choice of $\mathbf{z}$ as an optimal solution to $\operatorname{Primal}(\mathbf{x})$, and the bound $\mathbf{c}^{\top}\mathbf{w}>-\mathbf{g}^{\top}\mathbf{r}$, we get $\varphi(\mathbf{t})>\varphi(\mathbf{x})-\varepsilon\mathbf{g}^{\top}\mathbf{r}=\varphi(\mathbf{x})+\left\langle\mathbf{g},\mathbf{t}-\mathbf{x}\right\rangle$, a contradiction to $\mathbf{g}\in\partial^{\wedge}\varphi(\mathbf{x})$.
+∎
+Consider a generalized flow instance defined in a graph $G=(V,E)$, sink $t\in V$, capacities $\mathbf{h}\in\mathbb{R}_{\geq 0}^{E}$, gain factors $\bm{\gamma}\in\mathbb{R}_{>0}^{E}$ and supply set $S={\mathcal{M}}$.
+For the rest of this section, let $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ be a generalized network utility function defined in the form $\operatorname{GenflowLP}(\mathbf{x})$ in Section 2.
+Using Lemma D.1 and complementary slackness, we get the following characterization of subgradients of generalized network utility functions.
+Note that for a generalized network utility function, the domain is $\mathcal{D}=\mathbb{R}_{\geq 0}^{\mathcal{M}}$, because $\mathbf{f}=0$ is always a feasible solution to $\operatorname{GenflowLP}(\mathbf{x})$, and the definition assumes that the objective is always bounded.
+Lemma D.2.
+Let $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{+}$ be a generalized network utility function defined in the form $\operatorname{GenflowLP}(\mathbf{x})$ for $S={\mathcal{M}}$.
+For a given supply vector $\mathbf{x}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, let $\mathbf{f}$ be an optimal solution to $\operatorname{GenflowLP}(\mathbf{x})$.
+Then, $\mathbf{g}\in\partial^{\wedge}u(\mathbf{x})$ if and only if $\mathbf{g}=\bm{\pi}|_{\mathcal{M}}$ for a vector $\bm{\pi}\in\mathbb{R}_{\geq 0}^{V}$ that satisfies
+$\gamma_{e}\pi_{w}-\pi_{v}$ $=0\,,\quad\forall e=(v,w)\in E:h_{e}>f_{e}>0\,$, (10)
+$\gamma_{e}\pi_{w}-\pi_{v}$ $\leq 0\,,\quad\forall e=(v,w)\in E:f_{e}=0\,$,
+$\gamma_{e}\pi_{w}-\pi_{v}$ $\geq 0\,,\quad\forall e=(v,w)\in E:f_{e}=h_{e}\,$,
+$\pi_{v}$ $=0\,,\quad\forall v\in{\mathcal{M}}:\operatorname{net}_{\mathbf{f}}(v)>-x_{v}\mbox{ or }v\in V\setminus({\mathcal{M}}\cup t):\operatorname{net}_{\mathbf{f}}(v)>0\,$,
+$\pi_{t}$ $=1\,$,
+$\bm{\pi}$ $\geq\mathbf{0}\,.$
+In accordance with this lemma, we say that $\mathbf{f}\in\mathbb{R}^{E}$ and $\bm{\pi}\in\mathbb{R}^{V}$ form a fitting pair with respect to $\mathbf{x}$, if $\mathbf{f}$ is a feasible generalized flow with supply $\mathbf{x}$, and $\bm{\pi}$ satisfies (10).
+Note that this implies that $\mathbf{f}$ and $\bm{\pi}$ are primal and dual optimal solutions to $\operatorname{GenflowLP}(\mathbf{x})$.
+Further, $\mathbf{g}\in\partial^{\wedge}u(\mathbf{x})$ if and only if $\mathbf{g}$ is the restriction of $\bm{\pi}$ to $\mathcal{M}$} for a fitting pair $(\mathbf{f},\bm{\pi})$ with respect to $\mathbf{x}$.
+For every arc $e=(v,w)\in E$, let us define the reverse arc $\overleftarrow{e}=(v,w)$ with gain factor $\gamma_{\overleftarrow{e}}=1/\gamma_{e}$.
+By sending $\alpha$ units of flow on $\overleftarrow{e}$ we mean decreasing the flow value on $e$ by $\gamma_{e}\alpha$.
+Given a fitting pair $(\mathbf{f},\bm{\pi})$ with respect to $\mathbf{x}$, we define the auxiliary network $G^{\textrm{aux}}_{\mathbf{f},\bm{\pi}}=(V,E^{\textrm{aux}}_{\mathbf{f},\bm{\pi}})$ by
+$E^{\textrm{aux}}_{\mathbf{f},\bm{\pi}}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\left\{e=(v,w)\in E\,:\,\gamma_{e}\pi_{w}-\pi_{v}=0\mbox{ and }f_{e}<h_{e}\right\}\cup\left\{\overleftarrow{e}\,:\,e\in E\,:\,\gamma_{e}\pi_{w}-\pi_{v}=0\mbox{ and }0<f_{e}\right\}\,.$
+In the proof of Theorem 1.5, we use two flow subroutines in the generalized flow network.
+The first one starts with a fitting pair $(\mathbf{f},\bm{\pi})$ with respect to a demand vector $\mathbf{x}$, and outputs a flow $\mathbf{f}^{\prime}$ for a decreased demand vector $\mathbf{x}^{\prime}$ such that the flow value $\operatorname{net}_{\mathbf{f}^{\prime}}(t)$ is decreased to the smallest possible amount while maintaining $(\mathbf{f}^{\prime},\bm{\pi})$ as a fitting pair.
+Input: $\mathbf{x}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, and a fitting pair $(\mathbf{f},\bm{\pi})$ with respect to $\mathbf{x}$
+Output: A supply vector $\mathbf{0}\leq\mathbf{x}^{\prime}\leq\mathbf{x}$ and a feasible generalized flow $\mathbf{f}^{\prime}$ such that $(\mathbf{f}^{\prime},\bm{\pi})$ is a fitting pair with respect to $\mathbf{x}^{\prime}$, and subject to this, $\operatorname{net}_{\mathbf{f}^{\prime}}(t)$ is minimal.
+Algorithm 1 Pullback-Flow
+Lemma D.3.
+Pullback-Flow can be formulated as a linear program.
+For the output flow $\mathbf{f}^{\prime}$, the auxiliary graph $E^{\textrm{aux}}_{\mathbf{f}^{\prime},\bm{\pi}}$ does not contain any directed paths from the sink $t$ to $\operatorname{supp}(\mathbf{x}^{\prime})$.
+Proof.
+When fixing $\bm{\pi}$, the requirement that $(\mathbf{f}^{\prime},\bm{\pi})$ be a fitting pair can be formulated by linear constraints on $\mathbf{f}^{\prime}$.
+For example, for every arc $e=(v,w)$, if $\gamma_{e}\pi_{w}-\pi_{v}<0$ then we require $f^{\prime}_{e}=0$, and if $\gamma_{e}\pi_{w}-\pi_{v}>0$ then $f^{\prime}_{e}=h_{e}$; if $\gamma_{e}\pi_{w}-\pi_{v}=0$ then any value $0\leq f^{\prime}_{e}\leq h_{e}$ is allowed.
+For the second part, assume for a contradiction that $E^{\textrm{aux}}_{\mathbf{f}^{\prime},\bm{\pi}}$ contains a directed path $P$ from $t$ to a node $j\in{\mathcal{M}}$ with $x^{\prime}_{j}>0$.
+Then, we can send a small amount of flow along this path by maintaining the flow balance on every intermediate node, and decreasing $\operatorname{net}_{\mathbf{f}^{\prime}}(t)$ and $x^{\prime}_{j}$ by positive amounts.
+This contradicts the minimality of $\operatorname{net}_{\mathbf{f}^{\prime}}(t)$.
+∎
+Remark D.4.
+Pullback-Flow can be in fact formulated as a (classical) maximum flow problem.
+To see this, one can use the standard generalized flow relabeling technique, first introduced by Glover Glover and Klingman (1973), see also e.g., Goldberg et al.
+(1991), Olver and Végh (2020)).
+That is, we multiply the flow at node $v$ by $\pi_{v}$ and replacing $\gamma_{e}$ by $\gamma_{e}\pi_{w}/\pi_{v}$ for every $e=(v,w)\in E$.
+On every arc, we fix $f^{\prime}_{e}$ to the upper or lower bound unless $\gamma_{e}\pi_{w}-\pi_{v}=0$, in which case the relabeled gain factor is 1.
+Our second subroutine aims to increase the supply at a designated node $k\in{\mathcal{M}}$ while the supply at all other goods is non-increasing, and the flow value $\operatorname{net}_{\mathbf{f}}(t)$ is non-decreasing.
+The goal is again to maintain a fitting pair $(\mathbf{f}^{\prime},\bm{\pi})$ for a demand vector $\mathbf{x}^{\prime}$.
+Input: $\mathbf{x}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, and a fitting pair $(\mathbf{f},\bm{\pi})$ with respect to $\mathbf{x}$, and a good $k\in{\mathcal{M}}$ such that $\pi_{k}>0$.
+Output: A supply vector $\mathbf{x}^{\prime}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $x^{\prime}_{k}\geq x_{k}$ and $x^{\prime}_{j}\leq x_{j}$ if $j\in{\mathcal{M}}\setminus\{k\}$, and a feasible generalized flow $\mathbf{f}^{\prime}$ such that $(\mathbf{f}^{\prime},\bm{\pi})$ is a fitting pair with respect to $\mathbf{x}^{\prime}$, $\operatorname{net}_{\mathbf{f}}(t^{\prime})\geq\operatorname{net}_{\mathbf{f}}(t)$, and subject to this, $x^{\prime}_{k}$ is maximal.
+Algorithm 2 Reroute-Flow
+The following lemma can be derived similarly to Lemma D.3.
+Lemma D.5.
+Reroute-Flow can be formulated as a linear program.
+For the output flow $\mathbf{f}^{\prime}$, the auxiliary graph $E^{\textrm{aux}}_{\mathbf{f}^{\prime},\bm{\pi}}$ does not contain any directed path from $k$ to the set $(\operatorname{supp}(\mathbf{x}^{\prime})\setminus\{k\})\cup\{t\}$.
+From Lemma 4.2, we get the following.
+Lemma D.6.
+Consider a generalized network utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$, $b>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$, and $\mathbf{x}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$.
+Then, $\mathbf{x}\in\mathcal{GD}^{u}(\mathbf{q},b)$ if and only if there exists a fitting pair $(\mathbf{f},\bm{\pi})$ with respect to $\mathbf{x}$ such that $u(\mathbf{x})=\operatorname{net}_{\mathbf{f}}(t)$ and
+$\pi_{j}\leq\frac{q_{j}}{b}\cdot u(\mathbf{x})\,\quad\forall j\in{\mathcal{M}}\,,\,\text{ with equality for }j\in\operatorname{supp}(\mathbf{x})\,.$ (11)
+Theorem 1.5 follows by showing the two properties in Definition 2.14 in Lemmas D.7 and Lemma D.8 below. We start by property (ii).
+Lemma D.7.
+Let $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ be a generalized network utility function, $b>b^{\prime}>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$, and let $\mathbf{x}\in\mathcal{GD}^{u}(\mathbf{q},b)$.
+Then, there exists $\mathbf{x}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$ such that $\mathbf{x}^{\prime}\leq\mathbf{x}$.
+Proof.
+The proof is given by a finite algorithm, shown as Algorithm 3.
+We start from fitting pair $(\mathbf{f},\bm{\pi})$ with respect to $\mathbf{x}$ as in Lemma D.6.
+Thus, the inequality (11) holds for $\bm{\pi}$, $\mathbf{x}$, $\mathbf{q}$ and $b$.
+If $u(\mathbf{x})=0$, then we must have $\pi_{j}=0$ for all $j\in{\mathcal{M}}$ and thus (11) holds also when replacing $b$ by $b^{\prime}$, and therefore we can return $\mathbf{x}^{\prime}=\mathbf{x}$.
+For the rest of the proof, let us assume $u(\mathbf{x})>0$.
+Similarly, we can assume that $\mathbf{q}\neq\mathbf{0}$.
+Note that $\pi_{j}>0$ for every $j\in\operatorname{supp}(\mathbf{x})\cap\operatorname{supp}(\mathbf{q})$.
+When replacing $b$ by $b^{\prime}$, all inequalities with in (11) with $q_{j}>0$ become strict; the ratio of the right and left hand sides will be $\alpha\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}b/b^{\prime}>1$ for all $j\in\operatorname{supp}(\mathbf{x})\cap\operatorname{supp}(\mathbf{q})$.
+Also, recall $u(\mathbf{x})=\operatorname{net}_{\mathbf{f}}(t)$.
+Algorithm 3 alternates between two operations: changing $\mathbf{x}$ and $\mathbf{f}$ while keeping $\bm{\pi}$ unchanged; and changing the dual $\bm{\pi}$ while keeping $\mathbf{x}$ and $\mathbf{f}$ unchanged.
+Throughout, it maintains a fitting pair with respect to a decreasing vector $\mathbf{x}^{\prime}$, such that the required equalities in (11) are gradually restored.
+In the first step, we call $\textnormal{{Pullback-Flow}}(\mathbf{x},\mathbf{f},\bm{\pi})$ to get demands $\mathbf{x}^{\prime}\leq\mathbf{x}$ and a corresponding fitting pair $(\mathbf{f}^{\prime},\bm{\pi})$.
+We have $u(\mathbf{x}^{\prime})=\operatorname{net}_{\mathbf{f}^{\prime}}(t)\leq\operatorname{net}_{\mathbf{f}}(t)=u(\mathbf{x})$.
+Note that $\frac{q_{j}}{b^{\prime}}u(\mathbf{x}^{\prime})/\pi_{j}$ is the same value $\alpha^{\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\alpha u(\mathbf{x}^{\prime})/u(\mathbf{x})$ for every $j\in\operatorname{supp}(\mathbf{x}^{\prime})\cap\operatorname{supp}(\mathbf{q})$.
+Case I:
+If $\alpha^{\prime}\leq 1$, then we choose $\lambda\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}(\alpha-1)/(\alpha-\alpha^{\prime})\in[0,1]$ such that $\lambda\alpha+(1-\lambda)\alpha^{\prime}=1$.
+Then, $\mathbf{x}^{\prime\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\lambda\mathbf{x}^{\prime}+(1-\lambda)\mathbf{x}$ and the flow $\mathbf{f}^{\prime\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\lambda\mathbf{f}^{\prime}+(1-\lambda)\mathbf{f}$ satisfy that $(\mathbf{f}^{\prime\prime},\bm{\pi})$ is a fitting pair with respect to $\mathbf{x}^{\prime\prime}$, and (11) holds for $\bm{\pi}$, $\mathbf{x}^{\prime\prime}$ and $b^{\prime}$.
+Thus, we can terminate by outputting $\mathbf{x}^{\prime\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$.
+Case II:
+If $\alpha^{\prime}>1$, then we update the dual $\bm{\pi}$ to $\bm{\pi}^{\prime}$ as follows.
+Let $T\supseteq\operatorname{supp}(\mathbf{x}^{\prime})$ be the set of nodes that can reach $\operatorname{supp}(\mathbf{x}^{\prime})$ in the auxiliary graph $(G,E^{\textrm{aux}}_{\mathbf{f}^{\prime},\bm{\pi}})$.
+By Lemma D.3, $t\notin T$.
+Let us define $\bm{\pi}^{\prime}$ such that $\pi^{\prime}_{v}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\pi_{v}$ for $v\notin T$ and $\pi^{\prime}_{v}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\delta^{\prime}\pi_{v}$ for $v\in T$ for the largest value $1\leq\delta^{\prime}\leq\alpha^{\prime}$ such that $(\mathbf{f}^{\prime},\bm{\pi}^{\prime})$ is a fitting pair.
+We claim that $\delta^{\prime}>1$.
+Recall that fitting pairs are required to satisfy (10).
+The sign of $\gamma_{e}\pi^{\prime}_{w}-\pi^{\prime}_{v}$ is the same as the sign of $\gamma_{e}\pi_{w}-\pi_{v}$ for all arcs $e=(v,w)$ with both endpoints inside $T$ or outside $T$.
+There is no arc $e=(v,w)$ entering or leaving $T$ with $0<f^{\prime}_{e}<u_{e}$, since $e,\overleftarrow{e}\in E^{\textrm{aux}}_{\mathbf{f}^{\prime},\bm{\pi}}$ for all such arcs.
+For every arc $e=(v,w)$ entering $T$ with $f^{\prime}_{e}=0$, we must have $\gamma_{e}\pi_{w}-\pi_{v}<0$ so that $e\notin E^{\textrm{aux}}_{\mathbf{f}^{\prime},\bm{\pi}}$.
+Similarly, for every arc $e=(v,w)$ leaving $T$ with $f^{\prime}_{e}=u_{e}$, we must have $\gamma_{e}\pi_{w}-\pi_{v}>0$.
+Hence, we can always choose $\delta^{\prime}>1$.
+If $\delta^{\prime}=\alpha^{\prime}$, then $\bm{\pi}^{\prime}$, $\mathbf{x}^{\prime}$, $\mathbf{q}$ and $b^{\prime}$ satisfy (11), since the right hand side for all $j\in\operatorname{supp}(\mathbf{x}^{\prime})$ has increased by a factor $\alpha^{\prime}$, and therefore equality holds for all these indices.
+We can therefore conclude $\mathbf{x}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b)$ and terminate by outputting $\mathbf{x}^{\prime}$.
+In case $\delta^{\prime}<\alpha^{\prime}$, then (11) holds with strict inequality for all $j$ for $\bm{\pi}^{\prime}$, $\mathbf{x}^{\prime}$, $\mathbf{q}$, and $b^{\prime}$.
+Moreover, the ratio of the right hand side and the left hand side is the same for all $j\in\operatorname{supp}(\mathbf{x}^{\prime})\cap\operatorname{supp}(\mathbf{q})$, namely, $\alpha^{\prime}/\delta^{\prime}<\alpha$.
+We now restart the above procedure with $(\mathbf{f}^{\prime},\bm{\pi}^{\prime})$ and $\mathbf{x}^{\prime}$ in place of $(\mathbf{f},\bm{\pi})$ and $\mathbf{x}$.
+We claim that the algorithm terminates in a finite number of iterations.
+This holds because the linear program Pullback-Flow only depends on the sign pattern of $\gamma_{e}\pi_{w}-\pi_{v}$ for the arcs $e=(w,v)$ and on $\operatorname{supp}(\mathbf{x})$.
+There are $3^{|E|}$ possible sign patterns; the set $\operatorname{supp}(\mathbf{x})$ is monotone decreasing and therefore there are at most $|{\mathcal{M}}|$ different possibilities.
+The same configuration of arc sign patterns and $\operatorname{supp}(\mathbf{x})$ cannot occur twice, because $u(\mathbf{x}^{\prime})=\operatorname{net}_{\mathbf{f}^{\prime}}(t)$ is strictly decreasing in every iteration of the algorithm.
+Hence, after a finite number of iterations we arrive at $\bm{\pi}^{\prime}$ and $\mathbf{x}^{\prime}$ satisfying (11) for $b^{\prime}$.
+Since $\mathbf{x}^{\prime}$ is decreasing throughout, we have $\mathbf{x}^{\prime}\leq\mathbf{x}$ for the input supply vector $\mathbf{x}$, as required.
+∎
+Input: $\mathbf{q}\in\operatorname{pdom}(u)$, $b,b^{\prime}>0$, $\mathbf{x}\in\mathcal{GD}^{u}(\mathbf{q},b)$.
+Output: $\mathbf{0}\leq\mathbf{x}^{\prime}\leq\mathbf{x}$ such that $\mathbf{x}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$.
+1 Find a fitting pair $(\mathbf{f},\bm{\pi})$ with respect to $\mathbf{x}$ such that $u(\mathbf{x})=\operatorname{net}_{\mathbf{f}}(t)$ and (11) holds;
+2 if $u(\mathbf{x})=0$ or $\mathbf{q}=\mathbf{0}$ then return $\mathbf{x}$;
+3 $\bar{\mathbf{x}}\leftarrow\mathbf{x}$ ; $\alpha\leftarrow b/b^{\prime}$ ;
+4 while $\alpha>1$ do
+5 $(\mathbf{x}^{\prime},\mathbf{f}^{\prime})\leftarrow\textnormal{{Pullback-Flow}}(\bar{\mathbf{x}},\mathbf{f},\bm{\pi})$ ;
+6 $\alpha^{\prime}\leftarrow\alpha\frac{u(\mathbf{x}^{\prime})}{u(\mathbf{x})}$ ;
+7 if $\alpha^{\prime}\leq 1$ then
+8    $\lambda\leftarrow\frac{\alpha-1}{\alpha-\alpha^{\prime}}$ ;
+9    $\mathbf{x}^{\prime\prime}\leftarrow\frac{\alpha-1}{\alpha-\alpha^{\prime}}\mathbf{x}^{\prime}+\frac{1-\alpha^{\prime}}{\alpha-\alpha^{\prime}}\bar{\mathbf{x}}$ ;
+10    return $\mathbf{x}^{\prime\prime}$
+11 else
+12    $T\leftarrow\left\{v\in V\,:\,v\text{ can reach }\operatorname{supp}(\mathbf{x}^{\prime})\text{ in }(G,E^{\textrm{aux}}_{\mathbf{f}^{\prime},\bm{\pi}})\right\}$  ;
+13    $\delta^{\prime}\leftarrow\min\left\{\alpha,\min\left\{\frac{\gamma_{e}\pi_{w}}{\pi_{v}}\,:\,e=(v,w)\in\partial^{\mathrm{in}}(T)\right\},\min\left\{\frac{\pi_{v}}{\gamma_{e}\pi_{w}}\,:\,e=(v,w)\in\partial^{\mathrm{out}}(T)\right\}\right\}$;
+14    if $\delta^{\prime}=\alpha^{\prime}$ then return $\mathbf{x}^{\prime}$;
+15    for $v\in T$ do $\pi_{v}\leftarrow\delta^{\prime}\pi_{v}$;
+16    $\alpha\leftarrow\alpha^{\prime}/\delta^{\prime}$ ; $\bar{\mathbf{x}}\leftarrow\mathbf{x}^{\prime}$ ;
+17
+18
+Algorithm 3 Budget-Decrease
+We next prove property (i) in Definition 2.14.
+Lemma D.8.
+Let $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ be a generalized network utility function and $\mathbf{q},\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$ such that $\mathbf{0}\leq\mathbf{q}^{\prime}\leq\mathbf{q}$, and let $\mathbf{x}\in\mathcal{GD}^{u}(\mathbf{q},b)$.
+Then, there exists $\mathbf{x}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ such that $x^{\prime}_{j}\leq x_{j}$ whenever $q^{\prime}_{j}=q_{j}$.
+Proof.
+Clearly, it suffices to show that the statement holds when $q^{\prime}_{k}<q_{k}$ for a single good $k\in{\mathcal{M}}$.
+Again, we start from fitting pair $(\mathbf{f},\bm{\pi})$ with respect to $\mathbf{x}$ as in Lemma D.6.
+Similarly to Lemma D.7, we may assume that $u(\mathbf{x})>0$ and $\mathbf{q}\neq\mathbf{0}$.
+Let us now replace $\mathbf{q}$ by $\mathbf{q}^{\prime}$.
+The condition (11) can only be violated for $j=k$.
+If it is not violated, then we may return $\mathbf{x}^{\prime}=\mathbf{x}$; this is always the case when $\pi_{k}=0$.
+For the rest of the proof, assume that the (11) is violated for $k$ with the cost $q^{\prime}_{k}$.
+Note that $\pi_{j}=\beta q^{\prime}_{j}$ for every $j\neq k$, where $\beta\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}u(\mathbf{x})/b$; we now have $\pi_{k}>\beta q^{\prime}_{k}\geq 0$.
+This value of $\beta$ will be fixed throughout the argument.
+We first consider the case $q^{\prime}_{k}>0$.
+The proof uses the subroutine $\textnormal{{Reroute-Flow}}(\mathbf{x},\mathbf{f},\bm{\pi})$ with the special node $k$.
+This subroutine finds $\mathbf{x}^{\prime}$ and $\mathbf{f}^{\prime}$ such that $(\mathbf{f}^{\prime},\bm{\pi})$ is a fitting pair with respect to $\mathbf{x}^{\prime}$, and further $x^{\prime}_{k}\geq x_{k}$, $x^{\prime}_{j}\leq x_{j}$ for $j\neq k$, and $u(\mathbf{x}^{\prime})=\operatorname{net}_{\mathbf{f}}(t^{\prime})\geq\operatorname{net}_{\mathbf{f}}(t)=u(\mathbf{x})$.
+Consider the set of nodes $T$ reachable from $k$ in $E^{\textrm{aux}}_{\mathbf{f}^{\prime},\bm{\pi}}$.
+According to Lemma D.5, $t\notin T$ and $T\cap\operatorname{supp}(\mathbf{x}^{\prime})=\{k\}$.
+Similarly as Case II in the proof of Lemma D.7, we can decrease $\pi_{v}$ multiplicatively on $T$ to obtain $\bm{\pi}^{\prime}$ such that $(\mathbf{f}^{\prime},\bm{\pi}^{\prime})$ is a fitting pair with respect to $\mathbf{x}^{\prime}$.
+We choose the largest multiplicative decrease such that $(\mathbf{f}^{\prime},\bm{\pi}^{\prime})$ remains a fitting pair, and that $\pi_{k}^{\prime}\geq\beta q_{k}^{\prime}$ is maintained.
+We then repeat this procedure for $\mathbf{x}^{\prime},\mathbf{f}^{\prime}$, and $\bm{\pi}^{\prime}$ in the place of $\mathbf{x},\mathbf{f}$, and $\bm{\pi}$.
+We can again argue that this algorithm terminates in a finite number of iterations.
+Note that $\operatorname{supp}(\mathbf{x}^{\prime})$ is monotone decreasing throughout, and $\pi^{\prime}_{j}=\beta q^{\prime}_{j}$ is maintained for all $j\in\operatorname{supp}(\mathbf{x}^{\prime})\setminus\{k\}$.
+For goods with $x^{\prime}_{j}=0$, $\pi^{\prime}_{j}$ may decrease, and therefore $\pi^{\prime}_{j}\leq\beta q_{j}$ for these goods.
+At termination, we have $\pi^{\prime}_{j}=\beta q^{\prime}_{j}$ for all $j\in\operatorname{supp}(\mathbf{x}^{\prime})$ — including $j=k$ — and $\pi^{\prime}_{j}\leq\beta q^{\prime}_{j}$ whenever $x^{\prime}_{j}=0$.
+Moreover, note that $x^{\prime}_{j}\leq x_{j}$ for all $j\neq k$.
+[…]
+Consider now the case $q^{\prime}_{k}=0$.
+We proceed as above, but our goal is to eventually set $\pi^{\prime}_{k}=0$.
+As in the previous case, we repeatedly call $\textnormal{{Reroute-Flow}}(\mathbf{x},\mathbf{f},\bm{\pi})$ and then decrease $\pi^{\prime}_{v}$ multiplicatively on $T$ by the largest possible factor such that $(\mathbf{f}^{\prime},\bm{\pi}^{\prime})$ remains a fitting pair.
+If this largest factor is unbounded, then we can set $\pi^{\prime}_{v}=0$ for all $v\in T$.
+Again, the algorithm should terminate in a finite number of iterations, using also the assumption that $\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$.
+Once $\pi^{\prime}_{k}=0$ is reached, we again have $\pi^{\prime}_{j}=\beta q^{\prime}_{j}$ for all $j\in\operatorname{supp}(\mathbf{x}^{\prime})$, including $j=k$, and $\pi^{\prime}_{j}\leq\beta q^{\prime}_{j}$ whenever $x^{\prime}_{j}=0$.
+The rest of the proof is the same as in the case $q^{\prime}_{k}>0$.
+∎
+This concludes the proof of Theorem 1.5, namely, that all generalized network utilities are Gale-substitutes.
+Nevertheless, they are not $\Sigma$-Gale-substitutes, and so Theorem 1.3 is only applicable in the case when the generalized network utilities are non-satiable.
+We conclude by showing Theorem 1.6.
+This is a modification of the argument in Section B; we omit some simple details.
+Proof of Theorem 1.6.
+In the proof of Theorem 1.3 in Section B, property ($\Sigma$) is only used in the proof of Lemma 5.2.
+We show that for generalized network utilities, if $\mathbf{x}\in{\mathcal{D}}^{u}(\mathbf{p},b)$ for $\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ and further $\mathbf{x}$ minimizes $\left\langle\mathbf{p},\mathbf{x}\right\rangle$ in ${\mathcal{D}}^{u}(\mathbf{p},b)$, then there exists $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{p},b)$ such that $\mathbf{y}\leq\mathbf{x}$.
+From this property, we can derive Lemma B.1 for thrifty competitive equilibria; the rest of the proof remains identical to the proof of Theorem 1.3.
+Consider now (6) in the proof of Lemma 5.2.
+If $\lambda>0$, the same argument is applicable.
+Consider now the case $\lambda=0$.
+This means that the constraint $\left\langle\mathbf{p},\mathbf{x}\right\rangle\leq b$ is not active in the program defining ${\mathcal{D}}^{u}(\mathbf{p},b)$; that is, $\max_{\mathbf{z}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\{u(\mathbf{z})\,:\,\left\langle\mathbf{p},\mathbf{z}\right\rangle\leq b\}=\max_{\mathbf{z}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}u(\mathbf{z})=v^{\star}$.
+In particular, $u(\mathbf{x})=v^{\star}$.
+The budget minimization property means that $\mathbf{x}$ is an optimal solution to the program
+[…]
+$\operatorname{net}_{\mathbf{f}}(t)$ $\geq v^{\star}$
+$\operatorname{net}_{\mathbf{f}}(v)$ $\geq-x_{v}\quad\forall v\in S$
+$\operatorname{net}_{\mathbf{f}}(v)$ $\geq 0\quad\forall v\in V\setminus(S\cup\{t\})$
+$\mathbf{0}$ $\leq\mathbf{f}\leq\mathbf{h}\,.$
+From here, using LP duality and Lemma D.1, we can see that there exists a supergradient $\mathbf{g}\in\partial^{\wedge}u(\mathbf{x})$ and $\mu\in\mathbb{R}_{\geq 0}$ such that
+$\mu g_{j}\leq p_{j}\,,\quad\forall j\in{\mathcal{M}}\,,\,\mbox{with equality if }x_{j}>0\,.$
+If $\mu>0$, then we can use the same argument for $\lambda=1/\mu$.
+If $\mu=0$, given that $\mu$ acts as the dual variable of $\operatorname{net}_{\mathbf{f}}(t)\geq v^{\star}$, which is no longer active, the optimal value of (12) becomes 0 since $\mathbf{0}$ is a feasible allocation of $u(\cdot)$ according to our definition.
+This implies the existence of a zero-cost optimal allocation $\mathbf{x}$: $u(\mathbf{x})=v^{*}$ and $\mathbf{p}^{\top}\mathbf{x}=0$, which indicates that $\mathbf{x}$ belongs to both ${\mathcal{D}}^{u}(\mathbf{p},b)$ and $\mathcal{GD}^{u}(\mathbf{p},b)$.
+Thus, the result follows.
+∎
+Appendix E Examples
 In this section, we present multiple examples mentioned in the paper.
-A.1 Connectivity of equilibria with bounded linear utilities
+E.1 Connectivity of equilibria with bounded linear utilities
 We first investigate the (dis)connectivity of competitive equilibria with bounded linear utilities.
-We begin by demonstrating that the thrifty competitive equilibria with bounded linear utilities might not be connected.
-Example A.1 (Disconnectivity of thrifty competitive equilibrium).
-There are two agents and two goods.
-Each agent has a unit budget.
-The first agent’s utility is $u_{1}(x_{1})=1.3\min\{x_{11},0.8\}+0.45x_{12}$ and the second agent’s utility is $u_{2}(x_{2})=0.01\min\{x_{21},0.3\}+2\min\{x_{22},0.8\}$.
-First, we show there are at least two equilibria.
-The first equilibrium is with $\mathbf{p}=(1,1)$.
-Under this price, $x_{11}=0.8$, $x_{12}=0.2$, $x_{21}=0.2$, $x_{22}=0.8$.
-Another competitive equilibrium is with $\mathbf{p}=(1.3,0.45)$.
-Under this price, $x_{11}=0.7$, $x_{12}=0.2$, $x_{21}=0.3$, $x_{22}=0.8$.
-Now, we show the second competitive equilibrium is isolated under the thrifty assumption.
+Our initial example illustrates that the set of thrifty competitive equilibria may be disconnected, even when agents have bounded linear utility functions, where “thrifty” denotes that, given the prices, each agent selects the least costly bundle from the set of utility-maximizing bundles.
+Example E.1 (Disconnectivity of thrifty competitive equilibrium).
+There are two agents and two goods. Each agent has a unit budget. The agents’ utility functions are
+[…]
+[…]
+First, we show that there are at least two equilibria. The first equilibrium is
+$p_{1}=1\,,\quad p_{2}=1\,,\quad x_{11}=0.8\,,\quad x_{12}=0.2\,,\quad x_{21}=0.2\,,\quad x_{22}=0.8\,.$
+Another competitive equilibrium is
+$p_{1}=1.3\,,\quad p_{2}=0.45\,,\quad x_{11}=0.7\,,\quad x_{12}=0.2\,,\quad x_{21}=0.3\,,\quad x_{22}=0.8\,.$
+We show the second competitive equilibrium is isolated under the thrifty assumption.
 First, we show the price is isolated.
-Suppose there is a nearby price $\mathbf{p}^{\prime}=(1.3+\epsilon_{1},0.45+\epsilon_{2})$.
+Suppose there is a nearby equilibrium price $\mathbf{p}^{\prime}=(1.3+\varepsilon_{1},0.45+\varepsilon_{2})$.
 For this price, the thrifty demand of the second agent remains the same: $x_{21}=0.3$, $x_{22}=0.8$.
 To maintain equilibrium, the first agent should receive the same allocation.
-As $1.3\times 0.7+0.45\times 0.2=1$, $\epsilon_{1}$ and $\epsilon_{2}$ can not be both positive or negative.
-However, if one $\epsilon$ is negative, the first agent will prefer this item more at $\mathbf{p}^{\prime}$.
-This implies the allocation of the other item of the first agent is 00.
+As $1.3\times 0.7+0.45\times 0.2=1$, $\varepsilon_{1}$ and $\varepsilon_{2}$ cannot be both positive or negative.
+However, if $\varepsilon_{j}<0$, the first agent will prefer good $j$ more at $\mathbf{p}^{\prime}$.
+This implies the allocation of the other good of the first agent is 0.
 This leads to a contradiction.
 Second, we show the allocation is isolated.
-Suppose there is a nearby equilibrium allocation of the second agent $x^{\prime}_{2}=(x_{21}-\epsilon_{1},x_{22}-\epsilon_{2})$ with at least one of $\epsilon_{1}$ and $\epsilon_{2}$ is strictly positive.
-In this case, both agents should spend all their budgets as they are non-satiable.
+Note that the second agent is currently satiated.
+Suppose there is a nearby equilibrium allocation.
+Agent 2 must have a different allocation, which can only be smaller.
+Thus, $\mathbf{x}^{\prime}_{2}=(x_{21}-\varepsilon_{1},x_{22}-\varepsilon_{2})$ with $\varepsilon_{1},\varepsilon_{2}\geq 0$, and at least one of them is strictly positive.
+In the allocation $\mathbf{x}^{\prime}$, both agents are non-satiated, and therefore they are spending their full budgets.
 Therefore, the equilibrium prices satisfy $p^{\prime}_{1}+p^{\prime}_{2}=2$.
-For the first agent, the allocation is $x^{\prime}_{11}=0.7+\epsilon_{1}$, $x^{\prime}_{12}=0.2+\epsilon_{2}$.
-Note that both $\epsilon_{1}$ and $\epsilon_{2}$ are very tiny.
-This implies $p^{\prime}_{1}:p^{\prime}_{2}=1.3:0.45$.
-Therefore, $p^{\prime}_{1}=\frac{2}{1.3+0.45}\times 1.3$ and $p^{\prime}_{2}=\frac{2}{1.3+0.45}\times 0.45$.
+For the first agent, the allocation is $x^{\prime}_{11}=0.7+\varepsilon_{1}$, $x^{\prime}_{12}=0.2+\varepsilon_{2}$.
+Note that both $\varepsilon_{1}$ and $\varepsilon_{2}$ are very tiny.
+This implies $p^{\prime}_{1}:p^{\prime}_{2}\approx 1.3:0.45$.
+Therefore, $p^{\prime}_{1}\approx\frac{2}{1.3+0.45}\times 1.3$ and $p^{\prime}_{2}\approx\frac{2}{1.3+0.45}\times 0.45$.
 However, the first agent cannot afford these allocations under these prices.
-The next example shows that competitive equilibria are not connected without the thrifty assumption.
-Example A.2 (Disconnectivity of competitive equilibria).
-There are three agents and three goods.
-Each agent has a unit budget.
-The first agent’s utility is $u_{1}(x_{1})=1.3\min\{x_{11},0.8\}+0.45x_{12}$, the second agent’s utility is $u_{2}(x_{2})=0.01\min\{x_{21},0.3\}+2\min\{x_{22},0.8\}+10^{-10}x_{23}$, and the third agent’s utility is $u_{3}(x_{3})=x_{33}$.
-First, we show there are at least two equilibrium.
-The first equilibrium is with $\mathbf{p}=(1,1,1)$, and the allocation is $x_{1}=(0.8,0.2,0)$, $x_{2}=(0.2,0.8,0)$, and $x_{3}=(0,0,1)$.
-The second equilibrium is with $\mathbf{p}=(1.3,0.45,1.25)$ and the allocation is $x_{1}=(0.7,0.2,0)$, $x_{2}=(0.3,0.8,0.2)$, and $x_{3}=(0,0,0.8)$.
+Nevertheless, it remains possible that thrifty competitive equilibria are connected through intermediate equilibria that are not thrifty.
+However, with a slight modification of the previous example, the next example shows that competitive equilibria are not connected without the thrifty assumption.
+Example E.2 (Disconnectivity of competitive equilibria).
+There are three agents and three goods. Each agent has a unit budget. The utility functions are
+[…]
+[…]
+$u_{3}(\mathbf{x}_{3})$ $=x_{33}\,.$
+First, we show there are at least two equilibria.
+The first equilibrium is with $\mathbf{p}=(1,1,1)$, and the allocation is $\mathbf{x}_{1}=(0.8,0.2,0)$, $\mathbf{x}_{2}=(0.2,0.8,0)$, and $\mathbf{x}_{3}=(0,0,1)$.
+The second equilibrium is with $\mathbf{p}=(1.3,0.45,1.25)$ and the allocation is $\mathbf{x}_{1}=(0.7,0.2,0)$, $\mathbf{x}_{2}=(0.3,0.8,0.2)$, and $\mathbf{x}_{3}=(0,0,0.8)$.
 In order to show our results, we first state the following properties: for any equilibrium,
@@
 First, we demonstrate that the price is isolated.
-Suppose there is a nearby equilibrium price $\mathbf{p}^{\prime}=(1.3+\epsilon_{1},0.45+\epsilon_{2},1.25+\epsilon_{3})$.
-For this price, the allocation of the second agent on the first two items is the same.
+Suppose there is a nearby equilibrium price $\mathbf{p}^{\prime}=(1.3+\varepsilon_{1},0.45+\varepsilon_{2},1.25+\varepsilon_{3})$.
+For this price, the allocation of the second agent on the first two goods is the same.
 Therefore, the first agent should receive the same allocation.
-As $1.3\times 0.7+0.45\times 0.2=1$, $\epsilon_{1}$ and $\epsilon_{2}$ can not be both positive or negative.
-Similar to the previous example, if one of $\epsilon_{1}$ and $\epsilon_{2}$ is negative, then the first agent will prefer this item more at this nearby price.
-This implies the allocation of the other item of the first agent is 00.
-This leads to the conclusion $\epsilon_{1}=\epsilon_{2}=\epsilon_{3}=0$.
+As $1.3\times 0.7+0.45\times 0.2=1$, $\varepsilon_{1}$ and $\varepsilon_{2}$ cannot be both positive or negative.
+Similar to the previous example, if one of $\varepsilon_{1}$ and $\varepsilon_{2}$ is negative, then the first agent will prefer this good more at this nearby price.
+This implies the allocation of the other good of the first agent is 0.
+This leads to the conclusion $\varepsilon_{1}=\varepsilon_{2}=\varepsilon_{3}=0$.
 Second, we show the allocation is isolated.
@@
 Additionally, $p^{\prime}_{1}\times x^{\prime}_{11}+p^{\prime}_{2}\times x^{\prime}_{12}=1$.
-This implies $p_{1}^{\prime}$ is very close to $p_{1}$ and $p_{2}^{\prime}$ is very close to $p_{2}$: if $\|x^{\prime}-x\|_{\infty}\leq\epsilon$, then $|p_{1}^{\prime}-p_{1}|\leq O(\epsilon)$ and $|p_{2}^{\prime}-p_{2}|\leq O(\epsilon)$.
+This implies $p_{1}^{\prime}$ is very close to $p_{1}$ and $p_{2}^{\prime}$ is very close to $p_{2}$: if $\|\mathbf{x}^{\prime}-\mathbf{x}\|_{\infty}\leq\varepsilon$, then $|p_{1}^{\prime}-p_{1}|\leq O(\varepsilon)$ and $|p_{2}^{\prime}-p_{2}|\leq O(\varepsilon)$.
 [^8] Note that $p^{\prime}_{1}:p^{\prime}_{2}=1.3:0.45=p_{1}:p_{2}$.
@@
 We have a similar result for $p_{1}^{\prime}-p_{1}$.
-This also implies $|p_{3}^{\prime}-p_{3}|\leq O(\epsilon)$ as $p^{\prime}_{1}+p^{\prime}_{2}+p^{\prime}_{3}=3$.
+This also implies $|p_{3}^{\prime}-p_{3}|\leq O(\varepsilon)$ as $p^{\prime}_{1}+p^{\prime}_{2}+p^{\prime}_{3}=3$.
 Since price is isolated, $p^{\prime}_{1}=p_{1}$, $p^{\prime}_{2}=p_{2}$, and $p^{\prime}_{3}=p_{3}$.
 Under this price, $\mathbf{x}$ is the only allocation.
-A.2 Non-monotonicity of Gale Demands
+E.2 Non-monotonicity of Gale Demands
 The next example shows the somewhat counterintuitive property that increasing the prices may lead to an increase in the Gale utility.
-Example A.3 (Non-monotonicity of Gale demands).
-There are two goods and a single agent $i$ with $b_{i}=1$.
-The agent’s utility function is $u_{i}(x_{i})=\min\{x_{i1}+x_{i2},2+0.1\cdot x_{i1}+0.2\cdot x_{i2}\}$.
-When $\mathbf{q}=\left(\frac{0.1}{2/0.9+0.2\cdot 0.01},\frac{0.2}{2/0.9+0.2\cdot 0.01}\right)$, the associated Gale demand is $\mathcal{GD}_{i}(b_{i},\mathbf{q})=\{(2/0.9,0.01)\}$ and the utility is $u_{i}(\mathcal{GD}_{i}(b_{i},\mathbf{q}))=2/0.9+0.2\cdot 0.01$.
-When $\mathbf{q}^{\prime}=\left(\frac{1}{0.01+2/0.85},\frac{0.2}{0.01+2/0.85}\right)$, the Gale demand becomes $\mathcal{GD}_{i}(b_{i},\mathbf{q}^{\prime})=\{(0.01,2/0.85)\}$ and $u_{i}(\mathcal{GD}_{i}(b_{i},\mathbf{q}^{\prime}))=2/0.85+0.01$.
-Clearly, $\mathbf{q}<\mathbf{q}^{\prime}$, yet $u_{i}(\mathcal{GD}_{i}(b_{i},\mathbf{q}))<u_{i}(\mathcal{GD}_{i}(b_{i},\mathbf{q}^{\prime}))$.
-A.3 Tight Example on Price of Anarchy
-The next example shows that the bound in Theorem 1.1 is tight.
-Example A.4 (Tight Example on Price of Anarchy).
-We consider the case with two agents and two goods.
-The first agent has a budget of $1$, and desires the first good: $u_{1}(x_{1})=x_{11}$.
-The second agent has a budget of $\mathrm{e}-1$ and has a utility function $u_{2}(x_{2})=\varepsilon x_{21}+\min\{x_{22},1\}$, where $\varepsilon>0$ is sufficiently small.
-There is a competitive equilibrium with prices $p_{1}=\mathrm{e}$ and $p_{2}=0$, and allocations $x_{1}=(\frac{1}{\mathrm{e}},0)$ and $x_{2}=(\frac{\mathrm{e}-1}{\mathrm{e}},1)$.
-The Nash welfare objective for this allocation is $\left(\frac{1}{\mathrm{e}}(1+\varepsilon\cdot\frac{\mathrm{e}-1}{\mathrm{e}})^{\mathrm{e}-1}\right)^{\frac{1}{\mathrm{e}}}$.
-In the allocation maximizing Nash welfare, we have $y_{1}=(1,0)$ and $y_{2}=(0,1)$, where the Nash welfare objective is $1$.
-When $\varepsilon\rightarrow 0$, the difference of these two scenarios converges to $\left(\frac{1}{\mathrm{e}}\right)^{\frac{1}{\mathrm{e}}}$.
-A.4 Utility Difference in Different Equilibria
-The following example shows agents might receive significantly different utilities in different equilibria.
-Example A.5 (Utility Difference in Different Equilibria).
-Consider a scenario involving $n$ agents and $n$ goods, where each agent has a unit budget.
-For $11\leq i\leq n-$ 1, agent $i$’s utility is defined as $u_{i}(x_{i})=\min\{x_{ii},1\}+\epsilon x_{in}$, while agent $n$’s utility is given by $u_{i}(x_{i})=x_{nn}$.
-There exist at least two competitive equilibria.
-In the first competitive equilibrium, all goods are priced at $1$, and each good $i$ is allocated to agent $i$ for $1\leq i\leq n$, resulting in a utility of $1$ for agent $n$.
-However, in the second competitive equilibrium, prices are set to 00 for the first $n-1$ goods and the price of the $n$-th good is $n$.
-Consequently, goods $1$ to $n-1$ are allocated to agents $1$ to $n-1$ respectively, while the $n$-th good is uniformly allocated to all agents.
-In this equilibrium, agent $n$ can only attain a utility of $\approx\frac{1}{n}$.
-A.5 Tight Example on Demand-Approximate CE
-The following example shows Theorem 1.2 is tight.
-Example A.6 (Tight Example on Demand-Approximate CE).
-Consider the case with two agents and one good.
-Both agents have unit budgets.
-The first agent’s utility on this good is linear: $u_{1}(x_{1})=x_{1}$.
-The second agent’s utility is defined as follows: $u_{2}(x_{2})=\alpha x_{2}$ for $x_{2}\leq\frac{1}{\alpha}$ and $u_{2}(x_{2})=1+\beta(x_{2}-\frac{1}{\alpha})$ for $x_{2}>\frac{1}{\alpha}$, where $\alpha$ is a large number and $\beta\leq\alpha$.
-When $\beta=\frac{1-\frac{1}{\alpha}}{1-\frac{2}{\alpha}}$, the Nash welfare maximizing allocation is $x_{1}=1-\frac{1}{\alpha}$ and $x_{2}=\frac{1}{\alpha}$, with the dual variable $q$ equal to $\frac{1}{x_{1}}=\frac{1}{1-\frac{1}{\alpha}}$.
-Under this price, agent $2$ can obtain $x^{\prime}_{2}=1-\frac{1}{\alpha}$ with his/her budget, resulting in a utility of $1+\beta(x^{\prime}_{2}-\frac{1}{\alpha})=2-\frac{1}{\alpha}$.
-In contrast, in the Nash welfare maximizing allocation, agent $2$ only receives a utility of $1$.
-A.6 (In)-approximate Nash Welfare utility by Equilibrium utility
-In this paper, we show for $\Sigma$-Gale-substitues utilities, the utility of any agent at the Nash welfare maximizing allocation is at least half the maximum utility they may get at any CE.
-It’s natural to ask, inversely, if the utility agent can attain at any CE can approximate the utility they receive at Nash welfare maximizing allocation.
+Example E.3 (Non-monotonicity of Gale demands).
+Consider a market with two goods and a single agent with a unit budget, $b=1$. The agent’s utility function is
+$u(\mathbf{x})=\min\{x_{1}+x_{2},2+0.1\cdot x_{1}+0.2\cdot x_{2}\}.$
+Consider the price vector:
+$\mathbf{q}=\left(\frac{0.1}{2/0.9+0.2\cdot 0.01},\frac{0.2}{2/0.9+0.2\cdot 0.01}\right).$
+At this price, the agent’s Gale demand and corresponding utility are
+$\mathcal{GD}(\mathbf{q},b)=\{(2/0.9,0.01)\},\quad u(\mathcal{GD}(\mathbf{q},b))=2/0.9+0.2\cdot 0.01.$
+Now, consider a higher price vector:
+$\mathbf{q}^{\prime}=\left(\frac{1}{0.01+2/0.85},\frac{1}{0.01+2/0.85}\right).$
+Under price $\mathbf{q}^{\prime}$, the agent’s Gale demand and utility become
+$\mathcal{GD}(\mathbf{q}^{\prime},b)=\{(0.01,2/0.85)\}\quad u(\mathcal{GD}(\mathbf{q}^{\prime},b))=2/0.85+0.01.$
+Clearly, $\mathbf{q}<\mathbf{q}^{\prime}$ component-wise, yet we observe that $u(\mathcal{GD}(\mathbf{q},b))<u(\mathcal{GD}(\mathbf{q}^{\prime},b))$.
+It is also worth noting that such a case cannot arise under separable utilities (see the proof of Lemma C.2 for further details).
+E.3 (In)-approximate Nash Welfare utility by Equilibrium utility
+Theorem 1.3 shows that for $\Sigma$-Gale-substitutes utilities, $u_{i}(\mathbf{y}_{i})\geq\frac{1}{2}u_{i}(\mathbf{x}_{i})$ for the Nash welfare maximizing allocation $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$ and CE allocation $(\mathbf{x}_{i})_{i\in{\mathcal{A}}}$.
+One may ask the converse: namely, for any agent $i$, is there a CE $(\mathbf{x}_{i})_{i\in{\mathcal{A}}}$ such that $u_{i}(\mathbf{x}_{i})\geq C\cdot u_{i}(\mathbf{y}_{i})$ for some constant C>0C>0999We note that the utility achieved by an agent in a Nash welfare maximizing allocation is unique, since the logarithm of the Nash welfare is a strictly concave function.?
 We show a constant approximation is not achievable by providing an example in which the CE is unique and there exists an agent whose utility at this CE is much worse than his/her utility at Nash welfare maximizing allocation.
-Example A.7.
-Consider a scenario with $n+2$ agents and $n+1$ goods.
-Set $\epsilon<1$.
-For $1\leq i\leq n$, let $u_{i}(x_{i})=\min\{x_{ii},0.5\}+\epsilon x_{i(n+1)}$.
-Additionally, set $u_{n+1}(x_{n+1})=x_{(n+1)(n+1)}$ and $u_{n+2}(x_{n+2})=\sum_{j=1}^{n}x_{ij}$.
-We first demonstrate the unique competitive equilibrium is achieved when $p_{j}=\frac{2}{n}$ for $1\leq j\leq n$ and $p_{n+1}=n$.
-In this equilibrium, $x_{ii}=0.5$ and $x_{i(n+1)}=\frac{n-1}{n^{2}}$ for $1\leq i\leq n$; $x_{(n+1)(n+1)}=\frac{1}{n}$; and $x_{(n+2)j}=0.5$ for $1\leq j\leq n$.
-We now show this is the only competitive equilibrium.
-First, for all $j:1\leq j\leq n$, $p_{j}>0$.
-Otherwise, agent $n+2$ would receive infinite goods.
-Given $p_{j}>0$ for any $1\leq j\leq n$, $x_{(n+2)j}\geq 0.5$.
-Consequently, the first $n$ goods share the same prices and $10.5p_{j}n\leq\sum_{j}x_{(n+2)j}p_{j}=$ 1 which implies $p_{j}\leq\frac{2}{n}$.
-If $p_{j}<\frac{2}{n}$ for $1\leq j\leq n$, then $p_{n+1}=\sum_{i}b_{i}-\sum_{j}p_{j}>n$.
-In this case, $x_{ii}=0.5$ for $1\leq i\leq n$ and $10.5p_{j}n=$ 1 for $1\leq j\leq n$, which implies the result.
-Therefore, agent $n+1$ receives utility $\frac{1}{n}$ in the unique competitive equilibrium. However, for a small enough $\epsilon$, this agent receives $1$ in the maximal Nash welfare solution.
-A.7 (In)-approximation for general utilities
+Example E.4.
+Consider a market with $n+2$ agents and $n+1$ goods. Let $\varepsilon<1$ be a small constant. For each agent $1\leq i\leq n$, the utility function is given by:
+$u_{i}(\mathbf{x}_{i})=\min\{x_{ii},0.5\}+\varepsilon x_{i(n+1)}.$
+For the remaining two agents, define their utilities as:
+$u_{n+1}(\mathbf{x}_{n+1})=x_{(n+1)(n+1)},\quad\text{and}\quad u_{n+2}(\mathbf{x}_{n+2})=\sum_{j=1}^{n}x_{(n+2)j}\,.$
+We claim that the unique competitive equilibrium occurs with prices:
+$p_{j}=\frac{2}{n}\quad\text{for }1\leq j\leq n,\qquad p_{n+1}=n.$
+At this equilibrium, the allocation is as follows:
+  • for agent $i$ such that $1\leq i\leq n$: $x_{ii}=0.5$ and $x_{i(n+1)}=\frac{n-1}{n^{2}}$;
+  • for agent $n+1$: $x_{(n+1)(n+1)}=\frac{1}{n}$;
+  • for agent $n+2$: $x_{(n+2)j}=0.5$ for $1\leq j\leq n$.
+We now show that this is the only competitive equilibrium.
+First, for all $1\leq j\leq n$, $p_{j}>0$, since otherwise agent $n+2$ could obtain an unbounded amount of good $j$, violating feasibility.
+Given $p_{j}>0$, and since $x_{(n+2)j}\geq 0.5$ for all $1\leq j\leq n$, the first $n$ goods share the same prices, and, additionally, the total spending of agent $n+2$ satisfies: $1=\sum_{j=1}^{n}x_{(n+2)j}p_{j}\geq 0.5\cdot p_{j}\cdot n$, which implies $p_{j}\leq\frac{2}{n}$.
+We now show $p_{j}=\frac{2}{n}$.
+If $p_{j}<\frac{2}{n}$, then, given $\sum_{j}p_{j}=\sum_{i}b_{i}$, $p_{n+1}=\sum_{i}b_{i}-\sum_{j=1}^{n}p_{j}>n$.
+In this case, agents $1\leq i\leq n$ receive $x_{ii}=0.5$, and therefore $x_{(n+2)i}=1-0.5=0.5$ for all $1\leq i\leq n$.
+This implies $0.5\cdot p_{j}\cdot n=1$, which leads back to $p_{j}=\frac{2}{n}$.
+Therefore, this pricing and allocation define the unique competitive equilibrium.
+Finally, observe that agent $n+1$ receives utility $\frac{1}{n}$ in this equilibrium.
+In contrast, under the allocation that maximizes Nash welfare, for sufficiently small $\varepsilon$, agent $n+1$ would receive utility $1$.
+This illustrates that the utility an individual agent receives in a Nash welfare maximizing allocation can be substantially higher than in the corresponding market equilibrium.
+E.4 (In)-approximation for general utilities
 Our next example demonstrates if the utility function is not Gale substitutes, then agents’ competitive equilibrium utilities can be much higher than the Nash welfare maximizing solution.
-Example A.8.
-There are three types of agents (A, B, C): $1$ type-A agent; $n^{2}$ type-B agents; $n$ type-C agents; and four types of goods (O, P, R, S): $1$ type-O good; $n$ type-P goods; $n^{2}$ type-R goods; $n$ type-S goods.
+Example E.5.
+We consider an economy with three types of agents (A, B, C):
+11 type-A agent, $n^{2}$ type-B agents, $n$ type-C agents;
+and four types of goods (O, P, R, S):
+11 type-O good, $n$ type-P goods, $n^{2}$ type-R goods, $n$ type-S goods.
 In total, there are $N=n^{2}+n+1$ agents and $M=n^{2}+2n+1$ goods.
-The type-A agent is interested in the type-O good and type-P goods.
-Type-A agent has a bounded linear utility with equal preferences over these type-O and type-P goods: $v_{ij}=1$ for $i=A$ and $j\in O\cup P$.
-This agent has an upper bound of 1 on the utility of each good: $u_{A}(x_{A})=\sum_{j\in O\cup P}\min\{x_{Aj},1\}$.
-For type-B agents, let’s label each of these agents by $(s,t)$ for $s\in[1,n]$ and $t\in[1,n]$ and also label each of type-R goods by $(s,t)$ for $s\in[1,n]$ and $t\in[1,n]$.
-Agent $(s,t)$ is interested in the corresponding $(s,t)$ good in R and the $s$-th good in P.
-The utility function of agent $i=(s,t)$ is $\min\{x_{iP_{s}}+\frac{n}{n+1},x_{iR_{s,t}}\}$, where $x_{iP_{s}}$ is agent $i$’s allocation of $P_{s}$ and $x_{iR_{s,t}}$ is agent $i$’s allocation of $R_{s,t}$.
-For type-C agents, the $s$-th type-C agent is only interested in the $s$-th type-S good and all type-R good with label $(s,\cdot)$.
-This agent also has a bounded linear utility with equal preference ($v_{ij}=1$) over these goods and an upper bound, $l_{ij}=1$, on the utility of each good.
-We first consider the maximal Nash welfare solution. Here, $k$ is a parameter, which will be specified later.
-The type-A agent receives type-O good and $k+\frac{1}{n+1}$ portion of type-P goods, his/her utility will be $1+nk+\frac{n}{n+1}$.
+The type-A agent is interested in the type-O good and type-P goods with equal preferences: $v_{ij}=1$ for $i=A$ and $j\in O\cup P$.
+However, the utility from each good is capped at $1$, so the utility is given by:
+$u_{A}(\mathbf{x}_{A})=\sum_{j\in O\cup P}\min\{x_{Aj},1\}.$
+The type-B agents are indexed by pairs $(s,t)$ where $s,t\in[n]$, for a total of $n^{2}$ agents. Similarly, the type-R goods are indexed by $(s,t)$. Each type-B agent $(s,t)$ is interested in
+  • the $s$-th type-P good, and
+  • the type-R good with index $(s,t)$.
+The utility function of type-B agent with index $i=(s,t)$ is
+$u_{i}(\mathbf{x}_{i})=\min\{x_{iP_{s}}+\frac{n}{n+1},x_{iR_{s,t}}\},$
+where $x_{iP_{s}}$ is agent $i$’s allocation of the $s$-th type-P good and $x_{iR_{s,t}}$ is agent $i$’s allocation of the type-R good with index $(s,t)$.
+The type-C agents are indexed by $s\in[n]$.
+Each type-C agent with index $s$ is only interested in the $s$-th type-S good and all type-R goods with index $(s,\cdot)$, and the utility from each good is capped at $1$.
+The utility function type-C agent with index $i=s$ is
+$u_{i}(\mathbf{x}_{i})=\sum_{j\in{S_{s}}\cup{R_{(s,\cdot)}}}\min\{x_{ij},1\},$
+where $S_{s}$ represents the $s$-th type-S good and $R_{(s,\cdot)}$ represents the set of type-R goods with index $(s,\cdot)$.
+The maximal Nash Welfare solution
+Let $k$ be a parameter, which will be specified later.
+The type-A agent receives the type-O good and $k+\frac{1}{n+1}$ portion of type-P goods.
+This results in a total utility of $1+nk+\frac{n}{n+1}$.
 Each type-B agent obtains $1-\frac{k}{n}$ portion of corresponding type-R good and $\frac{1}{n+1}-\frac{k}{n}$ portion of corresponding type-P good, yielding a utility of $1-\frac{k}{n}$.
@@
 The goal is to maximize $(1+nk+\frac{n}{n+1})(1-\frac{k}{n})^{n^{2}}(1+k)^{n}$.
-By calculation, the optimal $k$ is found to be no larger than $n^{-\frac{1}{2}}$.
+By calculation, the optimal $k$ is at most $n^{-\frac{1}{2}}$.
 [^10] First, we take the log and calculate the gradient, which is $\frac{n}{1+nk+\frac{n}{n+1}}+\frac{n}{1+k}-\frac{n}{1-\frac{k}{n}}$.
@@
 Therefore, the type-A agent receives at most $1+\sqrt{n}+\frac{n}{n+1}$ utility in the maximal Nash welfare solution.
-On the other hand, let’s consider the (thrifty) competitive equilibrium where all agents have unit budgets.
-The prices of type-O, P, and S goods are 00.
-and prices of type-R goods are $\frac{n+1}{n}$.
+Thrifty competitive equilibrium
+At the thrifty competitive equilibrium, the prices of type-O, P, and S goods are 0, while the prices of type-R goods are $\frac{n+1}{n}$.
 In this case, the type-A agent will get the type-O good and type-P goods, resulting in a utility of $n+1$.
@@
 Comparing the type-A agent’s utility in the two scenarios, the type-A agent achieves at most $1+\sqrt{n}+\frac{n}{n+1}$ utility in the maximal Nash welfare solution, whereas the agent attains $n+1$ utility in the (thrifty) competitive equilibrium.
-A.8 Tight Example of Theorem 1.3
-The following example shows the bound in Theorem 1.3 is tight.
-Example A.9.
-Consider the case that there are $n$ agents and $n$ goods.
-All agents have bounded linear utilities.
-Additionally, all agents have the same preferences ($v_{ij}=1$) on all goods except for the first good.
-For the first good, only the first agent likes the first good ($v_{11}=0$) and all other agents don’t like this good $v_{i1}=0$ for $i\neq 1$.
-Additionally, we assume the upper bound on the utility of each good of each agent is $1$, $l_{ij}=1$ for all $i$ and $j$.
-In this case, for the maximal Nash welfare solution, the utilities of all agents equal $1$.
-However, there exists a competitive equilibrium, at which the price of the first good equals 00 and the prices for all other goods equal $\frac{n}{n-1}$.
-Then, the first agent can get the first good and $\frac{n-1}{n}$ fraction of the second good, and achieve a utility of $1+\frac{n-1}{n}$ which tends to $2$ in the case $n\rightarrow\infty$.
-A.9 Example with Leontief-free utility functions
+E.5 Example with Leontief-free utility functions
 Not all Leontief-free utility functions satisfy property ($\Sigma$).
 The subsequent example illustrates a scenario where an agent can attain significantly more utility in competitive equilibrium compared to the Nash welfare maximizing allocation when agents have Leontief-free utilities.
-Example A.10.
+Example E.6.
 Consider the scenario with $n$ goods and $n+1$ agents.
-Agent $i=1,2,\cdots,n$ only likes good $i$, has budget $b_{i}=1$ and a linear utility function $u_{i}(x_{i})=x_{ii}$.
-Agent $n+1$ has a substantially larger budget $b_{k+1}=B$, and a utility function $\min\{\sum_{i=1}^{k}x_{(n+1)i},(n-1)(1-\epsilon)\}$ with a small $\epsilon>0$.
-There is a competitive equilibrium where the prices are $p_{1}=1$ and $p_{2}=p_{3}=\cdot=p_{n}=\frac{1}{\epsilon}$.
-Agent $1$ obtains $x_{11}=1$, the other agents get $x_{ii}=\epsilon$ for $i=2,\cdots,n$, and agent $n+1$ gets $x_{(n+1)1}=0$ and $x_{(n+1)j}=1-\epsilon$ for $j=2,\cdots,n$.
-The competitive equilibrium holds if $B>\frac{(n-1)(1-\epsilon)}{\epsilon}$.
-On the other hand, in the Nash welfare maximizing allocation, when $B$ is big enough, agent $n+1$ receives (approximately) $\frac{(1-\epsilon)(n-1)}{n}$ of each good, with the remaining goods shared among the remaining agents.
-Agent $1$ receives a utility of $\epsilon+\frac{1-\epsilon}{n}$, which is significantly less than the utility received in the competitive equilibrium.
-In this case, agent $n+1$’s utility doesn’t satisfy property ($\Sigma$).
-When $\mathbf{p}=\mathbf{0}$, $x_{n+1}=(0,1-\epsilon,1-\epsilon,\cdots,1-\epsilon)\in\mathcal{GD}_{n+1}(\mathbf{p},B)$.
-However, when $\mathbf{p}=(1,\frac{1}{\epsilon},\frac{1}{\epsilon},\cdots,\frac{1}{\epsilon})$, for any $x_{n+1}\in\mathcal{GD}_{n+1}(\mathbf{p},B)$, $x_{(n+1)1}>0$, which violates property ($\Sigma$).
-A.10 Example with matching utility functions
-The following example illustrates that matching utilities are not Gale-substitutes in general.
-Example A.11.
-Consider the scenario with $3$ goods and an agent with unit budget.
-The agent’s utility is a matching utilities, with preferences for the goods being $1$, $2.5$, and $0.7$ respectively.
-The initial prices for the goods are $0.3$, $1.2$, and $0.1$.
+Agent $i=1,2,\cdots,n$ only likes good $i$, has budget $b_{i}=1$ and a linear utility function $u_{i}(\mathbf{x}_{i})=x_{ii}$.
+Agent $n+1$ has a substantially larger budget $b_{n+1}=B$, and a utility function $\min\{\sum_{j=1}^{n}x_{(n+1)j},(n-1)(1-\varepsilon)\}$ with a small $\varepsilon>0$.
+There is a competitive equilibrium where the prices are
+$p_{1}=1\text{ and }p_{2}=p_{3}=\cdot=p_{n}=\frac{1}{\varepsilon}\,.$
+Agent $1$ obtains $x_{11}=1$, the other agents get $x_{ii}=\varepsilon$ for $i=2,\cdots,n$, and agent $n+1$ gets $x_{(n+1)1}=0$ and $x_{(n+1)j}=1-\varepsilon$ for $j=2,\cdots,n$.
+The competitive equilibrium holds if $B>\frac{(n-1)(1-\varepsilon)}{\varepsilon}$.
+On the other hand, in the Nash welfare maximizing allocation, when $B$ is big enough, agent $n+1$ receives (approximately) $\frac{(1-\varepsilon)(n-1)}{n}$ of each good, with the remaining goods shared among the remaining agents.
+Agent $1$ receives a utility of $\varepsilon+\frac{1-\varepsilon}{n}$, which is significantly less than the utility received in the competitive equilibrium.
+In this case, agent $n+1$’s utility doesn’t satisfy property ($\Sigma$). When $\mathbf{p}=\mathbf{0}$,
+$\mathbf{x}_{n+1}=(0,1-\varepsilon,1-\varepsilon,\cdots,1-\varepsilon)\in\mathcal{GD}_{n+1}(\mathbf{p},B)\,.$
+However, when $\mathbf{p}=(1,\frac{1}{\varepsilon},\frac{1}{\varepsilon},\cdots,\frac{1}{\varepsilon})$,
+$\text{for any }\mathbf{x}_{n+1}\in\mathcal{GD}_{n+1}(\mathbf{p},B),\ \ x_{(n+1)1}>0,$
+which violates property ($\Sigma$).
+Observe that we can make the budgets of all agents the same by creating $B$ copies of agent $n+1$, each with the utility function $\min\{\sum_{j=1}^{n}x_{(n+1)j},(n-1)(1-\varepsilon)/B\}$.
+E.6 Example with matching utility functions
+The following example shows that matching utility functions (i.e., those with a unit-demand constraint) are not Gale-substitutes in general.
+This also implies that utility functions with more general constraints (as in Garg et al.
+(2022b), Jalota et al.
+(2023)) need not satisfy the Gale-substitutes property either.
+Example E.7.
+Consider a market with $3$ goods and an agent with unit budget.
+The agent has a matching utility function (i.e., a unit-demand constraint), with valuations of $1$, $2.5$, and $0.7$ for goods 1, 2, and 3, respectively.
+The initial prices vector is $(0.3,1.2,0.1)$.
 In this case, the Gale demand is
-${\arg\max}_{x=(x_{1},x_{2},x_{3}):x_{1}+x_{2}+x_{3}\leq 1}\{\log(x_{1}+2.5x_{2}+0.7x_{3})-(0.3x_{1}+1.2x_{2}+0.1x_{3})\},$
-which gives the solution $(0,103/198,95/198)$.
-Now we keep the prices of the first two goods unchanged and increase the price of the third good from $0.1$ to $0.6$.
-By calculation, under the new prices, the new Gale demand is $(5/9,4/9,0)$.
-The Gale demand of good $2$ decreases, which implies that it is not substitutes.
-Appendix B Proof of Lemma 8.1
-See 8.1
-Proof.
-First, assume that $\pi\in\mathbb{R}_{\geq 0}^{\ell}$ is an optimal solution to $\operatorname{Dual}(x)$ for a given $x\in\mathbb{R}^{m}$.
-We need to show that $g=B^{\top}\pi$ is a supergradient at $x$, that is, $\varphi(t)\leq\varphi(x)+\left\langle g,t-x\right\rangle$ for any $t\in\mathcal{D}$.
-The programs $\operatorname{Dual}(x)$ and $\operatorname{Dual}(t)$ have the same feasible region, with linear objective functions $\mathbf{B}x+b$ and $\mathbf{B}t+b$, respectively.
-By weak duality, the cost of every feasible dual solution is at least the primal optimum value, thus, $\varphi(t)\leq t^{\top}\mathbf{B}\pi+b^{\top}\pi=g^{\top}t+b^{\top}\pi$.
-The statement follows as
-$\varphi(t)\leq g^{\top}t+b^{\top}\pi=g^{\top}(t-x)+x^{\top}\mathbf{B}\pi+b^{\top}\pi=\varphi(x)+g^{\top}(t-x)\,,$
-noting also that $\varphi(x)=x^{\top}\mathbf{B}\pi+b^{\top}\pi$ since $\pi$ is an optimal solution to $\operatorname{Dual}(x)$, and $\varphi(x)$ is the optimum value.
-Conversely, we need to show that if $g\in\partial^{\wedge}\varphi(x)$, then there exists a dual optimal solution $\pi$ such that $g=\mathbf{B}^{\top}\pi$.
-Let $z$ be a primal optimal solution to $\operatorname{Primal}(x)$.
-Using complementary slackness, we can turn $\operatorname{Dual}(x)$ into a feasibility system.
-Namely, $g=\mathbf{B}^{\top}\pi$ for a dual optimal solution if and only if
-$\mathbf{B}^{\top}\pi$ $=g$ (12)
-$\mathbf{A}_{i}^{\top}\pi$ $=c_{i}\,,\quad\forall i:z_{i}>0$
-$\mathbf{A}_{i}^{\top}\pi$ $\geq c_{i}\,,\quad\forall i:z_{i}=0$
-$\pi_{j}$ $=0\,,\quad\forall j:(\mathbf{A}z)_{j}<(\mathbf{B}x+b)_{j}$
-$\pi_{j}$ $\geq 0,\quad\forall j:(\mathbf{A}z)_{j}=(\mathbf{B}x+b)_{j}\,.$
-Here, $\mathbf{A}_{i}$ denotes the $i$-th column of $\mathbf{A}$. By Farkas’s lemma, if (12) is not feasible, then the following system has a feasible solution:
-$(\mathbf{A}w)_{j}+(\mathbf{B}r)_{j}$ $\leq 0\,,\quad\forall j:(\mathbf{A}z)_{j}=(\mathbf{B}x+b)_{j}$
-$w_{i}$ $\geq 0\,,\quad\forall i:z_{i}=0$
-$c^{\top}w+g^{\top}r$ $>0\,.$
-For a small enough $\varepsilon>0$, $z+\varepsilon w$ is feasible to the system $\operatorname{Primal}(x-\varepsilon r)$.
-Hence, for $t=x-\varepsilon r$, $\varphi(t)\geq c^{\top}(z+\varepsilon w)=c^{\top}z+\varepsilon c^{\top}w$.
-Using $c^{\top}z=\varphi(x)$ by the choice of $z$ as an optimal solution to $\operatorname{Primal}(x)$, and the bound $c^{\top}w>-g^{\top}r$, we get $\varphi(t)>\varphi(x)-\varepsilon g^{\top}r=\varphi(x)+\left\langle g,t-x\right\rangle$, a contradiction to $g\in\partial^{\wedge}\varphi(x)$.
-∎
+${\arg\max}_{\mathbf{x}:x_{1}+x_{2}+x_{3}\leq 1}\{\log(x_{1}+2.5x_{2}+0.7x_{3})-(0.3x_{1}+1.2x_{2}+0.1x_{3})\},$
+which yields the solution $(0,103/198,95/198)$.
+Now, keeping the prices of the first two goods unchanged, we increase the price of the third good from $0.1$ to $0.6$.
+Under these new prices, the Gale demand becomes $(5/9,4/9,0)$.
+Since the demand for good $2$ decreases as the price of good $3$ increases, this shows that the utility function does not satisfy the Gale-substitute property.